A Complete Bibliography of the Journal of Business & Economic Statistics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 October 2023 Version 1.06

Title word cross-reference

(0, 1, 1)_{12} [Bel87]. 1/2 [GI11]. 4 [Fre85]. α [McC97]. F(d) [CdM19]. ℓ_{1} [PX21]. F [WS20]. k [PW22, SW20]. L [BC11]. M [TZ19]. N [AW07, CSZZ17, GK14]. ⊖ = 1 [Bel87]. P [Han97, Ray90, SU22]. R [CW96]. R^{2} [MRW19]. T [AW07, GK14, KW22, RSW18, FSC03, IM10, JS10, MV91, NS96, TM19]. U [TZ19].

-Distributed [WS20]. -Estimators [TZ19]. -Means [PW22]. -Penalized [PX21]. -Processes [TZ19]. -R [Fre85]. -Squared [CW96]. -Statistic [IM10, TM19]. -Statistics [BC11]. -values [SU22].

100 [ACW97]. 11 [McK84, dBM90]. 19 [KNW22].

225 [ACW97]. 2SLS [Lee18].

500 [ACW97].

a.m [HRV20]. Ability [CM09, Han05, HD21, Son12, SU22, dNS23]. Abnormal [MlM87]. Absorbing [KK15]. Abuse [Die15a]. Accelerated [KLS15]. Access [AKO12]. Accessibility [FM97]. Accident [GFdH87]. Accommodating [KDM20]. Accountability [BDJ12, LPS22]. Accounting [HPSW88, HW84, MPTK91]. Accounts [Che12]. Accumulated [DM01]. Accumulation [Cag03, DM01]. Accuracies [Fom86]. Accuracy [BP22, DM95a, DM02, Die15a, McN86a, NP90, PG93, SW22b, Wec89]. Accurate [McK18]. Achievement [GRTZ22]. Acidic [PR87]. Acknowledgment [Ano93a]. Acreages [HF87]. Across [PG93, PP11]. Act [Fre85]. Action [GW85]. Active [ZCZ^{+}21]. Activities [AC03, But85]. Activity [FS93, JLR11, Kol20, Li11, RS94b, ZCZ^{+}21]. Actually [CO19]. Adaptive [BZ20, CZ14, CHL18, CNT20, CL17, HHK11, HF19, Tsa93b, ZS15]. Addiction [BG01]. Additional [LZ95]. Additive [Abe99, FZ18, FXMFY20, FH94, Lia12, OHS14]. Additively [Koz20]. Addressing [Mel18]. Adequacy [HM89]. Adjust [Nev03]. Adjusted [AW84, LD19, MS03, NTP94, Rad83, Rub86, Son10]. Adjusting [vdHH86]. Adjustment [BH84a, BH02, BCP19, BW84, CW86b, DM84, DL87, EG98, ES01a, FMB^{+}98a, Ghy90, GGS96a, Ghy97, HR91, HKR97, Hil85, Jai89, Jor99, Lin86, LHM20, McE17, ML90, McK84, Pat86, Pfe91a, Pie83, Pie84, PGC84, Sim85, Tay02, dBM90]. Adjustment-Models [Pie84]. Adjustments [DFPS11, Lit88a, OV84, Ver12]. Administrative [But85, CMW23, Car85, HT20, JS85, MAB^{+}23, Nor85, Wai85, ZW93, ZHB23]. Administrative-List [ZW93]. Advance [CHLR92]. Advertising [GKP00]. AFDC [GOV84]. Affect [CLR14, HW04]. Affecting [NG89]. Affects [RS94b]. Affine [CDJW10, IRS15, dJ00]. Affluence [Dag87]. Afriat [FW05]. Africa [YSD03]. After [Gra19, Kim01, Lec99, McC19, MA23, Mel90]. Against [BSS01, Hal90, LM00, LLW90, Qu11, SK91, TS01]. Age [DMS99]. Agencies [GB87, SvD06]. Agency [CM93]. Agglomeration [CRZ19]. Aggregate [BS87a, BC01, BM93, CO19, Che12, EH90, Haa96, Har83, HH11b, Jac00, JS87, MP83, PPL94, RDP95, Sle98, SW90]. Aggregated [Erm91, Lüt84]. Aggregates [Aiz90, GKMV09, McG95, Pie83, PGC84]. Aggregation [AR91, CL87, DW85, Jor99, JSS88, Mar99, Nic89, RS95]. Aging [Ran88]. Agricultural [BFR97, CHL88]. Agriculture [Gor89]. Ahead [Knü15, Ott90, Tau93b, ZH92]. Ahn [KT18]. Aid [GOV84, PPT92]. AIM [KOS94a]. Ain’t [Mor87]. Air [LA84]. Airport [HHPS07]. Alan [Pel21]. Alcoholic [AES86]. Algebra [Gre84]. Algorithm [CILY17, DS98, HSA87, KJGV09, MM03, WHW22]. Algorithms [DM95b, GK96, Kor19, Sil01]. Allocation [AW00, BK11, Blo19, EC06, FFX16, Joh87, LHYZ22]. Allocations [Wri83]. Allowing [Wer99]. Almost [Cha87, IP12, Ram94]. Alpha [Fri22]. Alpha-Stable [Fri22]. Alphas [MLST20]. Alternative [BSS01, BN12, Cal91, Can92, DL92, DMMN83, Fom86, Gar89, GH22, Gbu89, HHY96, Har91, HO84, KWD13, McK18, McN86a, Oom06b, RS94a, Sen03, SB89a, TU90, WTC84, WW96, Wri00a]. Alternatives [Hal90]. Amazon [PSD14]. Ambiguity [TV15]. American [Rog87]. Amount [RFG20]. Amplitude [vdHH86]. Analysis [AL84, And00, Ano84c, ACL19, Bäu13, BR89, BO01, BZ99, Bie00, BH84c, BCG85, CKY85, CST18, CM88, CL93, CW01, CdH96, Das08, DFPS11, DM16, DH88, DV04, Era01, EHM98, Eri98, FLSVP04, FM90, FKK99, FL98, FG00, FSTO04, GFdH87, GSC03, GH91, Goi85, Goo93, GK03, GJP02, HH21, Hag83, HM89, HL16, HP11, HH19a, Har91, HK01, HKYZ20, HPZZ21, Hof91, Hor98, HMI95, HT14a, HLM16, HS89, IRS15, JPR94a, JPR02, Joc18, JJR18, KS98, Koh83a, Koo91, KOS94a, KS94, KP03, KS96a, LeS91, LS94, LW09, LLPZ20, Li11, LSCX22, LQZ19, LLCL22, Lin86, LLL19, LdV99, LSS16, MN88, MK01, Mar99, MM20, ML88, McG95, MT09, Mor92, Mou87, MN20, NWK10, NW13, NP97, NSR88, NG89, Nef93, NP85, Nie07, OV84, Ott90, PRM99, PJ20]. Analysis [PRF08, PT00, RW86a, RT01, Rus19, She92, TGCD16, Tsa00, VP92, VZ11, WlMCP98, Wan03, WJS08, Wat00, kWM90, WLZT22, ZH92, ZHCB16, ZL22, ZPWW22, dA88, vHPP04]. Analytic [BC91]. Analytical [Bra07]. Analyzing [FB97, KMV10, MSW83, Mal83, OS20, Tsa16]. Anatomy [BS98]. and/or [IP12]. Andrews [Bon20, Tam20, NS10]. Anniversary [GH02b, GH02a]. Announcement [And07, Ano86a, Ano86b, Ano87a, Ano87b, Ano97a, GH04a, TW98, Woo98, Woo01]. Announcements [Ano84a, Ano86c, BS87b, HRV20, IOS90, LeS92, RW96]. Annual [SW88]. Anomalies [Rav06]. Ante [KK15, Cle14]. Antipoverty [JR03]. Apple [PSD14]. Appliances [LA84]. Application [Agu10, ACL19, ACD09, BGS22, BK05, BCCP18, BL05, BBC04, BCHK16, BB05, Bey01, BK91, Bie00, BL88b, BBCL20, BV05, Cal91, CT16, CsL94, Cha15, CF99, CLLM03, CHLR92, Das08, DHS19, DP22, EN94, Era01, FK86, FS91, FLM^{+}19, FHDS11, FB99, Fuh92, GLL93, GZ84, HMS17, HS16, HP10, HPP00, HLR^{+}22, Hos18, JLR11, Kal22, KW19, KLGS10, Kum87, LD19, LS20b, LSCX22, LLCL22, LM94b, Lou07, LSS16, MMG^{+}06, Mar83, MLR19, Mor85, OV84, Ott90, PRC84, PV92a, PB91, PFS98, Pro08, Ram94, Rit86, RS94b, San17, SL20, She19, Sta08, Sto90, SMU13, Tar07, Wan96, XZL20, Xie22, ZH92, ZLY17, dBM90, vdKK03]. Applications [ABS99, AAMW20, App85, Ber01, BRT12, CT93, CW96, CNR20, CI03, DS13, FZ13, FMT10, GW14, GGW14, HH19b, HLW14, KLL20, Kak84a, LMS09, MQZ22, MOT22, MT09, Tsa93c, XP11, YF22]. Applied [ACL11, CPP18, JVC94, Jar87, KW97]. Approach [AGJT14, AJ05, AWX20, AG05, Ant83, AFO21, ABD19, BH94, BK15, BFS94, BCR90, BHL92, BH09b, Blo19, Bru92, BH93, Byk23, CCÖU19, CGL20, Can92, CK22, Che18, CFM11, CT18, CI87, CHL88, CDF15, CHNY22, CEJ^{+}12, Das91, DIW96, Due99, FS22, Fis92, FB97, FH98, GFdH87, GG90, Gor89, GMP85, Gre89, GR12, GHY92, GL13, GW88, Ham91b, HHK19, HW90, Hil86, Ioa95, JLS16, KKB94, Kli22, KYK^{+}21, KR88, Kot84, KGM91, LCZT22, LeS91, LM92, Li99, LT00, LSX15, LTY^{+}22, Lit86c, MRW12, MHB19, NP86, PFA00, PR98, PT11, Phi87, PGC84, RS92, Rat12, Sel89, She93, SZ20, Shr89, SW95, TK10, Tay19, TM86, TH89, TA89, TMH^{+}23, Tse89, kTY12, Val88, WlMCP98, Wes15, ZZZ20]. Approaches [ABS99, Bel04, WTC84]. Appropriate [EN94]. Approximate [CH14, Han97, Kap10, Mac94, PPT92]. Approximately [AC94, WZ20a]. Approximation [BBK12, Chr90a, Chr90b, FS97, McG90]. Approximations [CT90, NP86, SS06]. APT [BDW96]. Arbitrage [BFZ98a, CG00, FKK99, HPS96, MB88]. Arbitrage-Free [BFZ98a]. Arbitrage-Pricing [HPS96]. Arbitration [Swa03b]. ARCH [BH97, Cai94, EGR91, HS99b, LK93, LT00, LTY19, Wei86]. Area [FMT97, Gra97, Isa90, LSZ14, PB91, VKR06]. Areas [PFS98]. Argentina [Mel90]. ARIMA [CT90, FMB^{+}98a, FP01, Gri93, Kan86, KP95, NP90, SOK01, kWM90]. ARIMA-Model-Based [FP01]. ARIMA-Noise [kWM90]. Arising [FG00]. ARMA [Hog19, KK15, Lüt84, Lüt86]. ARMA-GARCH [Hog19]. Aroma [MT03]. Around-the-Clock [MKL07]. Arrival [BMP10]. Arrivals [BJ99]. Arriving [CCP13]. Article [LB16, Mil84]. Articles [Ano83b, Ano84e]. Artificial [SW95]. ASA [Ano83c]. Asaf [Pel21]. Ascending [BGG22]. Assess [NG89, PV22]. Assessing [ACD09, BMM20, Cle17, DH89, JLG04, SW95, Wec89, WLL18]. Assessment [GOV84, Ker85, PB91, Pre95, TV15]. Asset [AGU04, BK11, BHMR19, BCP19, Bur94, CZ19, DG96, EC06, EH19, FOTY20, GSA04, GM20, GR12, GMS87, HS96a, HS96b, HV03, Ing90a, Jac00, Lux08, MV18, PRPS22, Var16, Yog08]. Asset-Pricing [Bur94, GMS87, HS96a, HV03]. Assets [DTU06, DU12, GL16, MLST20, MS18, Mor93, Ser88b]. Assignment [FJ08, GRTZ22]. Assistants [Ano93a]. Assisted [ST22b, XZW22]. Associate [Ano93a, Ham94, Chr96, Eri98]. Associated [HS93]. Assumption [DHL14, RH04, vdKK03]. Assumptions [Cal91, HW04, JJK20, Pie84]. AST [GdXPE16]. Asymmetric [BGG22, EG98, HS96b, IRR22, NZ20, SV16, SS13, Tay19]. Asymmetries [Bru92, CK03, KV13, Knü09, KP99]. Asymmetry [ACJK22, FB99, LS20c, SL01, SvD06]. Asymptotic [GHR87, HK14, Han97, HJNR98, KD00, Mac94, MNW19, MV18, Pan91, SW21, TM19]. Asymptotically [FS97, JL22, McC19, WS20]. Asymptotics [Yan20b]. Asynchronous [BBCL20]. ATT [DHL14]. Attainment [Jef08, Rou95]. Attendance [FLM^{+}19]. Attenuation [AKLS20]. Attrition [HR19, RW86b]. Auction [AK13, AGV19, LW18, WJS08]. Auctioning [ELLV97]. Auctions [BGG22, CCP13, CE08, ELLV97, LPV03, Rav06, WV11]. Audit [Pre95, TH83]. Auditing [BGR93, Cla94, Sto90, WNG89]. Augmented [BG23, CL95, DS13, JS20, Nor99, Phi10, QZL15]. Austral [Mel90]. Australia [DHZ10, dBE98]. Australian [AV07, TD84]. Author [Ano83b, Ano84e, Ano93b, TR93]. authorship [Don22, JJKL21, JJKL22, LC22, MLZ22, ZK22]. Auto [KM08].

Autocontours [GRSY11]. Autocorrelated [LS02, Mül14a]. Autocorrelation [BHL92, Cha20b, ZTCW15]. Autocorrelations [Ali84, Ali87, PFS98, Zha15]. Automatic [ELZ13]. Automobile [Mor85, OG86, TN92]. Automotive [BM93]. Autoregression [KMR19, LL19, LLL22, Lob03, NSS07, PR98, Ray90, Sal95]. Autoregressions [BB20, CNR20, Cla11, ELZ13, Kew20, KD00, Knü09, LS02, LL10, Lit86b, LS22, PY17, Run87b, Run02]. Autoregressive [AC93, ABJT22, Ali84, ACR19, AT14, AC94, BL05, BCD17, BG96, CS86, CL17, DL20, DP87, DP02, Gre89, HL16, HT22, Hos18, HF19, Kim01, LYY22, LLZ16, LZLL17, LW19a, LGG22, LTvD03, LRS99, MG84, MT06, NP97, NS05, OS89, RF01, SCS20, SK91, Sta08, SQ17, TT06, Tsa89, TT02, kTY12, Tuz22, YL17, YLN17, ZK05, ZLWW20b, ZLY17]. Autoregressive-Moving-Average [Ali84]. Auxiliary [GdXPE16, HF87, Nev03]. Auxiliary-to-Study [GdXPE16]. Available [Nev03]. Average [AI11, AHL15, Ali84, BCD17, BCFFL19, BW22, CS05, DL87, DW93, DBS84, FHLZ20, FFL13, Gos02, GN15, GIR20, GS03, HS96a, Hei22, HW20a, LTU15, LRW09, SK91, Sta08, ST22b, TT06, Tsa89, ACD09]. Averaging [CG19, DL20, FLTS19, FLYZ20, KLP08, KW19, RLYZ22, SL20, YLL19, ZWZ12, ZWZL22, ZZWZ23]. Aversion [ACJK22, NRW01, Web00, ZO04]. Avoidance [Paa88]. Award [Ano00c, Ano01, Ano02d, Ano03a, Ano03b, Ano04b, Ano04c, Ano05c, Ano05d, Ano06b, Ano06c, Ano07c, Ano07d, Ano08c, Ano08d, Ano09c, Ano10c, Ano11c, Ano13c, Ano13d, Ano15c, Ano96a, Woo98]. Awareness [YLYJ22]. Axiom [MRZ92a]. Axiomatic [CDF15]. Axioms [Mel05].

B [TTW95]. Back [BGR15]. Background [Des07]. Backlog [OV84]. Backsolving [Ing90b, Sim90]. Backtesting [EO10]. Bagged [KP22]. Balanced [Che12]. Balancing [FIL^{+}23, Hei22]. Band [LZZ20]. Bands [LY18]. Bandwidth [CGZ19, LLR13, SL11]. Bandwidths [Yan20b]. Bang [MN95a]. Bank [AGO^{+}14a, CLR14, DS90, Due99, FB99, LSS19, Yil03, vUS87, AGO^{+}14a]. Bank-Deposit-Rate [DS90]. Banking [ABS99, GPP98, KBZ89, WW09]. Bankruptcy [San17]. Bargaining [ELLV97, FB97, FPAB02]. Bargaining-Comparing [FPAB02]. Baroett [BH94]. Barseghyan [CMM23, Gua23, Hon23, Mor23]. Baseball [Kro83]. Based [Ant83, AKKL21, AMRT22, BO01, BC11, BCS08, BFZ98b, BOP23, BRT12, BJ06, BS08, BE96, BT01, CCKR17, CF02, Che07a, Cla94, CGT14, DF11, DM01, DL20, DHS19, FFX16, FHWZ18, FP01, FW03, FSK08, GRW84, Gos02, GK03, GMS87, GSS07, Hal94b, HP13, Hil85, Hog22, IM10, JMZ18, JS10, KL90, Kew20, KD00, Knü09, Kot85, Lab90, LK93, LS15, LW18, MPTK91, MABF08, MB13, Nie04, NP86, NEV15, PT12a, PHPY20, PF19, PGC84, QW20, ST89, SK99, SRdCFB18, SX20, Tay02, Tay19, WHKCS14, Wri00c, XHL20, HSS20, LQZ19, TXLT22]. Basic [DM84, Rei99]. Baskets [CG00]. Bass [BF05]. Battlefront [Poi89, Poi92]. Bayes [AES86, AC93, All90, FMT97, GK15, GGW14, KW04b, LMN22a, MHB19, PV03, Wan96]. Bayesian [AW00, BW15, BFS94, BW19, BGR15, BCIK22, BS08, BN12, CCM16, CST18, CI03, Cha20a, CG19, CGZ19, CT18, Cla11, DW93, DIW96, DL20, FLSVP04, FS19, FB97, FKK99, FSTO04, GCC11, GK08, GSC96, GM20, GM83, HHL18, Ham91b, HL16, HK01, HHKO22, Hil86, HMI95, HF19, JPR94a, JPR02, Joh87, KWD13, KLP08, KJGV09, KS98, KK15, Kli22, KYK^{+}21, Koc10, Koo91, KOS94a, KP03, KPS13, LeS91, LS87, LM90, Li11, Lit86b, Liu21, LYZ22, LMN22b, LNS11, LdV99, NW13, NP97, NS05, NSS07, Oli87, PFA00, PR98, PRF08, Poi91a, PT00, Qia22, RAL99, SK20, SL99, Str03, Str22, Swa03b, TK10, TGCD16, TM86, VP92, VDP00, WZ00, Wat00, WV11, WHKCS14, WLL18, dA88]. Be [CO99, GI19, JSSvN20, LMT96]. Bear [MM00, MMS12]. Beef [FB92]. Beef-Cattle [FB92]. Before [MA23, Mel90]. Behavin [Mor87]. Behavior [Agu10, DP06, Fuh92, Gew88b, HP89, JS87, Kea97, KD08, Lie98, Luc00, MS97, Mis95, ME04, PSD14, RS94a, TM19, Wat00, dNS23]. Behavioral [Cal91, CMHM19, HW04, LLCL22]. Being [AJL96, MN88, OS20]. Beliefs [DGJ20, McK18]. Bell [Sim85]. Benchmarking [She92, TH89]. Benefit [JR03, LZ95, McC95]. Benefits [MVVZ19]. Benford [BCCP18]. Best [EKM87, KS08]. Beta [BR08, LS87]. Beta-Binomial [LS87]. Betas [Kal22]. Better [Ste87]. Between [APW20, BG93, BHL92, Coc89, Dag87, DZ04, DTU06, DMMN84, DU12, Gas85, Kim93, KR90, KP95, LT19, Met98, MS97, NV12, OF97, PRW06, PW10, PPL94, RF15, Rot15, Ser88a, SH07, Vin85a]. Between-Group [Rot15]. Beverages [AES86]. Beyond [BDJ12, DP13, McG95, Nor99, Pow22, Wil09]. Bias [AI11, All90, AR91, AS96a, AT17, BRW12, BH09b, BCF94, BG05, Bra07, BC05, CGJ19, DE19, DW85, DF20, FS88, GIR19, HS96c, HI01, Koh83a, MNO85, Ran88, RS09, Sol86]. Bias-Corrected [AI11, BC05, DE19]. Biases [KLN85, Kot84]. Bidder [ACJK22]. Bidders [BGG22, CCP13]. Big [BFT16, Tsa16, MN95a]. Bilinear [Wei86]. Bilinearity [BH97]. Bimodal [SSBD15]. Binary [Bel04, Car01, CSZZ17, GLL93, GK96, HP11, LNB10, Mal83, MMO22, NWK10]. Binary-Choice [GLL93]. Binomial [LS87, Sta08]. Birth [AD08]. Births [Spl00]. Birthweight [AD08]. Bivariate [CGT17, GE12, LR03, LG20, Wat00, Wat03, Zha20]. BivarSate [Lie98]. Black [Ren07, ZH19]. Black-Box [ZH19]. Black-White [Ren07]. Blacks [Gas85, Vin85a]. Blanchard [Hau96]. Block [HK94]. Blocked [XZW16]. Blockmodels [MHCP22]. Blockwise [BC89a]. BLS [Nor85]. Board [Ano12a, Ano13a, Ano14a, Ano15a, Ano16a, Ano17a, Fre86a]. Bond [BFZ98a, HRV20, ZC17]. Bonds [KW90]. Book [Ano93a, BG88, CRL^{+}85, CRTO86, CBC^{+}87, DCW^{+}86, DJA88, DC85, KP87, Mil85, MOJ^{+}86, NNH87, NNW88, PAC85, SB88a, WCM^{+}86]. Booms [ACS15]. Boosted [YQZ18]. Bootstrap [BS92, Cla94, DM10, DC86, GHLT19, GPD15, Gos02, GJ01, GHR87, JL22, Kim01, LXY22, MNW19, NS96, PPT92, PF84, Psa16, SB19, WKX23, Xu00]. Bootstrap-After-Bootstrap [Kim01]. Bootstrap-Based [Gos02].

Bootstrapping [BK91, BN02, CNR20, FP84, GHPS22, Ray90]. Borrowing [Hau91]. Both [AAW11, DNW98, GK14, HK17, Pak11]. Bound [Aiz91, Cla94, LTMZ19]. Boundaries [FXMFY20]. Boundary [And02, YF20]. Bounded [Kra86]. Bounded-Influence [Kra86]. Bounding [SW90]. Bounds [BCFFL19, Bur94, CF15, GNK86, GIR20, HL22, LW09, MM03, MPTK91, MRW19, PT12a, Rus19, Vol22]. Box [ZH19, Col91, HT83a, KFL03, SY94, ZS14]. Bracketology [KB10]. Branch [KBZ89]. Brand [AL95, BG05, CP98, JVC94, MHB19]. Brand-Choice [BG05, JVC94]. Brand-Level [MHB19]. Brazil [Dur98]. Break [BLS92, Chr92, KMN05, RS94a, Sen03]. Breakdown [AK06]. Breaking [HM07, Ser95]. Breaks [BT03, CWW20, DRV20, HM16, INT21, KW22, KNW22, LLF22, Smi22, Tim01]. Bridge [CH14]. Britain [BZ99]. Broken [Poi91b]. Brokerage [HH11a]. Bryan [Pel21]. Bubble [CNR20, Fri22]. Bubbles [PJ20]. Buckle [MRZ92a]. Budget [Meg87, Met98, Mof86, PRC98]. Budgeting [JSS88]. Buffered [ZLY17]. Building [PRC84]. Buildings [TH83]. Bulk [PRC98]. Bull [MM00, MMS12]. Bumps [ABMW99]. Bureau [But85, Jae97, Wai85]. Business [AVV01, Ano00c, Ano01, Ano02d, Ano03a, Ano03b, Ano04b, Ano04c, Ano05c, Ano05d, Ano06b, Ano06c, Ano07c, Ano07d, Ano08c, Ano08d, Ano09c, Ano10c, Ano11c, Ano13c, Ano13d, Ano15c, ABS93, ADS09, BTZ04, Bar85, BJOS99, CD09, CP08, CdH96, CK03, ERT88, Fil94, Ghy94, GH02a, Goo93, Ham94, HI97, HS93, HS96c, Knü09, LSS19, McK18, Nef93, Pop84, Sic94, SvD06, Tau93b, Tho20, TR93, VKR06, VLK04, Zue03]. Business-Cycle [BJOS99, CdH96, Fil94, Goo93, HI97]. Business-Related [VLK04]. Businesses [Pre95]. Busts [ACS15]. Butterworth [Góm01]. BVAR [KCR17].

C [Sim85]. Cake [HO84]. Calculating [Cha93a, Tar07, Yit91]. Calculation [GA10]. Calculations [DBCB19]. Calibration [DIW96, GS91b, Knü15]. California [Fre86a, KT95, Kra07]. Campaigns [WLL18]. Can [AFFT22, CO99, FV95, HS89, JSSvN20, LMT96, McK18, NW91, RW96, Wat00]. Canada [Fer97, GH95, Kar95]. Canadian [DM84, DMS99]. Canonical [Ott90, Pos00, ZH92]. CAP [Erm91]. Capabilities [FMB^{+}98a]. Capacity [Mor85, OTZ91]. Capital [CFM11, HK86, Mor93, PvP06, Wil09]. Capital-Labor [CFM11]. Car [GFdH87]. Card [KD08]. Care [CW96, Das08, DMT06, DMMN83, GH91, HO84, HLR87, Kow16]. Careers [Ren07]. Carlo [AS96b, BC89a, CW01, DF15, GO98, HM89, Har91, HLR87, Lum95, Ray90, Ril91, Sch89, vdKH89]. CARMA [TT06, BDY11]. Carrots [AvdBL20]. Cars [Oht87]. Case [CH14, DH89, Fre85, Fre86a, Ghy90, HT83a, HS96c, KW90, Kar95, KKPS22, KN89, Kro83, LeS92, LXHT22, Pow22, Sle98, SB89a]. Case-Control [LXHT22]. Cases [FG00]. Catastrophe [BK16]. Catch} [Lav20]. Categorical [HLW14, LR08, LRW09, MN20, Sta88b, Swa03a]. Category [MHB19]. Category-Level [MHB19]. Cattle [FB92]. Cauchy [DH12].

Causal [BMM20, CM88, FGK20, Hub15, Kli16, Lec09, LZ19, MQZ22, MOT22, ZH19]. Causality [BRT12, CT16, GR15, Kan85, RF15, ST18, ST20]. Cause [JJK20]. Cautionary [JJK20]. CAViaR [EM04]. CCE [SW22b]. CDF [LR08]. CDS [OP18]. Ceilings [PW10]. Cell [VP92]. Cellular [Hau99]. Cement [Das91]. Censored [BFV22, HW90, Ioa95, dRJH11, Kow16, MT99, PZZ20, PX21, RS09]. Censoring [BCFFL19, Fra19, KST11]. Census [But85, McK84, ZW93, dBM90, Jae97, Wai85]. Central [AGO^{+}14a, CLR14, Yil03]. Century [CD09]. Chain [CW01, Dan92, Gre89]. Chains [Gra97]. Change [ACH03, BHZ20, BGR15, GK08, Han97, HMR20, JMZ18, LL90, ML90, Mos91, OG86, TZ19, WW09, XZW16, ZLLC23, dHK14]. Change-Point [BHZ20, BGR15, GK08, TZ19]. Change-Points [XZW16]. Changed [Li11]. Changepoint [FG00, HT22]. Changes [BS87a, Bor85, BS95, Fuh92, GD89, Ham94, HSS20, Inc93, IOS90, JS99, KP10, Kim93, Pat86, PY19, Pro08, Slo87, WZ00]. Changes-in-Changes [HSS20]. Changing [BK15, CW92, GPR13, HD85, Jus98, KD08, KN89, LTMZ19, MT18, Per90, PV92b, WHYC22]. Channel [BCG85]. Chaotic [Ser95]. Characteristic [JK02]. Characteristics [Bar85, BB05, Gra97, HS96c, Son10]. Characterization [Mar85]. Charity [Das08]. Checking [HK92, WS20]. Checks [CM12, LP12]. Chemicals [Mor93]. Chemistry [HJS20]. Chi [Rit86]. Chi-Squared [Rit86]. Childbearing [JJK20]. Children [GOV84, KSS09]. Chilean [Kas09, YCJK20]. Chinese [GGW14]. Choice [Abe99, AL95, BG05, Bor89, BCM10, Car01, CP98, Edw08, ES01b, GLL93, GO98, GK18, JVC94, Kas22, Kea97, KJGV09, KT19, KW19, LT12, LSS16, ME04, MT96, PPL94, RI12, SW94, SV88, Wer99]. Choice-With-Screening [LT12]. Choices [LT12]. Choosing [AAMW20, APW20, DMMN84, Hay89, PPT92]. Chow [KW92]. Chow-type [KW92]. Christopher [Ano02b]. Chronic [Dur98]. Circularly [Che22]. Circulation [Cra86, Goi85, vdKH89]. citation [Don22, JJKL21, JJKL22, LC22, MLZ22, ZK22]. Cities [BK05, PZ90]. Claim [Yan20a]. Claims [HS89]. Class [BL05, CGT17, Cre17, FMT10, GW85, GP98, HL22, HMR20, Kum87, LS16, VM89]. Classes [LZ95]. Classical [Bur94, DW93, HI01, Poi91a, VSV99]. Classification [DV04, FFX16, GGW14, Mey88, WTC84]. Classification-Error [Mey88]. Classified [WK91]. Classifier [HSZ22b]. Classifier-Lasso [HSZ22b]. Classroom [GRTZ22]. Clock [MKL07]. Closed [OLBvD20]. Closed-Form [OLBvD20]. Closings [CRZ19]. Cluster [CKMS22, PT18, PT23]. Cluster-Robust [PT18, PT23]. Clustered [PF19, ZCZ^{+}21]. Clustering [CGM11, FSK08, GR12, JS10, KM08, MNW19, Mel22, PW22, PHPY20]. Co [BKW19, Don22, HM07, JJKL21, JJKL22, LC22, MLZ22, ZK22]. Co-authorship [Don22, JJKL21, JJKL22, LC22, MLZ22, ZK22]. Co-Breaking [HM07]. Co-citation [Don22, JJKL21, JJKL22, LC22, MLZ22, ZK22]. Co-Movements [BKW19]. Coauthorship [WF22]. Cocitation [WF22]. Coding [AV05b]. Coefficient [BHL92, CFLS19, CGLS18, CGZ19, DMP19, DFP11, DGP20, HT22, LHLF02, LLCL22, Ost19, PHPY20, SWW88, SMU13, WZ15, XZW14, YZK19]. Coefficient-Probability [SWW88]. Coefficients [BG05, CHL88, Das91, DBS84, HL16, JVC94, Kli22, LGG22, NSR88, Tyr83, Wri00c]. Cohort [DMS99, Juo18]. Coincident [Tho20]. Coins [Goi85]. Cointegrated [CdM19, GG95, KP10, PV03]. Cointegrating [Bos96, CT18, Ell00, EJP05, HJNR98, HP10, LRS99, Pos00, SL00, Str03, Wri00c]. Cointegration [AK06, BK11, BZ20, Bre01, CGT17, CL93, CD98, CT18, ES01a, EHM98, Eri98, FH94, FL98, Gre94, GK03, Hau91, HW10, LL07, LV15, Mac94, Nie04, Nie07, She92, SRdCFB18, vH89]. Cointergrating [GJ01]. Coke [GKP00]. Collaborative [YLYJ22]. Collaborators [Ano83a, Ano84d, Ano85b, Ano86d, Ano87d, Ano88a, Ano89a, Ano90a, Ano91a, Ano92a, Ano93c, Ano94b, Ano95a, Ano96d, Ano97b, Ano98a, Ano99a, Ano00a, Ano02a, Ano04a, Ano05b, Ano06a, Ano07a, Ano08a, Ano09a, Ano10a, Ano11a, Ano12b, Ano13b, Ano14b, Ano15b, Ano16b, Ano17b, Ano19, Ano20]. Collection [BL93]. Collective [Blo19, DLP19, Luo02]. Colleges [Rou95]. Collinearity [Iwa91, WN89]. Combination [AGJT14, BK15, CT09, Die88b, FS91, GK05, GdXPE16, GW88, PP11, TH89, Wei86]. Combinations [DL20]. Combine [KCR17, MRW12]. Combined [ARvD18, Cla94, LSX15]. Combining [BB22, CGR07, CW86a, HH11b, Jan88a, LM92, SU22, VLS90]. Come [Don22]. Commemorative [GH02a]. Comment [Agu08, ASMZ06, Aka84, All96, And94a, ABFN06, AG14, And14, Ans83, Ans93, App85, AL18, Bai98, Baj08, Bal04, BR06, BNS06, Bel91, Bel96, BW88, Bla87, Bra13, Bro88, Buj96, Bur84, But85, Can09, Car85, CC14, CMM23, Cha92, CS09, fC88, CH16, Chr07, Cle98, CW08, Coo88, Cro12, DL84, Dan94, Del86, DL04, Dic88, Die88a, Die06, Duf09, Eas88, Eng94, Eri93, Fas84, Fau07, FG84, Fin83, Fin96, FS14, FZ14, Fra14, Fre89, Gal07, GP88, GK92, GM06, Gas85, Gew88a, Gew94, Gew98, GJ94, Ghy98, GS06a, Gia16, Gra13, Gra84, Gra86, Gra93, Gra98, Ham91a, Han86, Han93, HT15, HW14, HR94, Hay92, Hil86, Hil89, Hil91, HW20a, HfJL98, Hog83, Hog88, Hon91, Hon05, Hon08, HRvD12, HM14, Hyl96, Hyl98, Ino15, Jac13, Job88, Joh04, Ken14, Kha08, Kie14, Kil07, Kil15]. Comment [Kil22, KT13, Koe88, KM91, Lah12, Lin14, LZ14, Lit86a, MaC92, Mag08, Man13, Mar84, Mar98, McA14, Mik09, MR08, Mon88, MC98, Mül18, Nel94, New83, Nor85, Oom06a, Ord93, Pat15, PY06, Pie84, Por18, Qua93, RR94, RTW12, Rob98, Ros84, Ros86, Ros12, Ros14, RL84, Sab88, Sac13, San88, Sch88, SAW92, Sco14, SK94, Sim85, Sim87, Sim07, Smi83, Spi86, Sta88a, Sto16, Sun14, Sun18, SB88b, Swa16, Tsa93a, Tsa93b, Uhl94, Vog14, Vog18, Wai85, Wal98, Wal04, Wat87, Wat88, Wat96, Wec89, Wes12, Win89, Win05, WP87, Wri09, Wri15, Yao14a, Yao14b, Yog09, You88, YL14, ZC09, vdK05]. Comments [CFR05, CR84, DMP19, Mil84, RR22, ZH92]. Commerce [DHW17, KR88]. Commercial [WW09]. Committee [AT17, Pie83].

Commodity [vH89]. Common [BC91, BT22, BKT11, BJ99, CCM16, CHT07, Che22, CHN97, CP10, DE19, ES93, EK93b, GG95, HHPS07, HL16, KB04, LGLY16, Lu22, PS16, Qia22, Rit86, Urg07, WV11]. Common-Scaling [CP10]. Common-Stock-Return [BC91]. Communications [Gas85]. Community [BGL20, Rou95, XZW22]. Commuting [VG97]. Comovements [GMP11]. Company [Ham13]. Comparable [Tar07]. Comparative [AL89, DL92, MPTK91, Pow22]. Comparing [AG07, BP22, BMY04, DM95a, DM02, Die15a, DFK19, EMW09, FW09, FPAB02, GR11, HO84, HI97, MNO85, Pat19]. Comparison [ACL11, CP08, Chr90a, DZ04, DMMN83, DU12, Fom86, FB97, Gar89, Gbu89, GHR87, HLR87, KP95, LeS92, Mal83, McN86a, MG84, MS97, PGFF94, Qua19, SW88, TU90, WTC84, Wei86, WW96, Wüt20, Zil97, dBM90]. Comparisons [CM12, KS08]. Compensating [Lav20]. Compensation [LW09]. Competing [BH97, CZ18, McG91]. Competition [CRZ19, HJS20, Xie22]. Competitive [MF05]. Compliance [HS19a]. Component [CS20, Gar89, Gos02, HH21, HT14a, JJR18, Kim93, KHT05, LdV99, Mar87, PY09, Rei99, ZL22]. Components [BW84, CHN97, CDJW10, GG95, HK92, HpK23, KR90, LZ18, MMS12, Mar85, NPW93, OF97, RW86a, Ric93, Shi00, Slo87]. Composite [CJJ^{+}23, HH83, HZ22, LSS16, Phi87, Tuz22]. Composition [Rot15]. Compositional [AVV01, GD89, HHL^{+}22, LSCX22]. Compound [YQZ18]. Computation [Fdv00, RS92]. Computations [Swa03b]. Computers [BB05, Col93, NTP94]. Concatenation [MS03, Rub86]. Concave [Hen22, Tsi22]. Concepts [KS20]. Conceptual [RCW83]. Concerts [VSV99]. Concurrent [McK84]. Condition [CdM19, Zil97]. Conditional [AHL15, Aie13, ABJT22, AT14, AT11, BG93, BB89, BHZ20, BL05, BO16, BHL92, BG96, BRT12, Bru92, BH93, BBCL20, CM02, CJZZ18, Dav04, DE13, DNW98, Due99, Eng02, FHLZ20, FT12, FFK18, Fri22, GA10, GK05, Gre89, HV03, Hog19, HD21, HMI95, HLM16, KWD13, KN89, LR08, LLR13, LSX15, Lin97, LSU11, LSZ14, MLST20, MQZ22, MMO22, MT06, Mil87, MSU10, NEV15, NZ20, PRPS22, SW20, SCS20, SV88, SU09, Tao19, TMH^{+}23, TT02, kTY12, TW14, VP93, YLN17, ZK05, ZZZ20, ZLY17]. Conditional-Demand [HMI95]. Conditional-Variance [BB89]. Conditionally [DS98, FSC03, GIR20, ZL15]. Conditioning [Koz20, LA84]. Conditions [ADS09, CD09, CHL18, FIL^{+}23, FW03, GKR12, KFL03, KZ20, Mos99, SW21]. Conference [ERT88, HF86]. Confidence [AJL96, CPDH15, GIR19, Hog19, HL22, KSD98, LY18, Ros05, Wri00b, Wri00c]. Conflict [DP19]. Conformal [YYX20]. Confounders [DMP19, PPS19]. Connected [BKW19]. Connectivity [DL20]. Conscious [FHLW00]. Conscription [IV95]. Consensus [GSY01, Has87]. Consequences [AKO12, Mar99]. Conservation [DH88]. Consideration [BM23b, Gua23, Hon23, Mor23]. Consistency [HW84, SW21]. Consistent [AW07, BDB14, Cha20b, CSZZ17, GK96, Hen22, Lee18, LM94a, McC19, Pos00, Rac97, WLJ07]. Constant [CH95]. Constrained [BS95, Dol99, GRL13, PS92, Qia19, YCJK20]. Constraints [BSS01, Con90, FP97, FXMFY20, FB97, FPAB02, HL04, Hau91, HKYZ20, KRW^{+}14, LTU15, Meg87, Mof86, NC92, PvP06, Tah01, Wan03]. Construct [Rei99]. Constructed [HP11]. Constructing [KJGV09]. Construction [CJJ^{+}23, DF20, KSD98, NP86, YT07]. Consumer [AL95, AJL96, BS87a, Bra07, DM84, DO95, DW88, DP06, FLS07, Ham13, Hei83, Jai89, JS87, JW99, Kea97, KSM97, KD08, Ker85, Kok87, Kou13, MF05, MT96, Mos91, MS99, OG86, PS92, Ran88, RW98, SW94, TN92, Ver12]. Consumer-Choice [SW94]. Consumer-Demand [RW98]. Consumption [ABMW99, CM90, EN94, Erm91, FOTY20, GDG^{+}05, GK18, GMS87, Hau96, Jac00, LS94, Li05, LA84, LW97, Lus96, Nor94, OS89, PRW06, Sle98, Sto88a, SW88, Web00]. Consumption-Based [GMS87]. Consumption-Habit [Li05]. Contagion [CT16, CST18, FMT10]. Containing [DNW98, KSD98]. Contamination [CGQ20, MH17]. Contemporaneously [Lüt84]. Contemporary [Hof91]. Content [IOS90]. Contestants [VZ11]. Contested [JT95]. Context [CMM23]. Context-Dependent [CMM23]. Contingent [FLSVP04, Wer99]. Contingent-Valuation [Wer99]. Continuous [CR12, DG02, GMS87, HPS96, HLZ22, JK02, Lec99, LR08, LRW09, LLL20, LXY22, PdV06, RE05, Yan20b, ZZZ^{+}22, TT06]. Continuous-Time [CR12, DG02, GMS87, HPS96, JK02, RE05, Yan20b, TT06]. Contraception [AKO12]. Contrarian [CGK97]. Contrast [HHL^{+}22, MJY19]. Contrasts [DTU06]. Contribution [Spl00, Wil09]. Control [AR88, BCHK16, KKPS22, Kli22, LKFL22, LXHT22, McK18, Pow22, Sal95, WHT22]. Controlled [HK94]. Controls [BCW16]. Convergence [CHT07, CT22, Kim05]. Convex [DL20, Dol99, Hen22]. Conway [FG84, Mil84, SSBD15]. Cooking [Ghy00]. Coordinate [RS94b]. Coorection [FB99]. Copula [AS20, BRT12, Che07a, GKK20, LW15, LMS20, LG20, NEV15, OP18, OLBvD20, PF19, SV16, WW20, Yan20a, ZSZ20]. Copula-Based [BRT12, NEV15, PF19]. Copula-Linked [ZSZ20]. Copulae [GHS09]. Copulas [CMT19, OP15, SK20, ZT06]. Core [DG06]. Corporate [KW90, LTY^{+}22, San17]. Corps [CF15]. Correct [AFFT22, CH14]. Corrected [AI11, BC05, DE19, YT07]. Correcting [BRW12, PSW04a]. Correction [Ano87c, Ano94a, Ano96b, Ano96c, Ano05a, Ano21a, Ano21b, BBD20, FO05, GK03, Ilm90, Str03]. Corrections [Ano85a, GH22, MMS12, Mül14a, PV92b]. Correlated [BH09a, CSZZ17, CW01, DE19, HMS17, HpK23, May86, RSW18, WZ20a]. Correlation [Aie13, AT14, BHZ20, BGL20, BBCL20, BV05, BT01, CST18, CHNY22, DM99, Die88b, Eng02, GA10, HHPS07, IM10, JR22, KR92a, LTY19, MW22, NS10, Ng06, OT21, Ott90, RF15, Rod16, TXLT22, Vog98, YF22, ZH92]. Correlation-based [TXLT22]. Correlations [AT11, BO16, CKL11, DM16, EC06, OF97, Pos00, RE12, TT83, TT02, YO19]. Corridors [CPDH15]. Corrigendum [Ano84b, PT23]. Coskewness [AGU04]. CoSmo [Dol99]. Cost [ACJK22, BS87a, CP10, Fre88, GG90, GRW84, Had99, HH11a, HD85, IV95, JS99, KFL03, Kok87, KPS00, KOS94a, KSS09, ML90, Nor99, Oht87, Wri83]. Cost-of-Living [CP10, GG90, HD85, Kok87, Nor99]. Costing [RW86a]. Costly [Asa02]. Costs [Asa02, DBS84, DP06, LW19b, SJS18]. Cotrending [Bie00]. Could [AKO12]. Count [BW19, CW96, CW01, GT96, GE12, HF89b, JLR11, MABF08, Mon97, San17, Wec89, Win95, vO00]. Count-Data [Win95, vO00]. Count-Panel-Data [Mon97]. Count-Valued [BW19]. Counterfactual [HLL20, MM20]. Counterfactuals [McK18]. Countries [Gew88b, KOS00, MS97, PZ90]. Country [MM19]. Country-Specific [MM19]. Counts [KLS12]. CoVaR [FHWZ18]. Covariance [ABJT22, AS96a, BBCL20, Cha20a, Cha20b, Cla96, ELW19, FFX16, GR15, GJ22, KLZ20, LLT^{+}15, LFWT18, LLF22, LSCX22, LSU11, LSS16, PSSE20, Por02, SX20, Swa03a, TMH^{+}23]. Covariances [GL16, OJLD18, TW14]. Covariate [FIL^{+}23, Hei22, KST11, LS15, Lu15, TZ19, XZW22]. Covariate-Assisted [XZW22]. Covariates [DBCB19, FHWZ18, FH19, HK14, HHL^{+}22, Hil86, Kas22, KKPS22, LCZT22, LLL19, MHCP22, Rot15, Wes15]. Covariation [BHMR19, Var16]. Coverage [CHSS04, HL22]. COVID [KNW22]. COVID-19 [KNW22]. Cox [Col91, KFL03, ZS14]. CP1 [Rei99]. CPI [DF20, DG99, Hau99, Nor99]. CPS [KLN85]. Crash [Fri22, MN95a, ZA92, ZA02]. Crashes [BMP10, PJ20]. Credit [BGS20, CGT14, DBCB19, DM16, FP97, KD08, KLS12, PS92, RCW83]. Credit-Constrained [PS92]. Credit-Scoring [RCW83]. Crime [Hos18]. Crisis [AGO^{+}14a, KLS12]. Criteria [KP95, PGFF94, Pie84, ZWZ12]. Criterion [BMY04, KW19, Lu15, YLL19]. Critical [CL95, Jan88a]. Critique [BS87b, McC97, MR91a]. Cromwell [Ros08]. Crop [HF87]. Cross [All90, BGR15, Bor89, BCP19, CR88b, DC86, GDG^{+}05, GL13, HLM05, JR22, Kim22, May86, MKL07, Ng06, RF15, RSW18, SV16, TV15, TT83, WTC84, WK91, Wu19, dJ00]. Cross-Asset [BCP19]. Cross-Classification [WTC84]. Cross-Classified [WK91]. Cross-Correlated [RSW18]. Cross-Correlation [RF15]. Cross-Correlations [TT83]. Cross-Lagged [May86]. Cross-Listed [MKL07]. Cross-Section [BGR15, JR22, Ng06, TV15, dJ00]. Cross-Sectional [Bor89, GDG^{+}05, HLM05, SV16]. Cross-Sectionally [Kim22]. Cross-Validation [All90, CR88b, DC86, Wu19]. Crossing [JY21]. Culling [CSZ22]. Cure [DBCB19]. Currency [CG00, Guo98]. Current [AHI99, Bra90, CM90, Lus96]. Current-Quarter [Bra90]. Curvature [AJ05, KFL03, Mos99, Ram94, RW98]. Curve [App85, CGM16, JY21, Kak84a, KM09b, KYK^{+}21, KLGS10, RW09]. Curves [ARBS87, BdB18, CL17, CG02, FB97, FPAB02, HK01, LY18, McD87a, WP83]. Customer [KS96a, MS88a]. Customs [BCCP18]. Cuts [CJW86]. Cycle [BTZ04, BJOS99, Cag03, CGL20, CP08, CdH96, CK03, DL87, Fil94, Ghy94, Góm01, Goo93, Ham94, HI97, Knü09, MS89, Nef93, Pop84, Sic94, SvD06, Tho20, VKR06, Zue03]. Cycles [AVV01, CHT07, Har85]. Cyclical [FS93, Gew88b, Jef08, PS96].

D [Gas85, Mül18, ZSZ20]. D-Vines [ZSZ20]. Daily [AGK13, BB89, BB02, HHPS07, HP85, Hsi89, JLP23, KWD13, Mil87, RT00, Tuc92]. Dairy [KGM91]. Damages [GW85]. Dan [BR95]. Daniel [TTW95]. Danish [HS19a, Kon98]. DARM [McD87a]. Data [Abe99, AL84, ABA99, Alv04, And00, AB22, AGK13, ACL11, AHI99, BC91, BN05, BFT16, BL93, BMM20, BFHM11, BGS20, BT89, Bor89, Bra90, BBCL20, BC05, CW96, CGJ19, CMW23, Car01, CsL94, Cha92, CF99, CGL13, CSZZ17, Che18, CLQ20, Che22, CW01, CGR07, CM09, CG08, Cle17, CG17, CVF06, CO99, CHNY22, CEJ^{+}12, CHLR92, DF11, DL92, Deh03, Dev07, DGP20, Don22, DH89, EH90, EH23, ES01b, Erg19, FFX16, FHLZ20, FLTS19, FM97, FLM20, FLSVP04, FG00, GK14, GFHPZ87, GNK86, GGKS98, GGS96a, Ghy97, Ghy00, GRL13, GdXPE16, GK96, GE12, HGB^{+}12, HHPS07, Hal94b, Har83, HK17, HPP00, HMvdB12, Hos18, HLW14, HS89, HF87, HLM15, HT20, IR05, Ish22, JVC94, JS10, JLR11, Juo18, KW22, KL90, KR92a, KLN85, Kim22, KK12, KYK^{+}21]. Data [KS96a, KMR19, KP11, Kum87, Kum89, LZ87, LY18, LCZT22, Lec95, LW09, LM90, LH99, Li99, LR08, LRW09, LLL20, LZW^{+}22, LSCX22, LQZ19, LGG22, LS15, LLCL22, LTY19, Lit88a, Liu21, LS98a, Lou07, Lu22, Lus96, MQZ22, MN88, MTLJ97, MM09, Mal83, Man15, MP83, MRZ92a, MM20, MM92, MVVZ19, MRW12, Mel18, Mit90, MAB^{+}23, Mon97, NWK10, NP97, Ng06, NPF11, OS20, PHPY20, PRC98, PF19, Pfe91a, Phi10, Pic01, PM20, QZL15, RV19, RLYZ22, Rei99, RW86b, Rod16, San20b, SS84, ST89, SL01, Sle98, SW94, Sta88b, SWJ19, SL11, SX20, SSBD15, SW88, TM19, Tau86b, TR86, Tsa00, Tsa16, VLS90, WlMCP98, WJS08, Wil83, WK91, Win95, WZ20a, WSX22, XZW16, Yan20a, YZK19, ZW93, ZS15, ZWLL22, ZL22, ZLS19, Zho96, ZTCW15, ZLWW20b, ZPWW22, Zil97, ZHB23]. Data [dHK14, dJ87, vO00, CMO20, FS22, JSSvN20, SvD06]. Data-Analytic [BC91]. Data-Based [Hal94b]. Data-Dependent [LH99]. Data-Driven [SL11]. Data-Filtering [GGS96a]. Data-Rich [BGS20, Man15, MVVZ19]. Database [MN16]. Databases [Woo00a]. Dataset [FW09]. Datasets [BT22, CFS09, Kap10, Qia22]. Dating [CGL20, CP08]. DAX [ACW97]. Day [LA84, SH87]. Daytime [MN95a]. Deadliest [Lav20]. Deadweight [HPH87]. Dealing [CMT19]. Deaths [Spl00]. Debates [PR87]. Debiased [CKMS22, HW20a]. Decentralized [WHW22]. Deciphering [HRV20]. Decision [Agu10, AK13, BK91, GSC03, KS94, LNB10, Luo02, Sim90]. Decision-Making [Luo02]. Decision-Theoretic [KS94]. Decisions [AT17, Bar85, BFR97, KW04a]. Decline [HP05, JJK20]. Decomposing [ACH03, KL08, Rot15]. Decomposition [AG10, ACL19, Hub15, LHM20]. Decompositions [GL19]. DeCoster [RR94]. Decreasing [ZO04]. Decriminalization [DHZ10]. Deducing [BMP10]. Default [DM16, Kie10, KL08, KLS12, LTY^{+}22, LSZ14]. Deficit [Met98, Qui95]. Defined [Kan86]. Definition [HK86]. Deflators [LG89]. Delay [KP03].

Delinquency [LLPZ20]. Demand [AAW11, ACL11, BK05, Bar83b, BFS94, BV87, Bor85, Car06, Cha87, CK08, CO99, DG96, DW85, DW88, DMMN83, DH88, EN94, FID16, FHLW00, Fis92, GZ84, Har83, HO84, HLR87, HS85, Hei83, HW90, HK94, HMI95, HH83, JR90, Koe06, Kou13, LPS99, MTLJ97, Mel90, MF05, MS96, Mos91, Mos99, Oku11, Olm09, PS92, Ram94, Rog87, RW98, San20b, SY98, Son10, You88, vdKK03, vHPP04]. Demand-Systems [Fis92]. Demeaning [Wes18b]. Democratization [Rou95]. Demographic [But85, IM97]. Demographic-Specific [IM97]. Denmark [Jus98]. Dense [ZL22]. Densities [BL05, GW14, SV16, WW20]. Density [AGJT14, ARvD18, AG07, Ber01, ÇD22, CT22, Cla11, DF15, FE22, GP98, GR11, GHNS09, HCP11, IRR22, Knü15, LLN89, LS15, Liu21, MJY19, MOT22, OXM13, Wu19]. Density-Ratio [GP98]. Density-Tempered [DF15]. Department [CRZ19, KR88]. Depend [Wes01]. Dependence [AFO21, BO16, BR89, BH97, BS20a, CGQ20, Erg19, EH19, GHS09, HLM05, HM16, Hog22, KG15, Kea97, LT19, LT04, OP15, Pro12, Psa16, PV22, RT00, Seo20, SV16, WHT22, Win95, ZLWW20b, Zue03]. Dependence-Based [Hog22]. Dependencies [BJ99]. Dependency [ZL22]. Dependent [Ang01, BH93, Car84, CMM23, Das08, DM94b, Est98, GSA04, GOV84, KST11, Kim22, LY18, LH99, LK88, Lou07, LG20, PRM99, RS15, Swa03a, Tsa16, XZW16, Yog08, ZZZ20]. Deposit [DS90]. Deposition [PR87]. Depreciation [Ran88]. Deprivation [App85, Kak84a]. Deregulation [DS90]. Derivative [GS10]. Derivatives [CS05, Li11]. Descent [WHW22]. Design [FH19, HHK11, Kot84, LXHT22, PBC95, War89]. Designed [FG00]. Designing [HS93]. Designs [BBD20, BI20, Don19, FLM16, GI19, Gro86, KJGV09, PLCW22, QY19]. Desk [dRJH11]. Detecting [Bal93, BS20a, Gur91, HSZ22b, Wil93, XZW16]. Detection [BCCP18, BHZ20, BGL20, CCKR17, CT90, FL98, HT22, Inc93, Joh87, LLF22, LKFL22, XZW22]. Determinants [Cle85, HW84, KGM91]. Determination [AMS19, Ger88, Joh87, KP95, Pos00]. Determining [BN07, BK91, DP87, DP02, Kap10, MM92, Pro08, Rot15]. Deterministic [Gag90, Wil93]. Detrending [Kan85]. Deutsche [GR89]. Developing [RCW83]. Development [CM93, HHW91, LeS91]. Development/Firm [HHW91]. Deviance [BMY04, HH19a]. Deviations [Mil87, Ros05]. Diagnostic [HK92, PTZ86, SW21, Tse87, WS20]. Diagnostics [JMP16, Mou87, Qua88, Sch87, Sil01, ZK05]. Diagonality [LLT^{+}15]. Dichotomous [Est98, Wer99]. Dichotomous-Choice [Wer99]. Dickey [CL95]. Did [JJK20, OG86]. Diebold [Die15a]. Difference [Abr03, CK22, CE08, DHS19, LS20b, PX21, SH07, WZS22]. Difference-in-Differences [CK22, DHS19]. Differenced [LMT96, MV88]. Differences [AK84, BS20a, CK22, CS05, DHS19, KLL20, Öll86, Ren07, Rot15]. Differences-in-Differences [KLL20]. Differencing [DP87, DP02, SD87]. Different [DL20, KP95, Lee18]. Differentials [IM97, Lav20]. Differentiated [DHW17, LW19b]. Diffusion [BCS08, BF05, DG02, Era01, IRS15, Kim22, Kri17, Smi02, SW02, SC17, TY89, ZZZS19, ZC17]. Diffusion-Index [Kim22]. Digital [TGCD16]. Dimension [CKLGS12, LS20b, LZ19, MM92]. Dimensional [Bab22, BCHK16, CCKR17, CZZ22, FHWZ18, FHLZ20, FLM20, HHL^{+}22, HKYZ20, HLW14, KLZ20, Kor19, Koz20, LS20c, LKFL22, LSCX22, LTY19, MLST20, PSSE20, PX21, ST22b, TMH^{+}23, WKX23, LXHT22]. Dimensions [BOP23, KKS22, OP15, PX22a, PX22b]. Dips [AW84]. Direct [CD09, CW92, DHS19, FE22, HSS20]. Direction [Kli16]. Dirichlet [CG02, Kou13, MN20, WHKCS14]. Disaggregate [HH11b, NP90, PPL94]. Disaggregated [Ive01]. Disaggregating [Wil83]. Disaggregation [JMP16, Los85, dA88]. Disagreement [Str22]. Discerning [KS20]. Disciplining [DGJ20]. Disclosure [DL89, Paa88]. Discontinuity [BBD20, BI20, Don19, FLM16, FH19, GI19, HZ22, MJY19, OXM13, PLCW22, QY19]. Discord [Has87]. Discount [GSC96]. Discovery [MKL07, WHT22]. Discrepancy [Che12, LZGL23]. Discrete [Abe99, Agu10, BCR90, BCM10, CDS20, CR12, HLZ04, KS20, LLL20, MHCP22, Rit86, RE05, SSBD15, ZZZ^{+}22]. Discrete-Choice [Abe99]. Discrete-State [RE05]. Discrete-State-Space [BCR90]. Discrete-Time [CR12, HLZ04]. Discriminant [Hag83]. Discriminate [Gra97]. Discrimination [Ano84c, CR83, CR84, FG00, Gre84, Iwa91, MB84, Sch87, WN89]. Discussants [Ley85]. Discussion [Ano84c, Atk83, Bar83a, Bon20, CC83, Don22, Dun83, Ful83, Gua23, Hon23, KT18, LC22, MLZ22, Mor23, Par83, Pel21, PM22, Sin21, Tam20, Tsa00, WF22, Wes18a, XCK21, ZK22]. Disentangling [CP19]. Disequilibrium [BR89]. Dismissal [Fli97]. Disparity [PW10]. Dispersion [Win95, XZL20]. Distance [BGKS04, CPP18, Gas85, GN15, SJS18, Vin85a, YF22]. Distance-Weighted [BGKS04]. Distillation [ZHCB16]. Distort [Ver12]. Distortion [GIR19]. Distressed [Ham13]. Distributed [HL83, KC85, PRG^{+}22, Shr89, Tsa85, WS20]. Distributed-Lag [Shr89]. Distribution [AHI99, BV87, CT16, Che12, CLQ20, CG02, Goo96, GKR12, GL13, Hen22, KWD13, KBR95, LLW90, LC85, LLR13, LZW^{+}22, Lit83, LZGL23, Mac94, MM09, Man15, NG89, Rad83, Rus19, RS94b, SJS18, Sch83, Slo87, SS06, Tay19, VB95]. Distribution-Free [Goo96, GL13]. Distributional [AAW11, BZ13, CVF06, LPS22, She19]. Distributions [Ali84, ARBS87, BC91, Che18, CZZ22, CGR07, Dan92, EMW09, FZ13, GHNS09, HGB^{+}12, Ham91b, HCP11, HMN10, HPZZ21, JS10, LMS09, Pan91, PT03, Tuc92]. Disturbances [CS86, LK93, SK91]. Divergence [LS20b]. Diverging [YYX20]. Diversification [GHY92]. Diversion [Rou95]. Diversity [NG89]. Dividend [Lou07, MPV20, Wol00]. Do [AK84, BDJ12, BB87, CLR14, FL97, Gra97, HW04, KV13, MP05, PSvD09, Ver12, Whi98]. Documents [GGW14]. Does [FID16, Pow22, Ren07, SYA08, vUS87]. Dollar [KR90, LM00]. Domain [SRdCFB18]. Domestic [TN92]. Dominance [And18, AMRT22, BDB14, BZ13, BC11, CVF06, DE13, KZ20, ST10, SB19, ZFSC00]. Dominant [LYY22]. Doors [PW10]. Double [CKMS22, Hal94a, LLZ16, LZLL17, PJ20, Qia22, SK99, YL17]. Double-Question [PJ20]. Double/Debiased [CKMS22]. Doubly [Had99]. Dow [ACD09]. Downswing [Car06]. Draft [DHS19]. Drifting [CCM16]. Driven [AMNP18, BG92, BDY11, BBCL20, HH21, MPV20, OLBvD20, ST22a, SL11, TT06]. DSGE [BF15, KPS13]. Dual [Mor85]. Due [BMM20, TZ19]. Dummy [Ang01, Mel05, SH07]. Durable [Pak11]. Durables [JR90]. Duration [AHI99, BFV22, BCFFL19, BH84c, DW04, DM94b, LT04, MT06, San20a, kTY12, Win95, Zue03]. Duration-Dependent [DM94b]. Durations [HMvdB12]. Durbin [Ali87]. During [AGO^{+}14a, Cle85, MA23]. Dutch [vdKH89]. Dyadic [CT22, TM19]. Dyadic-Robust [TM19]. Dynamic [AW00, Agu10, Aie13, AW07, AT11, BW15, BRW12, Bau18, Bäu13, BFR97, BCIK22, BGS20, BT89, BCG85, BC05, CDS20, CWW20, CP98, Col93, CKL11, DE19, Die15c, Due05, Eng02, EC06, EK12, ELW19, FM90, FSC03, FS97, FB92, FJ08, GA10, GRSY11, HH21, HHK19, HS19a, HPP00, Isk20, dRJH11, JLR11, KLZ20, KPS20, KR90, KP99, KLGS10, KMV10, KLS12, Lie98, LR03, Lin86, LSU11, Lou07, MPV16, MHB19, MB13, NW13, OP18, OLBvD20, Phi10, RS07, San17, TMH^{+}23, Tsa85, Wat00, Wat03, WW96, YLPC20, Yao08, Zha20]. Dynamics [AH20, Alv04, AG10, BHLM07, BZ99, BBCL20, CKY85, CM02, CE05, CMHM19, DM91, DS90, DRV20, DH89, Dur98, ES01b, FOTY20, Fil94, FB99, GR89, GK15, HT20, KK15, LTMZ19, Mey88, MZ23, Nie04, Pis95, Pro12, RR94, SY98, Vel22, WJS08, Zha20]. Dynamites [MN95a]. Dystopia [APW20].

E-Commerce [DHW17]. Early [CG17, Mor87, Ren07]. Earnings [AMS19, AN91, BZ13, CMW23, FBB06, GJP02, HT20, Lec99, MA23, MRW12, MZ23, MAB^{+}23, Pis95, ZHB23]. East [KK92, Lec99]. Eat [HO84]. Echelon [LP96]. Echelon-Form [LP96]. Econometric [BS87b, DH88, ELLV97, FP84, Ghy00, HP11, Jar87, LS94, LSS16, MS18, Mor92, MT85, MT02, RW86a]. Econometrics [BG01, BF05, GP98, Ken95, Meg87, Mof86, dlV85]. Economic [ARvD18, AV05b, AJL96, Ano00c, Ano01, Ano02d, Ano03a, Ano03b, Ano04b, Ano04c, Ano05c, Ano05d, Ano06b, Ano06c, Ano07c, Ano07d, Ano08c, Ano08d, Ano09c, Ano10c, Ano11c, Ano13c, Ano13d, Ano15c, AV86, BR08, BCM22, BH84a, BH02, BL93, CW86a, CM96, Dag87, DBCB19, EHM98, Eri98, FS93, Gas85, GH02a, Góm01, Gra87, GW88, HT83a, Isa90, LLL22, LMT96, Lit86c, Man15, McE17, McG95, MN95b, MLR19, Mor93, MS96, Ost84, Qi99, Rac01, Rod16, Ros08, RS95, Rot15, Sea93, Sim85, Smi05, SW90, Tau93b, TR93, TW14, Vin85a, Wai85, XZL20, dJ87]. Economics [LLCL22, Mey95, Urg07, Yit96]. Economies [GHY92, OTZ91, RU01]. Economy [KNP04]. ECU [LM00]. Editor [ALN07, GH02b, GH02a, GH02c, GH03, GH04b, HW11, LNHW10, NL08, NL09, And05, And06, Ano86f, Chr96, Eri98, Gew91, Gew92, Ham94, Tau93a, Tau94, Tsa00, Woo00b]. Editorial [And07, Ano83a, Ano84d, Ano85b, Ano86d, Ano87d, Ano88a, Ano89a, Ano90a, Ano91a, Ano92a, Ano93a, Ano93c, Ano94b, Ano95a, Ano96d, Ano97b, Ano98a, Ano99a, Ano00a, Ano02a, Ano04a, Ano05b, Ano06a, Ano07a, Ano08a, Ano09a, Ano10a, Ano11a, Ano12a, Ano12b, Ano13a, Ano13b, Ano14a, Ano14b, Ano15a, Ano15b, Ano16a, Ano16b, Ano17a, Ano17b, Ano19, Ano20, GH04a, TW98, Woo01, Zel83]. Editors [Ano93a, CHT17]. Education [Jae97, LPS22]. Educational [IM97, Jef08, Rou95]. EEC [HV87]. Effciency [Por02]. Effect [AFO21, Coc89, CR12, DG99, GR12, Hei22, HLZ22, Ish22, JR90, KT95, Kli16, LTY19, Rot15, Rou95, San20a, Vos16, Yan22]. Effective [ERT88]. Effectiveness [AP06]. Effects [AI11, AD08, AHL15, AL95, AC03, AN91, AJK18, APR19, AvdBL20, ACD09, Bar85, BZ13, BH09a, BH09b, BMM20, BCFFL19, BGS20, BW22, BCP19, CJW86, CGL13, CF15, CSZZ17, CG17, CHNY22, Das91, DE19, DMS99, DW85, Des07, DHS19, FHLZ20, FFL13, FLM^{+}19, FH94, FM13, GPI13b, GMP11, GH95, GPP98, HMS17, HK86, HW20a, HS16, HS99b, HMvdB12, HSZ22a, HLL20, HS19b, HSS20, JLS16, Juo18, KW22, Kan85, KNW22, KLS12, Kra07, Lec99, LPS22, LRW09, LGG22, LW97, Lu15, Mon97, Mou87, NW95, Pak11, PF19, PT18, PT23, QY19, RS22, Rus19, She19, SZ20, Sol86, ST22b, TGCD16, TN92, Tuz22, VV99, WZ20a, WLL18, ZT06]. Efficacy [GH91]. Efficiency [BDK07, BS87b, BL93, CF02, CGK97, GIR20, GL16, HH19b, KS08, KOS94a, Kum89, Ril91, RS07, ST10, ST89, WW09, WNG89, Yan22]. Efficient [BN02, BS89, BT14, CT01, Fdv00, GK08, GdXPE16, HS96a, Hei22, HV03, HJL22, HSZ22b, KJGV09, LYY22, LRW09, Liu14, NV12, QZL15, XZ22, YSD03, Zil97]. Effort [MS88a]. EGARCH [BJ06, VDP00]. EGLS [Ril91]. Elastic [CZ14, CHL18]. Elasticities [AES86, AWX20, Atk83, Bor85, BM93, GDG^{+}05, GHR87, MTLJ97, MF05, MY18]. Elasticity [Bor89, BM93, CFM11, GB87, Koh83a, Kow16]. Elderly [LT12]. Electric [Cal91]. Electricity [CL17, Har83, HK94, Koh83a, LS94, LA84, Smi00]. Eliciting [BdB18, DGJ20, McK18]. Eligibility [Fer97]. Elliptical [CZZ22, KS94]. Elongation [BC91]. Elusive [CFM11]. Email [WLL18]. Embodied [ML90]. Empirical [BHMR19, Blo19, CKY85, CMO20, CL87, CSZ22, CT22, CP98, FMT97, FP84, GK15, GJP02, Hau96, HP13, IRS15, Imb02, Jai89, JK02, KW04b, MJY19, Mar99, OV84, PS92, PT00, Ser88a, Shr89, Tao19, TV15, VZ11, Wan96, XP11, Zil97]. Employee [BFHM11, HMvdB12]. Employer [BFHM11, HMvdB12]. Employer-Employee [BFHM11]. Employer-to-Employer [BFHM11]. Employment [Ano84c, BS20b, CR83, CR84, FG00, GJP02, Iwa91, Jef08, KW04a, KT95, Lec99, MB84, NW91, NW95, PR98, Pow22, Pro12, Ren07, Spl00, Wan96, WN89]. Empty [VP92]. Empty-Cell [VP92]. Encompassing [Bey01, HLN98, Wes01]. Encouragement [FLM^{+}19]. End [BFS94, HMI95, LW97]. End-of-Life [LW97]. End-Use [HMI95, BFS94]. Endogeneity [CMT19, FM13, Wüt20]. Endogenous [AMS19, AAW11, Ang01, CFLS19, Car01, CHW20, Cui20, HL16, Hos18, HpK23, JS20, KP03, LL94, Lew12, OHS14, VV99]. Enduring [RW09]. Energy [DH88, FRS83a, MTLJ97, TH83, TD84]. Energy-Audit [TH83]. Engel [HK01]. Engineering [DH88]. Engineering/Econometric [DH88]. England [MS96]. English [CCP13]. Enhancing [ZC17]. Enrollment [NW95]. Entrepreneurial [PvP06]. Entropic [AFAA22, KCR17]. Entropy [CKY85, Rod16]. Entry [CT09, HH11a]. Enumeration [ZW93]. Environment [ARvD18, BGS20, Man15, MVVZ19]. Environments [BFR97]. EOV [Ano12a, Ano13a, Ano14a, Ano15a, Ano16a, Ano17a]. Equal [Bat90, CM09, HD21, SW22b]. Equality [KT19, Seo20]. Equality-Minded [KT19]. Equalization [Fre86a]. Equation [Atk83, BCR90, IP12, MY18, Sal95, SL20, WL21]. Equations [AT14, DW85, DTU06, Est98, FH90, Kra86, Liu14, MK01, MF05, TR86, Whi98]. Equicorrelation [EK12]. Equilibrium [ALT08a, Fli97, Ing90a, MT99, Xie22]. Equity [AFT20, BJ99, CG17, ES93, HC94a, HC94b, IRS15, IOS90, KMN05, KR94a, KR94b, LTU15, MM09]. Equivalence [Bel87, JS87, YSD03]. Equivalent [Góm99, Ver12]. Error [AWX20, AV86, BD01, CCKR17, CK19, Cha20a, CGQ20, CHNY22, FO05, FB99, Gar89, GL19, GK03, Has87, HD21, HI01, HT20, KP11, LZGL23, MP83, Mar87, MPTK91, MRW12, Mey88, NPF11, PSW04a, Pis95, PY17, Sch87, Str03, TR86, Vol22, YLL19]. Error-Coorection [FB99]. Error-Correcting [PSW04a]. Error-Correction [FO05, GK03]. Errors [AV05b, AS12, BDK07, BH84c, Cla94, Dev07, DG99, DV04, GK18, GHR87, HS96a, HK01, HT06, HL22, JM91, LLP15, LG89, MO88, May86, MJY86, Mor87, PFS98, RS92, RS15, TR86]. Errors-in-Variables [Dev07]. Establishment [McG95, PBC95, Spl00]. Establishment-Survey [PBC95]. Estate [HH11a, Ive01]. Estimate [GD89, JT95, Lew12, McC97, ST89, Vog98]. Estimated [FG20, GNK86, HS99a, Lec95, OF97, PPL94, SW22b, Wes01]. Estimates [AES86, AK95, AJK18, Bar85, BFHM11, Bor89, CJW86, CR01, CG12, HKKP01, JLG04, KW04b, Kli01, KK92, Koh83a, KMMP22, Kow16, MB88, Mor87, NW91, NS05, PV03, Pis95, QW20, Rad83, TH83, Van08]. Estimating [AZ85, AHL15, AWX20, BK05, BHKN10, BFS94, BHMR19, BM93, CCÖU19, Che18, CFM11, CG02, CGR07, CHL88, DW93, Des07, Dol99, DG06, EH90, Ell00, FZ13, Fre88, GW85, GKP00, GMS87, HR91, Ham91b, HS16, HLWY14, HLM15, JR03, Kol20, KL08, Kor10, KGM91, Kut94, LT19, LT00, MOT22, MZ23, NPF11, Pak11, Sta88b, Sto90, SMU13, TN92, Tsi22, VV99, XZW14, YCJK20, ZTCW15]. Estimation [Abr03, Aie13, AL89, ABA99, AS96a, AW07, AS96b, AC94, And02, Ang01, Are02, AGV19, BW15, BLL16, BRW12, BKT11, BG92, BD01, Bra90, BDY11, BS89, BC05, BWH86, Byk23, CCP13, CHL18, CL87, CFG95, CILY17, Cha20b, CGL13, CSZZ17, CGLS18, CGZ19, Che20, CT22, CG22, CS05, DO95, Das91, DE19, DBS84, DM95b, DNW98, DP22, DG02, EH23, Erg19, EO10, EPFK15, FLW96, FQX14a, FZ18, FHLZ20, FIL^{+}23, FXMFY20, FMT97, FFK18, FSC03, FS19, FS23, FE22, Fuh92, GI11, GLL93, GA10, GK14, GDG^{+}05, GK18, GRW84, Góm01, GG95, GSC96, GN15, GdXPE16, GHNS09, Had99, Han19, HS96a, HHN08, HMN10, Har83, HS96b, HK01, HP13, HW90, Hen22, HV03, HK17, HMvdB12, Hos19, HSZ22a, HLL20, HH83, HSA87, HF87, IP12, Ioa95, Isa90]. Estimation [JR90, JMZ18, JK02, JY21, JLK14, JGL19, KW22, KW90, KC85, Kas22, KR92a, KW97, Kie10, Kim90, KK12, Kou13, Kum89, LLN89, LS20a, LFWT18, LL19, LV15, Lav20, LS87, LYY22, LPV03, LR08, LRW09, LW19a, LLL20, LSCX22, Lia12, LS15, LR03, LW19b, Liu14, LJ18, LMN22b, LSU11, Lou07, Lu15, LG20, Luc00, LW18, MQZ22, MTLJ97, MV88, Mar87, MH17, MHCP22, MY18, MJY86, Mon97, MB13, MT85, MT02, NSS07, NEV15, Oku11, OXM13, PZZ20, PX21, PPR03, PRC98, PP11, Pfe91a, PFS98, Phi10, PZ19, Por02, PM20, Pow22, PT18, PT23, QZL15, RF01, RS07, RS94b, SJS18, She19, Sil01, Sol86, SRdCFB18, Str03, SMU13, SWJ19, SX20, kTY12, Tuz22, UY23a, Val88, Val91, Var16, Vel22, Vos16, Wat03, WW20, WW09, WP83, Wu19, XP11, XZ22, Yan20a]. Estimation [Yan20b, Yan22, YSD03, YYX20, ZWL22, ZG22, ZWZ22, Zil97, ZC17, vdKH89, vHPP04]. Estimator [Cui20, DH12, Die15c, FFX16, FZ18, GJ22, HC94a, HC94b, Hor98, JS20, Kew20, KR94a, KR94b, Lee18, Lum95, MSU10]. Estimators [AI11, AIPR18a, AIPR18b, AFAA22, AG02, AL18, BG23, BCHL20, CH14, Cla96, DL87, Dev07, DHL14, Gar89, Gbu89, GHPS22, HHY96, Har91, Hof91, HH04, HLM16, Kal22, KT18, Kot84, Kra86, MTLJ97, PB91, Por02, Por18, Ril91, SWW88, SW21, SK99, SM83, TZ19, Tau86b, VM89, WW96, Yit91, ZZWZ23, Zil97]. EU [WGBM12]. Euler [BCR90, Whi98]. Euler-Equation [BCR90]. Euro [LSZ14, VKR06]. European [AGO^{+}14a, AFO21, Car06, KK92, VP93]. Evaiuation [BCF94]. Evaluating [AP06, Deh03, GLLS11, IRR22, IV95, JJK20, Knü15, MT06]. Evaluation [Ang95, BF15, Cav22, GH22, GK05, Jan88a, Lec09, Rod84]. Evaluations [Bel04]. Event [HLR^{+}22, LK88]. Event-Dependent [LK88]. Events [BFP06, Gra19, LTTL19]. Evidence [AFT20, AMRT22, AOS00, BLS92, BGR93, BHMR19, Boh91, CD09, CMW23, CL87, CRZ19, CM96, DHZ10, DHW17, DMP19, FP97, GMP11, GH95, GB87, HW95, Hau91, HP13, HK94, HP85, KSM97, KD00, LLW90, LW19b, Lus96, Mer99, OP18, Ost19, PZ90, Rav06, Ril91, RDP95, SW96, YSD03, ZA92, ZA02, dNS23, HW04]. Evolution [Byk23, KMF05]. Evolving [MM19]. Ex [Cle14, KK15]. Ex-Ante [KK15]. Exact [Cre17, JSS88, Lan97, Nic89, Shi00]. Examination [KR88, RCW83]. Examining [BZ13]. Example [EG98, GPP98, Sta88b]. Excess [GT96, Nor94]. Exchange [AOS00, BB89, BB02, Bek95, BB87, CKY85, CST18, Che93, CE05, FKO97, GR89, Hsi89, Joh02, Kim90, Kla05, KK92, LN20, MS18, Mil87, Nie04, NPW93, PRM99, VP93, Wol87, Yil03, Zho96, ZLY17, vUS87, MN95a]. Exchange-Rate [AOS00, GR89, Kim90, PRM99]. Exclusion [CK19]. Exist [AK84]. Exit [CT09]. Exogeneity [EHM98, Eri98, FFK18]. Exogenous [GIR20, HJ00, KR92a]. Expanded [Ing90b]. Expectation [Fuh92]. Expectations [AKLS20, Aru20, BBK12, BL88b, Boh91, Goo96, GMP85, GSS07, HV87, Hof91, HRV20, KL90, LZ87, MM03, MM10, PRM99, She92, Zar85, dHM90]. Expected [ABD15, Bek95, BWH86, CD09, GS06b, KC85, KP11, SM83, Tay19, TV15]. Expectile [XZW14, XHL20]. Expectiles [DPS22]. Expenditure [DMS99, Kow16, SYA08]. Expenditures [PRC98]. Experience [Kas09, Kro83, Lit86b]. Experience-Productivity [Kro83]. Experiences [AGO^{+}14a, GM97]. Experiment [DPES19, FLM^{+}19, GO98, HK94, HW04]. Experimental [AL84, GH22, HHK11, KKB94, MS99, RW86b, War89]. Experimentation [TGCD16]. Experiments [Mey95, VB95]. Expert [CJJ^{+}23, Kie10]. Experts [CT09]. Explain [FV95, Ren07, Wat00]. Explaining [BK05, Pic01, WJS08]. Explanation [Nor94]. Explanatory [Ash85, HK17, NK84, PZ19]. Exploratory [AL84]. Exploring [FLM^{+}19, KL90]. Explosive [CNR20]. Exponential [HK19, HH16, PTZ86, WD96]. Exponential-Family [WD96]. Exponential-Type [HK19]. Exponentially [Öll86]. Exponents [GI11]. Exports [BB87, Koe06, Kon98]. Exposure [Dan92]. Exposures [BGR15]. Extended [Gag90, TT83]. Extensions [PV92b]. External [BI20, BB22, DPES19, Kor10, KCR17]. Externalities [APR19, CRZ19]. Extracting [CEJ^{+}12]. Extraction [FP01, Ghy87]. Extremal [Hog22, SW20, XHL20]. Extreme [EH23, LW19a, SS06, Swa03a]. Extremes [BMP10, NZ20, XHL20].

Facilities [CLR14]. Factor [AW00, BN07, BW15, Bäu13, BG23, CH14, CR16b, CR16a, CWW20, Che20, CZ18, DS98, DGP20, DS13, FFX16, GPD15, GL13, GL20, GB87, HHL18, Han19, HKYZ20, JLR11, JS20, Kap10, KLS12, LD19, LS07, LSX15, LS20c, MLST20, MPV16, OP15, OLBvD20, PX22b, RE12, SX20, TMH^{+}23, UY23a, UY23b, War89, ZLWW20a]. Factor-Adjusted [LD19]. Factor-Augmented [DS13, JS20]. Factor-Based [FFX16, SX20]. Factor-Spline-GARCH [RE12]. Factorial [ABD19]. Factors [AW07, BG23, BGR15, CH14, Che22, CZ19, CR88b, CGM16, DBCB19, FG20, GSC96, HL16, Han19, Kap10, LGLY16, NG89, Oku11, PS16, PX22a, Yam16, Yao08, YYX20]. Factors* [SW22b]. Fail [BKW19, Whi98]. Failure [BS98, SH05]. Fair [FG84, Pre95]. Fairness [Ano84c, CR83, CR84, FG84]. Fake [HS19a]. False [LKFL22, WHT22]. Families [ARBS87, GOV84]. Family [Des07, PTZ86, Rad83, WD96]. Fan [MRZ92a, NP85]. Farm [HL04]. Farmland [PB91]. Farms [KGM91]. Fashion [AJL96]. Fast [Gra97, HHKO22, LMN22a, Str22]. Fast-Food [Gra97]. Fat [OJLD18]. Fat-Tailed [OJLD18]. Faulty [KRW^{+}14]. FAVAR [BLL16]. Feasible [Phi10]. Feature [HLW14, ZPWW22]. Features [DM95b, EK93b, HHPS07, HI97, Urg07]. Fed [DF11]. Federal [RT01, AGO^{+}14a, GG90, JLG04, Pie83, Sal95]. Federal-Funds [RT01]. Federated [WHW22]. Feedback [Ghy87, McG91]. Fees [MV18]. Fertility [DPES19]. Field [FLM^{+}19]. Fields [LB16]. File [MS03, Rub86]. Filter [GLLM20, QW20, SS06]. Filtered [ACR19]. Filtering [BWH86, GGS96a, Góm99, YLYJ22]. Filters [BW84, CW86b, Góm01]. Final [Swa03b]. Final-Offer [Swa03b]. Finance [Era01, HPSW88, Urg07]. Finances [DG99, KSM97]. Financial [AGO^{+}14a, Ana09, AG10, AGK13, BH94, BS20a, BK12b, CT16, Che07a, DU12, GCC11, GS16, Gra87, HL04, IOS90, NZ20, PJ20, Qi99, Rac01, RE05, SS06, Tah01, Tau86b, YF22, Zha20]. Financial-Firm [BH94]. Financially [Ham13]. Financing [Wan03]. Findings [JR90]. Finite [BCHL20, FS97, Góm99, HHY96, HS99b, HL22, HLM16, KW04b, KDM20, Lum95, SW21, Swa03a, Tuc92, Val88]. Finite- [Tuc92]. Finite-Sample [HHY96, HS99b, Lum95]. Finland [AES86]. Firm [AC03, BH94, BHLM07, CFG95, Gra19, HHW91, HMvdB12, KBZ89, Lou07, MF05]. Firm-Level [MF05]. Firm-Specific [CFG95, HMvdB12]. Firms [DBS84, HJL22, MlM87, PW10, VLK04]. First [AGV19, Bor89, Don22, LPV03, LW18, Ray90, Tau93b]. First-Choice [Bor89]. First-Order [Ray90]. First-Price [AGV19, LPV03, LW18]. Fit [BCCP18, CDF15, DSSW07a, Est98, Rit86, ZZZS19]. Fitting [PSSE20]. Five [Lit86b]. Fixed [AN91, Asa02, BH09b, CGL13, Das91, DP06, DTU06, JS20, KW22, LGG22, Mon97, PT18, PT23, RSW18, SW20]. Fixed- [SW20, RSW18]. Fixed-Effects [Das91]. Fixed-T [JS20]. Flat [Var16]. Flat-Top [Var16]. Flexible [Ant83, Bar83b, BBK12, Cha87, Cha20a, GE12, HMN10, HHKO22, KG15, KOS94a, Kum89, LZ87, Mos99, Ram94, RFG20, RW98, Ser88b, SY94, Swa03a]. Florida [GB87]. Flow [ZHCB16]. Flows [AZ85, BHLM07, BFHM11, GS16, HHPS07, HF86, Mer99, Sta88b]. Fluctuations [Coo98, ST22a]. Fluency [DV04]. Focused [KW19, YLL19, ZWZ12]. Fold [Lia12]. Follow [BL93]. Follow-up [BL93]. Followers [PS16]. Following [DHS19]. Food [DP19, GDG^{+}05, Gra97, HH83, PRC98]. Force [AZ85, GLL93, GD89, HF86, Sta88b]. Forecast [ACvD22, BK15, CT09, CL17, Cle14, GL19, GSY01, HLN98, HH11b, Ilm90, JS98, KV13, KZ20, Mor87, PT12a, PP11, Qua19, RI12, SV16, TH89, TR86, Vol22]. Forecaster [Has87]. Forecasters [AGK13, Bat90, EMW09, GW09, JS19]. Forecasting [AGO^{+}14a, AV07, And94b, AV08, BRY13, BCM22, BK15, BW19, BB87, BT89, BS20b, BJ06, Can92, CR16b, CR16a, CF99, CGLS18, Che07b, CD98, CG00, CG08, CR12, DM01, DK00, FK86, Fom86, FS91, Fre92, GFHPZ87, GCC11, GW09, GPR13, HT83a, HPP00, KLP08, KK12, Kor19, LTU15, LM90, Lit86b, Lit86c, LMS20, LSU11, Lüt84, Lüt86, MM09, Man09, Man15, Mar83, MPV16, McN86a, MVVZ19, MG84, Mit90, Oli87, OS89, Ott90, PFA00, PR98, PY19, PT17, Phi87, RW09, SS15, Smi00, SW22b, SW02, SW12, Tay19, TM86, TMH^{+}23, Tsa93b, WJS08, Wol87, Yam16, YF22, YYX20, ZH92, dA88]. Forecasts [AG07, Ash85, BK12a, Ber01, BP18, Cha93a, Cla11, CW86a, Die88b, Dua88, Due05, FW09, GK05, GR11, GW88, HH11b, IRR22, Kan86, Kim22, Knü15, KCR17, LM92, Liu21, MMO22, McN86a, NP85, Pat19, RT01, Ste87, Wec89, Wes01, Zar85]. Foreign [AOS00, Bek95, CKY85, Che93, Hsi89, Joh02, LN20, NPW93, YP08, Zho96]. Foreign-Exchange [Bek95, CKY85, Che93, Hsi89, Zho96]. Forests [FGK20]. Form [FW09, HHW91, KOS94a, Kum89, LP96, ML88, McK18, Olm09, OLBvD20, Sim90, Ter85]. Formalization [Bey01]. Formation [HHK19, HV87, Joc18, Kor10, Yog08]. Forms [GO98, HI01, Ram94, Ser88b, SY94]. Forward [NPW93]. Fractional [CNT20, CL93, LV15, Lou07, Nie04, Nie07]. Fractionally [CdM19, TWW13]. Fragmentary [FLTS19]. Frailty [KLS12]. Framework [CZ19, CM93, Deh03, DGJ20, HLZ22, HS19b, Joh87, KMF05, MSW85, ZG22]. Franc [vUS87]. Francesca [Gua23, Hon23, Mor23]. Fraud [BCCP18]. FRED [MN16]. FRED-MD [MN16]. Free [BFZ98a, Goo96, GL13, HH11a, HHW91]. Freedman [Hil86, SB89a]. Frequency [And00, Bab22, BG05, BK12b, BCP19, BBCL20, CDS20, CG22, CG08, CP19, DZ04, FFX16, Ghy00, GMP11, Kal22, MPV16, MH17, PM20, RE12, SS15, SRdCFB18, SX20, Tsa00, kTY12, Zho96, CMO20]. Frequentist [BF15, ZWZL22]. Friedman [Sto88a]. Frontier [CFG95, CHW20, FLW96, GHY92, Had99, HW20b, HJL22, HpK23, KGM91, ST89, SW22a, Wan03, Wil93, YZK19]. Frontiers [ABS99, Kum87, LYZ22, SS84, Tsi22]. FTSE [ACW97]. FTSE-100 [ACW97]. Fuel [Das91]. Full [HJ00, SM83, VDP00]. Fuller [CL95]. Function [Abr85, BP22, BH94, BH09b, BV05, CLQ20, CFM11, Chr90a, Chr90b, Col90, Gor89, GIR20, Had99, JT95, JK02, JLG04, dRJH11, KFL03, KOS94a, LT00, Lin97, NSR88, Tau90, Yan22]. Functional [CL17, CGQ20, GO98, HHW91, HLR^{+}22, KMR19, Kum89, ML88, Ram94, Ser88b, SY94, SMU13, WJS08, ZL22]. Functional-Form-Free [HHW91]. Functionals [Rus19, Yan20b]. Functions [AN91, BG92, CG22, DW88, Hay86, Hen22, IRR22, LR08, LLR13, Mac94, Ser95, Tao19, XP11, YCJK20]. Fund [JLP23, LD19, WHT22]. Funds [RT01]. Further [Boh91, CC97, GKR12, JSSvN20, SM83, ZA92, ZA02]. Future [BV87, But85, TM86, Wai85]. Futures [DF11, MT99]. Fuzzy [BI20, FLM16].

Gain [NP90]. Gains [PT03]. Gambling [Goo96]. Games [ALT08a, KS20, Xie22]. Gamma [JMZ18]. Gap [CLLM03, PRF08, QW20]. Gaps [CG12, Hub15]. GARCH [AT14, AT11, CG19, CS20, Due97, FQX14a, HK19, HK14, HH16, Hog19, Kar95, KSW22, LL90, LS07, LK93, LLL11, Lin97, Lum95, MT99, NC92, RE12, VDP00, Cav22]. GARCH-X [HK14]. Gas [LS94]. Gasoline [OG86, Oht87]. Gaussian [And01, BDK07, CZZ22, Cre17, Die15c, FS19, FH98, GA10, HLWY14, JS10, KL08, KLS15, LL19, OT21, RFG20, SV16]. GDP [AGJT14, ÇD22, CG17, FS22, Gew88b, HPP00, JSSvN20, KV13, KMMP22, LdV99, MPV16]. Geary [RS92]. Gender [CLLM03]. General [AT11, Cat20, Qua88, TMH^{+}23, VB95]. General-Liability [VB95]. Generalization [BGMP21]. Generalized [Abe99, Ali87, AFAA22, AB22, And02, Are02, BH94, BD09, BL05, BCW16, BLC20, BN02, CZ14, CPDH15, Chr96, GH02b, GJ22, HS99a, HW02, HH19a, HPZZ21, Imb02, JSW02, JLG04, Kew20, KGM91, LS07, NS10, Phi10, SWY02, SW12, Swa03a, Tau86b, TT02, Xu00, YLN17, ZK05, ZWL22]. Generalizing [MS88a, PR98]. Generally [RSW18]. Generated [CO99]. Generating [ARBS87, Bor89, GK96]. Generator [FB99]. Generic [BL93, KZ20]. Genetic [DM95b]. Gentkow [Tam20]. Gentzkow [Bon20]. German [Bey01, OF97, VLK04]. Germany [Lec99, LW09]. Getting [SvD06]. Giacomini [PM22, RR22, Kil22]. Gibbs [AC93, Due99]. Gibbs-Sampling [Due99]. Gini [BD09, BFZ98b, GNK86, SWW88, Xu00, Yit91]. Gini-Based [BFZ98b]. Give [FL97]. Glass [PW10]. Global [AGO^{+}14a, DPES19, DFP11, FFX16, Han19, LTU15, MM19, PSW04a, RKL22]. Globally [Cha87]. GLS [Wes18b]. GMM [Chr96, AS96a, AS96b, BE96, CHL18, CdH96, GKR12, HHY96, KM09b, LL19, LYY22, Mon97, SW21, SMU13]. GMM-Based [BE96]. GNP [Bru92, CC97, Chr92, DM94b, Mor87]. Goals [JS85]. Gold [CM88]. Good [Hay89, LB16]. Goodness [BCCP18, CDF15, Rit86, ZZZS19]. Goodness-of-Fit [BCCP18, Rit86, ZZZS19]. Goods [Kim05, Pak11]. Google [BS20b, FS22]. Got [Ost84]. Government [Hau91, JS99]. Gradient [HS16, WHW22, YQZ18]. Gradients [LLN89, MMG^{+}06]. Granger [BRT12, CT16, Kan85, ST18, ST20]. Graphical [GS16, KS96a, WLTL20]. Great [Cam07, MA23, ZA92, ZA02]. Greedy [CILY17]. Greenbook [FW09]. Grid [LXY22]. Gross [AZ85, Bra90, Ghy90, HF86, Sta88b, dL90]. Group [BBD20, CH14, CILY17, FZ18, Han19, LNB10, Mou87, Rot15, Sol86]. Group-Specific [BBD20, Han19]. Grouped [AB22, Che18, Che22, Das08, Deh03, Dev07, GNK86, HGB^{+}12, KYK^{+}21, Smi22]. Grouping [GZ84]. Growth [BCR90, BS20b, Cal91, Chr90b, Cle14, Coc89, Col90, DM94b, Gag90, HPP00, Ing90b, KOS00, Lab90, McG90, NW91, PT11, Sim90, Spl00, Tau90, TU90, dHM90]. Growths [LdV99]. Guest [CHT17]. Gun [BCHK16].

H [Gas85, KT18, Mül18]. Habit [AKO12, Kor10, Li05, Yog08]. HAC [Mül14a]. Half [CD09, Ros05]. Half-Life [Ros05]. Hamilton [QW20]. Hansen [Ano02c, Bur94]. Happens [Gra19]. Happiness [HS16]. HAR [Mül18, Wes18a, LLSW18a, LLSW18b, Vog18]. Hard [BC89b]. Having [Vog98]. Hawaiian [CF99]. Hayek [KB10]. Haynes [You88]. Hazard [FHDS11, JT95, LTY^{+}22, SC03]. Headed [Ost84]. Health [CW96, DMT06, Edw08, FHDS11, GH95, HLR87, LT12, LW97, SY98]. Health-Care [CW96, HLR87]. Heat [LN20]. Heavy [CKL11, DPS22, FQX14a, FZ13, LS20c, YL17, OJLD18]. Heavy-Tailed [CKL11, DPS22, FQX14a, FZ13, YL17]. Hedging [FID16, RW22]. Hedonic [BB05, DZ04, Mel05, Oht87, SH07]. Help [KV13, MP05]. Henderson [DB08]. Herd [CSZ22]. Here [Ost84]. Heterogeneity [AH20, AKLS20, AB22, BCM10, CGJ19, CMHM19, CHNY22, FLS07, FSTO04, GR12, GK15, Haa96, HSZ22b, IM10, Ioa95, JLS16, KDM20, Kea97, Luo02, LPS99, PW22, Pro12, San20a, Smi22]. Heterogeneous [AT17, APR19, CMM23, EH23, LZW^{+}22, SV88, TGCD16, VZ11, Wil09].

Heteroscedastic [CNT20, DH12, DS98, FSC03, Had99, HT22, KST11, LMS20, LS22, RS15, TS01, XHL20, ZL15]. Heteroscedasticity [AT14, BAT83, BL05, BHL92, BG96, BH93, CFG95, CHNY22, Dav04, DNW98, Due99, Gre89, Hsi89, Kim93, KW92, Lew12, LZZ20, ML88, MY18, TT02, VP93, Wes14, XZ22, ZK05, ZLY17]. Hidden [ABD19, BCD17]. Hierarchical [Deh03, Fdv00, KW04a, RAL99, WK91]. High [And00, Bab22, BCHK16, BOP23, BCP19, BBCL20, CCKR17, CDS20, CZZ22, CP19, FFX16, FHWZ18, FHLZ20, FLM20, GI19, Ghy00, HHL^{+}22, Kal22, KLZ20, KKS22, Kor19, Koz20, LCZT22, LZ95, LS20c, LKFL22, LSCX22, LXHT22, LTY19, MLST20, OP15, PX21, PM20, RE12, SX20, ST22b, TMH^{+}23, Tsa00, kTY12, WKX23, Zho96, CMO20]. High-Dimensional [Bab22, BCHK16, CCKR17, CZZ22, FHWZ18, FHLZ20, FLM20, HHL^{+}22, KLZ20, Kor19, Koz20, LS20c, LKFL22, LSCX22, LTY19, MLST20, PX21, ST22b, TMH^{+}23, WKX23, LXHT22]. High-Frequency [And00, BCP19, BBCL20, CDS20, FFX16, Ghy00, Kal22, SX20, Tsa00, kTY12, Zho96]. High-Order [GI19]. High-School [LZ95]. Higher [BH93, BT01, Cav22, CG22, Kew20, Vel22]. Higher-Order [BT01]. Hiilmer [Sim85]. Hilbert [DB08]. Hiring [FG00]. Historical [FM97, LTU15, MM09, Siz13]. Histories [MT96]. History [Woo00a]. Hoeffding [Cla94]. Homogeneity [FH90, GL20, LQZ19, SY94, ZWLL22]. Homogeneous [KLZ20, VZ11]. Homophily [Mel22]. Homothetic [Atk83, HH19b, Koh83a]. Homotheticity [Bar83a]. Horizon [CD98, PT11, PT12a, Qua19, Siz13]. Hospital [LC85, RAL99]. Hours [Xie22]. House [BGKS04]. Household [AKLS20, AHI99, DH89, Har83, HP89, JS87, KSM97, KSS09, MF05, MS97, MS89, PRC98, SYA08]. Household-Level [MF05]. Households [DLP19]. Housing [ACS15, BK05, GZ84, PRC84, Ran88]. HPV [HS19a]. Human [ZCZ^{+}21]. Humps [ABMW99]. Hurst [LJ18]. Hybrid [Cha22]. Hyperbolic [LLL11]. Hyperparameters [AAMW20]. Hypotheses [BLS92, KSW16]. Hypothesis [Boh91, BV05, FKO97, Goo96, HH19a, KJAG85, KN89, KB10, KS94, ML90, MS89, PdV06, PJ14, PT18, PT23, She92, TS01, WZS22, ZO04, ZA92, ZA02, lT85].

Ichimura [KT18]. Ideal [Cha87, FS97, Ram94]. Identifiability [Bos96]. Identification [Agu10, AN91, AKKL21, ALT08a, BW15, BKT11, BB20, BH09a, BdB18, BB22, BN12, CdM19, CR88b, DF11, DGJ20, DP06, DP22, DLP19, FLM16, FFL13, GPI13b, GIR20, GL20, JLS16, Kas22, Kea92, Kli16, KCF^{+}15, KPS13, KP95, KP11, Liu14, MLR19, MY18, PS16, SWY02, Tsa85, TWW13, Vel22, Vol22]. Identification-Robust [AKKL21]. Identified [AK13]. Identifiers [AV05b]. Identify [Lee18, Lew12]. Identifying [AT17, DU12, FL97, Gos10, JJK20, LLL22, MM00, San20b, SZ20]. Identities [McE17]. Ideological [AT17]. Idiosyncratic [GS06b]. IGARCH [Lum95]. Illusion [Smi05]. Illusions [KLN85]. Illustration [JJR18]. Illustrations [CT93]. Immigrants [DV04]. Impact [DP19, McC95, MM19, TR86].

Impacts [AAW11, MN95a]. Implementation [HP13]. Implementing [Sil86]. Implications [Atk83, BMP10, CP98, Cle85, CGK97, FOTY20, IM97, KBR95, Koh83a, Kou13, MS88a, Sea93, ST89]. Implicit [Guo98]. Implied [ABD15, FE22, GW14, Nie07]. Implied-Realized [Nie07]. Import [Gbu89, Kim90]. Importance [KLS15]. Imposing [KFL03, Mos99, RW98, SY94]. Improve [GI11, NW91, PG93]. Improved [BN02, Dev07, JSSvN20, SB19]. Improving [BL93, ÇD22, CG12, FM97, KW04b, RT01]. Impulse [CG22, GMP11, Koc10, KOS94b, Lin97, NSS07, Wri00b]. Imputation [KRW^{+}14, SH07]. Imputations [LCZT22, MS03, Rub86]. In-Patient [RAL99]. In-Sample [CR16b, CR16a]. Inactive [ZCZ^{+}21]. Incentive [TN92]. Incentives [BDJ12, LPS22]. Incidence [JR03, PSD14]. Incidental [JR22, Wes15]. Including [DBCB19, FH19]. Income [AES86, BS87a, BCF94, BM93, BM87, CM90, Che18, CPP18, CGR07, CHLR92, Dag87, DH89, EN94, FS88, GOV84, Gra97, GK15, HGB^{+}12, HS16, KJAG85, LS16, Lus96, MZ23, NG89, NPF11, Pak11, PRW06, Rad83, RS94a, RS94b, Sch83, Ser88a, Sho11, Slo87, ZO04]. Income-Happiness [HS16]. Income-Induced [BS87a]. Income-Share [BM93]. Incomparable [APW20, Tar07]. Incomplete [BFR97, Hos19, Tim01, VSV99]. Incorporating [DGJ20, FFX16, WLTL20]. Increase [GD89, KT95]. Increased [AKO12]. Increases [AAW11, Kas09, OG86]. Increasing [Hen22]. Increasingly [FHDS11]. Independence [BRT12, CT93, CJZZ18, Fra19, HLM16, KP11, MRZ92a, RF15, ZZZ20]. Independent [HH21]. Index [AIPR18a, AIPR18b, Ano83b, Ano84e, Ano85c, Ano86e, Ano87e, Ano88b, Ano89b, Ano90b, Ano91b, Ano92b, Ano93b, Ano93d, Ano94c, Ano95b, Ano96e, Ano97c, Ano98b, Ano99b, Ano00b, Ano07b, Ano08b, Ano09b, Ano10b, Ano11b, AL18, BCSS20, BGR93, Bla89, Bra07, CGL13, CLQ20, CJJ^{+}23, CP10, DM84, DO95, FHWZ18, Fom86, HD85, HF97, HW20a, HKKP01, JY21, KWD13, Ker85, KT18, Kim22, Kot84, LLY20, LTY^{+}22, LQZ19, LM94b, McC97, MOZ07, NWK10, NTP94, Nor99, Por18, Ran88, RS94b, Sel89, SCS20, Shr89, Son10, Tho20, TR93, XZL20, ZWL22, PT00, Ver12]. Indexes [ACW97, BB05, BO01, BFZ98b, DZ04, GFdH87, HT06, Jai89, Kok87, Kri17, Mel18, MS99, Oht87, Rei99, SH07, SW02, VM89, Val88, Val91, dHK14]. Indexing [GG90, JS99]. India [FP97, Tar07]. Indicator [KR88]. Indicators [GMP85, LM89, LeS91, MOZ07, PSvD09, Shr89]. Indices [Bar83b, BD09, GNK86, IRS15, Ive01, PQ10, Xu00]. Indirect [DHS19, GS03, HSS20]. Indirectly [Mos91]. Individual [HS19b, KW04b, LLL19]. Individual-Level [HS19b, KW04b]. Individual-Specific [LLL19]. Induced [BS87a, GP17, UY23a, UY23b]. Inducing [HKO20]. Industrial [ACD09, FFX16]. Industries [Mor93, PS96, YCJK20]. Industry [ABS99, Che12, FB92, HH11a, KLS12, Mor85, Rog87, TD84]. Inefficiency [AMRT22, CFG95, HH11a, KGM91]. Inequalities [BSS22, FW05].

Inequality [BD09, BSS01, BCF94, Bla89, BM87, Cla94, DFK19, FBB06, FS88, GNK86, MN88, NC92, Qia19, RS94a, Slo87, ZFSC00]. Inference [AC93, ACR19, And18, AKKL21, BLL16, BD09, BCHK16, BCW16, BFZ98b, BCHL20, BN02, CGM11, CR16b, CR16a, CNT20, CI03, CDY13, Cre17, DPS22, DH20, FSC03, GK14, GK08, Gos02, Gos10, GS91b, GS03, GM83, HGB^{+}12, HJNR98, HHKO22, HHPS19, HW20a, HL22, HW04, HLL20, HLQS17, IM10, JL22, Kal22, KK15, Kim22, KCF^{+}15, Koz20, Lan97, LL07, LLSW18a, LLSW18b, LSX15, LXHT22, LYZ22, LZ19, MJY19, MQZ22, MNW19, MM20, MOT22, MV18, MHB19, Mül18, MW22, MT85, MT02, NV12, OXM13, PRT00, Qia22, QY19, RS15, RSW18, SW20, She19, SK20, ST22b, TM19, TK10, UY23b, Vog18, VDP00, WHYC22, WV11, Wes18a, WZ20a, WZ15, WLL18, Xie22, Xu00, YL17, YYX20, Zha15, ZLS19, ZFSC00, ZWZ22]. Inferences [CFLS19]. Inferring [AHI99, Fuh92, GP17, MT96, BKW19]. Infinite [BCD17, HCP11, Li11, Tuc92]. Infinite-Variance [Tuc92]. Inflation [Aru20, Bie00, BH93, BWH86, CKP13, Cha15, CI87, CG12, Cle14, CMHM19, DG06, Dur98, GPR13, HW95, IM97, Jai19, Kim93, Los85, MVVZ19, Met98, PT11, SH05, Ver12, XKZ21, dBE98, lT85, Shr89]. Inflationary [Cle85, GMP85, KL90]. Influence [AJL96, GM97, Kra86, NG89, Sil01, WLZT22, ZK05]. Influences [MSW83]. Influential [Joh87, Peñ90]. Information [Bäu13, BMY04, BJ99, ÇD22, CFS09, HJ00, HH11b, Hos19, Isk20, IOS90, JMP16, KLP08, Kie10, KW19, Nev03, PG93, PR87, SS13, SM83, SW95, Tim01, TW14, YLL19, ZWZ12, dJ00]. Information-Theoretic [KLP08]. Informational [BS87b]. Infrastructure [MS96]. Inherited [Sho11]. Inhomogeneous [GHS09]. Initial [BZ13]. Innovation [AC03, GK08]. Innovations [CDJW10, FSC03, KHT05, TY89, WS20]. Input [HK86, MS96, She93]. Input-Output [She93]. Inputs [AD08, HpK23, SW22a]. Instability [Han92, Han02, PT17, Ros98, SW96]. Instantaneous [GR15]. Institutional [TH83]. Institutions [CO99]. Instrument [FGK20, FFL13, KM09b, She93]. Instrumental [ABA99, AK95, BFV22, Bre19, FZ18, HHN08, HMN10, HK01, HH04, Kli16, Kli01, Kow16, PPL94, ST22b, WW96, WHKCS14]. Instrumental-Variable [ABA99]. Instruments [BB22, HHL18, IP12, KR92a, KCF^{+}15, Lee18, Mon97, OP13, RS22, SWY02, TW14, Zil97]. Insurance [DMT06, Fer97, GM97, GH95, HJ00, McC95, VB95, WTC84, YQZ18, Yan20a]. Insurance-Market [VB95]. Integrated [Cui20, JLK14, LW18, NV12, PY09, Phi87, SX20, TWW13, Yan20b]. Integrated-Quantile-Based [LW18]. Integration [HMI95, INT21, LD96, RU01, TS01]. Interaction [BHL92, HHPS07, Hos19, LKFL22, RS22]. Interactions [BHKN10, HHK19, Pic01]. Interactive [CM88, CHNY22, HSZ22a, Juo18, KW22, KNW22, KS96a]. Interarrival [MABF08]. Interbank [CLR14]. Intercept [Knü09]. Interdependencies [PSW04a]. Interdistributional [BM87]. Interest [AB02, Aru20, Aud06, BHW09, BI15, Bie00, CG00, Coc89, Dua88, EG98, Haa96, HLZ04, KK15, KMV10, LM94b, LSS19, Mis95, RW96, She92, Smi02, XKZ21, ZC17]. Interest-Rate [BHW09, CG00, RW96, She92]. Intergenerational [AWX20, Sho11]. Interindustry [PR98]. Internal [Kor10, ML90]. Internal-Adjustment-Cost [ML90]. International [AFT20, BLS92, CsL94, ES93, GFHPZ87, GS16, HW95, JJR18, Kar95, Mit90, RS92]. Interpolating [Sil86]. Interpret [Sal95]. Interpretable [PX22a]. Interpretation [FL83, LM00, LLN89, SB89a]. Interpretations [ZH19]. Interpreting [Kli01, VV99]. Interregional [CE08]. Intertemporal [EH90, GK18, Koo91, NRW01, Pak11, Tah01, Web00]. Interval [Cha93a, FLSVP04, GRL13, YT07]. Interval-Valued [GRL13]. Intervals [GIR19, GPD15, Hog19, Kim01, KSD98, Ros05, SOK01, Wri00b]. Intervention [vdKH89, GPP98]. Interventions [Yil03]. Interview [Ano02b, Ano02c]. Intraday [DF11]. Intrahousehold [Blo19]. Intratemporal [Pak11]. Intrinsic [NSS07]. Introducing [GH95]. Introduction [Ang95, CHT17, Chr96, Eri98, GH02b, GH02a, Ham94, JW99, KW97, Ley85, Tsa00]. Intrumental [Are02]. Invalid [CHL18, FFL13, KCF^{+}15]. Invalidity [GHLT19]. Invariant [HK17, Shi00]. Inventories [Ros98, Sic94]. Inventory [GRW84, OV84]. Inverse [And01, And18, CZZ22, DHL14, FLM20, QZL15, Sal95]. Inverse-Control [Sal95]. Inversion [LMS20]. Inverted [HK94]. Invertible [AIPR18a, AIPR18b, AL18, KT18, Por18]. Investigation [CC97, CP98, Lum95, MS18, Ray90, Sch89]. Investment [Asa02, CHL88, DTU06, FP97, FB92, Gra19, HL04, KC85, Kas09, Mor93, Tah01, Wan03, Whi98]. Involved [BH84a, BH02, RCW83, Sim85]. Involving [CM88, Har91]. Inward [WLZT22]. IPO [HLR^{+}22]. Iron [TD84]. Irregular [DM95b]. Irreversible [FP97]. Isaiah [Bon20]. Israel [GMP85]. Issue [BFT16, GH02b, GH02a, KW97]. Issues [BH84a, BH02, DM84, Mel18, RCW83, Sim85]. Itô [KSW22]. Iterated [MB88, Phi10, Xu00]. Iterated-Bootstrap [Xu00]. Iteration [Chr90a, Chr90b, Col90]. Iterations [Tau90]. Iterative [PPR03]. IV [Bab22, DM10, GIR19, HHL18, SMU13].

J [KT18, Mül18, Tam20]. Jackknife [Yit91]. Jagannathan [Bur94]. Japanese [Mor92]. JBES} [Ang95, KW97, Ano95c, GH02b, Gal93, Gew93, Tau93c, Zel93]. Jenkins [HT83a]. Jesse [Bon20]. Jiashun [MLZ22]. Jin [MLZ22]. Job [BS20b, CF15, Dos05, Gar89, HMvdB12, Lan97, Lec99, VG97]. Job-Match [Gar89]. Job-Search [Lan97]. Joe [You88]. Joint [DRV20, FKO97, MOT22, Rus19, Zha20]. Jones [ACD09]. Journal [GH02a, LB16, Tau93b, TR93]. Journals [Ano83c]. Judgment [Man09]. Jump [BMP10, BLC20, CM02, DU12, HW10, IRS15, KS17, Kol20, LTTL19, ZZZS19]. Jump-Diffusion [IRS15]. Jumps [DNW98, DU12, DP13, HP13, HPS96, JLK14, LN20, PRM99, TT11, VP93]. Juvenile [LLPZ20].

Kalman [BWH86]. Keeping [BH84c]. Kelly [Pel21]. Kennedy [CJW86]. Kernel [CT22, CS05, DB08, GHNS09, HLWY14, RLYZ22, Var16, WW20, Wu19, YCJK20, Yan22, ZLS19, ZZWZ23]. Kernel-Weighted [RLYZ22, YCJK20]. Kernelized [LZGL23]. Kernels [LSS16]. Key [LLPZ20]. Keynesian [DSSW07a, KM09b]. Khamis [RS92]. Kind [FG84]. Kink [Han17, ZWZ22]. Kitagawa [Kil22, PM22, RR22]. Known [EJP05]. Kolmogorov [CPP18]. Korea [CPP18]. Kronecker [Cha20a]. Kurtosis [BN05, Cav22].

L [DHL14]. Labeling [HYSW16]. Labor [AZ85, AAW11, Blo19, BMM97, Bra90, CFM11, GLL93, GD89, GH95, HF86, JGL19, Koe06, Lec09, Mey88, Oku11, RW86b, Sta88b]. Labor-Force [AZ85, GLL93, Sta88b]. Labor-Marker [GH95]. Labor-Market [BMM97, Bra90, Mey88]. Laboratory [HW04]. LAD [WLJ07]. LAD-Lasso [WLJ07]. Lag [BCP19, CL95, HL83, LS20a, Shr89, Tsa85]. Lagged [BCP19, DM99, May86]. Lagrange [Hal90]. Land [MMG^{+}06]. Landsat [HF87]. Laplace [Cui20, Tay19]. Large [AW07, AB22, BT22, CCÖU19, CH14, CCM16, Cha20a, Cha22, CHNY22, DM16, ELW19, FFX16, FW09, GK14, GNK86, GL13, GL16, HKYZ20, Kap10, LLT^{+}15, LLF22, LHYZ22, Lit88a, MHCP22, NWK10, PX22a, PX22b, Yan20b, YLL19, vHPP04]. Large-Dimensional [HKYZ20]. Large-Portfolio [DM16]. Large-Scale [AB22]. Lars [Ano02c]. Laspeyres [VM89, Val88]. Lasso [HSZ22b, WLJ07, HJL22]. Last [Vos16]. LATE [CLT22]. Latent [BJ99, DG06, Ger88, Kas22, Kut94, LGLY16, MN20, NW13, SWJ19]. Latent-Variable [Ger88]. Later [Die15a]. LATEs [Lee18]. Laurent [Bar83b, Ram94]. Law [AB88, AL89, FZ13, BCCP18]. Layer [XZW22]. Lazarus [Mül18, Wes18a]. LDC [AOS00]. Lead [BCP19, PSvD09]. Lead-Lag [BCP19]. Leaders [PS16]. Leading [KR88, LM89, LeS91, MOZ07, PSvD09]. Leading-Indicator [KR88]. League [Kro83]. Learn [HS89]. Learning [BGS22, BCSS20, CKMS22, DHW17, LCZT22, McD87a, MVVZ19, WP83, WHW22, ZCZ^{+}21]. Least [Hag83, HLR87, Jan88a, Lab90, NV12, Phi10, YCJK20]. Least-Squares [Phi10]. Left [DMP19, PPS19]. LEHD [MA23]. Leisure [SY98, SW88]. Lender [Vos16]. Lending [CLR14]. Lens [JS19]. Leptokurtic [LLL22]. Less [BH93, KNP04]. Less-Volatile [KNP04]. Lessons [BdB18, Lav20]. Let [HO84]. Letter [Ano86f]. Level [ACL11, Bal93, CT90, DRV20, GK18, HS19b, HW10, INT21, KW04b, MF05, MHB19, NS96, PV92a, PQ10, Smi05, WZ00]. Level-Shift [CT90]. Levels [BH93, FBB06, FL83, McC95]. Leverage [Cat20, CR12, LS20c, Yan22]. Levon [Gua23, Hon23, Mor23]. Lévy [BDY11, TT06]. Lévy-Driven [BDY11, TT06]. Lewis [Mül18, Wes18a]. Li [MLZ22, SK91]. Liability [VB95]. Life [Cag03, LW97, MS89, Ros05].

Life-Cycle [MS89]. Lifetime [ABMW99, RR97]. Likelihood [AG07, CSZ22, CT22, Cre17, DG02, FQX14a, FSC03, FH98, GHPS22, GK03, Imb02, Lum95, MJY19, MB13, Por02, RH04, SM83, Tao19, TA89, Tuz22, Wu19, CMO20]. Likelihood-Based [GK03, MB13]. Likelihoods [FQX14a]. Limit [MT99]. Limited [Ang01, BM23b, CGR07, Gua23, Hon23, Mor23, PRM99]. Limited-Dependent [PRM99]. Limiting [GKR12]. Linear [AB22, BCW16, CS86, CGL13, CGQ20, Chr90a, Chr90b, CZ18, CHNY22, Dav04, DGJ20, Die15c, GK14, GGS96a, GJ01, Gre94, GPP98, GL13, HK19, HHL18, HHL^{+}22, HS96a, HH19a, JS20, KW90, KRW^{+}14, McC19, McG91, McG90, Meg87, Mof86, MW22, OV84, RW22, SY94, SMU13, SW95, Tyr83, VLS90, War89, WW96, Yit96]. Linear-in-Variance [HK19]. Linear-Quadratic [Chr90a, Chr90b, McG90]. Linearity [HP10]. Link [CR12, Kim93]. Linkage [Str22]. Linkages [BCP19, EH19]. Linked [BFHM11, ZSZ20, ZHB23]. Links [Wan96]. Liquidity [CLR14, Coc89, FID16]. List [ZW93]. Listed [MKL07]. Litigation [Fre85, GW85]. Litterman [Sil86]. Living [BS87a, CP10, GG90, HD85, JS99, Kok87, KPS00, Nor99]. LM [DRV20, HMS17]. Load [AL84, Koh83a, Smi00]. Loading [BG23]. Loadings [Kri17, OLBvD20]. Loan [PS92]. Local [BLC20, DPES19, DFP11, FFL13, GIR19, GL19, GHNS09, Hei22, HZ22, LT19, LTU15, Mos99, Nie07, PLCW22, PM20, RW98, SMU13, ST22b, Zha15]. Locally [Can98, DW20, LK93, OT21, WV22, XKZ21, Zha15]. Location [Bar85, BK91, FLM20, MJY86, RV19]. Location-Scale [RV19]. Log [HHL^{+}22, OV84]. Log-Contrast [HHL^{+}22]. Log-Linear [OV84]. Logistic [FMT97, Hag83, LXHT22]. Logit [AR91, BS89, CS86, Con90, JVC94, KW04b, LC85, SV88, Tse87, VSV99]. Lognormal [ZG22]. Lognormally [KC85]. London [MN95a]. Long [BK07, BK11, BJ99, Che93, Che20, CD98, CS20, DRV20, Erg19, FL97, FBB06, GG95, Gos10, GS91a, HW95, HM16, KLN85, Kla05, KOS94b, LL07, LS98a, LV00, MTLJ97, MM09, MPV20, ORT08, PQ10, Pic01, Qu11, RT00, Siz13, Smi05, TWW13]. Long- [Pic01]. Long-Horizon [CD98, Siz13]. Long-Memory [BK07, DRV20, GG95, LS98a, PQ10, Siz13]. Long-Range [Erg19, HM16, RT00]. Long-Run [BK11, BJ99, FL97, Gos10, GS91a, KOS94b, MTLJ97, MM09, MPV20, TWW13]. Long-Term [CS20, FBB06]. Longitudinal [AB22, BMM20, HMvdB12]. Look [Agu10, BC91, Tau93b, Wol00]. Looking [McG95]. Lorenz [ARBS87, BDB14, CG02, KYK^{+}21, SB19]. Loss [BP22, BG92, GHNS09, HPH87, IRR22, JMP16, MJY86, SM83]. Losses [WTC84]. Lost [Goi85]. Lottery [DHS19]. Lotto [FHLW00]. Low [GMP11, MH17, RE12]. Low-Frequency [GMP11, MH17]. Lower [Aiz91, LTMZ19, MPTK91]. Loyalty [AL95]. Lucas [KN89, MR91a, lT85]. Lumber [Lin86]. Lumpy [DFPS11].

M [Mül18, Tam20, SW21]. M.B.A. [RMS88]. Machine [AFFT22, BGS22, CKMS22, MVVZ19]. Macro [AR91, Cle17, DMS99, DBCB19, KLS12, RS94b]. Macro-Economic [DBCB19]. Macroeconometric [PSW04a]. Macroeconomic [AGO^{+}14a, AGK13, AV08, Bäu13, BGR15, Cam07, CG08, CGM16, CM96, DS13, DP22, FK86, GFHPZ87, GS91b, Har85, HRV20, Isk20, JS19, Kor19, LZ87, LTMZ19, LMS20, MN16, McN86a, MG84, Mit90, Nef93, Ott90, PT17, Poi91a, PY17, SV16, SW96, SW02, Yam16, Zar85, ZH92]. Macroeconomics [HW02]. Macromodel [Bey01]. Madness [KB10]. Magazine [Dan92]. Magazine-Exposure [Dan92]. Major [Kro83]. Making [ERT88, GSC03, LNB10, Luo02]. Male [CMW23, MA23, MZ23, MAB^{+}23]. Malpractice [HS89]. Management [Ber01, CM93, Luc00]. Manela [Pel21]. Manufacturing [Kon98, Mor92, MS96, NSR88, YCJK20]. Manuscripts [Ano83c]. Many [BSS22, BCW16, BW19, CHL18, HHL18, HHN08, HJL22, KCF^{+}15, KSD98, LP12, SW12]. Map [SW94]. Mapping [Lab90, LB16]. March [KB10]. Marginal [BH09a, FZ13, GKK20, KS08, MQZ22]. Marginalized [DF15, Qia22]. Marginals [MOT22]. Mariano [Die15a]. Marijuana [DHZ10]. Mark [GR89]. Marker [GH95]. Market [ACW97, ACS15, AMRT22, AFO21, ABD15, ACD09, BMP10, BAT83, BO16, BO01, BS98, BMM97, Bor89, Bra90, Car06, CL87, CM02, CP98, CE08, CP19, CGK97, CO99, CGT14, DF11, Due97, EH19, Goo96, GS06b, GSS07, HL06a, HW84, JLP23, dRJH11, JLR11, KNW22, KD08, Lec09, LS98a, LV00, MN95a, ML88, Mey88, PS92, PQ10, SW94, Tau86b, VB95, Woo00a]. Market-Based [CGT14, GSS07]. Market-Value [HW84]. Marketing [BCG85, HPSW88, WLL18]. Markets [AFT20, BFR97, Bek95, BR89, CST18, CM88, CHN97, CG17, ES93, GCC11, JJR18, LN20, MM00, MMS12, NTGK22, NZ20, RU01, WGBM12, Zha20]. Markov [AC93, ACR19, AMNP18, ABD19, BCD17, BP18, Cai94, CST18, Cav22, CE05, CW01, Dan92, Due97, DM94b, FO05, Goo93, Gre89, GM13, GLLM20, Kim93, Knü09, Lit83, LdV99, Lux08, NPZ01, PV03, Smi02, SLL98]. Markov-Chain [Dan92, Gre89]. Markov-Switching [ACR19, BCD17, CST18, FO05, Goo93, GM13, GLLM20, Kim93, Knü09, LdV99, Lux08, Smi02]. Markups [Mor92]. Martingale [FKO97, LS20b, PJ14, WZS22, Yil03]. Massive [Qia22, ZPWW22]. Match [Gar89, GIR20]. Matched [DZ04, HT20]. Matched-Model [DZ04]. Matching [AI11, AKKL21, BCHL20, GIR20, JR03, LZ19, MS03, Rod84, Rub86]. Math [LZ95]. Matrices [ABJT22, DL20, ELW19, HL16, KLZ20, LLF22, LSS16, RF15, TMH^{+}23]. Matrix [Cha20b, FFX16, GW88, LS20a, LLT^{+}15, LFWT18, LS20b, LSCX22, Por02, SX20]. Matter [BB87]. Matters [BK12b, CO19, Mer99]. Matthew [Bon20, PM22]. Max [CZ18]. Max-Linear [CZ18]. Maximum [CKY85, DG02, FQX14a, FSC03, FH98, GHPS22, Lum95, Por02, Rod16, SM83]. Maximum-Entropy [Rod16]. Maxwell [SSBD15]. McFadden [CP98]. MCMC [Era01, HMvdB12]. MD [MN16]. Mean [AC93, BDK07, Cha15, Cha20b, CJZZ18, Due97, Fre85, GL16, KP11, KSD98, LK93, Mar87, Per90, PV92b, ST18, Tay02, VP93, Vog98]. Mean-Reverting [Due97]. Mean-Structure [Cha20b]. Mean-Variance [BDK07, GL16]. Means [Bel87, PW22, Pfe91a, PB91]. Measure [BHLM07, Bla89, CPP18, Est98, HHPS19, HLQS17, IOS90, KCBB97, Qua88]. Measured [DMP19, PPS19]. Measurement [AWX20, ADS09, AV86, BI15, Bla89, BH84c, CHSS04, CK19, Cle85, CI87, FS88, Gas85, GK18, Gor89, HK01, HPH87, HP89, HD21, HI01, HT20, JSSvN20, JM91, KS08, Kok87, LG89, LK88, Los85, MRW12, MlM87, Mor87, Mor85, NPF11, OTZ91, Pis95, Rog87, Sch87, Shr89, SH05, Ver12, Vin85a]. Measurement-Error [Sch87]. Measures [ACD09, BCCL19, Cal91, CW96, CFG95, Cav22, DFK19, EN94, GSS07, HH16, Hil85, HW84, Hog19, MN88, PJ20, RS94a, ST89]. Measuring [AJ05, AL84, AV86, ACH03, BS95, CT97, CG00, Dag87, DG96, EH19, GM97, GHY92, GPP98, HI97, HF86, KPS00, Kro83, LS16, McC97, Rat12, RS22, ST18, ST20, Wri83]. Mechanism [KL90]. Mechanisms [FB99]. Median [AC94, Fre85, HCP11]. Median-Unbiased [AC94]. Mediation [HLM16, LLCL22]. Mediator [CHW20]. Medicaid [CHSS04]. Medical [DMMN83, HO84, HS89, Kow16]. Medicare [LT12]. Medigap [LW19b]. Memories [Che20]. Memory [AS12, BK07, Che93, Che20, Dav04, DRV20, GG95, HW95, LS98a, LV00, ORT08, PQ10, Qu11, Siz13, Smi05]. Men [LZ95]. Menu [KKB94]. Merchandising [AL95]. Message [BB89, BB02, Kor19, VP93]. Messy [ST89]. Meta [GSC03, VP92]. Meta-Analysis [GSC03, VP92]. Metal [CM88]. Metals [Mor93]. Meteor [LN20]. Metering [HMI95]. Method [And02, Are02, BSS22, BN02, CF02, Chr96, CT01, GH02b, GdXPE16, HS99a, HW02, HS85, Imb02, JLG04, KL90, Kew20, Mel05, RW98, SWY02, Tau86b, Xu00, YT07, Yit91, You88, ZWL22, ZLS19, ZWZL22, dBM90, vdHH86]. Method-of-Moments [JLG04, Tau86b]. Methodological [TY89]. Methodologies [Fom86]. Methodology [SL20, WD96]. Methods [BL88b, BCHL20, CZ14, CP08, DL87, DL92, DF20, DP22, FMB^{+}98a, Góm99, JSW02, Jan88a, JMP16, KKPS22, KS96a, LMN22a, MPTK91, MVVZ19, MKL07, PRC84, PPL94, PT18, PT23, RS94b, SW12, Tau90, TU90, TT06, Tsa16]. Metric [Ros08]. Michigan [MN88]. Micro [BGR93, GK18]. Micro-Level [GK18]. Microcomputers [dlV85]. Microdata [DL89, HW90, McG95, Paa88]. Microeconometric [DFPS11, MS97, Qia22]. Microeconomics [KW97]. Microstructure [ACD09, CP19, HL06a, JLK14, Woo00a]. MIDAS [GM13]. Migration [Fre92]. Military [BZ13]. Mimicking [BR08]. Mind [MO88]. Minded [KT19]. Minflex [Ram94]. Minimum [AAW11, CPP18, GN15, KT95, MJY19, Mar87, NW95, Pow22, RKL22, SJS18, SM83, WB04]. Minority [AK84]. Minority/White [AK84]. Mint [vdKH89]. Miracle [Chi85, Deh03]. Misclassification [OS20]. Mismeasured [CLT22, Lew12]. Mismeasurement [BBD20]. Missing [BMM20, Bre19, CLT22, CLQ20, GM83, KRW^{+}14, LCZT22, LLL19, Lit88a, Lüt86, MQZ22, Mol10, NPF11, NP86, QZL15, ZLWW20b]. Missing-Data [Lit88a]. Misspecification [GK14, GZ84]. Misspecified [CR16b, CR16a, Luc00, Pat19]. Mitchell [RR94]. Mixed [Bab22, CG08, FHDS11, GKK20, LR08, LRW09, LLL20, MPV16, Ng08, RV19, RLYZ22, SS15, WlMCP98, Yan20a, ZZZ^{+}22]. Mixed-Data [RV19, RLYZ22]. Mixed-Frequency [Bab22, CG08, MPV16, SS15]. Mixed-Poisson-Regression-Model [WlMCP98]. Mixing [KP11]. Mixture [DBCB19, GA10, GK08, HLWY14, KDM20, LeS92, LM92, LZLL17, Lie98, LR03, MRW12, RFG20, SSBD15, Wat00, Wat03, WD96]. Mixture-Amount [RFG20]. Mixture-Model [LM92]. Mixtures [CDS20, Ham91b, KW04b, KS94, Swa03a, WHKCS14]. Mobility [AMS19, AWX20, MLR19]. Mode [KPS20, Wec89]. Model [Abe99, Abr03, AB88, ABJT22, AS96b, And01, AT11, ABD19, BCR90, BKT11, BDW96, BH09a, BGR15, BCCL19, BF05, BCP19, BBCL20, BT14, BN12, Cai94, CFLS19, CHL18, CL87, CST18, Cat20, CKP13, Cha15, CL17, CSZZ17, CG19, CWW20, CHW20, Che20, CZ19, CP98, Chr90b, Col90, CKL11, Dan92, Dav04, DL20, DZ04, DMS99, DMMN84, DM94b, EYZ20, Erm91, FLTS19, FLYZ20, FT12, FP01, FB92, FO05, FP84, Fre86b, FHDS11, FSK08, FSTO04, Gag90, GA10, GKMV09, Ger88, GRW84, GP17, Goo93, GM20, GMS87, Gur91, HHL18, HM89, Hal94b, HS19a, HS96b, Hau96, Hil85, HPS96, HYSW16, HpK23, HW10, Ing90b, JVC94, JS10, Juh14, KWD13, KLP08, KDM20, Kar95, Kea92, KS98, KN89, KW19, KL08, KMV10, Kot85, KP11, KHT05, Lab90, LS20a, LL90]. Model [LS07, LLW90, LM92, LC85, LS87, LH99, LT12, LLZ16, LS20c, LTY^{+}22, LHYZ22, LGG22, LLCL22, Lin86, Lit83, LMN22b, LSU11, LdV99, Lux08, MV88, MPV16, McC19, McD87a, ML88, MB88, McG90, MRW12, Mel22, Mil87, MPV20, MS88a, Mos91, NS96, NC92, NTGK22, NP86, OP18, PZZ20, PX21, PRM99, PSW04a, PT17, PGC84, RLYZ22, RE12, RU01, Rog87, RAL99, Ros98, RE05, Sch83, SC03, SL20, SCS20, SK91, Sim90, SV16, SLL98, Son10, SV88, Str03, ST22b, SW95, Tau90, Ter85, Tsa85, Tse87, TT02, Tuz22, WlMCP98, WZ00, WL21, WW96, XZW14, YO19, YZK19, YLL19, YLN17, ZWZ12, Zha20, ZS14, ZLWW20b, ZWZL22, ZLLC23, ZT06, dHM90, vdKK03, vdKH89]. Model-Assisted [ST22b]. Model-Based [FSK08, GRW84, Hil85, JS10, Kot85, NP86, PGC84]. Model-Selection [LH99, SW95]. Modeling [AMS19, Abr85, AR88, AG10, ABD19, BGKS04, BO16, BFP06, Can92, CNR20, Cha15, CT18, CR12, CHLR92, DP22, GGKS98, GSC03, GHS09, GKK20, HHK19, HH16, HKR97, Hsi89, Ing90a, JR90, Jar87, KWD13, KG15, Kea97, KN89, KBR95, KMF05, Kon98, KOS00, LM00, LGLY16, Lie98, MSW85, MKL07, MHB19, OP15, PSD14, PSW04a, PY17, Pro12, SH87, SCS20, Smi00, Sta88b, SSBD15, Swa03b, TT83, TW14, VSV99, WV22, ZW93, ZS15, ZSZ20, dBE98, vO00]. Models [AGU04, AW00, AIPR18a, AIPR18b, AR91, AFFT22, ACR19, AT14, AAMW20, AT17, AB22, AC94, And94b, ACL11, Ang01, AMNP18, AL18, AK13, AGV19, AP06, BF15, BFZ98a, BN07, BW15, BLL16, BK07, BHKN10, BMP10, BR08, BHZ20, BAT83, BRW12, Bau18, BL05, BCD17, Bel87, BCHK16, BCW16, Bel04, BR89, BH97, BO01, BMY04, BC89b, BH09b, BGG22, BG05, BP18, BT89, BJ06, BB22, BCM10, BGL20, BC05, Bur94, BW84, CGT17, CCKR17, CMKSW03, CT93, CW96, CGJ19, CH14, CdM19, CR16b, CR16a, Car84, CNT20, Cav22, CKLGS12, CHT17, CS86, CT90, CF99, CGL13, CDY13, CGLS18, CGZ19, CMT19, CT01, Cla96, CM12, CG12, Col93, Col91, CS20, Con90, Cre17, CZ18, CHNY22, Dav04, DL20, DSSW07a, DS98, DGJ20, DK00, Die15c, DBCB19]. Models [DFP11, DGP20, DNW98, DW04, DMMN83, Due99, DS13, DV04, EH90, EGR91, Era01, Erg19, ELZ13, FLW96, FQX14a, FMT97, FLYZ20, FLSVP04, FFK18, FM90, FSC03, FS97, FRS83a, FH98, GLL93, GLLS11, GK14, GIR19, GK08, GS16, GPD15, GHPS22, Gos02, Gos10, GN15, GdXPE16, GS91b, Gre94, GK03, GGW14, GS03, GM13, GL13, GL20, GT96, GE12, HK19, HR91, HS99a, HK14, HL16, Han19, HHL^{+}22, HS96a, HL83, HP11, HH19a, Har91, HT83a, HF89b, HK92, HHKO22, HO84, HLR87, HM16, HW20a, HS16, HV03, Hof91, Hog19, HLZ04, HK17, HPP00, HW20b, HJL22, HT22, Hos18, Hos19, HSZ22a, HLZ22, HF19, HKO20, HLM15, IP12, IRS15, Ish22, Isk20, JPR94a, JPR02, JMZ18, Jor99, JLR11, Juo18, KW04a, KW22, KW04b, KW90, Kap10, Kas22]. Models [KR92a, KP10, KST11, KT18, Kim93, Kim01, KK15, KSW22, KKS22, KSW16, Koc10, KS94, KP03, KLGS10, KPS13, KB04, KLS12, KLS15, KP95, Kor10, Koz20, Kra86, LZ87, Lan97, LL07, LeS92, Lec95, Lec09, LYY22, Lew12, LD96, LT00, LHLF02, Li11, LLL11, LSX15, LZLL17, LW19a, Lia12, LQZ19, Lie98, LR03, Lin97, Liu14, Liu21, LMN22a, LNS11, Lou07, Lu22, LSS19, LXY22, Lum95, LTvD03, LW18, LZGL23, LP96, MLST20, MSW83, Mar85, Mar87, May86, McG91, MABF08, MT06, Mey88, MRW19, Mon97, MW22, MT85, MT02, NW13, NS05, NP90, NEV15, Olm09, OJLD18, OLBvD20, OS89, OV84, PSSE20, PPR03, PHPY20, PX22b, PRC84, PRPS22, PF19, Phi10, Pie84, PZ19, Por18, PT18, PT23, Qia19, Qia22, RCW83, RS15, RT01, RFG20, San17, SW22a, Smi02]. Models [SOK01, SL99, Sta08, SW22b, SMU13, SQ17, SWJ19, SL11, SSBD15, Swa03a, SW95, TU90, TM86, TT06, TR86, Tsa85, Tsa89, Tsa93c, Tse89, Tuc92, Tyr83, UY23a, UY23b, VLS90, Vel22, VV99, VDP00, WLW19, WS20, WZS22, War89, Wat00, Wat03, Wei86, Wer99, WK91, Wil93, Win95, Wol87, kWM90, WZ20a, WSX22, Wüt20, XHL20, XZ22, YL17, YQZ18, ZZZS19, ZWL22, ZWLL22, ZG22, Zha15, ZH19, ZLS19, ZL15, ZLY17, ZZZ^{+}22, ZC17, dJ00, dBM90, vO00, BS89]. Moderation [Cam07]. Modified [LH99, QW20]. Molinari [CMM23, Mor23, Gua23, Hon23]. Moment [Ant83, BSS22, CHL18, Cha92, Che07a, Dav04, FS19, GW14, GKR12, HKYZ20, Isk20, JJR18, SW21, Tao19, Zil97]. Moment-Based [Ant83, Che07a]. Moment-Condition [Zil97]. Moment-Implied [GW14]. Moments [And02, Are02, BLC20, BN02, CZ14, CF02, Cav22, CSZ22, Chr96, CT01, Fri22, GH02b, HS99a, HW02, Imb02, JSW02, JLG04, Kew20, Knü15, SWY02, Tau86b]. Momentum [Joh02, LL94, Yao08]. Monetary [AJK18, BH94, BI15, Car06, DG96, DM91, FO05, GKMV09, GSS07, Jus98, LM00, MP83, Pie83, PGC84, RR94, RDP95, Ser88b, SW90, VP93]. Monetary-Asset [DG96]. Money [Bar83b, BS87b, BI15, BF89, Car06, CK08, Coc89, Cra86, EN94, Fis92, LeS92, Mel90, Poi91a, Rad83, RW96, RDP95, Ser88a, Ser95, SW88]. Money-Demand [Fis92]. Money-Supply [LeS92]. Monitoring [Ana09, GH91, SBK18]. Monotone [RH04, Tsi22]. Monotonic [BS08, Dol99]. Monotonicity [FG20, Mel05]. Monte [AS96b, BC89a, CW01, DF15, GO98, HM89, Har91, HLR87, Lum95, Ray90, Ril91, Sch89, vdKH89]. Monthly [Bel87, BWH86, Gbu89, KMMP22, MN16]. Moreira [Pel21]. Mortality [BK16, FHDS11]. Mortgage [FB99]. Mortgage-Rate [FB99]. Most [LK93, Shi00]. Movements [BKW19, CM88]. Moves [MT99]. Moving [Ali84, BCD17, DL87, DW93, Gos02, GN15, GS03, HS96a, MM03, SK91, Sta08, TT06, Tsa89]. Muitinomial [MPTK91]. Muitinomial-Based [MPTK91]. Multi [BR08, BCP19, DMMN84, HR19, Knü15, NWK10, OLBvD20, PT11, PT12a, Qua19, WL21, XZW22, ZWZ22]. Multi-Asset [BCP19]. Multi-Beta [BR08]. Multi-Factor [OLBvD20]. Multi-Horizon [PT11, PT12a, Qua19]. Multi-Index [NWK10]. Multi-Kink [ZWZ22]. Multi-Layer [XZW22]. Multi-Part [DMMN84]. Multi-Period [HR19]. Multi-Step-Ahead [Knü15]. Multi-Threshold [WL21]. Multichoice [VSV99]. MulticolIinearity [SH87]. Multidimensional [WD96, XZL20]. Multifactor [YLL19]. Multifractal [Lux08]. Multifrequency [LZZ20]. Multifrequency-Band [LZZ20]. Multilateral [DF20, HT06]. Multiline [CRZ19]. Multinomial [BN12, KW04b, Kas22, Kea92, LC85, LMN22a, LMN22b, Tse87, WK91]. Multioutput [KBZ89]. Multiperiod [Dua88, KC85]. Multiple [APW20, Aud06, CG12, Coo98, DM95b, FSK08, GFdH87, GK08, GMP85, GPP98, Inc93, KW22, KP10, KRW^{+}14, KW19, LD19, LPS22, LD96, LLF22, Lüt89, MSW83, MS03, PZ19, Pro08, Rub86, Shr89, SW22a, TT83, TA89, WZ00, WHYC22, Wil93, YL17]. Multiple-Indicators [Shr89]. Multiple-Threshold [YL17]. Multiple-Valued [PZ19]. Multiple-Vintage [CG12]. Multiplicative [CS20, KP11, VM89, WV22]. Multiplier [Hal90]. Multipower [Kol20]. Multiprocess [LeS92]. Multiproduct [DBS84]. Multivalued [ACL19]. Multivariate [AT11, BHZ20, BL05, BDK07, CPDH15, CKL11, Dan92, DS90, DFP11, EYZ20, ELZ13, FSC03, GS10, GL16, GM83, JLR11, Kar95, LS07, LS20b, LGLY16, LNS11, LSU11, LZGL23, MN88, MB88, McE17, MRW19, NEV15, Ott90, PG06, RV19, RTW12, RF15, RFG20, SU22, TT02, WZS22, YO19, YLN17, ZH92, ZZZS19, ZSZ20, vdKH89]. Multiway [CGM11, CKMS22, MNW19]. Multiyear [Ott90, ZH92]. Multiyear- [ZH92]. Multiyear-Ahead [Ott90]. Museum [FLM^{+}19]. Mutual [LD19, WHT22]. My [Poi91b, Gal93, Gew93, Tau93c, Zel93].

N [SW21]. NAIRU [DG06]. Naïve [GGW14]. NAPM [DL92]. Narrative [GKR22a, GKR22b, Kil22, PM22, RR22]. National [Bra90, CGR07, Ghy90, GH95, dL90, vUS87]. Natural [DPES19, Mey95]. NBA [HJS20]. NBD [MS88a]. NBER [FM97]. Near [RF01, SK91, Wri00b]. Near-Regression [SK91]. Nearly [Gos02, Gos10, Pan91]. Negative [BDY11]. Neglected [BK07, HMS17, HS85, You88]. Neither [BC01]. Nelson [BR95, KMV10, TTW95]. Neoclassical [SW94]. NERC [NP85]. Nested [BS89, CM12, SW22b]. Net [CZ14, CHL18]. Netherlands [IV95]. Network [BKW19, BGL20, GS16, HHK19, HJS20, HYSW16, Joc18, LFWT18, LLPZ20, Liu14, PHPY20, WLZT22, WHW22, ZTCW15]. Network-Based [PHPY20]. Networks [Byk23, Don22, GPI13b, JJKL21, JJKL22, LC22, MLZ22, Mel22, RW22, SW95, WF22, XZW22, YLYJ22, ZK22]. Neural [RW22, SW95]. Neutrality [McG91]. Newcomb [BCCP18]. News [BCM22, HS19a, ST22a, ZHCB16]. News-Driven [ST22a]. Newsy [Tho20]. Next [JS85]. Nikkei [ACW97]. No [AR91, JY21]. No-Crossing [JY21]. Nodal [MHCP22]. Nodewise [CCÖU19]. Noise [BRY13, CP19, HL06a, HP13, JLK14, LZZ20, Mar87, NV12, PY09, WKX23, kWM90]. Noisy [BBCL20, LJ18, SX20, Var16]. Nominal [Poi91a]. Non [BDK07, BDY11, CZZ22, Cre17, FH98, JS10, KL08, KLS15, LL19]. Non-Gaussian [BDK07, CZZ22, Cre17, FH98, JS10, KL08, KLS15, LL19]. Non-Negative [BDY11]. Nonacceptable [RAL99]. Nonadaptive [PPR03]. Nonadmissibility [FP01]. Nonaffine [IRS15]. Noncausal [CNR20, Fri22]. Nonclassical [AWX20]. Noncompliance [BCFFL19, CF15]. Nonconstant [GR15]. Nondecreasing [FPAB02]. Nondifferentiability [DM95b]. Nonexperimental [MRZ92a]. Nonhomothetic [SW90]. Nonignorable [BMM20, CLQ20, HR19, Sta88b]. Noninvertibie [Tsa93c]. Noninvertible [Gos02, GN15]. Nonlinear [ACW97, BH97, BH09b, Bie00, Bre01, Cla96, Col93, Das91, DM91, DH20, ELZ13, EH19, FLYZ20, Fuh92, GGS96a, HK17, KLS15, LT19, Lob03, Mar83, MB88, Olm09, Qi99, Rac01, RR94, Sim90, ST18, TU90, XZ22]. Nonlinearities [KV13, ML88, dL98]. Nonlinearity [BK07, HM16, HP85, JS98, VFL99]. Nonmarket [KSS09]. Nonnested [CK08, EN94]. Nonnormal [CDJW10, LL10, MJY86]. Nonparametric [AJ05, Aiz91, AFAA22, AWX20, Ana09, AFO21, BDB14, BKT11, BH09a, BFV22, Bie00, BRT12, BCM10, CT16, CI03, CGLS18, CGZ19, DH20, DU12, FZ18, FFK18, Gre89, JGL19, Juh14, Kal22, KSW16, LW09, LMS09, LR08, LLR13, LLL20, LZW^{+}22, LW19b, LG20, Oku11, OHS14, PZZ20, PRPS22, PT92, Rac97, RV19, San20a, SW22a, SS13, SB19, SW88, TK10, TGCD16, WW09, WHKCS14, Wil93, WZ15, XP11, Yan20a, YYX20, ZWLL22]. Nonparametrsc [BZ99]. Nonpecuniary [BKT11]. Nonresponse [BD01, Fre86a, Hil86, HS96c, Ker85, Kot87, Sta88b, SB89a]. Nonrevisable [dHK14]. Nonseasonal [OF97, Tay02]. Nonseparability [GK18]. Nonseparable [FFK18]. Nonspurious [Yam16]. Nonstationarities [dL98]. Nonstationarity [DGP20, EJP05, Mis95, Ng08, PV92a]. Nonstationary [AT14, BT22, BZ20, FS23, Góm99, Gos10, HK14, MM20, SL01, SL11, WHYC22]. Nonstochastic [FL83]. Nonstrong [BG23]. Nonsynchronous [Var16]. Nontraded [Kim05]. Nonzero [ZWZ12]. Norm [CCKR17, FLM20]. Normal [ACS15, AS20, And01, Ham91b]. Normality [BN05, Hal90, KD00, PTZ86, SW21, vdKK03]. Normalization [LRS99]. Normalized [ABA99, BGKS04, DW88, ZFSC00]. Northwest [DH88]. Norwegian [MS89]. Note [Bel87, Car84, CHT07, Has87, HO84, Iwa91, Kea92, Kot85, Luc00, MS03, NPW93, Oht87, Öll86, OTZ91, PRG^{+}22, Sel89]. Notes [PF84]. Novel [ZWL22]. Nowcast [FG20, FS22]. Nowcasting [AGJT14, ARvD18, BGS22, ÇD22]. Nowcasts [KCR17]. Nuclear [RR97]. Nuisance [BSS01, GS03]. Number [AW07, BN07, CCP13, CH14, GL16, Kap10, YYX20]. Numbers [HF97, Sel89, Tho20]. Numerical [DG02]. Numerically [KLS15]. NYSE [BS98].

Observation [OLBvD20]. Observation-Driven [OLBvD20]. Observational [BMM20]. Observations [Ali84, GM83, HF89b, HKR97, HLR^{+}22, Joh87, LJ18, Lüt86, Mil84, MT99, NP86, Peñ90, Wec89]. Observed [CG22, DGP20]. Obtained [Tau86b]. Obtaining [MPTK91]. Occur [RW96]. Odds [Fri22, GW88, Inc93]. Odds-Matrix [GW88]. OECD [Gew88b, Ost84, PZ90]. Off [Lec99]. Off-the-Job [Lec99]. Offer [Swa03b]. Offers [JT95]. offs [lT85]. Oil [ACvD22, BK12a, BK15, KV13, ZA92, ZA02]. Oil-Price [ZA92, ZA02]. Old [Jae97, Wol00]. OLS [GI11]. Omitted [CS20]. One [CW86b, LP12, Ott90, PRG^{+}22, ZH92]. One-Sided [CW86b]. One-Step [PRG^{+}22]. One-Year-Ahead [Ott90, ZH92]. Online [PSD14, WJS08]. Only [BSS01]. Open [dRJH11]. Opening [Xie22]. Openings [CRZ19]. Operational [BL88b]. Operational-Subjective [BL88b]. Opinion [CJJ^{+}23]. Opportunities [BFZ98a]. Optima [DM95b]. Optimal [Ash85, BOP23, BP18, BT89, Can98, EJP05, FIL^{+}23, GK18, Gro86, KJGV09, KBR95, Lan97, LPV03, LLR13, LNB10, Luc00, Nie04, RLYZ22, RR97]. Optimality [Phi87]. Optimization [HL22]. Optimizing [EKM87]. Option [AFFT22, AFAA22, AFT20, BCSS20, BCCL19, FE22, HLM15, LMS09, PRT00]. Options [DTU06, Guo98]. Order [BC11, BT01, Cav22, CL95, DP87, DP02, GI19, Hay86, INT21, KP95, MT96, PLCW22, Ray90, SS06, SRdCFB18, Vel22]. Order-Determination [KP95]. Ordered [ARBS87, DV04, HS16, Tuz22]. Ordering [Hen22]. Orderings [ZFSC00]. Ordinal [BCF94, KW90]. Ordinary [Hag83, HLR87]. Orleans [CE08]. Orthogonal [CILY17, LS07]. Other [Bar85, BR89, DM95b, Ghy97, GK96, HI01, KLL20, Siz13]. Otter [ZH92]. OU-Gamma [JMZ18]. Our [Sle98]. Out-of-Sample [HLZ04, JS98, RI12, Wol87]. Outcome [CG22, MHB19, PT03]. Outcomes [BFV22, BMM20, BCFFL19, KKPS22, KP11, McK18, Rot15, San20a, VB95]. Outlier [FL98, GH91]. Outliers [FH94, SS06, VFL99, Wil93]. Outlook [Gra87, Ost84]. Output [AP06, CsL94, CG12, Cle14, Coo98, DG06, Gor89, HP05, Koo91, KOS00, Kut94, LG89, PT11, PRF08, Poi91a, QW20, RDP95, She93, lT85]. Output-Inflation [lT85]. Output-Substitution [Gor89]. Outputs [SW22a, Wil93]. Outward [WLZT22]. Over-Identification [AN91]. Overall [LB16]. Overcoming [FP01]. Overdifferencing [Bel87]. Overdispersed [WK91]. Overdispersion [Gur91]. Overnight [KSW22, MN95a]. Overstatement [MPTK91]. Overview [HF86, MS99]. Owner [Ver12]. Owners [HS96c, McK18]. Ownership [LA84].

P [KT18, ZH92, BS08]. P-Splines [BS08]. Pacific [DH88]. Pairwise [Abr03, PX21, YO19]. Pairwise-Difference [Abr03]. Pandemic [Cav22]. Panel [ABS99, ABA99, Alv04, AW07, ACL11, BCHK16, BH09b, BG23, BC05, CGJ19, CH14, Car01, Cha92, CGL13, CSZZ17, CHNY22, DGP20, DH89, ES01b, Erg19, GK14, GDG^{+}05, HL16, HLM05, HK17, HPP00, HSZ22a, HpK23, Ish22, JVC94, JGL19, JS10, Juo18, KW22, KR92a, KOS00, KL08, Kor10, KS96a, Kum87, Kum89, LL07, Lec95, Li99, LQZ19, LGG22, Liu21, Lou07, Lu22, Lus96, MTLJ97, May86, MZ23, Mon97, MS89, MW22, NP97, Ng06, PHPY20, Phi10, RW86b, SS84, Sta88b, SWJ19, Tsa00, Tuz22, Wes14, Wes15, Wes18b, WZ20a, WSX22, XZW16, YZK19, ZWLL22, ZLS19, Zil97]. Panel-Data [HPP00, KR92a, Li99]. Panels [DE19, DH12, GK03, HR19, JS20, KNW22, LL07, MB13, MP05, Ng08, RSW18, Val91, YLL19]. Parallel [Col93]. Parameter [AAMW20, And02, BP22, Cha22, FLM20, GA10, GS03, Han92, Han02, HHKO22, HKO20, Ioa95, KN89, Kim90, LeS92, Li15, LJ18, MSW85, MH17, PT03, RU01]. Parameterization [Tsa89]. Parameterized [MM03]. Parameterizing [dHM90]. Parameters [AL89, BSS01, Cha15, CFM11, CO99, DW93, EJP05, Fuh92, Gri93, Ham91b, HLM15, JR22, KMV10, MJY86, Str22, Tau86b, Vog98, Wes01, Wol87, Yit91]. Parametric [CT93, DH20, EO10, FE22, HW20b, KMF05, LMS09, PM20, Rat12, Ril91]. Pareto [HPZZ21, SCS20]. Parimutuel [Goo96]. Parities [RS92]. Parity [BBC04, CL93, HLM05, HP10, Kim05, MP05, PV92a, Pro08, Ros05]. Parsimonious [Tsa89]. Part [DMMN84]. Partial [BdB18, BGL20, DM01, FFL13, HR91, HJNR98, HW20a, Jor99, Lin86, MLR19, OT21, She19, TA89, YF22]. Partial-Likelihood [TA89]. Partially [AK13, CFLS19, CGL13, DGP20, HH19a]. Participation [GLL93]. Particular [Sim90]. Parts [Bel87]. Pass [GLLM20, Kim90]. Pass-Through [Kim90]. Passing [Kor19]. Patent [WlMCP98]. Path [Gag90]. Patient [RAL99]. Pattern [CF99, Ram94]. Patterns [AB22, CH95, DMS99, FS93, LS94, LLL19, Oli87, ZCZ^{+}21]. Pay [BF89]. Payroll [LM90, NW91]. PC [FW03]. PC-Based [FW03]. PCDID [CK22]. Peak [AL84, Koh83a, Zha20]. Peak-Load [Koh83a]. Peaks [PSvD09]. Pearson [ZL15]. Pearson-Type [ZL15]. Peer [GPI13b, Kra07]. Penalized [CSZ22, WHKCS14, PX21]. Penalty [BH09b, Lia12]. Penetration [PFA00].

Pengsheng [MLZ22]. Pepsi [GKP00]. Perceived [Jai19]. Perception [Ham13]. Perceptions [HK86]. Performance [And94b, BCHL20, CP08, EKM87, Gra19, HV87, HS99b, HLZ04, HH04, HLR^{+}22, HLM16, JS98, KCBB97, MPTK91, MV18, Mor93, MS96, OS89, PvP06, PY19, PT92, VLK04, Wol87]. Period [AK06, HR19, Mer99]. Periodic [BG96, Ghy94, HHPS07, Kot87, OF97, OS89, WHYC22]. Permanent [CM90, KJAG85, KHT05, Lus96, MSW83, ZO04]. Permutation [DFK19]. Persistence [AKO12, CM96, Jai19, LL90, MN95b]. Persistent [CR12, GP17]. Personal [AV05b, BB05, Die15a, NTP94]. Perspective [Die15a, FZ18, GK15, Liu21, Luc00, Nor85, SBK18]. Perturbation [ZS15]. Peter [Ano02c]. Petroleum [WGBM12]. Pharmaceutical [BGR93]. Phase [ABS93]. Phases [Fil94, Sic94]. Phillips [KM09b, KLGS10]. Physician [Das08]. Piece [VZ11]. Piecewise [Meg87, Mof86]. Piecewise-Linear [Meg87, Mof86]. Pileups [DW93]. Pilot [PRG^{+}22]. Pitfalls [Ghy90, Hub15, KLL20, MV88, NK84, Pro08]. Plaintiff [Fre86a]. Plan [LT12]. Plants [KM08, RR97]. Player [LLPZ20]. Ploberger [NS10]. Plug [YT07]. Plug-in [YT07]. Point [AK13, BHZ20, BGR15, Car85, EMW09, Fra19, GK08, Hor98, HMR20, TZ19, ZCZ^{+}21]. Point-Estimator [Hor98]. Points [HW20b, LeS91, LD96, XZW16]. Pointwise [Shi00]. Poisson [AMNP18, Gur91, SSBD15, WlMCP98, YQZ18]. Poisson-Driven [AMNP18]. Police [GB87]. Policies [FLSVP04, MT09, Vos16]. Policy [Ang95, AJK18, BI15, Col90, EHM98, Eri98, FO05, FRS83a, GSS07, HS89, IOS90, Lou07, McG95, RT01, Sal95, Sea93]. Policy-Function [Col90]. Political [DHS19]. Polynomial [PLCW22]. Polynomials [GI19, Tyr83]. Pool [BC01]. Pooled [DE19, GFHPZ87, Har83, Mit90]. Pooling [Bra90, HPP00, LM90]. Poorly [DMP19, PPS19]. Popular [LYY22]. Population [BM93, KSD98, Pfe91a, Val88]. Populations [Dag87, MPTK91]. Portfolio [AW00, BOP23, DM16, Edw08, FFX16, GA10, KP22, KW19, Li15, LHYZ22, LSS16, PT00, RKL22]. Portfolios [AMRT22, BR08, BHW09, CCÖU19]. Portmanteau [WS20]. Positive [Gur91, LG20]. Possibilities [Gor89]. Possible [INT21]. Possibly [BDK07, BG23, GN15, Luc00, PV03, Pat19, SL01, WHYC22]. Post [BCW16, DS90, LXHT22, ZW93, Cle14]. Post-Deregulation [DS90]. Post-Enumeration [ZW93]. Post-Selection [BCW16, LXHT22]. Posterior [BW22, Inc93, KOS94b, MRZ92a]. Posteriors [DW93]. Postwar [Ghy90]. Potency [FB97]. Potential [DG06, EJP05, Kut94, MHB19, Pro08]. Potentially [CK22]. Poverty [BD09, Bla89, NPF11, Tar07]. Powe [RR97]. Powell [KT18, Sil01]. Power [Aiz91, ALT08a, BBC04, CL93, EJP05, HLM05, HP10, Kim05, MP05, NS96, PV92a, PTZ86, Pro08, Ray90, Ros05, RW09, Son12]. Powerful [BV05, LK93, Shi00]. PPP [Li99]. Practical [CPP18, DP22]. Practice [LLSW18a, LLSW18b, Mül18, Vog18, Wes18a]. Preaveraging [HP13]. Preaveraging-Based [HP13]. Prebisch [BV05]. Prebisch-Singer [BV05]. Precautionary [Cag03, FP97, Nor94]. Precious [CM88]. Precision [HPH87, Wri83]. Predetermined [Mon97, MB13, Zil97]. Predict [CL17]. Predictability [AG05, Ana09, BHW09, Cam07, DH20, FOTY20, GP12, GP17, IR05, Jac00, LZ18, MV18, MRW19, PT11, Qi99, Rac01, RU01]. Predictable [BH93, CHN97]. Predicting [BJOS99, CG17, LD96, RKL22, RAL99, WTC84]. Prediction [Byk23, Col91, DW20, FLTS19, Fri22, GKMV09, GK83, GPD15, Has87, Kim01, LS87, LZW^{+}22, LNB10, Lüt89, MV91, Qua88, SOK01, Val88, Wan96, XHL20, YQZ18]. Predictions [ACL11, EMW09, KW04b]. Predictive [BP22, CDY13, CM09, CM12, DH20, DM95a, DM02, Die15a, GHLT19, GP12, GP17, Han05, HV87, HD21, Juh14, LLP15, Mal83, MRW19, NP90, PG93, PT92, PT03, Son12, SU22, YLPC20, ZLLC23]. Predictors [SW12, WLTL20, YF22, ZZZ^{+}22]. Preference [BM23b, DLP19, Gua23, Hon23, LPS99, Mor23, Mos91, Sal95, VSV99]. Preferences [Bar83a, CMM23, DHS19, GK18, GSA04, HD85, SW90, Yog08]. Pregnancies [AKO12]. Pregnant} [KLL20]. Preintervention [KKPS22]. Preliminary [Bra90, Gro86, MP83, NW91, TR86]. Preliminary-Data [MP83]. Premia [BR08, BGR15, Che07b, HRV20, NPW93]. Premium [AFT20, FID16, Guo98, HK19, KMN05, LTU15, LZ18, LP03, YQZ18]. Premiums [BG93, BTZ04]. Preparation [Ano83c]. Preparations [BGR93]. Prescription [BGR93]. Preselection [FS22]. Presence [BRY13, BP22, BT01, CFG95, CF15, GKMV09, Ghy87, GR15, HP13, INT21, NPF11, OHS14, PRC98, PT03, VFL99]. Present [BSS01, Pak11]. Press [MV91]. Pretest [Hal94b]. Pretested [KP22]. Pretesting [Ilm90]. Prevalence [DHZ10]. Price [ACvD22, AES86, AL95, AGV19, Atk83, BS87a, BK12a, BK15, BB05, BGR93, BKW19, Bra07, CM88, CL17, CE08, DM84, DO95, DZ04, DW85, DFPS11, FLS07, FE22, GKP00, Gra97, Hay89, HK86, HT06, HLM15, HW10, Jai89, JR90, JW99, Ker85, KBR95, Koh83a, Kon98, Kot84, Kow16, LPV03, LW18, MM92, McG91, Mel18, MF05, MKL07, MY18, MPV20, MT99, MS99, NTP94, OG86, Oht87, OV84, PZ90, Ran88, Rav06, Rog87, SJS18, Slo87, Son10, VM89, Val88, Val91, Ver12, WV11, WGBM12, ZA92, ZA02, dHK14]. Price-Induced [BS87a]. Prices [BGKS04, Bau18, BHMR19, BCG85, CDS20, DP19, ELLV97, FE22, GR89, HSZ22b, Ing90a, KV13, Kim90, Lie98, LT04, NV12, OG86, Olm09, Pat86, RS92, Ric93, RDP95, RE05, Ser88a, Shi00, Tim01, Var16, WJS08, Yog08, ZC17, vH89]. Prices-Statistical [BHMR19]. Pricing [AGU04, AL84, AFFT22, AFAA22, AFT20, BFZ98a, BCCL19, Bur94, CZ19, DS90, FOTY20, FLSVP04, Fre86b, GS10, GR12, GMS87, GL13, HS96a, HPS96, HV03, KKB94, KBR95, Koh83a, Li11, MB88, PRT00, PRPS22]. Primal [Mor85]. Primary [Mor93]. Primary-Metals [Mor93]. Prime [Due99]. Primitive [BN07]. Principal [FLM^{+}19, HT14a, ZL22]. Prior [AAMW20, Bäu13, GP98, Koc10, PG93, Rod16, Ros08]. Priori [Ter85]. Priors [Kou13]. Privacy [BS92]. Private [DMT06, MS96, MNO85]. Probabilistic [MM10]. Probabilities [ACR19, HL22, Joh87, KLGS10, MRZ92a]. Probability [BBK12, BV87, DHL14, EMW09, LTY^{+}22, QZL15, SWW88, TA89]. Probit [BN12, Kea92, Lec95, LMN22a, LMN22b, Ter85, Tuz22, WZ20a]. Problem [Hil86, JR22, KS08, Sch87, VP92, Wes15, Wol00]. Problems [DM95b, HS93, HL22, Kok87, Los85, PF84, QZL15, SH87]. Procedure [KW04b, Kap10, LD19, Sil86, Wan96, ZS15]. Procedures [CLLM03, Hil85, Jai89, Mal83, MG84]. Process [Ali84, AR88, Ano95c, CFS09, FS23, GGS96a, MT96, Pan91, Pop84, Qui95, SL00, Sen03, WHKCS14]. Processes [Agu10, BGKS04, BCS08, BDY11, DP87, DP02, Due97, DP22, DG02, Fli97, FS19, Fri22, Gri93, Han92, Han02, HLWY14, JK02, KG15, Lüt84, Lüt86, LRS99, MT99, PG06, RTW12, RFG20, TZ19, Tay03, ZK05, ZCZ^{+}21, ZZZ20]. Procurements [ACJK22]. Produce [SW94]. Producer [BGR93]. Product [BB05, Bra90, DBS84, Ghy90, PFA00, dL90]. Production [Ant83, BH94, CFM11, Das91, FLW96, Fom86, HPSW88, HV87, KSW16, KCBB97, Kum87, KGM91, Lin86, LYZ22, NSR88, OV84, Ros98, San20b, SS84, ST89, SY98, Tsi22, YCJK20]. Production-Frontier-Based [ST89]. Production-Smoothing [Ros98]. Productively [BDJ12]. Productivity [ACH03, BT14, Cal91, Kro83, Mor85, NSR88, Van08, WW09, Wil09, YP08]. Products [DHW17, Hau99, LW19b]. Professional [EMW09, GW09, JS19]. Profile [CLQ20]. Profiler [Gar89]. Profiles [BT14]. Profit [BH94, BG92]. Profit-Driven [BG92]. Profitability [CGK97, HW84]. Profits [Yao08]. Program [Ang95, ACL19, FMB^{+}98a, GOV84, HS19a, JR03, Mel90]. Programs [Deh03, JS99, Lec09, RMS88, TN92, Wai85]. Progress [TD84]. Progressivity [BFZ98b]. Projected [Che22]. Projecting [CHLR92]. Projection [Lab90]. Projections [GIR19, GL19, Woo00a]. Projections-IV [GIR19]. Projective [ZZZ20]. Promise [KLL20]. Propensity [BCHL20, CLQ20, CHLR92, FIL^{+}23, HHK11, JR03, MT09]. Propensity-Score [JR03]. Proper [IRR22]. Properties [Aie13, BE96, BT01, CsL94, CdH96, Chr96, Cla96, Dav04, GW14, GS91a, HHY96, HC94a, HC94b, Hof91, Koo91, KOS94b, KR94a, KR94b, LMS09, LS98a, Lum95, McG91, MRW19, Oom06b, Sch83, Tau86b]. Property [Yil03]. Proportional [FHDS11, SC03, Tse89]. Proportionality [Li99]. Prospects [APW20]. Prototypal [KS98]. Provision [Das08, FID16]. Provisional [BT89]. Proxies [GKR22a, GKR22b, Kil22, PM22, RR22]. Proximate [PX22a]. Proxy [CO19, CK19, JL22, LS22, dNS23]. Pseudo [AKKL21, GDG^{+}05, Juo18]. Pseudo-Matching [AKKL21]. Pseudo-Panel [GDG^{+}05]. PSID [Pis95]. Publiations [Ano86g]. Public [AC03, Hay89, MS96, MNO85]. Public-Good [Hay89]. Publications [Ano85d, Ano85e, Ano85f, Ano85g, Ano86h, Ano86i, Ano86j, Ano87f, Ano87g, Ano87h]. Pulpwood [Lin86]. Purchase [BG05, CP98, MT96, PSD14, SC03, SW94]. Purchase-Frequency [BG05]. Purchases [BFP06, MHB19, MS88a, PRC98]. Purchasing [BBC04, CL93, HLM05, HP10, Kim05, MP05, PV92a, Pro08, Ros05]. Pure [HW10, Son10]. Pure-Jump [HW10]. Pursuit [LQZ19, ZWLL22]. Puzzle [AKLS20, MP05, RW09, Sto88a]. Puzzling [MNO85].

QML [LYY22]. QMLE [HK14, ZL15]. Quadrant [LG20]. Quadratic [Chr90a, Chr90b, DW88, Gre94, HP13, McG90, MJY86, Var16]. Quadrature [Tau90]. Qualitative [DL92, Due99, Due05, HF89b, KL90, Tse89]. Quality [CM93, DH89, Gar89, Ger88, GH91, GOV84, LB16, ME04, NTP94, OG86, Son10, dHK14]. Quality-Adjusted [NTP94]. Quantal [Ter85]. Quantifying [ACvD22, DL92, Ham13]. Quantile [BGG22, BCFFL19, CPDH15, FXMFY20, FGH19, FM13, GLLS11, GCC11, GK05, GR11, GM20, HZ22, Ish22, JY21, KB10, Kow16, LR08, LLR13, LW19a, LLL20, Lia12, LW18, MQZ22, NEV15, PRG^{+}22, QY19, TK10, TXLT22, WLTL20, WW09, WZ15, Wüt20, XKZ21, YT07, ZWLL22, ZWZ22]. Quantile-Weighted [GR11]. Quantiles [DH22, SW20, ST20]. Quantitative [MR91a]. Quantity [BS95, Kon98]. Quarter [Bra90]. Quarterly [Alv04, NP86, SW88]. Quasi [FQX14a, GH22, GHPS22, Ham91b, LYZ22, Mey95]. Quasi-Bayesian [Ham91b, LYZ22]. Quasi-Experimental [GH22]. Quasi-Experiments [Mey95]. Quasi-Maximum [FQX14a, GHPS22]. Question [PJ20]. Questionnaire [PBC95]. Questions [BBK12, Jae97].

R [Fre85, Sim85]. R&D [AC03]. Race [Gra97]. Racial [BK05]. Raffaella [PM22]. Rail [RW86a]. Railroads [Kum87]. Rallies [BMP10, MMS12]. Random [BHL92, BC89b, BH09a, BG05, Bre19, CT90, CSZZ17, DBS84, DLP19, HMS17, HL16, HT22, HpK23, JVC94, Jor99, KW22, KMF05, Lit83, PRT00, PF19, Ser95, Tse89, Tuz22, WZ20a]. Random-Coefficient [BHL92]. Random-Coefficients [BG05, JVC94]. Random-Time [Jor99]. Random-Variance [PRT00]. Randomization [Ken95, Seo20]. Randomness [Ali87]. Range [BJ06, Erg19, HM16, RT00]. Range-Based [BJ06]. Rank [Abr03, BDW96, Bre01, CMKSW03, CT18, DFP11, GI11, HJNR98, LV15, LTTL19, LPS99, Pos00, SL00]. Rank- [GI11]. Rankings [LB16, VSV99]. Ranks [FFK18, Wri00a]. Rao [BGMP21]. Rate [AOS00, BHW09, CKY85, CST18, CG00, DS90, Due99, FB99, GR89, GD89, Haa96, HLZ04, Kim90, KK15, LCZT22, LKFL22, Mil87, Mis95, Nie04, PRM99, RW86a, RT01, RW96, RAL99, She92, SMU13, VZ11, Wol87, ZC17]. Rates [AB02, Aru20, Aud06, BB89, BB02, BMP10, BHW09, BI15, BB87, CCKR17, Che93, CE05, CT22, CLR14, Coc89, Dua88, EG98, FKO97, Fre92, HW95, HK94, HS96c, HPP00, Hsi89, IM97, KC85, Kim90, Kim93, Kim05, Kla05, KK92, KMV10, LM94b, LSS19, Smi02, VP93, XKZ21, Yil03, Zho96, ZLY17, vUS87]. Ratings [CGT14]. Ratio [AG07, CsL94, CR01, FKO97, GP98, Li05, MOT22, MPV20, RH04, Tao19, Wri00a]. Rational [BG01, BL88b, Boh91, HL83, Hof91, PRM99, Zar85, dJ87]. Rational-Expectations [Boh91, Hof91, PRM99]. Rationality [Ing90a, JM91, PT12a, SvD06]. Raw [Knü15]. Reaction [JLG04, dRJH11, KNW22]. Reactions [ZHCB16]. Read [Kil22, PM22, RR22]. Reagan [CJW86]. Real [AGJT14, ACvD22, ACW97, ADS09, Aru20, AP06, BK12a, BK15, Bru92, CP08, CM09, Cla11, CG12, Cle17, CFS09, DTU06, GKMV09, Gew88b, Ghy90, GPR13, HH11a, IM97, IR05, Ive01, KV13, KK15, Koo91, LMS20, LS98a, MM92, Mis95, PRW06, Poi91a, QW20, SS15, SZ20]. Real-Interest-Rate [Mis95]. Real-Time [ACW97, ADS09, BK12a, CP08, CM09, Cla11, CG12, Cle17, CFS09, GKMV09, GPR13, IR05, LMS20, MM92, QW20, SS15]. Reality [CM12, JJK20, Run87b, Run02]. Realized [ABJT22, ACD09, BRY13, BCCL19, CGLS18, DH22, HL06a, HH16, LSS16, Nie07, Oom06b, OJLD18, Var16, YO19, YLN17, ZS14]. Really [AK84, JJK20]. Realtime [FW09]. Reanalyzing [FPAB02]. Reappearance [Bar83a]. Reappraisal [TY89]. Reassessing [AL95]. Received [Ano85d, Ano85e, Ano85f, Ano85g, Ano86g, Ano86h, Ano86i, Ano86j, Ano87f, Ano87g, Ano87h]. Receivers [Dag87]. Recession [MA23]. Recessions [Cav22, RW09, Sta08]. Recipes [Ghy00]. Recipiency [McC95]. Recommendations [Jae97, LLSW18a, LLSW18b, Mül18, Vog18, Wes18a]. Reconciled [KMMP22]. Reconciliation [JSSvN20]. Reconciling [CMW23, GMP11, Jae97, MAB^{+}23]. Reconsidered [Phi87, Zue03]. Record [Str22]. Record-Specific [Str22]. Records [ABS93, But85, Car85, JS85, Wai85]. Recovered [CO99]. Recovery [HHL^{+}22]. Recreational [GT96]. Recurrent [NTGK22]. Recursive [BLS92, IR05, Lab90, OV84, PPR03, Tay02]. Redirecting [Gol84]. Reduce [Pow22]. Reduced [BDW96, CMKSW03, FW09, Olm09, Ter85]. Reduced-Form [Olm09, Ter85]. Reduced-Rank [BDW96]. Reduction [BH09b, Bra07, LS20b, LZ19]. Reemployment [TA89]. Reexamination [Sto88a, Tuc92]. Refereeing [Ano95c]. Reference [CGL20, Yog08]. Reference-Dependent [Yog08]. Reflections [And00, KLL20]. Reform [LW09, VB95]. Reforms [HS89]. Refreshment [HR19]. Regime [AB02, AvdBL20, BTZ04, Cai94, CHT17, GP12, Ham94, Kim93, KK15, NPZ01]. Regime-Specific [GP12]. Regime-Switching [KK15]. Regimes [Aud06, BJOS99, Coo98, Jus98]. Region [MM19]. Region-Specific [MM19]. Regional [BZ99, CHN97, Fom86, KPS00, LM90, PR98, PSW04a]. Regions [HL22]. Register [LW09, MRW12]. Regression [Bab22, BBD20, BDW96, BI20, BFV22, BG23, BCIK22, BS08, BT01, CCÖU19, CMKSW03, CW96, CPDH15, CDY13, CGQ20, CZZ22, Col91, CR83, CR84, Das91, DM10, DM99, DC86, DZ04, Don19, FXMFY20, FMT97, FLM16, FLYZ20, FSC03, FH19, GLLS11, GSC03, GI19, GHLT19, Gol84, GRL13, Gre84, GLLM20, Gur91, GE12, GM83, Hag83, HHL18, Han17, HHKO22, HS85, HW90, HZ22, HF87, Ioa95, Ish22, Iwa91, JY21, dRJH11, Juh14, KP10, KPS20, KW19, KB10, LLT^{+}15, LP12, LK93, LLP15, LLL20, LLY20, LZW^{+}22, LHYZ22, Lia12, LLL19, LXHT22, MQZ22, MSW85, McC19, MB88, Mel05, MB84, Mou87, MW22, NK84, OHS14, PZZ20, PX21, PRG^{+}22, PLCW22, PRC84, PF84, QY19, Qua88, Rac97, RS15, SK91, SK20, TK10, Tay02, Tsa85, Tyr83, VLS90, WlMCP98, WLJ07, WLTL20, Wes01]. Regression [WN89, WHKCS14, WZ15, XHL20, XKZ21, YT07, YLPC20, You88, YF20, ZWLL22, ZLS19, ZLWW20a, ZWZ22, ZZZ^{+}22, ZLLC23]. Regression-Based [BT01, Tay02]. Regressions [BGS22, BLC20, DH20, FGH19, GP12, Han92, Han02, HSA87, RLYZ22, RW22, Siz13, Yit96]. Regressor [GP17, Lew12]. Regressors [Ang01, CFLS19, Car01, Hos18, JS20, LP12, MK01, MV91, Mis95, MB13, OHS14, RS09]. Regular [EYZ20]. Regularized [LHM20]. Regulation [RW86a]. Reinterpreting [Erm91]. Reject [BC89b]. Rejoinder [AV05a, AIPR18a, AGO^{+}14b, AGS20a, ALT08b, BM23a, CR16b, CR84, DSSW07b, Die15b, FQX14b, FRS83b, GKR22b, GPI13a, HL06b, HC94b, HT14b, JJKL22, KM09a, KR94b, LLSW18b, Mül14b, PT12b, PSW04b, WZ20b]. Related [Bla89, GNK86, Hof91, MS18, VLK04, VLS90, Wil83]. Relating [Bor85]. Relation [CO19, Coc89, FV95, Nie07]. Relations [BG93, GK96, HP10, Kon98, SW96]. Relationship [BBC04, HHW91, KR90, Kro83, Met98, NV12, Ser88a]. Relationships [CM88, EPFK15, HPH87, PRW06]. Relative [App85, HHPS19, Kak84a, KJAG85, McG91, NG89, Ril91, Slo87, ZO04, vH89]. Relatives [KC85]. Releases [CG17]. Relevant [Sle98]. Reliability [dL90]. Reliable [FL97, QW20]. Remaining [JR22]. Remarks [Gal93, Gew93, Tau93c, Zel93]. Remembered [BR95]. Rent [Ver12]. Rényi [HMR20]. Repeated [BFP06, Pfe91a, kWM90]. Reply [BH84b, BH85, BL88a, Cha93b, CR88a, DM94a, EK93a, FMB^{+}98b, Fre86c, GGS96b, HF89a, HS88, Jab85, JPR94b, Jan88b, Kak84b, Kak85, KR92b, KS96b, KR94a, KR94b, Lit88b, LS98b, MRZ92b, McD87b, McN86b, MR91b, MS88b, Ott92, Pfe91b, Poi91b, Run87a, SB89b, Sto88b, Tau86a, Vin85b, Wri85]. Report [And05, And06, ALN07, ERT88, Gew91, Gew92, GH02c, GH03, GH04b, HW11, LNHW10, NL08, NL09, Poi89, Tau93a, Tau94, Woo00b, Pie83]. Reported [FL83]. Reporting [KSM97, SvD06]. Reports [GOV84, MO88]. Representation [PPT92]. Representations [RS94b]. Reproducing [DB08]. Research [AGS20b, Bon20, HHW91, JW99, MN16, McG95, PBC95, Tab20, Tam20, Woo00a]. Researchers [Jae97]. Reservation [ELLV97]. Reserve [AGO^{+}14a, LPV03, Pie83, JLG04, Sal95]. Residence [AHI99]. Residential [GGKS98, HK94, LS94]. Residual [BAT83, KD00, Nie04]. Residual-Based [KD00, Nie04]. Residuals [DM99, LL10, San17]. Resolution [LZW^{+}22]. Resolution-Wise [LZW^{+}22]. Resolve [Wes15]. Resort [Vos16]. Resource [DLP19]. Respond [SYA08]. Respondent [BL93]. Response [AL84, ACL11, Bel04, BD01, CSZZ17, CG22, Due99, DV04, GM83, HT83b, HS16, Koh83b, LA84, Lin97, MMG^{+}06, MO88, NWK10, RW86b, RW96, Ter85, Tse89, dNS23]. Responses [DL92, GMP11, HRV20, Koc10, KOS94b, MN20, NSS07, Wri00b]. Restaurant [Xie22]. Restocking [SBK18]. Restricted [Ell00, MB88]. Restriction [Bar83a, BN12, CK19, Xie22]. Restrictions [Bau18, BB22, Cha92, CDJW10, DLP19, FL97, GKR22a, GKR22b, Gos10, GJ01, GPP98, Kil22, KLGS10, PM22, Ram94, RR22, Tao19, Ter85]. Results [FL97, FP84, GKR12, Jai89, MAB^{+}23, Sil86, SM83, VP92]. Retail [AJL96, CRZ19, KT95, KBR95, PSD14]. Retailer [SBK18]. Rethinking [Wes15]. Retirement [Agu10]. Retrospection [Ana09]. Retrospective [KSM97, MO88, NP85]. Return [AG10, AK95, BC91, CO19, Coc89, Des07, EPFK15, FOTY20, HCP11, KC85, LL94, LZ18, MM09, MN95a, PQ10, Poi92, Siz13, SMU13]. Returns [AB88, Ana09, AV07, ABD15, BAT83, BKT11, Bek95, BCSS20, CD09, CM02, Che07a, Con90, CR88b, DH20, Dos05, DP13, GSA04, GM20, GS06b, HS96b, HP85, Jac00, JLP23, Joh02, KWD13, Kar95, Kli01, LLW90, LK88, Lux08, MM00, MV18, MPV20, MlM87, OJLD18, Qi99, Rac01, Rit86, TV15, Tuc92, Wat00, Wol00, dL98]. Revealed [Sal95]. Revenue [CJW86, Gor89]. Revenue-Function [Gor89]. Reverse [CR83, CR84, Gol84, Gre84, Iwa91, MB84, WN89]. Reversible [Asa02]. Reversion [VP93]. Reverting [Due97]. Review [Ano93a]. Reviews [BG88, CRL^{+}85, CRTO86, CBC^{+}87, DCW^{+}86, DJA88, DC85, KP87, Mil85, MOJ^{+}86, NNH87, NNW88, PAC85, SB88a, WCM^{+}86]. Revised [Gbu89]. Revision [Cle17, CFS09, JSSvN20, KK12]. Revisions [CG17, DL87, dJ87]. Revisited [AJK18, DP19, LN20, Ser88a, Wol00]. Rewards [BDJ12]. Rice [BCG85]. Rich [BGS20, Man15, MVVZ19]. Ridge [DC86, ZLS19]. Right [AS20, BFV22, DMP19, Hay89, PPS19, SvD06]. Risk [ACvD22, AKLS20, AFT20, ACJK22, AFO21, BG93, BR08, BTZ04, BM23b, BK16, Bau18, Ber01, BGR15, CO19, Che07b, CZ19, CG00, DPS22, DBCB19, DL89, Edw08, EO10, EPFK15, EH19, FID16, GLLS11, GA10, GCC11, GSA04, Gua23, GJP02, Guo98, HK19, HHPS19, HPZZ21, Hog19, Hon23, HRV20, HLQS17, KL08, LZ18, LP03, LK88, Luc00, LSZ14, Mor23, NRW01, OP18, Paa88, Tay19, Web00, XZW14, ZO04, Zha20]. Risk-Aversion [ACJK22]. Risk-Management [Luc00]. Risk-Return [CO19, EPFK15]. Riskiness [Sho11]. Riverside [Deh03]. Roberts [FG84, Mil84]. Robust [And18, AKKL21, BGMP21, BV05, CGM11, CKMS22, CK08, CHNY22, FXMFY20, HHL^{+}22, HH04, IM10, Jan88a, Kim22, KM09b, LSCX22, MH17, MW22, OP13, PRC84, PT18, PT23, RI12, TM19, Tay03, WLJ07, WHYC22, Wes14, WW09, Wu19]. Robustification [Son12]. Robustness [GP98]. ROC [LY18]. Role [Che07b, FHLW00, GR12, HPSW88, KV13, LN20, MNO85, Pak11, Ros08, Rot15]. Rolling [AG02]. Rolling-Sample [AG02]. Root [BLS92, BBC04, BK12b, Can98, CSZZ17, CC97, DH12, DK00, EG98, Ghy90, Hal94b, KS94, LM94a, Mac94, NPZ01, Pan91, PGFF94, Per90, PV92b, RSW18, RF01, SW21, Sen03, SL99, TS01, Tay02, Wes14, Wes15, Wes18b, Wri00b, ZA92, ZA02]. Root- [CSZZ17]. Root-N [SW21]. Roots [BT01, FH94, Hal94a, Sch89, Sch02, SK99]. Rotating [Val91]. Rotation [Sol86]. Rounding [MM10]. Roy [BKT11]. Rubin [MS03]. Rule [BK91, Ros08, Sim90, Web00]. Rule-of-Thumb [Web00]. Rules [FS91, GR11, IRR22, Ing90a, LH99]. Run [BK11, BJ99, Coc89, FL97, Gos10, GS91a, Har83, KOS94b, MTLJ97, MM09, MPV20, Mor92, Pic01, TWW13]. Running [BBD20]. Ruud [KT18].

S&P [ACW97, PdV06, PT00]. Sacrifice [CR01]. Sales [AJL96, ACL11, GGKS98, LSCX22, PSD14, TN92]. Sample [Ali84, All90, AS96a, AG02, AK95, BCHL20, BE96, CR16b, CR16a, CsL94, CF15, CdH96, Chr96, CLLM03, Cla96, Don19, DMMN84, GH22, GNK86, GKR12, Gro86, HHY96, HLR87, Hof91, HS99b, HLZ04, HL22, Hos19, HLM16, JS98, Jar87, Kot87, Lum95, Mer99, Nev03, PZZ20, Pre95, PT18, PT23, RW86b, RI12, TT83, Val91, Vos16, Wol87, ZT06, vdKK03]. Sample-Audit [Pre95]. Sample-Selection [DMMN84, HLR87]. Sampled [KMR19, ZTCW15]. Samplers [DF15]. Samples [Bra07, CT97, HR19, HKKP01, SM83]. Sampling [AC93, ABS93, Cui20, DZ04, Due99, Hor98, KLS15, Kot84, NP90, Oom06b, PRG^{+}22, SWW88, Sch83, TM86, Val88, WNG89]. Sargan [Are02]. Saving [FP97, HP89, Nor94]. Savings [BFR97, BMM97, Cag03, Cle85]. Say [FID16, RW96]. Scalable [LMN22b, ZWZL22]. Scale [AB22, Con90, EN94, OTZ91, RV19]. Scaled [ZLWW20a]. Scales [JS87, YSD03]. Scaling [CP10, WD96]. Scan [CZ19]. Scanner [Mel18, Rei99, dHK14]. Scatterplot [Dol99]. Schemes [Oom06b]. Scholl [Fer97]. Scholl-to-Work [Fer97]. School [LZ95, NW95]. Schooling [AK95, Des07, Kli01, PvP06, SMU13]. Schools [BDJ12, ERT88, MNO85]. Science [LZ95]. SCLS [Sil01]. Score [BCHL20, BBCL20, CLQ20, FIL^{+}23, HH21, HHK11, JR03, LK93, LLL11, MPV20]. Score-Driven [BBCL20, HH21, MPV20]. Scores [MT09]. Scoring [GR11, IRR22, RCW83]. Screening [HLW14, LT12, ZPWW22]. Search [BS20b, DHW17, KD08, KS98, Kou13, Lan97, LW19b, LNS11, SJS18, VG97]. Searching [Chr92]. Seasonal [Abr85, BH84a, Bel87, BH02, Bey01, BW84, BT01, Can98, Can92, CH95, CF99, CW86b, DM84, DL87, DH88, FMB^{+}98a, Ghy87, Ghy90, GGS96a, Ghy97, HKR97, Hil85, Jai89, LHM20, McE17, McK84, Oli87, Öll86, OF97, OS89, Pfe91a, Pie83, Pie84, PGC84, Sim85, TS01, Tay02, Tay03, dBM90, vdHH86]. Seasonal-Adjustment [FMB^{+}98a]. Seasonalities [GK83]. Seasonality [Alv04, BH03, Hei83]. Seasonally [AW84, vdHH86]. Second [BC11, Bor89, EKM87, SD87, SS06, WV11]. Second-Choice [Bor89]. Second-Order [SS06]. Second-Price [WV11]. Secret [ELLV97]. Section [BGR15, Chr96, Eri98, JW99, JR22, Ng06, TV15, dJ00]. Sectional [Bor89, GDG^{+}05, HLM05, SV16]. Sectionally [Kim22]. Sections [GL13]. Sector [LM89, MF05, PRC84, VLK04]. Secular [Gew88b]. Security [BAT83, Car85, CR88b, DG99, LK88]. Seemingly [FH90, HSA87, MB88]. Segregation [BK05, CT97, LS16, Rat12]. Selecting [CH14, DK00, MV91]. Selection [BCW16, BCFFL19, BOP23, CHL18, CF15, CLLM03, DMP19, DMT06, Don19, DMMN84, EKM87, FHLW00, GA10, GH22, Hal94b, HLR87, Hos19, KP22, KMM21, Kra07, LS20a, LD19, LH99, LLR13, Li15, LNS11, Lu15, McC19, Nev03, Ost19, PZZ20, Pel21, PT00, She93, Sin21, SL11, SW95, TXLT22, Vos16, WLJ07, WLW19, WHT22, XCK21, ZWZ12, ZS15, ZT06, vdKK03, LXHT22]. Selectivity [Har91, MNO85, vO00]. Semi [FE22, LCZT22, LMS09]. Semi-Nonparametric [LMS09]. Semi-Parametric [FE22]. Semi-supervised [LCZT22]. SEMIFAR [BO01]. Semimartingale [PdV06]. Seminonparametric [Bru92, GLL93, Hal90]. Semiparametric [ABS99, AJK18, AGV19, Bel04, BGG22, CG19, CWW20, Das08, DW04, EGR91, EPFK15, FLW96, Goo96, GL20, Hal94a, Hos18, INT21, Joc18, KWD13, LT00, LHLF02, LPV03, LLY20, Lia12, LGG22, Liu21, LSU11, MSU10, Mos91, NEV15, PZZ20, Ril91, SK99, SL11, Tay19, Yan20b, YZK19, YSD03]. Seniority [Dos05]. Sensitivity [AV05b, AL95, BS87a, Cal91, GMP11, Jef08, KBR95, Nor94, Van08]. Separability [Atk83, BC89a, FW03, Koh83a, SW88, TD84]. Separable [Koz20, Mos91]. Separating [HS85, HS19b, You88]. Separation [BT14, Joh87]. Sequential [BLS92, Cha92, DF15, HLM16, Lec09, Li11, Rav06, WLW19, ZLWW20a]. Serial [BR89, BV05, BT01, CHNY22, DM99, Die88b, KR92a, LTY19, NS10, SV16, Vog98]. Serial-Correlation [DM99, Vog98]. Serially [LY18, May86]. Series [Abr85, AC93, AR88, And94b, AW84, AV86, AP06, BGS22, BN05, Bal93, BH84a, Bel87, BH02, BW19, BS20a, BK12b, Byk23, Can92, CJW86, CNT20, CT90, CW86b, CM96, DPS22, DM01, DW20, DRV20, Due99, EH90, ELZ13, FK86, FH90, FZ13, FSK08, GFdH87, GDG^{+}05, GK83, Góm99, Góm01, GHPS22, GS91a, GM20, GPP98, HM89, Hal94b, HV87, HP11, HT83a, Har85, HF89b, HK92, HS85, HM16, Hog22, JS98, Kim93, KYK^{+}21, KL08, LT19, LS02, LMT96, LZLL17, LS15, LHM20, LTY19, Lit83, LdV99, LZGL23, Lüt89, MSW83, Mar83, Mar85, MG84, Mer99, MSU10, Mül14a, NP97, Nef93, Ott90, PV03, Peñ90, PPT92, Per90, PV92b, Pic01, RS15, RF15, RS95, RF01, SH87, SD87, SL01, Sil86, Sim85, Smi05, SS06, SW96, Tay02, Tay03, TM86]. Series [TT83, TMH^{+}23, WZ00, WS20, WHYC22, WZS22, Wec89, Wei86, Wil83, kWM90, XZW16, Yam16, You88, ZH92, ZSZ20, dJ00, dBM90, vdHH86]. Serniarid [FP97]. Service [BZ13, HK86, LM89, VLK04]. Services [BH94]. Set [JLS16, Koz20]. Sets [Lus96, NWK10, Wri00c]. Setting [LLT^{+}15]. Sex [AKO12, Sch87]. Shape [BP22, FXMFY20, FB97, FPAB02, HCP11, LP03, YCJK20]. Shape-Constrained [YCJK20]. Shapes [Nef93]. Shapiro [Bon20, Tam20]. Share [Atk83, BM93]. Shared [PZ90]. Shares [AKLS20, CPP18, DLP19]. Sharing [ZO04]. Sharp [CCKR17, QY19, Rus19]. Shift [CT90, Tyr83, Vog98]. Shifts [AC93, Bal93, BTZ04, BT03, KK15, PV92a, PY09, PQ10, Qui95, RW96, SL00, Smi05]. Shock [Ing90b, ZA92, ZA02]. Shocks [BN07, BGS20, CG22, CM96, FHDS11, GMP11, LLL22, Lu22, MN95b, MT18, San20b, SZ20]. Short [AS12, AK06, Coc89, CGK97, Gas85, Haa96, Har83, MTLJ97, MPV16, Mor92, Pic01, Smi00, Smi02]. Short-Run [Coc89, Har83, MTLJ97, Mor92, Pic01]. Short-Term [CGK97, Haa96, MPV16, Smi00, Smi02]. Shortfall [Tay19]. Should [AGK13, GI19]. Shower [LN20]. Shrinkage [BOP23, FZ18, FS22, HHL18, Han19, HF19, HKO20, Kli22, MTLJ97, SW12, WLJ07]. Shrinkage-Based [BOP23]. Siable [AB88]. Siable-Law [AB88]. Side [Chi85]. Sided [CW86b]. Siegel [KMV10]. Sieve [Hor98, SWJ19, WNG89]. Sign [BB22, FLM20, LKFL22]. Signal [CEJ^{+}12, FP01, HHL^{+}22]. Significance [FW05, FL83, Gro86, MR91a, Rac97]. Signs [Wri00a]. Silver [CM88]. Similarity [DLP19]. Simple [AIPR18a, AIPR18b, ALT08a, AL18, FFX16, GI11, GK96, KT18, Ng08, PT92, Por18, RW98, WS20]. Sims [Ano02b]. Simulated [GS91b]. Simulation [AKKL21, BCS08, CF02, KD00, LS15, TT06, vdKH89]. Simulation-Based [AKKL21, BCS08, CF02, LS15]. Simultaneous [CHL18, IP12, Kra86, Liu14, Lu22, MY18, TR86]. Simultaneous-Equations [Kra86, TR86]. Singer [BV05]. Single [CGL13, CLQ20, FHWZ18, HW20a, HS16, JY21, LTY^{+}22, LQZ19, ZWL22]. Single-Index [CGL13, HW20a, JY21, LTY^{+}22]. Single-Index-Based [FHWZ18]. Single-Variable [HS16]. Singular [ABJT22, LHM20]. Singularity [CS86]. SIPP [CHSS04, CMW23]. Site [BK91]. Size [GIR19, HPH87, HHW91, KJAG85, Rad83, RI12, Slo87, Sto90]. Skeptic [Ric93]. Skew [BL05, JS10]. Skew- [JS10]. Skewness [BC91, BN05, FT12, Joh02]. Skilled [WHT22]. Slaves [CE08]. Sliced [CZZ22]. Slope [CGJ19, CHNY22]. Small [All90, AS96a, BF15, Bra07, BE96, CT97, CsL94, CdH96, Chr96, Cla96, FMT97, GMP11, Hof91, HKKP01, Isa90, Kot84, MlM87, PB91, PFS98, PT18, Rat12, Yan20b, PT23]. Small-Area [FMT97, Isa90, PB91]. Small-Sample [All90, AS96a, BE96, CsL94, CdH96, Chr96, Cla96, Hof91, PT18]. Smirnov [CPP18]. Smoking [DHZ10, Kra07]. Smooth [KDM20, LHLF02, LW15, LTvD03, RV19, VFL99, YZK19]. Smoothed [ACR19, Öll86]. Smoother [DB08, Dol99]. Smoothing [FGH19, HLWY14, Ros98]. Social [Car85, CP10, DG99, GPI13b, Hos19, HYSW16, LS16, Mel22, RS22, YLYJ22]. Socioeconomic [FHDS11]. Software [dlV85]. Solution [KS20, TU90]. Solve [MP05]. Solving [BCR90, Chr90b, Col90, Col93, Gag90, Ing90b, Lab90, McG90, Sim90, Tau90, TU90, dHM90]. Some [And00, BSS01, CJW86, CM96, FP84, Ghy00, HM89, HHY96, HP11, HH04, HK17, Jai89, Lec95, Mar99, Mel18, Mil84, PF84, PS96, PB91, Pis95, Ril91, RW98, Sch83, Tsa16]. Somewhere [APW20]. Source [Yao08]. Sources [CP19, Isk20, KOS00]. South [Smi00, YSD03]. Sovereign [LSZ14]. Space [BCR90, Cre17, CEJ^{+}12, DB08, Ing90b, Kli22, KLGS10, KB04, KLS15, MKL07, MPV20, Qia19, RS94b, WLW19]. Spacings [Ng06]. Spain [GFdH87]. Spanish [Ram94]. Spanning [AHPT19, GL16]. Sparse [Kri17, Li15, PX22a, VLS90, WLTL20, WZ20a, ZL22, ZLWW20a]. Sparsely [KMR19]. Sparsity [HKO20, UY23a, UY23b]. Sparsity-Induced [UY23a, UY23b]. Spatial [BGKS04, DL20, HL16, HHK19, Hos18, LS20a, LC85, LYY22, LGG22, Lu22, MMG^{+}06, MW22, SQ17, ZL22, ZTCW15, ZLWW20b]. Spatially [Ive01]. Spatio [GGKS98]. Spatio-Temporal [GGKS98]. Speaking [DV04]. Special [BFT16, Chr96, Eri98, HF86, JW99, Kas22, KW97]. Specialty [LB16]. Specific [BBD20, CFG95, GP12, Han19, HMvdB12, IM97, LLL19, MM19, Str22]. Specification [BBC04, BR89, BCS08, Cat20, DP22, ELZ13, FM90, GRSY11, HL83, Hay89, Hog22, HLZ22, Lec95, LP96, MK01, MT09, PRPS22, SQ17, ZZZ^{+}22]. Specifications [IRS15, LW15, PG93, PPL94, WSX22]. Spectral [CKY85, LM94b, MHCP22]. Spectral-Temporal [LM94b]. Spectrum [AW84]. Speculative [FV95]. Speed [HR91, Pat86]. Speeds [Lin86]. Spell [KLN85]. Spells [TA89]. Spike [Gra19]. Spike-Investment [Gra19]. Spillover [FLM^{+}19, HS19b]. Spillovers [LN20]. Splicing [HF97]. Spline [RLYZ22, RE12]. Splines [BS08, WHKCS14]. Split [AK95]. Split-Sample [AK95]. Splitting [Jar87]. Spot [BHMR19, HLZ04]. Spreads [ABD15, OP18]. Spurious [LS98a, ORT08, Qu11]. Squared [CW96, Mar87, Rit86]. Squares [Hag83, HLR87, Jan88a, Lab90, NV12, Phi10, Qua88, YCJK20]. Stability [CH95, Car06, DMP19, HS99a, HLR^{+}22, Lüt89, Ost19]. Stabilize [vUS87]. Stabilized [Hor98]. Stable [AL89, CF99, CR88b, Fri22, LLW90, Li15, McC97, Oli87]. Stable-Law [AL89]. Stage [CLLM03, EKM87, GHPS22, JSS88, Kra86, MHB19, Val91]. Stages [Pop84]. Standard [FFX16, GHR87, HT06, KKPS22, RS92]. State [AP06, BCR90, BH93, Car84, Cre17, CEJ^{+}12, FE22, Fre86a, Fre92, GSA04, Kea97, Kli22, KLGS10, KB04, KLS15, MKL07, MPV20, PX22b, Pro12, Qia19, RE05, TH83]. State-Dependent [BH93, Car84, GSA04]. State-Space [KLS15, Qia19]. State-Switching [AP06]. State-to-State [Fre92]. State-Varying [PX22b]. Statement [Zel83]. States [Bar85, KK15, ZCZ^{+}21, Bar85, Bie00, BFHM11, Cra86, Fer97, GD89, Hau91, HP89, HH83, Kar95, Kim90, KM08, KMMP22, MLR19, RS94a]. Static [BHKN10]. Stationarity [BT03, GHLT19, HLM05, LMT96, LH99, Tay03, Wri00c]. Stationary [Ali84, AMNP18, BGKS04, DW20, GS91a, HK14, HW20b, Nie07, Pan91, PY09, Sen03, Shi00, WV22, XKZ21, Zha15]. Statistic [IM10, TM19]. Statistical [AR88, Ano84c, BD09, BHMR19, BL88b, Che12, CM93, Ghy90, GH91, Goi85, HHPS19, HC94a, HC94b, HYSW16, JS85, KR94a, KR94b, Lit86c, MS03, Nef93, NTGK22, PRT00, Rod84, Rod16, Rub86, SvD06, Tau86b, Wri83, ZLS19, ZFSC00]. Statisticians [Don22, JJKL21, JJKL22, LC22, MLZ22, WF22, ZK22]. Statistics [AV05b, BC11, CHLR92, ERT88, FMT97, Gbu89, HPSW88, HMR20, Isa90, MV91, Nor85, RMS88, Tar07, Ano00c, Ano01, Ano02d, Ano03a, Ano03b, Ano04b, Ano04c, Ano05c, Ano05d, Ano06b, Ano06c, Ano07c, Ano07d, Ano08c, Ano08d, Ano09c, Ano10c, Ano11c, Ano13c, Ano13d, Ano15c, Tau93b, GH02a, TR93]. Status [CHW20, DKT17, FHDS11, McK18]. Steel [Rog87, TD84]. Stein [FS91, LZGL23]. Step [BSS22, Hof91, Hos19, KW90, Knü15, MT85, MT02, PRG^{+}22]. Stephen [You88]. Steven [Sim85]. Stickiness [FLS07]. Sticks [AvdBL20]. Sticky [GR89]. Stochastic [ABS99, AS96b, And01, And18, Ant83, AHPT19, BCR90, BMY04, BC11, BB20, BCCL19, BT14, CCP13, Cat20, Cha15, CK22, CHW20, Chr90b, Cla11, CI87, Col90, CHL88, DE13, DP13, FLW96, FT12, FH98, Gag90, GSC96, GS10, Had99, HS96b, Hen22, HPS96, HW20b, HJL22, HpK23, HLM15, Ing90b, Ioa95, JPR94a, JPR02, JMZ18, KB04, Lab90, LS20c, LNS11, MK01, MPV16, McG90, MHCP22, NSR88, NTGK22, PG06, RS92, RTW12, ST10, Sel89, SW22a, Sim90, Smi02, SLL98, SL99, Tau90, TU90, TT06, Tsi22, Wan03, YO19, YZK19, ZFSC00, ZS14, dHM90]. Stochastic-Coefficients [CHL88, NSR88]. Stock [ACW97, AB88, AKLS20, AV07, BC91, BMP10, BO01, CD09, Car06, CM02, CHN97, DF11, DH20, Due97, GS06b, HCP11, HP85, JJR18, JLR11, KWD13, KNW22, Kar95, LL94, LLW90, LZ18, LS98a, LV00, LT04, MM00, MN95a, MM92, MY18, MPV20, Mül18, NTGK22, Pat86, PQ10, Qi99, Rac01, RT00, Ric93, Rit86, RU01, Shi00, Tim01, Tuc92, Wes18a, Wol00, ZHCB16, dL98]. Stock-Market [ACW97, BO01, Due97, JLR11, LS98a, LV00, MN95a]. Stock-Price [MM92]. Stock-Return [LL94, MN95a]. Stockouts [SBK18]. Stocks [ACD09, HHPS07, MKL07]. Stone [You88]. Store [ACL11]. Store-Level [ACL11]. Stores [CRZ19]. Strategic [BHKN10]. Strategies [CGK97, GKP00]. Strategy [BL93, BN12]. Stratification [FLM^{+}19, GRW84, Kot85]. Strictly [KR92a]. Strike [BGR15]. String [AJK18]. Strong [BV05, IP12]. Strongly [ARBS87, LS02, Mül14a, Pos00]. Structural [AT17, And94b, AGS20b, Bäu13, BB20, Bon20, BS20a, BT03, Cal91, CR01, CWW20, HS99a, Han97, HL83, HT83a, HW04, KW22, KNW22, KW97, KP10, Kew20, KMN05, Koc10, LM00, LL90, Lan97, LL10, LL19, LLL22, LLF22, Lüt89, Mar85, Mel22, Olm09, PPR03, Pro08, Qui95, RS94a, Ros98, SL00, SL20, Smi22, SW96, Tab20, Tam20, Tau86b, Tim01, Vel22, WZ00, WL21, WV11, Wol87, ZLLC23, dBM90]. Structural-Change [Han97]. Structure [ACW97, Ant83, BMP10, BTZ04, BHZ20, BRW12, Bau18, BC89b, Cha20a, Cha20b, CP98, Cla96, CHL88, Die15c, DGP20, EG98, Ghy94, GL20, Isk20, KLZ20, KBZ89, KMV10, LFWT18, Seo20, Ser95, She92, SW95, WLTL20, YLL19, dJ00]. Structured [Aud06, Jus98]. Structures [AS96a, Aru20, LZ18, SWJ19, Swa03a, WHYC22]. Student [FSC03, GRTZ22]. Studies [Pow22]. Study [AL89, AS96b, BC89a, BMM20, CKY85, Coc89, DL92, DH89, ELLV97, Fre85, Fre86a, GdXPE16, HM89, HV87, HT83a, Hau96, MZ23, PS92, Pis95, SB89a]. Subannual [HS93]. Subdomain [Sto90]. Subject [AC93, Cle17, KK12, PZZ20, TR93]. Subjected [GPP98]. Subjective [BBK12, BdB18, BL88b, EMW09, OS20]. Suboptimal [Ash85]. Subsampling [Qia22, ZPWW22]. Subset [EKM87, GK96]. Subsidies [AC03]. Substitutability [Ser88b]. Substitution [CFM11, DG96, DF20, EH90, Gor89, GB87, HK86, NRW01, Pak11, TD84, Web00]. Substitutions [Pak11]. Sufficient [LZ19, YYX20]. Suggestion [Mil84]. Sum [Qua88]. Summary [FMT97, GO98, NP85, Pie83]. Sums [BGKS04]. Sup [CCKR17]. Sup-Norm [CCKR17]. Superior [Han05]. Superpopulation [Kot84]. supervised [LCZT22]. Supplier [KM08, SBK18]. Supply [AAW11, Bar83b, Blo19, Chi85, CO99, HS85, JGL19, LeS92, Oku11, PS92, RW86b, Rog87, You88]. Supply-Side [Chi85]. Surfaces [MMG^{+}06]. Survey [ÇD22, CMW23, DL92, GOV84, HV87, HS96c, HT20, KL90, KSM97, Kot87, LZ87, MRW12, MAB^{+}23, MN20, PT11, PJ20, PFS98, PBC95, RMS88, SYA08, SWY02, ZW93, ZHB23, dNS23]. Surveys [ABS93, BL93, HS93, Lit88a, MM10, Pfe91a, Tar07, TH83]. Survival [BdB18, LW97]. Suspensions [BS98]. Sustainability [Qui95]. SVAR [BB22, Gos10]. SVARs [JL22, Vol22]. Swedish [DO95]. Swings [Kla05]. Swiss [vUS87]. Switches [AB02]. Switching [ACR19, AP06, BCD17, BP18, Cai94, CST18, Cav22, CHT17, CE05, Due97, FO05, Goo93, GM13, GLLM20, Kim93, KK15, Knü09, KHT05, LdV99, Lux08, NPZ01, Smi02, SLL98]. Switching-Regime [Cai94]. Symmetric [BH94, SD87, SK99]. Symmetrically [ABA99]. Symmetry [CJZZ18, Con90, Li99, Psa16, PV22]. Symposium [Ang95]. Synthetic [CLQ20, KKPS22, Kli22, LHYZ22, Pow22]. System [Bar83b, Cha87, Che12, Erg19, GG90, HH83, Ram94, RW86a, Smi00, TWW13, VP93, ZW93]. Systematic [HK86]. Systematically [Lüt86]. Systemic [OP18]. Systems [DG96, DW85, DW88, Fis92, GG95, HJNR98, HW90, LPS99, MY18, Mos99, OV84, RW98, vHPP04].

T [JS20]. Tacoma [MT03]. Tail [AFT20, BS20a, DPS22, EH19, FID16, GI11, Hog19, HKKP01, KK92, LLY20, Luc00, McC97, SCS20, Zha20]. Tail-Index [HKKP01]. Tailed [AS12, CKL11, DPS22, FQX14a, FZ13, OJLD18, YL17]. Tails [LS20c]. Tale [BB89, JJK20]. Target [CT01, PRM99]. Target-Zone [CT01]. Targeting [MP83, YF22]. Tariffs [Kas09]. Tastes [AT17, BS87a, Luo02, OG86]. Tax [ABS93, BFZ98b, CJW86, CHLR92, GG90, Pre95]. Taxes [Bar85, Blo19]. Taxonomy [McE17, PY17]. Teacher [GRTZ22]. Teacher-to-Classroom [GRTZ22]. Teaching [RMS88]. Team [HJS20]. Technical [Kum89, ST89, TD84]. Techniques [Col91, DG02, FK86, LM90, McN86a]. Technological [ML90]. Technology [GMP11, KBZ89, YP08]. Teen [AKO12, FLM^{+}19]. Teenage [JJK20]. Telephone [Hau99]. Tell [MRW19]. Temperature [LA84, Sea93]. Temperature-Trend [Sea93]. Tempered [DF15]. Temporal [GGKS98, JMP16, LM94b, Mar99, RS95]. Temporally [Erm91]. Temporary [Kas09, Ric93]. Tender [JT95]. Tension [LS16]. Tensor [BCIK22]. Tenure [BT14, Gar89]. Term [Aru20, BTZ04, BRW12, Bau18, BC89b, CGK97, CS20, Die15c, EG98, FBB06, Gal93, Gew93, Haa96, KMV10, LZ18, MPV16, She92, Smi00, Smi02, SW95, Tau93c, WLW19, Zel93, dJ00]. Term-Structure [dJ00]. Test [BGMP21, BZ13, BCS08, BRT12, CT16, Can98, CH95, CL95, CW92, FLM20, FH90, FW03, Gar89, GHLT19, Gre89, HMS17, Han05, Hay89, HHW91, HS99b, HMR20, JS98, JLP23, Juh14, LK93, LM94a, LLP15, MF05, NS96, Ng08, ORT08, OP13, PGFF94, PT92, PTZ86, Psa16, Qu11, RV19, Rit86, RU01, SD87, Shi00, SQ17, Tse87, WS20, ZZZS19, ZZZ20]. Testing [AJ05, AGU04, Aiz90, AG05, Ant83, BSS22, BK07, BCCP18, BT22, Ber01, Boh91, Bos96, BK12b, BZ20, Bre19, CGJ19, CJZZ18, Che18, CPP18, CK08, CT01, CS20, DE13, DH12, DHL14, DTU06, DNW98, EYZ20, EJP05, EK93b, EC06, ES01b, ELZ13, FW05, FG20, Fra19, GR15, GRSY11, Goo96, GJ01, GS91b, GS91a, Gre94, GSY01, GL13, Hal94b, HS99a, HJ00, HH11a, HH19a, HD21, Hog22, HP10, HLR^{+}22, JR22, Kap10, KP10, KSW16, Knü09, KFL03, LT19, LLT^{+}15, LD19, LPS22, Li99, LTY19, Lob03, LG20, LZGL23, LRS99, MLST20, MS89, Mos91, NS10, Ng06, Nic89, PW22, PRPS22, Per90, PV92b, PY09, PY19, PdV06, PJ14, PT18, PT23, Rac97, RH04, SL00, San17, SW21, ST10, SK91, SL99, Son12, SRdCFB18, SU22, SU09, SS13, SC17]. Testing [Tsa93c, VFL99, Vog98, WZS22, Wes15, Wes18b, WP83, WSX22, ZZZ^{+}22, ZLLC23, lT85, vdKK03]. Tests [Aiz91, Ali87, AG07, AS12, AK06, AN91, AMRT22, BN05, BR08, BLS92, BC89a, BDB14, BDK07, BBC04, BSS01, BR89, BC11, BFZ98b, Bre01, BV05, Bur94, BE96, BT01, BH03, BT03, CMKSW03, CT93, CsL94, Che07a, CM09, DM10, DM99, DK00, Die15a, DRV20, DFP11, DFK19, DU12, EN94, EG98, FGK20, FFK18, FKO97, FH94, FG00, FMT10, Ghy90, GZ84, Gro86, GL16, Gur91, Hal94a, HM89, Hal90, Han92, Han97, Han02, HLM05, HLN98, HLZ22, INT21, IR05, JM91, Kan85, Ken95, KD00, KM09b, KW92, Lec95, LH99, LLL11, LTTL19, LW15, LZZ20, Lüt89, Mac94, McC19, McG91, ML90, Mis95, NS10, NPZ01, Nie04, Pan91, PT12a, RF15, RI12, San20a, Sch89, Sch02, Sen03, Seo20, SK99, SL01, SW22b, SB19, SW88, Tao19, TS01]. Tests [Tay02, Tay03, WKX23, War89, Wes01, Wes14, Wri00a, Wri00c, YLPC20]. Text [GGW14, KMM21, Pel21, Sin21, WLW19, XCK21]. Text-Term [WLW19]. Theil [Bla89]. Their [Fil94, LSS16, McK18, Sea93, TR86, WJS08]. Them [HO84]. Theorem [All90]. Theoretic [KLP08, KS94]. Theoretical [BG93, HPH87]. Theory [AJK18, BLS92, Bäu13, BHMR19, CM96, DMP19, DTU06, FP97, GNK86, HK14, HP13, HH19b, JMZ18, KD00, Kot84, MB88, Nic89, Ost19, Sal95, She19, Sle98, SW94, VZ11, dNS23]. There [AR91, Deh03, KS17, Nef93]. Thesis [Ano96a, Ano00c, Ano01, Ano02d, Ano03a, Ano03b, Ano04b, Ano04c, Ano05c, Ano05d, Ano06b, Ano06c, Ano07c, Ano07d, Ano08c, Ano08d, Ano09c, Ano10c, Ano11c, Ano13c, Ano13d, Ano15c, Woo98]. Thick [AS12]. Thick-Tailed [AS12]. Thickness [McC97]. Third [Hay86]. Third-Order [Hay86]. Thirteen [AOS00]. Thoughts [Tab20]. Three [Góm99, GLLM20, Sic94]. Three-Pass [GLLM20]. Threshold [AT11, BBC04, CGT17, CCKR17, CHT17, CILY17, CMT19, ES01a, FKK99, GR11, GP17, GS03, Han17, HS16, KP03, LS02, LLZ16, NW13, SC17, TZ19, WL21, YL17, YF20, ZWZ12, Zha20, ZLLC23]. Threshold- [GR11]. Thumb [Ing90a, Web00]. Tidwell [SY94]. Ties [BKW19]. Tilted [XP11]. Tilting [GdXPE16, KCR17]. Timber [ELLV97]. Time [AGJT14, ACvD22, ACW97, Abr85, AC93, AR88, AAMW20, And94b, AK06, AHI99, AW84, ADS09, AFO21, AV86, AP06, BGS22, BN05, Bal93, BRY13, BK12a, Bek95, BH84a, Bel87, BH02, BW19, BGR15, BS20a, BK12b, BH84c, Byk23, CCP13, Can92, CH95, CJW86, Car84, CL87, CNT20, CKLGS12, Cha15, Cha22, CP08, CT90, CGLS18, CGZ19, CT18, CM09, Cla11, CG12, Cle17, CFS09, CR12, CKL11, CW86b, CM96, DPS22, DW20, DBCB19, DRV20, Due99, DG02, EH90, ELZ13, FK86, FH90, FZ13, FSK08, Fuh92, GFdH87, GDG^{+}05, GKMV09, GCC11, GK83, GHS09, Góm99, Góm01, GHPS22, GS91a, GM20, GPR13, GMS87, GPP98, GL16, HM89, Hal94b, HL16, HV87, HP11, HT83a, Har85, HF89b, HK92, HHKO22, HS85, HM16, HPS96, Hog22, HLZ04]. Time [HK17, HKO20, IR05, JS98, JMZ18, JK02, Joh02, Jor99, Kal22, KN89, Kim93, Kli22, KYK^{+}21, KL08, KMV10, KSS09, LT19, LS02, LeS92, LMT96, Li11, LZLL17, LGG22, LS15, LA84, LHM20, LTY19, Lit83, LMS20, LdV99, LXY22, LTvD03, LZGL23, Lüt89, MLST20, MSW83, MSW85, Mar83, Mar85, MM92, MG84, Mel05, Mer99, MSU10, ME04, Mül14a, NP97, Nef93, NK84, OP18, Ott90, PSSE20, Peñ90, PPT92, Per90, PV92b, Pic01, QW20, RS15, RF15, RU01, RS95, RF01, RE05, SH87, SS15, SD87, SL01, Sil86, SH07, Sim85, Smi05, SS06, SW96, SWJ19, Tay02, Tay03, TM86, TT83, TMH^{+}23, TT02, VG97, WZ00, WS20, WHYC22, WZS22, Wec89, Wei86, Wol87, kWM90, WZ15, XZW16, Yam16, Yan20b, You88, ZH92, ZSZ20, dJ00, dBM90, vdHH86, TT06]. Time-Changed [Li11]. Time-Dependent [Car84]. Time-Dummy [Mel05]. Time-Invariant [HK17]. Time-of-Day [LA84]. Time-Series [AR88, HV87, HM16, LTY19]. Time-Variant [ME04]. Time-Varying [ACvD22, AAMW20, BRY13, BGR15, Cha15, Cha22, CGLS18, CGZ19, CT18, CKL11, DBCB19, GCC11, GHS09, GL16, HL16, HK17, HKO20, Joh02, KMV10, LeS92, LGG22, LdV99, LTvD03, MLST20, OP18, PSSE20, SWJ19, TT02, Wol87, WZ15, HHKO22]. Time-Varying-Parameter [KN89]. Time-Varying-Parameters [Fuh92]. Timely [MOZ07]. Times [ACS15, Cle85, Cui20, MABF08, RE05]. Timing [CP98, JLP23, SC03]. Tobit [HSA87, ZWZ12]. Together [KKPS22]. Tolerance [AKLS20]. Too [BKW19]. Top [CPP18, Var16]. Toru [PM22]. Total [DBS84, MPTK91]. Tourism [CF99]. Tournaments [VZ11]. Trace [BN12]. Tracking [VKR06]. Tracy [MLZ22]. Trade [AOS00, KT95, KR90, MT09, lT85]. Trade-offs [lT85]. Traded [Kim05, MS18]. Trading [AG05, BS98, BJ99, FV95, JLR11, LL94, Lie98, LV00, SH87, Wat00]. Trading-Day [SH87]. Traditional [Jan88a]. Training [Lec99]. Transaction [HW10]. Transaction-Level [HW10]. Transactions [Cra86, Rog87, RE05]. Transfer [Abr85, YP08]. Transformation [Abr03, GHNS09, KST11, KKS22, WW20, ZS14]. Transformation-Kernel [WW20]. Transformations [Dol99, Ghy97]. Transformed [HSZ22a, HF87, YT07]. Transient [MSW83]. Transition [Fer97, KDM20, KS17, LTvD03, Mor93, RU01, VFL99]. Transitional [Fil94]. Transitions [BMM97, DM94b]. Transitory [KHT05]. Translating [PG93]. Translog [Hay86, Ser88b]. Transmission [MT18, Sho11, WGBM12]. Transmissions [Kar95]. Transparency [AGS20b, Bon20, Tab20, Tam20]. Traps [DKT17]. Travel [vdKK03]. Treatment [AI11, AHL15, APR19, AvdBL20, BCFFL19, CLT22, CF15, CHW20, FHLZ20, FFL13, FLM^{+}19, FJ08, FM13, Hei22, HLL20, HLZ22, HS19b, Ish22, JLS16, KT19, LRW09, Lu15, McK18, Rus19, San20a, ST22b, TGCD16, VV99, Vos16, ZT06, dHK14]. Treatments [ACL19, Mol10]. Tree [Aud06, YQZ18]. Tree-Boosted [YQZ18]. Tree-Structured [Aud06]. Trend [BLS92, BV05, CKP13, CW92, DL87, DMS99, Góm01, NS96, PY09, PFS98, RS94a, Sea93, Sen03, Ser95, WZ00, Wes15, WSX22]. Trend-Break [BLS92, Sen03]. Trend-Cycle [DL87]. Trending [RSW18]. Trends [AJL96, BT22, CMW23, CHT07, CK22, FBB06, Fdv00, GK83, Har85, IM97, MZ23, MAB^{+}23, PV03, PZ90, ZHB23, BS20b]. Trendy [AJL96]. Trimmed [CVF06]. Trinomial [Ter85]. Triple [ZW93]. Triple-System [ZW93]. Triples [PV22]. Trips [GT96]. Trivariate [ZT06]. Troughs [PSvD09]. True [BS87a, HD85, ORT08]. Truncated [BMM20]. Truncation [BCF94, FS88]. Turkey [Met98]. Turning [LeS91, LD96]. Turning-Points [LeS91]. Turnover [Dos05, HS96c]. TV [JJK20]. Tweedie [YQZ18]. Twentieth [GH02b, GH02a]. Twenty [Die15a]. Two [ABS93, BSS22, CLLM03, FB97, GHPS22, GK96, Hof91, Hos19, JSS88, KW90, Kra86, LT19, Lia12, Lus96, MPTK91, MHB19, MT85, MT02, RF15, Sil86, Val91, Wüt20]. Two-Fold [Lia12]. Two-Phase [ABS93]. Two-Stage [CLLM03, GHPS22, JSS88, Kra86, Val91]. Two-Step [BSS22, Hof91, Hos19, KW90, MT85, MT02]. Type [HK19, HMR20, ZL15, KW92]. Types [BM23b, DL20, DTU06, Gua23, Hon23, Mor23]. Typical [HCP11].

U.K. [GKMV09, MF05, OS89]. U.S. [AVV01, AP06, BGR15, BJOS99, Bra07, BCG85, ÇD22, CR01, Che12, CRZ19, Cle14, CG17, Coo98, CMHM19, DM94b, EN94, FB92, Ghy90, Hei83, HP05, KV13, KNP04, KR90, KP99, KLS12, Kum87, KGM91, LM00, LM94b, McN86a, MAB^{+}23, Mor85, Mor92, Mor93, MS99, MT18, NSR88, OG86, Oht87, OF97, PZ90, RS94b, SV16, Sta08, Wan96, WW09, Yam16, dL90]. Ultimatum [FB97, FPAB02]. Ultrahigh [HLW14]. Unbiased [AC94, IP12]. Uncertain [ARvD18, CC97, CEJ^{+}12, LW97]. Uncertainties [Sea93]. Uncertainty [ACvD22, Cam07, Cle14, Cle17, GKMV09, Haa96, JS19, Kim93, Li15, MN95b, MT18, MM19, Rod16, SZ20, ST22a]. Uncertainty- [Cle14]. Unconditional [FM13, GR15]. Unconfoundedness [DHL14]. Uncorrelated [San17, WS20]. Uncorrelatedness [SU09]. Uncovering [ACW97]. Undermined [HS19a]. Undersized [SM83]. Understanding [CO19]. Unemployment [AVV01, AK84, AH20, BMM20, BZ99, BH84c, Fer97, GD89, GM97, KLN85, KP99, LW09, MO88, McC95, Mer99, OF97, Sol86, TA89]. Unidentified [PT03]. Unification [Lec99]. Unified [Che18, HLZ22, YLPC20, ZG22, ZL22]. Uniform [AMRT22, CDY13, CT22, LLP15, McC19, QY19]. Unimodal [ARBS87]. Unintended [AKO12]. Union [Car06]. Unionization [Bar85]. Unit [BLS92, BBC04, BK12b, BT01, Can98, CC97, DH12, DK00, EG98, FH94, Ghy90, Hal94a, Hal94b, KS94, LM94a, Mac94, NPZ01, Pan91, PGFF94, Per90, PV92b, RSW18, Sch89, Sch02, Sen03, SK99, SL99, TS01, Tay02, Wes14, Wes15, Wes18b, Wri00b, ZA92, ZA02]. Unit-Root [BLS92, BBC04, Can98, DK00, EG98, Ghy90, KS94, Mac94, NPZ01, Pan91, PGFF94, Sen03, SL99, TS01, ZA92, ZA02]. United [Bar85, Bie00, BFHM11, Cra86, Fer97, GD89, Hau91, HP89, HH83, Kar95, Kim90, KM08, KMMP22, MLR19, RS94a]. Units [CT97, LYY22, Rat12]. Univariate [Kan86, MG84, MRW19, NC92, SD87, TM86, Wes15, Wri00b, ZSZ20]. Unknown [CSZZ17, Han17, PS16, Psa16, Tao19]. Unobservable [DMP19, Ost19, Rog87]. Unobserved [AB22, BB05, BCM10, BWH86, BW84, GK15, HK92, HSZ22b, Ioa95, JLS16, KW90, KDM20, Kim93, KHT05, LdV99, Mar87, NPW93, PW22, Pro12, dNS23]. Unobserved-Component [Kim93, KHT05, LdV99]. Unobserved-Components [BW84, HK92]. Unparallel [CK22]. Unrelated [FH90, HSA87, MB88]. Unresolved [Mel18]. Unspanned [CGM16]. Updating [PRG^{+}22, VP92]. Upon [NW91]. Urban [LLN89, MMG^{+}06]. Use [AGK13, Ash85, BDJ12, BT89, DMT06, DC86, Die15a, Góm01, HMI95, IOS90, KLN85, LZ87, NG89, NK84, NP86, TH83, Tyr83, Wes18b, BFS94]. Used [GI19, Oht87, RT01]. Useful [DMP19, DK00, FS22, MM09, MOZ07, PPS19, Sil86]. Uses [Car85, HW84, JS85]. Using [ABA99, BGKS04, Bäu13, BBK12, BHLM07, BFHM11, Bor89, BS20b, Bre19, BWH86, ÇD22, CKY85, Car84, CG02, CGR07, CMT19, CT01, CR88b, CW86b, CHLR92, DL92, DH12, DV04, EH90, EG98, FMT97, FW09, FLM^{+}19, GFHPZ87, GR11, HHK11, HGB^{+}12, HJ00, Har83, Hil86, HLM15, Inc93, JLP23, KLP08, KKPS22, Knü15, KYK^{+}21, Kol20, KS94, KMV10, Kor19, KCR17, Kum89, Lab90, LeS91, LW09, LM90, Lew12, LS15, LHM20, LSS16, LZ19, MTLJ97, MV91, McG91, MRW12, MF05, MKL07, Mit90, MHB19, MT96, MS03, NSR88, Nev03, Ng06, PPT92, PRM99, PSW04a, Pfe91a, Psa16, PV22, Qi99, Rac01, Rei99, Rub86, RFG20, Sal95, SJS18, ST89, SW94, SW02, SW12, SSBD15, SW95, Tau90, Tay19, Val88, VP92, Vel22, WJS08]. Using [Wes15, Wil83, Wri00a, Xu00, Yit96, ZCZ^{+}21, ZHB23, ZT06, dBM90]. Utilities [Cal91, Ver12]. Utility [Hay86, KMF05, Tse89]. Utilization [CW96, LC85, Mor85, OTZ91, RAL99]. Utopia [APW20].

Vaccine [HS19a]. Valid [JL22, LRS99, Str03]. Validation [All90, CR88b, DC86, Pis95, Wu19]. Validity [Aiz90, BI20, DPES19, FGK20, Mal83]. Valuation [FLSVP04, LMS09, Wer99]. Value [Chr90a, Chr90b, EH23, EO10, GLLS11, GA10, GCC11, HW84, IOS90, LHM20, PR87, Swa03a, Tau90, Tay19, WV11, XZW14]. Value-at-Risk [EO10, GLLS11, GCC11]. Value-Function [Chr90a, Chr90b, Tau90]. Valued [BW19, GRL13, PZ19]. Values [CL95, Han97, KRW^{+}14, KSD98, MMG^{+}06, PB91, Ray90, SU22]. Valuing [EC06]. VAR [AV08, CdM19, Jus98, KK12, Koc10, RT01, SL00, SS15]. Variability [Fis92, Hil85]. Variable [ABA99, Ang01, BH94, BBD20, Con90, EN94, FZ18, Ger88, HK01, HS16, Kow16, Kut94, Lin86, NK84, TXLT22, WLJ07]. Variable-Profit-Function [BH94]. Variables [AK95, Are02, Ash85, Bre19, CG22, Das08, Dev07, DG06, Due05, Est98, GM83, HHN08, HMN10, HH04, HK17, HF87, Kan86, KW90, Kli16, Kli01, Koz20, Lou07, Man15, PPL94, PT17, PZ19, Qi99, Rac01, Siz13, SV16, ST22b, Swa03a, WW96, WHKCS14]. Varian [FW03]. Variance [AC93, BB89, BDK07, BT03, Car84, CsL94, DO95, FKO97, FS93, GL19, GL16, HK19, HL06a, IRS15, Inc93, KN89, KB04, LL90, Lee18, LZ18, Mil87, MSU10, Oom06b, PRT00, PT18, PT23, RKL22, SWW88, Tuc92, Val91, Vol22, WZ00, Wri00a, XZW16, Yit91]. Variance-Ratio [CsL94, FKO97, Wri00a]. Variances [BG93]. Variant [ME04]. Variation [AVV01, Bek95, EYZ20, HP13, Kal22, Kol20, SH87, Tah01, vdHH86]. Variational [LMN22a]. Various [Pop84, Slo87]. VARMA [AV08, DS13, DP22, LP96, Vel22]. VARs [CCM16, Cha20a, Cha22]. Varying [ACvD22, AAMW20, AFO21, BRY13, BGR15, CFLS19, CKLGS12, Cha15, Cha22, CGLS18, CGZ19, CT18, CKL11, DBCB19, DFP11, DGP20, Fuh92, GCC11, GHS09, GGW14, GL16, HL16, HK17, HKO20, Joh02, KN89, Kim90, Kli22, KMV10, LeS92, LGG22, LLCL22, LdV99, LTvD03, MLST20, MSW85, OP18, PSSE20, PHPY20, PX22b, RU01, SWJ19, TT02, Wol87, WZ15, XZW14, HHKO22]. Varying-Coefficient [DFP11, DGP20, XZW14]. Varying-Parameter [Kim90]. Vast [LSS16, PSSE20]. Vector [BB20, Cla11, ELZ13, FK86, FO05, GK03, HF19, KPS20, Kew20, LL10, LL19, LLL22, Lit86b, Lüt84, Lüt86, LS22, LRS99, NS05, NSS07, PR98, PPT92, PY17, Run87b, Run02, Sal95, Tsa89, WLW19]. Vectors [Bos96, Ell00, EJP05, GJ01, LRS99, She92]. Velocity [Ser95]. Versus [AvdBL20, DH20, EN94, Fre85, IRS15, Ing90a, LTU15, NRW01, Olm09, RS94b, ACL11, AV08, BBC04]. Very [FHWZ18]. Via [LCZT22, Pos00, YF22, AC93, AG07, AG10, ACD09, CILY17, CG19, DPS22, DHL14, FS22, GLLS11, HPZZ21, HSA87, HSZ22b, JK02, LFWT18, LLL22, LGLY16, LNS11, PW22, PG06, RTW12, YQZ18]. Vietnam [DHS19]. View [Car85, Poi91a, Ric93]. Vines [ZSZ20]. Vinod [Gas85]. Vintage [CG12]. Violations [FW05]. Visit [PSD14]. Volatile [KNP04]. Volatilities [CKL11, LGLY16, RT00, YO19]. Volatility [AS96b, AV07, And01, AG02, AOS00, ABD15, ABD19, ACD09, BRY13, BHZ20, BO16, Bek95, BO01, BMY04, BB20, BJ99, BCCL19, BZ20, BJ06, Cam07, CMW23, CCM16, Cat20, Cav22, Cha15, CGLS18, Che07b, CDJW10, Cla11, CR12, Cui20, Due97, DP13, ES93, FID16, FT12, FV95, FS23, FH98, GS10, Guo98, GS06b, HH16, HS96b, HP05, HPS96, HT14a, HLM15, JPR94a, JPR02, JMZ18, JLK14, Joh02, KG15, Kar95, KSW22, LN20, LL94, LS20b, LS20c, Lin97, LNS11, MPV16, MA23, MS18, MZ23, MAB^{+}23, NTGK22, Nie07, NV12, PQ10, PG06, RTW12, Sho11, Siz13, Smi02, SLL98, SL99, TT06, TT11, kTY12, WV22, XP11, YO19, Yan20b, YLN17, ZS14, Zho96, ZHB23]. Volatility-Related [MS18]. Volume [Ano83b, Ano84e, Ano85c, Ano86e, Ano87e, Ano88b, Ano89b, Ano90b, Ano91b, Ano92b, Ano93d, Ano94c, Ano95b, Ano96e, Ano97c, Ano98b, Ano99b, Ano00b, Ano07b, Ano08b, Ano09b, Ano10b, Ano11b, BJ99, Cra86, FV95, LL94, Lie98, LV00, Wat00]. Volume-Volatility [FV95]. Volumes [Ano93b, TR93]. vs [Fre86a].

W [Mül18, ZH92]. Wage [AAW11, CF15, CLLM03, Fli97, Gar89, Hub15, IM97, KT95, Lav20, NW95, Pow22, WB04, PW10]. Wage-Tenure [Gar89]. Wald [BE96]. Wales [Smi00]. Walk [BC89b, Lit83]. Walks [Ser95]. Wanshan [MLZ22]. WARP [GK96]. Warrant [Pat86]. Warranty [Fre88]. Was [Deh03]. Water [Fre86b, Olm09]. Watson [Wes18a, Ali87, Mül18]. Waves [LN20]. Way [BF89, DTU06, GI11, Wol00]. Weak [BC89a, DP22, FLM16, FW03, IP12, KM09b, OP13, PV22, RS22, SWY02, SW88, UY23a, UY23b]. Weakly [RS15]. Wealth [Cag03, CO19, KSM97, Li05, PRW06, Ren07]. Wealth-Consumption [Li05]. Weekly [BS87b, HKR97, PGC84, VP93]. Weibull [MABF08]. Weight [HL16, LS20a]. Weighted [AG07, BGKS04, DHL14, GR11, QZL15, RLYZ22, YCJK20]. Weighting [BCHL20]. Weights [CG00, HT06, MS03, Nev03, Rub86]. Welfare [BD09, BM23b, Bel04, BS95, Gua23, Hon23, Mor23, Yit96, vHPP04]. Well [MN88, OS20]. Well-Being [MN88, OS20]. We’re [Ost84]. Wheat [BB87]. Where [Ost84]. Which [CZ19]. White [AK84, LZZ20, Ren07, WKX23, MV91]. White-Noise [WKX23]. Whites [Gas85, Vin85a]. Whittle [Nie07]. Who [LLPZ20, SYA08, Ste87]. Wholesale [WGBM12]. WIA [ACL19]. Wild [DM10, MNW19]. William [Sim85]. Windfall [KJAG85]. Window [Poi91a, Poi91b, RI12]. Windows [PP11]. winner [Ano96a]. Wise [LZW^{+}22]. Wishart [ABJT22, GS10, PG06, RTW12, YLN17]. Within [PW10, PTZ86]. Without [CK19, DLP19, FPAB02, HKYZ20, Ter85, Vog98, Xie22]. Women [LZ95]. Wooldridge [HW20a]. Words [Tho20]. Work [AR91, Fer97, MO88]. Worker [BHLM07, HMvdB12]. Worker- [HMvdB12]. World [BK15, GPR13, Sea93]. Worth [MS88a].

X [HK14, FMB^{+}98a, McK84, dBM90]. X-11 [McK84, dBM90]. X-12-ARIMA [FMB^{+}98a].

Year [Ott90, ZH92]. Years [Die15a, JS85, Lit86b, Tau93b]. Yield [CGM16, RW09]. Yielding [HW10]. Yields [HRV20, Wol00]. York [AGO^{+}14a]. Young [LZ95]. Youth [Kra07].

Zealand [Cle85]. Zellner [Ano96a, Ano00c, Ano01, Ano02d, Ano03a, Ano03b, Ano04b, Ano04c, Ano05c, Ano05d, Ano06b, Ano06c, Ano07c, Ano07d, Ano08c, Ano08d, Ano09c, Ano10c, Ano11c, Ano13c, Ano13d, Ano15c, SM83, Woo98]. Zero [FS23, KSD98, LTMZ19, LSS19]. Zero-Process [FS23]. Zeros [GT96, Wer99]. Zheng [MLZ22]. Zhu [HW20a]. Zone [CT01]. Zones [PRM99].

References

Amir-Ahmadi:2020:CPH

[AAMW20] Pooyan Amir-Ahmadi, Christian Matthes, and Mu-Chun Wang. Choosing prior hyperparameters: With applications to timevarying parameter models. Journal of Business & Economic Statistics, 38(1):124–136, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1459302.

Ahn:2011:DIM

[AAW11] Tom Ahn, Peter Arcidiacono, and Walter Wessels. The distributional impacts of minimum wage increases when both labor supply and labor demand are endogenous. Journal of Business & Economic Statistics, 29(1):12–23, 2011. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/jbes.2010.07076.

Akgiray:1988:SLM

[AB88] Vedat Akgiray and G. Geoffrey Booth. The siable-law model of stock returns. Journal of Business & Economic Statistics, 6(1):51–57, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509636.

Ang:2002:RSI

[AB02] Andrew Ang and Geert Bekaert. Regime switches in interest rates. Journal of Business & Economic Statistics, 20 (2):163–182, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102317351930.

Ando:2022:LSG

[AB22] Tomohiro Ando and Jushan Bai. Large-scale generalized linear models for longitudinal data with grouped patterns of unobserved heterogeneity. Journal of Business & Economic Statistics, 41(3): 983–994, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2097913.

Alonso-Borrego:1999:SNI

[ABA99] César Alonso-Borrego and Manuel Arellano. Symmetrically normalized instrumental-variable estimation using panel data. Journal of Business & Economic Statistics, 17(1):36–49, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1999.10524795.

Atilgan:2015:IVS

[ABD15] Yigit Atilgan, Turan G. Bali, and K. Ozgur Demirtas. Implied volatility spreads and expected market returns. Journal of Business & Economic Statistics, 33(1):87–101, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.923776.

Augustyniak:2019:NAV

[ABD19] Maciej Augustyniak, Luc Bauwens, and Arnaud Dufays. A new approach to volatility modeling: The factorial hidden Markov volatility model. Journal of Business & Economic Statistics, 37(4):696–709, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1415910.

Abe:1999:GAM

[Abe99] Makoto Abe. A generalized additive model for discretechoice data. Journal of Business & Economic Statistics, 17 (3):271–284, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524817.

Andersen:2006:C

[ABFN06] Torben G. Andersen, Tim Bollerslev, Per Houmann Frederiksen, and Morten Ørregaard Nielsen. Comment. Journal of Business & Economic Statistics, 24(2):173–179, 2006. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500106000000134. See [HL06a, HL06b].

Alfelt:2022:SCA

[ABJT22] Gustav Alfelt, Taras Bodnar, Farrukh Javed, and Joanna Tyrcha. Singular conditional autoregressive Wishart model for realized covariance matrices. Journal of Business & Economic Statistics, 41(3):833–845, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2075370.

Attanasio:1999:HBL

[ABMW99] Orazio P. Attanasio, James Banks, Costas Meghir, and Guglielmo Weber. Humps and bumps in lifetime consumption. Journal of Business & Economic Statistics, 17(1):22–35, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1999.10524794.

Abraham:1985:STS

[Abr85] Bovas Abraham. Seasonal time series and transfer function modeling. Journal of Business & Economic Statistics, 3 (4):356–361, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509469.

Abrevaya:2003:PDR

[Abr03] Jason Abrevaya. Pairwise-difference rank estimation of the transformation model. Journal of Business & Economic Statistics, 21(3):437–447, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619070.

Armstrong:1993:TPS

[ABS93] John Armstrong, Clayton Block, and K. P. Srinath. Two-phase sampling of tax records for business surveys. Journal of Business & Economic Statistics, 11(4):407–416, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509974.

Adams:1999:SAS

[ABS99] Robert M. Adams, Allen N. Berger, and Robin C. Sickles. Semiparametric approaches to stochastic panel frontiers with applications in the banking industry. Journal of Business & Economic Statistics, 17(3):349–358, 1999. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1999.10524823.

Albert:1993:BIG

[AC93] James H. Albert and Siddhartha Chib. Bayes inference via Gibbs sampling of autoregressive time series subject to Markov mean and variance shifts. Journal of Business & Economic Statistics, 11(1):1–15, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509929.

Andrews:1994:AMU

[AC94] Donald W. K. Andrews and Hong-Yuan Chen. Approximately median-unbiased estimation of autoregressive models. Journal of Business & Economic Statistics, 12(2): 187–204, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10510007.

Almus:2003:EPR

[AC03] Matthias Almus and Dirk Czarnitzki. The effects of public R&D subsidies on firms’ innovation activities. Journal of Business & Economic Statistics, 21(2):226–236, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500103288618918.

Awartani:2009:AMM

[ACD09] Basel Awartani, Valentina Corradi, and Walter Distaso. Assessing market microstructure effects via realized volatility measures with an application to the Dow Jones Industrial Average stocks. Journal of Business & Economic Statistics, 27(2):251–265, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2009.0018.

Atkinson:2003:MDP

[ACH03] Scott E. Atkinson, Christopher Cornwell, and Olaf Honerkamp. Measuring and decomposing productivity change. Journal of Business & Economic Statistics, 21(2):284–294, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500103288618963.

Aryal:2022:PBA

[ACJK22] Gaurab Aryal, Hanna Charankevich, Seungwon Jeong, and Dong-Hyuk Kim. Procurements with bidder asymmetry in cost and riskaversion. Journal of Business & Economic Statistics, 41(4):1143– 1156, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2115497.

Andrews:2011:CSR

[ACL11] Rick L. Andrews, Imran S. Currim, and Peter S. H. Leeflang. A comparison of sales response predictions from demand models applied to store-level versus panel data. Journal of Business & Economic Statistics, 29(2):319–326, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2010.07225.

Ao:2019:MTD

[ACL19] Wallice Ao, Sebastian Calonico, and Ying-Ying Lee. Multivalued treatments and decomposition analysis: an application to the WIA program. Journal of Business & Economic Statistics, 39(1):358–371, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1660664.

Alvarez:2019:IFS

[ACR19] Rocio Alvarez, Maximo Camacho, and Manuel Ruiz. Inference on filtered and smoothed probabilities in Markov-switching autoregressive models. Journal of Business & Economic Statistics, 37(3):484–495, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1380032.

Agnello:2015:BBN

[ACS15] Luca Agnello, Vitor Castro, and Ricardo M. Sousa. Booms, busts, and normal times in the housing market. Journal of Business & Economic Statistics, 33(1):25–45, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.918545.

Aastveit:2022:QTV

[ACvD22] Knut Are Aastveit, Jamie L. Cross, and Herman K. van Dijk. Quantifying time-varying forecast uncertainty and risk for the real price of oil. Journal of Business & Economic Statistics, 41(2): 523–537, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2039159.

Abhyankar:1997:UNS

[ACW97] A. Abhyankar, L. S. Copeland, and W. Wong. Uncovering nonlinear structure in real-time stock-market indexes: The S&P 500, the DAX, the Nikkei 225, and the FTSE-100. Journal of Business & Economic Statistics, 15(1):1–14, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524681.

Abrevaya:2008:EBI

[AD08] Jason Abrevaya and Christian M. Dahl. The effects of birth inputs on birthweight. Journal of Business & Economic Statistics, 26(4):379–397, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500107000000269.

Aruoba:2009:RTM

[ADS09] S. Boragan Aruoba, Francis X. Diebold, and Chiara Scotti. Realtime measurement of business conditions. Journal of Business & Economic Statistics, 27(4):417–427, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.07205.

Ahtola:1986:BEP

[AES86] Juha Ahtola, Anders Ekholm, and Antti Somervuori. Bayes estimates for the price and income elasticities of alcoholic beverages in Finland from 1955 to 1980. Journal of Business & Economic Statistics, 4(2):199–208, 1986. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1986.10509515.

Almeida:2022:NOP

[AFAA22] Caio Almeida, Gustavo Freire, Rafael Azevedo, and Kym Ardison. Nonparametric option pricing with generalized entropic estimators. Journal of Business & Economic Statistics, 41(4):1173– 1187, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2115499.

Almeida:2022:CMC

[AFFT22] Caio Almeida, Jianqing Fan, Gustavo Freire, and Francesca Tang. Can a machine correct option pricing models? Journal of Business & Economic Statistics, 41(3):995–1009, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2099871.

Ascorbebeitia:2021:EDE

[AFO21] Jone Ascorbebeitia, Eva Ferreira, and Susan Orbe. The effect of dependence on European market risk. A nonparametric time varying approach. Journal of Business & Economic Statistics, 40 (2):913–923, 2021. CODEN RFSCFJ. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2021.1883439.

Andersen:2020:PTR

[AFT20] Torben G. Andersen, Nicola Fusari, and Viktor Todorov. The pricing of tail risk and the equity premium: Evidence from international option markets. Journal of Business & Economic Statistics, 38(3):662–678, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1564318.

Andreou:2002:RSV

[AG02] Elena Andreou and Eric Ghysels. Rolling-sample volatility estimators. Journal of Business & Economic Statistics, 20 (3):363–376, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618504.

Anatolyev:2005:TAT

[AG05] Stanislav Anatolyev and Alexander Gerko. A trading approach to testing for predictability. Journal of Business & Economic Statistics, 23(4):455–461, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000640.

Amisano:2007:CDF

[AG07] Gianni Amisano and Raffaella Giacomini. Comparing density forecasts via weighted likelihood ratio tests. Journal of Business & Economic Statistics, 25(2):177–190, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000332.

Anatolyev:2010:MFR

[AG10] Stanislav Anatolyev and Nikolay Gospodinov. Modeling financial return dynamics via decomposition. Journal of Business & Economic Statistics, 28(2):232–245, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2010.07017.

Andreou:2014:C

[AG14] Elena Andreou and Eric Ghysels. Comment. Journal of Business & Economic Statistics, 32(2):168–171, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.902238. See [GPI13b,

GPI13a].

Aastveit:2014:NGR

[AGJT14] Knut Are Aastveit, Karsten R. Gerdrup, Anne Sofie Jore, and Leif Anders Thorsrud. Nowcasting GDP in real time: a density combination approach. Journal of Business & Economic Statistics, 32(1):48–68, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.844155.

Andreou:2013:SMF

[AGK13] Elena Andreou, Eric Ghysels, and Andros Kourtellos. Should macroeconomic forecasters use daily financial data and how? Journal of Business & Economic Statistics, 31(2):240–251, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2013.
767199.

Alessi:2014:CBM

[AGO^{+}14a] Lucia Alessi, Eric Ghysels, Luca Onorante, Richard Peach, and Simon Potter. Central bank macroeconomic forecasting during the global financial crisis: The European Central Bank and Federal Reserve Bank of New York experiences. Journal of Business & Economic Statistics, 32(4):483–500, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.959124. See [Ken14,

Sco14, HM14, Ros14] and rejoinder [Mül14b].

Alessi:2014:R

[AGO^{+}14b] Lucia Alessi, Eric Ghysels, Luca Onorante, Richard Peach, and Simon Potter. Rejoinder. Journal of Business & Economic Statistics, 32(4):514–515, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.958920. See [AGO^{+}14a].

Andrews:2020:R

[AGS20a] Isaiah Andrews, Matthew Gentzkow, and Jesse M. Shapiro. Rejoinder. Journal of Business & Economic Statistics, 38 (4):731, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1791886.

Andrews:2020:TSR

[AGS20b] Isaiah Andrews, Matthew Gentzkow, and Jesse M. Shapiro. Transparency in structural research. Journal of Business & Economic Statistics, 38(4):711–722, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1796395.

Adesi:2004:TAP

[AGU04] Giovanni Barone Adesi, Patrick Gagliardini, and Giovanni Urga. Testing asset pricing models with coskewness. Journal of Business & Economic Statistics, 22(4):474–485, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000244.

Aguirregabiria:2008:C

[Agu08] Victor Aguirregabiria. Comment. Journal of Business & Economic Statistics, 26(3):283–289, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500108000000114. See [ALT08a,

ALT08b].

Aguirregabiria:2010:ALI

[Agu10] Victor Aguirregabiria. Another look at the identification of dynamic discrete decision processes: an application to retirement behavior. Journal of Business & Economic Statistics, 28(2):201–218, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1198/jbes.2009.07072.

Aryal:2019:SEF

[AGV19] Gaurab Aryal, Maria F. Gabrielli, and Quang Vuong. Semiparametric estimation of first-price auction models. Journal of Business & Economic Statistics, 39(2):373–385, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1665530.

Ahn:2020:HUD

[AH20] Hie Joo Ahn and James D. Hamilton. Heterogeneity and unemployment dynamics. Journal of Business & Economic Statistics, 38(3):554–569, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1530116.

Anily:1999:IDH

[AHI99] Shoshana Anily, Jacob Hornik, and Miron Israeli. Inferring the distribution of households’ duration of residence from data on current residence time. Journal of Business & Economic Statistics, 17(3):373–381, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524826.

Abrevaya:2015:ECA

[AHL15] Jason Abrevaya, Yu-Chin Hsu, and Robert P. Lieli. Estimating conditional average treatment effects. Journal of Business & Economic Statistics, 33(4):485–505, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.975555.

Arvanitis:2019:SS

[AHPT19] Stelios Arvanitis, Mark Hallam, Thierry Post, and Nikolas Topaloglou. Stochastic spanning. Journal of Business & Economic Statistics, 37(4):573–585, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1391099.

Abadie:2011:BCM

[AI11] Alberto Abadie and Guido W. Imbens. Bias-corrected matching estimators for average treatment effects. Journal of Business & Economic Statistics, 29(1):1–11, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.07333.

Aielli:2013:DCC

[Aie13] Gian Piero Aielli. Dynamic conditional correlation: On properties and estimation. Journal of Business & Economic Statistics, 31(3):282–299, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.771027.

Ahn:2018:RSE

[AIPR18a] Hyungtaik Ahn, Hidehiko Ichimura, James L. Powell, and Paul A. Ruud. Rejoinder for “Simple Estimators for Invertible Index Models”. Journal of Business & Economic Statistics, 36(1):22–23, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1392313. See [AIPR18b].

Ahn:2018:SEI

[AIPR18b] Hyungtaik Ahn, Hidehiko Ichimura, James L. Powell, and Paul A. Ruud. Simple estimators for invertible index models. Journal of Business & Economic Statistics, 36(1):1–10, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.

1379405. See discussion [KT18, Por18, AL18] and rejoinder [AIPR18a].

Aizcorbe:1990:TVA

[Aiz90] Ana M. Aizcorbe. Testing the validity of aggregates. Journal of Business & Economic Statistics, 8(4):373–383, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1990.10509808.

Aizcorbe:1991:LBP

[Aiz91] Ana M. Aizcorbe. A lower bound for the power of nonparametric tests. Journal of Business & Economic Statistics, 9 (4):463–467, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509873.

Abrevaya:2005:NAM

[AJ05] Jason Abrevaya and Wei Jiang. A nonparametric approach to measuring and testing curvature. Journal of Business & Economic Statistics, 23(1):1–19, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000316.

Angrist:2018:SEM

[AJK18] Joshua D. Angrist, Òscar Jordà, and Guido M. Kuersteiner. Semiparametric estimates of monetary policy effects: String theory revisited. Journal of Business & Economic Statistics, 36(3):371–387, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1204919.

Allenby:1996:ETB

[AJL96] Greg M. Allenby, Lichung Jen, and Robert P. Leone. Economic trends and being trendy: The influence of consumer confidence on retail fashion sales. Journal of Business & Economic Statistics, 14(1):103–111, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524633.

Abowd:1984:DMW

[AK84] John M. Abowd and Mark R. Killingsworth. Do minority/white unemployment differences really exist? Journal of Business & Economic Statistics, 2(1):64–72, 1984. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1984.10509370.

Angrist:1995:SSI

[AK95] Joshua D. Angrist and Alan B. Krueger. Split-sample instrumental variables estimates of the return to schooling. Journal of Business & Economic Statistics, 13(2):225–235, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524597.

Andrews:2006:TCB

[AK06] Donald W. K. Andrews and Jae-Young Kim. Tests for cointegration breakdown over a short time period. Journal of Business & Economic Statistics, 24(4):379–394, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000297.

Aryal:2013:PDP

[AK13] Gaurab Aryal and Dong-Hyuk Kim. A point decision for partially identified auction models. Journal of Business & Economic Statistics, 31(4):384–397, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.794731.

Akaike:1984:C

[Aka84] Hirotugu Akaike. Comment. Journal of Business & Economic Statistics, 2(4):321–322, 1984. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1984.10509399.

Antoine:2021:IRI

[AKKL21] Bertille Antoine, Lynda Khalaf, Maral Kichian, and Zhenjiang Lin. Identification-robust inference with simulation-based pseudomatching. Journal of Business & Economic Statistics, 41(2):321– 338, 2021. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.2019046.

Ameriks:2020:HER

[AKLS20] John Ameriks, Gábor Kézdi, Minjoon Lee, and Matthew D. Shapiro. Heterogeneity in expectations, risk tolerance, and household stock shares: The attenuation puzzle. Journal of Business & Economic Statistics, 38(3):633–646, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2018.1549560.

Arcidiacono:2012:HPT

[AKO12] Peter Arcidiacono, Ahmed Khwaja, and Lijing Ouyang. Habit persistence and teen sex: Could increased access to contraception have unintended consequences for teen pregnancies? Journal of Business & Economic Statistics, 30(2):312–325, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2011.652052.

Aigner:1984:MPL

[AL84] Dennis J. Aigner and Lee A. Lillard. Measuring peak load pricing response from experimental data: an exploratory analysis. Journal of Business & Economic Statistics, 2 (1):21–39, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509367.

Akgiray:1989:ESL

[AL89] Vedat Akgiray and Christopher G. Lamoureux. Estimation of stable-law parameters: a comparative study. Journal of Business & Economic Statistics, 7(1):85–93, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509716.

Allenby:1995:RBL

[AL95] Greg M. Allenby and Peter J. Lenk. Reassessing brand loyalty, price sensitivity, and merchandising effects on consumer brand choice. Journal of Business & Economic Statistics, 13 (3):281–289, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524602.

Aradillas-Lopez:2018:CSE

[AL18] Andres Aradillas-Lopez. A comment on “Simple Estimators for Invertible Index Models”. Journal of Business & Economic Statistics, 36(1):18–21, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1379406. See [AIPR18b, AIPR18a].

Ali:1984:DSA

[Ali84] Mukhtar M. Ali. Distributions of the sample autocorrelations when observations are from a stationary autoregressive-moving-average process. Journal of Business & Economic Statistics, 2(3):271–278, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509394.

Ali:1987:DWG

[Ali87] Mukhtar M. Ali. Durbin–Watson and generalized Durbin–Watson tests for autocorrelations and randomness. Journal of Business & Economic Statistics, 5(2):195–203, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509578.

Allenby:1990:CVB

[All90] Greg M. Allenby. Cross-validation, the Bayes theorem, and small-sample bias. Journal of Business & Economic Statistics, 8(2):171–178, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509788.

Allenby:1996:C

[All96] Greg M. Allenby. Comment. Journal of Business & Economic Statistics, 14(1):126–128, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524635.

Andersen:2007:ER

[ALN07] Torben G. Andersen, Arthur Lewbel, and Serena Ng. Editors’ report 2006. Journal of Business & Economic Statistics, 25(4):503, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500107000000377.

Aradillas-Lopez:2008:IPE

[ALT08a] Andres Aradillas-Lopez and Elie Tamer. The identification power of equilibrium in simple games. Journal of Business & Economic Statistics, 26(3):261–283, 2008. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/073500108000000105. See comments [Agu08, Baj08,

Hon08, Kha08, Mag08, MR08, CW08] and rejoinder [ALT08b].

Aradillas-Lopez:2008:R

[ALT08b] Andres Aradillas-Lopez and Elie Tamer. Rejoinder. Journal of Business & Economic Statistics, 26(3):307–310, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500108000000169. See

[ALT08a].

Alvarez:2004:DSQ

[Alv04] Javier Alvarez. Dynamics and seasonality in quarterly panel data. Journal of Business & Economic Statistics, 22(4): 443–456, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000424.

Anzarut:2018:PDS

[AMNP18] Michelle Anzarut, Ramsés H. Mena, Consuelo R. Nava, and Igor Prünster. Poisson-driven stationary Markov models. Journal of Business & Economic Statistics, 36(4):684–694, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1251441.

Anyfantaki:2022:EUI

[AMRT22] Sofia Anyfantaki, Esfandiar Maasoumi, Jue Ren, and Nikolas Topaloglou. Evidence of uniform inefficiency in market portfolios based on dominance tests. Journal of Business & Economic Statistics, 40(3):937–949, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1888741.

Abowd:2019:MEM

[AMS19] John M. Abowd, Kevin L. McKinney, and Ian M. Schmutte. Modeling endogenous mobility in earnings determination. Journal of Business & Economic Statistics, 37(3):405–418, 2019.

ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1356727.

Angrist:1991:ITE

[AN91] Joshua D. Angrist and Whitney K. Newey. Over-identification tests in earnings functions with fixed effects. Journal of Business & Economic Statistics, 9(3):317–323, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509857.

Anatolyev:2009:NRM

[Ana09] Stanislav Anatolyev. Nonparametric retrospection and monitoring of predictability of financial returns. Journal of Business & Economic Statistics, 27(2):149–160, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.0010.

Andersen:1994:C

[And94a] Torbers G. Andersen. Comment. Journal of Business & Economic Statistics, 12(4):389–392, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10524554.

Andrews:1994:FPS

[And94b] Rick L. Andrews. Forecasting performance of structural time series models. Journal of Business & Economic Statistics, 12 (1):129–133, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10509996.

Andersen:2000:SRA

[And00] Torben G. Andersen. Some reflections on analysis of highfrequency data. Journal of Business & Economic Statistics, 18(2):146–153, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524857.

Andersson:2001:NIG

[And01] Jonas Andersson. On the normal inverse Gaussian stochastic volatility model. Journal of Business & Economic Statistics, 19(1):44–54, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
07350010152472607.

Andrews:2002:GMM

[And02] Donald W. K. Andrews. Generalized method of moments estimation when a parameter is on a boundary. Journal of Business & Economic Statistics, 20(4):530–544, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102288618667.

Andersen:2005:ER

[And05] Torben G. Andersen. Editor’s report 2004. Journal of Business & Economic Statistics, 23(4):495, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500105000000117.

Andersen:2006:ER

[And06] Torben G. Andersen. Editor report 2005. Journal of Business & Economic Statistics, 24(4):505, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000512.

Andersen:2007:EA

[And07] Torben G. Andersen. Editorial announcement. Journal of Business & Economic Statistics, 25(1):1, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000611.

Andrews:2014:C

[And14] Beth Andrews. Comment. Journal of Business & Economic Statistics, 32(2):191–193, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.875921. See [FQX14a, FQX14b].

Andreoli:2018:RII

[And18] Francesco Andreoli. Robust inference for inverse stochastic dominance. Journal of Business & Economic Statistics, 36(1):146–159, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1137758.

Angrist:1995:IJS

[Ang95] Joshua D. Angrist. Introduction to the jbes symposium on program and policy evaluation. Journal of Business & Economic Statistics, 13(2):133–136, 1995. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1995.10524587.

Angrist:2001:ELD

[Ang01] Joshua D. Angrist. Estimation of limited dependent variable models with dummy endogenous regressors. Journal of Business & Economic Statistics, 19(1):2–28, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/07350010152472571.

Anonymous:1983:EC

[Ano83a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 1(4):348–349, 1983. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1983.10509363.

Anonymous:1983:IAA

[Ano83b] Anonymous. Index to articles, by author, volume 1 (1983). Journal of Business & Economic Statistics, 1(4):350, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1983.10509364.

Anonymous:1983:PMA

[Ano83c] Anonymous. Preparation of manuscripts for ASA journals. Journal of Business & Economic Statistics, 1(1):86–89, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1983.10509327.

Anonymous:1984:A

[Ano84a] Anonymous. Announcements. Journal of Business & Economic Statistics, 2(3):290, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509397.

Anonymous:1984:C

[Ano84b] Anonymous. Corrigendum. Journal of Business & Economic Statistics, 2(4):413, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/

07350015.1984.10509418. The article [DMMN83] is in the public domain, not copyrighted.

Anonymous:1984:DSA

[Ano84c] Anonymous. Discussion of the statistical analysis of fairness and employment discrimination. Journal of Business & Economic Statistics, 2(2):110, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509375.

Anonymous:1984:EC

[Ano84d] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 2(4):414–415, 1984. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1984.10509419.

Anonymous:1984:IAA

[Ano84e] Anonymous. Index to articles, by author, volume 2 (1984). Journal of Business & Economic Statistics, 2(4):416–417, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1984.10509420.

Anonymous:1985:C

[Ano85a] Anonymous. Corrections. Journal of Business & Economic Statistics, 3(1):100, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509435.

Anonymous:1985:EC

[Ano85b] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 3(4):415, 1985. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1985.10509483.

Anonymous:1985:IV

[Ano85c] Anonymous. Index to volume 3 (1985). Journal of Business & Economic Statistics, 3(4):416–417, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1985.10509484.

Anonymous:1985:PRa

[Ano85d] Anonymous. Publications received. Journal of Business & Economic Statistics, 3(1):99, 1985. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1985.10509434.

Anonymous:1985:PRb

[Ano85e] Anonymous. Publications received. Journal of Business & Economic Statistics, 3(2):178, 1985. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1985.10509449.

Anonymous:1985:PRc

[Ano85f] Anonymous. Publications received. Journal of Business & Economic Statistics, 3(3):291, 1985. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1985.10509461.

Anonymous:1985:PRd

[Ano85g] Anonymous. Publications received. Journal of Business & Economic Statistics, 3(4):414, 1985. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1985.10509482.

Anonymous:1986:Ab

[Ano86a] Anonymous. Announcement. Journal of Business & Economic Statistics, 4(3):395, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509536.

Anonymous:1986:Ac

[Ano86b] Anonymous. Announcement. Journal of Business & Economic Statistics, 4(4):495, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509550.

Anonymous:1986:Aa

[Ano86c] Anonymous. Announcements. Journal of Business & Economic Statistics, 4(1):145, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509509.

Anonymous:1986:EC

[Ano86d] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 4(4):496–497, 1986. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1986.10509551.

Anonymous:1986:IV

[Ano86e] Anonymous. Index to volume 4 (1986). Journal of Business & Economic Statistics, 4(4):498–500, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509552.

Anonymous:1986:LE

[Ano86f] Anonymous. Letter to the Editor. Journal of Business & Economic Statistics, 4(1):143, 1986. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1986.10509508.

Anonymous:1986:PRa

[Ano86g] Anonymous. Publiations received. Journal of Business & Economic Statistics, 4(1):141, 1986. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1986.10509507.

Anonymous:1986:PRb

[Ano86h] Anonymous. Publications received. Journal of Business & Economic Statistics, 4(2):278–279, 1986. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1986.10509524.

Anonymous:1986:PRc

[Ano86i] Anonymous. Publications received. Journal of Business & Economic Statistics, 4(3):394, 1986. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1986.10509535.

Anonymous:1986:PRd

[Ano86j] Anonymous. Publications received. Journal of Business & Economic Statistics, 4(4):493, 1986. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1986.10509549.

Anonymous:1987:Aa

[Ano87a] Anonymous. Announcement. Journal of Business & Economic Statistics, 5(1):163, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509574.

Anonymous:1987:Ab

[Ano87b] Anonymous. Announcement. Journal of Business & Economic Statistics, 5(4):551, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509625.

Anonymous:1987:C

[Ano87c] Anonymous. Correction. Journal of Business & Economic Statistics, 5(2):313, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509593.

Anonymous:1987:EC

[Ano87d] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 5(4):552–553, 1987. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1987.10509626.

Anonymous:1987:IV

[Ano87e] Anonymous. Index to volume 5 (1987). Journal of Business & Economic Statistics, 5(4):554–556, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509627.

Anonymous:1987:PRa

[Ano87f] Anonymous. Publications received. Journal of Business & Economic Statistics, 5(1):162, 1987. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1987.10509573.

Anonymous:1987:PRb

[Ano87g] Anonymous. Publications received. Journal of Business & Economic Statistics, 5(2):311, 1987. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1987.10509592.

Anonymous:1987:PRc

[Ano87h] Anonymous. Publications received. Journal of Business & Economic Statistics, 5(4):550–551, 1987. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1987.10509624.

Anonymous:1988:EC

[Ano88a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 6(4):521–522, 1988. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1988.10509702.

Anonymous:1988:IV

[Ano88b] Anonymous. Index to volume 6 (1 988). Journal of Business & Economic Statistics, 6(4):523–525, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509703.

Anonymous:1989:EC

[Ano89a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 7(4):504–505, 1989. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1989.10509764.

Anonymous:1989:IV

[Ano89b] Anonymous. Index to volume 7 (1989). Journal of Business & Economic Statistics, 7(4):504–505, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509765.

Anonymous:1990:EC

[Ano90a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 8(4):481–482, 1990. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1990.10509820.

Anonymous:1990:IV

[Ano90b] Anonymous. Index to volume 8 (1 990). Journal of Business & Economic Statistics, 8(4):481–482, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1990.10509821.

Anonymous:1991:EC

[Ano91a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 9(4):475–476, 1991. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1991.10509875.

Anonymous:1991:IV

[Ano91b] Anonymous. Index to volume 9 (1991). Journal of Business & Economic Statistics, 9(4):475–476, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509876.

Anonymous:1992:EC

[Ano92a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 10(4):475–476, 1992. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1992.10509927.

Anonymous:1992:IV

[Ano92b] Anonymous. Index to volume 10 (1992). Journal of Business & Economic Statistics, 10(4):477–478, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1992.10509928.

Anonymous:1993:AEA

[Ano93a] Anonymous. Acknowledgment of the editorial assistants, book review Editors, and associate Editors. Journal of Business & Economic Statistics, 11(4):432–433, 1993. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1993.10509981.

Anonymous:1993:AIV

[Ano93b] Anonymous. Author index, volumes 1–10 (1983–1992). Journal of Business & Economic Statistics, 11(4):437–481, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1993.10509983.

Anonymous:1993:EC

[Ano93c] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 11(4):483, 1993. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1993.10509984.

Anonymous:1993:IV

[Ano93d] Anonymous. Index to volume 11. Journal of Business & Economic Statistics, 11(4):485–486, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509985.

Anonymous:1994:C

[Ano94a] Anonymous. Correction. Journal of Business & Economic Statistics, 12(1):139, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10509999.

Anonymous:1994:EC

[Ano94b] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 12(4):493–494, 1994. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1994.10524572.

Anonymous:1994:IV

[Ano94c] Anonymous. Index to volume 12 (1994). Journal of Business & Economic Statistics, 12(4):495–496, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10524573.

Anonymous:1995:EC

[Ano95a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 13(4):489–490, 1995. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1995.10524622.

Anonymous:1995:IV

[Ano95b] Anonymous. Index to volume 13 (1995). Journal of Business & Economic Statistics, 13(4):491–492, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524623.

Anonymous:1995:NRP

[Ano95c] Anonymous. New refereeing process at the JBES. Journal of Business & Economic Statistics, 13(4):493, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524624.

Anonymous:1996:ZTA

[Ano96a] Anonymous. 1995 Zellner Thesis Award winner. Journal of Business & Economic Statistics, 14(2):259, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524654.

Anonymous:1996:Ca

[Ano96b] Anonymous. Correction. Journal of Business & Economic Statistics, 14(1):133, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524638.

Anonymous:1996:Cb

[Ano96c] Anonymous. Correction. Journal of Business & Economic Statistics, 14(2):257, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524653.

Anonymous:1996:EC

[Ano96d] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 14(4):497–498, 1996. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1996.10524679.

Anonymous:1996:IV

[Ano96e] Anonymous. Index to volume 14 (1 996). Journal of Business & Economic Statistics, 14(4):499–500, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524680.

Anonymous:1997:A

[Ano97a] Anonymous. Announcement. Journal of Business & Economic Statistics, 15(1):109, 1997. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1997.10524693.

Anonymous:1997:EC

[Ano97b] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 15(4):493–494, 1997. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1997.10524727.

Anonymous:1997:IV

[Ano97c] Anonymous. Index to volume 15 (1997). Journal of Business & Economic Statistics, 15(4):495–496, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524728.

Anonymous:1998:EC

[Ano98a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 16(4):508, 1998. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1998.10524790.

Anonymous:1998:IV

[Ano98b] Anonymous. Index to volume 16 (1998). Journal of Business & Economic Statistics, 16(4):509–510, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524791.

Anonymous:1999:EC

[Ano99a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 17(4):505, 1999. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1999.10524840.

Anonymous:1999:IV

[Ano99b] Anonymous. Index to volume 17 (1999). Journal of Business & Economic Statistics, 17(4):506–507, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524841.

Anonymous:2000:EC

[Ano00a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 18(4):512–513, 2000. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.2000.10524890.

Anonymous:2000:IV

[Ano00b] Anonymous. Index to volume 18 (2000). Journal of Business & Economic Statistics, 18(4):514–515, 2000. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.2000.10524891.

Anonymous:2000:ZTA

[Ano00c] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 18(2):263, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524868.

Anonymous:2001:ZTA

[Ano01] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 19(2):254, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500101316970476.

Anonymous:2002:EC

[Ano02a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 20(4):546–547, 2002. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500102288618685.

Anonymous:2002:ICS

[Ano02b] Anonymous. Interview with Christopher A. Sims. Journal of Business & Economic Statistics, 20(4):448–449, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102288618586.

Anonymous:2002:ILP

[Ano02c] Anonymous. Interview with Lars Peter Hansen. Journal of Business & Economic Statistics, 20(4):442–447, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102288618577.

Anonymous:2002:ZTA

[Ano02d] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 20(1):145–146, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102753410453.

Anonymous:2003:ZTAa

[Ano03a] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 21(2):336–337, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288618990.

Anonymous:2003:ZTAb

[Ano03b] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 21(4):581–582, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619250.

Anonymous:2004:EC

[Ano04a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 22(4):489–490, 2004. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500104000000460.

Anonymous:2004:ZTAa

[Ano04b] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 22(1):126–127, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619340.

Anonymous:2004:ZTAb

[Ano04c] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 22(4):486–487, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000451.

Anonymous:2005:C

[Ano05a] Anonymous. Correction. Journal of Business & Economic Statistics, 23(3):363, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500105000000045.

Anonymous:2005:EC

[Ano05b] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 23(4):496–497, 2005. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500105000000126.

Anonymous:2005:ZTAa

[Ano05c] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 23(1):130–131, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000659.

Anonymous:2005:ZTAb

[Ano05d] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 23(4):493–494, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500105000000108.

Anonymous:2006:EC

[Ano06a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 24(4):506–507, 2006. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500106000000503.

Anonymous:2006:ZTAa

[Ano06b] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 24(1):125–126, 2006. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500105000000298.

Anonymous:2006:ZTAb

[Ano06c] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 24(4):503–504, 2006. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000521.

Anonymous:2007:EC

[Ano07a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 25(4):504–505, 2007. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500107000000386.

Anonymous:2007:IV

[Ano07b] Anonymous. Index to volume 25 (2007). Journal of Business & Economic Statistics, 25(4):506–507, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500107000000359.

Anonymous:2007:ZTAa

[Ano07c] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 25(1):121–122, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000666.

Anonymous:2007:ZTAb

[Ano07d] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 25(4):501–502, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500107000000368.

Anonymous:2008:EC

[Ano08a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 26(4):558–559, 2008. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500108000000196.

Anonymous:2008:IV

[Ano08b] Anonymous. Index to volume 26 (2008). Journal of Business & Economic Statistics, 26(4):560–561, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500108000000204.

Anonymous:2008:ZTAa

[Ano08c] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 26(1):128–129, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500107000000395.

Anonymous:2008:ZTAb

[Ano08d] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 26(4):555–556, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500108000000178.

Anonymous:2009:EC

[Ano09a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 27(4):567–568, 2009. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/jbes.2009.274ec.

Anonymous:2009:IV

[Ano09b] Anonymous. Index to volume 27 (2009). Journal of Business & Economic Statistics, 27(4):569–570, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.274in.

Anonymous:2009:ZTA

[Ano09c] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 27(4):564–565, 2009. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.274za.

Anonymous:2010:EC

[Ano10a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 28(4):575–576, 2010. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/jbes.2010.284ec.

Anonymous:2010:IV

[Ano10b] Anonymous. Index to volume 28. Journal of Business & Economic Statistics, 28(4):577–578, 2010. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2010.284in.

Anonymous:2010:ZTA

[Ano10c] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 28(4):572–573, 2010. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2010.284za.

Anonymous:2011:EC

[Ano11a] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 29(4):598–599, 2011. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/jbes.2011.294ec.

Anonymous:2011:IV

[Ano11b] Anonymous. Index to volume 29 (2011). Journal of Business & Economic Statistics, 29(4):600–602, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2011.294in.

Anonymous:2011:ZTA

[Ano11c] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 29(4):595–596, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2011.294za.

Anonymous:2012:EBE

[Ano12a] Anonymous. Editorial board EOV. Journal of Business & Economic Statistics, 30(4):??, 2012. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2012.739058.

Anonymous:2012:EC

[Ano12b] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 30(4):592–594, 2012. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2012.721674.

Anonymous:2013:EBE

[Ano13a] Anonymous. Editorial board EOV. Journal of Business & Economic Statistics, 31(4):??, 2013. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2013.835593.

Anonymous:2013:EC

[Ano13b] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 31(4):560–561, 2013. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2013.848082.

Anonymous:2013:ZTAa

[Ano13c] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 31(1):122–123, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.750219.

Anonymous:2013:ZTAb

[Ano13d] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 31(4):562–563, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.848080.

Anonymous:2014:EBE

[Ano14a] Anonymous. Editorial board EOV. Journal of Business & Economic Statistics, 32(4):??, 2014. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2014.972254.

Anonymous:2014:EC

[Ano14b] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 32(4):606–608, 2014. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2014.972253.

Anonymous:2015:EBE

[Ano15a] Anonymous. Editorial board EOV. Journal of Business & Economic Statistics, 33(4):610, 2015. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2015.1103118.

Anonymous:2015:EC

[Ano15b] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 33(4):607–609, 2015. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2015.1103119.

Anonymous:2015:ZTA

[Ano15c] Anonymous. The Zellner Thesis Award in Business and Economic Statistics. Journal of Business & Economic Statistics, 33(1):165–166, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1004896.

Anonymous:2016:EBE

[Ano16a] Anonymous. Editorial board EOV. Journal of Business & Economic Statistics, 34(4):??, 2016. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2016.1221617.

Anonymous:2016:EC

[Ano16b] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 34(4):689–692, 2016. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2016.1221618.

Anonymous:2017:EBE

[Ano17a] Anonymous. Editorial board EOV. Journal of Business & Economic Statistics, 35(4):??, 2017. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2017.1357927.

Anonymous:2017:EC

[Ano17b] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 35(4):642–645, 2017. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2017.1357925.

Anonymous:2019:EC

[Ano19] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 37(4):771–774, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2019.1670479.

Anonymous:2020:EC

[Ano20] Anonymous. Editorial collaborators. Journal of Business & Economic Statistics, 38(4):951–954, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1826254.

Anonymous:2021:Ca

[Ano21a] Anonymous. Correction. Journal of Business & Economic Statistics, 39(4):1080, 2021. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2021.1970572.

Anonymous:2021:Cb

[Ano21b] Anonymous. Correction. Journal of Business & Economic Statistics, 40(1):467, 2021. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2021.1971536.

Ansley:1983:C

[Ans83] Craig F. Ansley. Comment. Journal of Business & Economic Statistics, 1(4):307–309, 1983. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1983.10509356.

Ansley:1993:C

[Ans93] Craig F. Ansley. Comment. Journal of Business & Economic Statistics, 11(2):136–137, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509939.

Antle:1983:TSS

[Ant83] John M. Antle. Testing the stochastic structure of production: a flexible moment-based approach. Journal of Business & Economic Statistics, 1(3):192–201, 1983. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1983.10509339.

Arize:2000:ERV

[AOS00] Augustine C. Arize, Thomas Osang, and Daniel J. Slottje. Exchange-rate volatility and foreign trade: Evidence from thirteen LDC’s. Journal of Business & Economic Statistics, 18 (1):10–17, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524843.

Ashley:2006:EES

[AP06] Richard A. Ashley and Douglas M. Patterson. Evaluating the effectiveness of state-switching time series models for U.S. real output. Journal of Business & Economic Statistics, 24 (3):266–277, 2006. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500105000000216.

Apps:1985:CRD

[App85] Patricia Apps. Comment on “The Relative Deprivation Curve and Its Applications”. Journal of Business & Economic Statistics, 3(2):169–171, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509446. See [Kak84a].

Arduini:2019:TEH

[APR19] Tiziano Arduini, Eleonora Patacchini, and Edoardo Rainone. Treatment effects with heterogeneous externalities. Journal of Business & Economic Statistics, 38(4):826–838, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1592755.

Anderson:2020:SBU

[APW20] Gordon Anderson, Thierry Post, and Yoon-Jae Whang. Somewhere between utopia and dystopia: Choosing from multiple incomparable prospects. Journal of Business & Economic Statistics, 38(3):502–515, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1515765.

Alwan:1988:TSM

[AR88] Layth C. Alwan and Harry V. Roberts. Time-series modeling for statistical process control. Journal of Business & Economic Statistics, 6(1):87–95, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509640.

Allenby:1991:TNA

[AR91] Greg M. Allenby and Peter E. Rossi. There is no aggregation bias: Why macro logit models work. Journal of Business & Economic Statistics, 9(1):1–14, 1991. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1991.10509822.

Arnold:1987:GOF

[ARBS87] Barry C. Arnold, Christopher A. Robertson, Patrick L. Brockett, and Boo-Yau Shu. Generating ordered families of Lorenz curves by strongly unimodal distributions. Journal of Business & Economic Statistics, 5(2):305–308, 1987. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1987.10509590.

Arellano:2002:SIV

[Are02] Manuel Arellano. Sargan’s intrumental variables estimation and the generalized method of moments. Journal of Business & Economic Statistics, 20(4):450–459, 2002. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500102288618595.

Aruoba:2020:TSI

[Aru20] S. Boragan Aruoba. Term structures of inflation expectations and real interest rates. Journal of Business & Economic Statistics, 38(3):542–553, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1529599.

Aastveit:2018:CDN

[ARvD18] Knut Are Aastveit, Francesco Ravazzolo, and Herman K. van Dijk. Combined density nowcasting in an uncertain economic environment. Journal of Business & Economic Statistics, 36(1):131–145, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1137760.

Altonji:1996:SSB

[AS96a] Joseph G. Altonji and Lewis M. Segal. Small-sample bias in GMM estimation of covariance structures. Journal of Business & Economic Statistics, 14(3):353–366, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524661.

Andersen:1996:GES

[AS96b] Torben G. Andersen and Bent E. Sørensen. GMM estimation of a stochastic volatility model: a Monte Carlo study. Journal of Business & Economic Statistics, 14(3):328–352, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524660.

Amsler:2012:TSM

[AS12] Christine Amsler and Peter Schmidt. Tests of short memory with thick-tailed errors. Journal of Business & Economic Statistics, 30(3):381–390, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.669668.

Amengual:2020:NCR

[AS20] Dante Amengual and Enrique Sentana. Is a normal copula the right copula? Journal of Business & Economic Statistics, 38(2):350–366, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1505631.

Asano:2002:CRI

[Asa02] Hirokatsu Asano. Costly reversible investment with fixed costs. Journal of Business & Economic Statistics, 20(2): 227–240, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102317351976.

Ashley:1985:OUS

[Ash85] Richard Ashley. On the optimal use of suboptimal forecasts of explanatory variables. Journal of Business & Economic Statistics, 3(2):129–131, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509440.

Ait-Sahalia:2006:C

[ASMZ06] Yacine Aït-Sahalia, Per A. Mykland, and Lan Zhang. Comment. Journal of Business & Economic Statistics, 24(2): 162–167, 2006. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000152. See [HL06a, HL06b].

Audrino:2011:GMT

[AT11] Francesco Audrino and Fabio Trojani. A general multivariate threshold GARCH model with dynamic conditional correlations. Journal of Business & Economic Statistics, 29(1):138–149, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2010.08117.

Amado:2014:CCM

[AT14] Cristina Amado and Timo Teräsvirta. Conditional correlation models of autoregressive conditional heteroscedasticity with nonstationary GARCH equations. Journal of Business & Economic Statistics, 32(1):69–87, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.847376.

An:2017:ISM

[AT17] Yonghong An and Xun Tang. Identifying structural models of committee decisions with heterogeneous tastes and ideological bias. Journal of Business & Economic Statistics, 35(3):452–469, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1084309.

Atkinson:1983:DIH

[Atk83] Scott E. Atkinson. Discussion: The implications of homothetic separability for share equation price elasticities. Journal of Business & Economic Statistics, 1(3):211–214, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1983.10509341.

Audrino:2006:TSM

[Aud06] Francesco Audrino. Tree-structured multiple regimes in interest rates. Journal of Business & Economic Statistics, 24 (3):338–353, 2006. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000053.

Ashley:1986:MME

[AV86] Richard Ashley and David Vaughan. Measuring measurement error in economic time series. Journal of Business & Economic Statistics, 4(1):95–103, 1986. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1986.10509498.

Abowd:2005:R

[AV05a] John M. Abowd and Lars Vilhuber. Rejoinder. Journal of Business & Economic Statistics, 23(2):162–165, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500104000000712. See

[AV05b, Win05, vdK05, Hon05, CFR05].

Abowd:2005:SES

[AV05b] John M. Abowd and Lars Vilhuber. The sensitivity of economic statistics to coding errors in personal identifiers. Journal of Business & Economic Statistics, 23(2):133–152, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500104000000677. See

comments [Win05, vdK05, Hon05, CFR05] and rejoinder [AV05a].

Anderson:2007:FVA

[AV07] Heather M. Anderson and Farshid Vahid. Forecasting the volatility of Australian stock returns. Journal of Business & Economic Statistics, 25(1):76–90, 2007. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500106000000440.

Athanasopoulos:2008:VVV

[AV08] George Athanasopoulos and Farshid Vahid. VARMA versus VAR for macroeconomic forecasting. Journal of Business & Economic Statistics, 26(2):237–252, 2008. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000313.

Arni:2020:TVR

[AvdBL20] Patrick Arni, Gerard J. van den Berg, and Rafael Lalive. Treatment versus regime effects of carrots and sticks. Journal of Business & Economic Statistics, 40(1):111–127, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1784744.

Abbring:2001:BCC

[AVV01] Jaap H. Abbring, Gerard J. Van Den Berg, and Jan C. Van Ours. Business cycles and compositional variation in U.S. unemployment. Journal of Business & Economic Statistics, 19 (4):436–448, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
07350010152596682.

Ansley:1984:DSS

[AW84] Craig F. Ansley and William E. Wecker. On dips in the spectrum of a seasonally adjusted time series. Journal of Business & Economic Statistics, 2(4):323–324, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1984.10509400.

Aguilar:2000:BDF

[AW00] Omar Aguilar and Mike West. Bayesian dynamic factor models and portfolio allocation. Journal of Business & Economic Statistics, 18(3):338–357, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524875.

Amengual:2007:CEN

[AW07] Dante Amengual and Mark W. Watson. Consistent estimation of the number of dynamic factors in a large N and T panel. Journal of Business & Economic Statistics, 25(1):91–96, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000585.

An:2020:NNM

[AWX20] Yonghong An, Le Wang, and Ruli Xiao. A nonparametric nonclassical measurement error approach to estimating intergenerational mobility elasticities. Journal of Business & Economic Statistics, 40(1):169–185, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1787176.

Abowd:1985:EGL

[AZ85] John M. Abowd and Arnold Zellner. Estimating gross laborforce flows. Journal of Business & Economic Statistics, 3 (3):254–283, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509457.

Babii:2022:HDM

[Bab22] Andrii Babii. High-dimensional mixed-frequency IV regression. Journal of Business & Economic Statistics, 40(4):1470– 1483, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1933501.

Baillie:1998:C

[Bai98] Richard T. Baillie. Comment. Journal of Business & Economic Statistics, 16(3):273–276, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524764.

Bajari:2008:C

[Baj08] Patrick Bajari. Comment. Journal of Business & Economic Statistics, 26(3):289–292, 2008. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500108000000123. See [ALT08a, ALT08b].

Balke:1993:DLS

[Bal93] Nathan S. Balke. Detecting level shifts in time series. Journal of Business & Economic Statistics, 11(1):81–92, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1993.10509934.

Baltagi:2004:C

[Bal04] Badi H. Baltagi. Comment. Journal of Business & Economic Statistics, 22(2):163–164, 2004. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/073500104000000028. See [PSW04a] and rejoinder

[PSW04b].

Barnett:1983:DRR

[Bar83a] William A. Barnett. Discussion: The recent reappearance of the homotheticity restriction on preferences. Journal of Business & Economic Statistics, 1(3):215–218, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1983.10509342.

Barnett:1983:NIM

[Bar83b] William A. Barnett. New indices of money supply and the flexible Laurent demand system. Journal of Business & Economic Statistics, 1(1):7–23, 1983. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1983.10509317.

Bartik:1985:BLD

[Bar85] Timothy J. Bartik. Business location decisions in the United States: Estimates of the effects of unionization, taxes, and other characteristics of states. Journal of Business & Economic Statistics, 3(1):14–22, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509422.

Barone-Adesi:1983:MMH

[BAT83] Giovanni Barone-Adesi and Prem P. Talwar. Market models and heteroscedasticity of residual security returns. Journal of Business & Economic Statistics, 1(2):163–168, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1983.10509335.

Batchelor:1990:AFE

[Bat90] R. A. Batchelor. All forecasters are equal. Journal of Business & Economic Statistics, 8(1):143–144, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1990.10509784.

Baurle:2013:SDF

[Bäu13] Gregor Bäurle. Structural dynamic factor analysis using prior information from macroeconomic theory. Journal of Business & Economic Statistics, 31(2):136–150, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2012.747839.

Bauer:2018:RRP

[Bau18] Michael D. Bauer. Restrictions on risk prices in dynamic term structure models. Journal of Business & Economic Statistics, 36(2):196–211, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1164707.

Bessler:1987:FWE

[BB87] David A. Bessler and Ronald A. Babula. Forecasting wheat exports: Do exchange rates matter? Journal of Business & Economic Statistics, 5(3):397–406, 1987. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1987.10509604.

Baillie:1989:MDE

[BB89] Richard T. Baillie and Tim Bollerslev. The message in daily exchange rates: a conditional-variance tale. Journal of Business & Economic Statistics, 7(3):297–305, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509739.

Baillie:2002:MDE

[BB02] Richard T. Baillie and Tim Bollerslev. The message in daily exchange rates. Journal of Business & Economic Statistics, 20(1):60–68, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102753410390.

Benkard:2005:HPI

[BB05] C. Lanier Benkard and Patrick Bajari. Hedonic price indexes with unobserved product characteristics, and application to personal computers. Journal of Business & Economic Statistics, 23(1):61–75, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000262.

Bertsche:2020:ISV

[BB20] Dominik Bertsche and Robin Braun. Identification of structural vector autoregressions by stochastic volatility. Journal of Business & Economic Statistics, 40(1):328–341, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1813588.

Braun:2022:ISM

[BB22] Robin Braun and Ralf Brüggemann. Identification of SVAR models by combining sign restrictions with external instruments. Journal of Business & Economic Statistics, 41(4):1077–1089, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2104857.

Bec:2004:TUR

[BBC04] Frédéric Bec, Mélika Ben Salem, and Marine Carrasco. Tests for unit-root versus threshold specification with an application to the purchasing power parity relationship. Journal of Business & Economic Statistics, 22(4):382–395, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000389.

Buccheri:2020:SDC

[BBCL20] Giuseppe Buccheri, Giacomo Bormetti, Fulvio Corsi, and Fabrizio Lillo. A score-driven conditional correlation model for noisy and asynchronous data: an application to high-frequency covariance dynamics. Journal of Business & Economic Statistics, 39(4):920–936, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1739530.

Bartalotti:2020:CRD

[BBD20] Otávio Bartalotti, Quentin Brummet, and Steven Dieterle. A correction for regression discontinuity designs with group-specific mismeasurement of the running variable. Journal of Business & Economic Statistics, 39(3):833–848, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1737081.

Bellemare:2012:FAS

[BBK12] Charles Bellemare, Luc Bissonnette, and Sabine Kröger. Flexible approximation of subjective expectations using probability questions. Journal of Business & Economic Statistics, 30(1): 125–131, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic).

URL http://www.tandfonline.com/doi/full/10.1198/jbes.
2011.09053.

Barnett:1989:MCS

[BC89a] William A. Barnett and Seungmook Choi. A Monte Carlo study of tests of blockwise weak separability. Journal of Business & Economic Statistics, 7(3):363–377, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509745.

Berger:1989:WRW

[BC89b] Allen N. Berger and Roger Craine. Why random walk models of the term structure are hard to reject. Journal of Business & Economic Statistics, 7(2):161–167, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509724.

Badrinath:1991:DAL

[BC91] S. G. Badrinath and Sangit Chatterjee. A data-analytic look at skewness and elongation in common-stock-return distributions. Journal of Business & Economic Statistics, 9 (2):223–233, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509848.

Bonham:2001:APN

[BC01] Carl S. Bonham and Richard H. Cohen. To aggregate, pool, or neither. Journal of Business & Economic Statistics, 19 (3):278–291, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500101681019936.

Bun:2005:BCE

[BC05] Maurice J. G. Bun and Martin A. Carree. Bias-corrected estimation in dynamic panel data models. Journal of Business & Economic Statistics, 23(2):200–210, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000532.

Berrendero:2011:TSO

[BC11] José R. Berrendero and Javier Cárcamo. Tests for the second order stochastic dominance based on L-statistics. Journal of Business & Economic Statistics, 29(2):260–270, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2010.07224.

Bormetti:2019:SVM

[BCCL19] Giacomo Bormetti, Roberto Casarin, Fulvio Corsi, and Giulia Livieri. A stochastic volatility model with realized measures for option pricing. Journal of Business & Economic Statistics, 38(4):856–871, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1604371.

Barabesi:2018:GFT

[BCCP18] Lucio Barabesi, Andrea Cerasa, Andrea Cerioli, and Domenico Perrotta. Goodness-of-fit testing for the Newcomb–Benford Law with application to the detection of customs fraud. Journal of Business & Economic Statistics, 36(2):346–358, 2018.

ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1172014.

Bauwens:2017:AMA

[BCD17] Luc Bauwens, Jean-François Carpantier, and Arnaud Dufays. Autoregressive moving average infinite hidden Markov-switching models. Journal of Business & Economic Statistics, 35(2):162– 182, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2015.1123636.

Bishop:1994:TBO

[BCF94] John A. Bishop, Jong-Rong Chiou, and John P. Formby. Truncation bias and the ordinal evaiuation of income inequality. Journal of Business & Economic Statistics, 12(1): 123–127, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10509995.

Blanco:2019:BAQ

[BCFFL19] German Blanco, Xuan Chen, Carlos A. Flores, and Alfonso Flores-Lagunes. Bounds on average and quantile treatment effects on duration outcomes under censoring, selection, and noncompliance. Journal of Business & Economic Statistics, 38 (4):901–920, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1609975.

Brorsen:1985:DAP

[BCG85] B. Wade Brorsen, Jean-Paul Chavas, and Warren R. Grant. A dynamic analysis of prices in the U.S. rice marketing channel. Journal of Business & Economic Statistics, 3(4): 362–369, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509470.

Belloni:2016:IHD

[BCHK16] Alexandre Belloni, Victor Chernozhukov, Christian Hansen, and Damian Kozbur. Inference in high-dimensional panel models with an application to gun control. Journal of Business & Economic Statistics, 34(4):590–605, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1102733.

Bodory:2020:FSP

[BCHL20] Hugo Bodory, Lorenzo Camponovo, Martin Huber, and Michael Lechner. The finite sample performance of inference methods for propensity score matching and weighting estimators. Journal of Business & Economic Statistics, 38(1):183–200, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1476247.

Billio:2022:BDT

[BCIK22] Monica Billio, Roberto Casarin, Matteo Iacopini, and Sylvia Kaufmann. Bayesian dynamic tensor regression. Journal of Business & Economic Statistics, 41(2):429–439, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2032721.

Briesch:2010:NDC

[BCM10] Richard A. Briesch, Pradeep K. Chintagunta, and Rosa L. Matzkin. Nonparametric discrete choice models with unobserved heterogeneity. Journal of Business & Economic Statistics, 28 (2):291–307, 2010. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.07219.

Barbaglia:2022:FEN

[BCM22] Luca Barbaglia, Sergio Consoli, and Sebastiano Manzan. Forecasting with economic news. Journal of Business & Economic Statistics, 41(3):708–719, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2060988.

Buccheri:2019:HFL

[BCP19] Giuseppe Buccheri, Fulvio Corsi, and Stefano Peluso. Highfrequency lead-lag effects and cross-asset linkages: a multi-asset lagged adjustment model. Journal of Business & Economic Statistics, 39(3):605–621, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2019.1697699.

Baxter:1990:SSG

[BCR90] Marianne Baxter, Mario J. Crucini, and K. Geert Rouwenhorst. Solving the stochastic growth model by a discrete-state-space, Euler-equation approach. Journal of Business & Economic Statistics, 8(1):19–21, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509767.

Bhardwaj:2008:SBS

[BCS08] Geetesh Bhardwaj, Valentina Corradi, and Norman R. Swanson. A simulation-based specification test for diffusion processes. Journal of Business & Economic Statistics, 26(2):176–193, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500107000000412.

Bernales:2020:LIO

[BCSS20] Alejandro Bernales, Gonzalo Cortazar, Luka Salamunic, and George Skiadopoulos. Learning and index option returns. Journal of Business & Economic Statistics, 38(2):327–339, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1505629.

Belloni:2016:PSI

[BCW16] Alexandre Belloni, Victor Chernozhukov, and Ying Wei. Postselection inference for generalized linear models with many controls. Journal of Business & Economic Statistics, 34(4):606–619, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1166116.

Bollinger:2001:ERE

[BD01] Christopher R. Bollinger and Martin H. David. Estimation with response error and nonresponse. Journal of Business & Economic Statistics, 19(2):129–141, 2001. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500101316970368.

Barrett:2009:SIG

[BD09] Garry F. Barrett and Stephen G. Donald. Statistical inference with generalized Gini indices of inequality, poverty, and welfare. Journal of Business & Economic Statistics, 27(1):1–17, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2009.0001.

Barrett:2014:CNT

[BDB14] Garry F. Barrett, Stephen G. Donald, and Debopam Bhattacharya. Consistent nonparametric tests for Lorenz dominance. Journal of Business & Economic Statistics, 32(1):1–13, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2013.
834262.

Bissonnette:2018:ESS

[BdB18] L. Bissonnette and J. de Bresser. Eliciting subjective survival curves: Lessons from partial identification. Journal of Business & Economic Statistics, 36(3):505–515, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2016.1213635.

Bacolod:2012:BID

[BDJ12] Marigee Bacolod, John DiNardo, and Mireille Jacobson. Beyond incentives: Do schools use accountability rewards productively? Journal of Business & Economic Statistics, 30 (1):149–163, 2012. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.637868.

Beaulieu:2007:MTM

[BDK07] Marie-Claude Beaulieu, Jean-Marie Dufour, and Lynda Khalaf. Multivariate tests of mean-variance efficiency with possibly non-Gaussian errors. Journal of Business & Economic Statistics, 25(4):398–410, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000468.

Bekker:1996:AMR

[BDW96] Paul Bekker, Pascal Dobbelstein, and Tom Wansbeek. The APT model as reduced-rank regression. Journal of Business & Economic Statistics, 14(2):199–202, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524646.

Brockwell:2011:ENN

[BDY11] Peter J. Brockwell, Richard A. Davis, and Yu Yang. Estimation for non-negative Lévy-driven CARMA processes. Journal of Business & Economic Statistics, 29(2):250–259, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2010.08165.

Burnside:1996:SSP

[BE96] Craig Burnside and Martin Eichenbaum. Small-sample properties of GMM-based Wald tests. Journal of Business & Economic Statistics, 14(3):294–308, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524658.

Bekaert:1995:TVE

[Bek95] Geert Bekaert. The time variation of expected returns and volatility in foreign-exchange markets. Journal of Business & Economic Statistics, 13(4):397–408, 1995. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1995.10524614.

Bell:1987:NOE

[Bel87] William Bell. A note on overdifferencing and the equivalence of seasonal time series models with monthly means and models with (0, 1, 1)_{12} seasonal parts when ⊖ = 1. Journal of Business & Economic Statistics, 5(3):383–387, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509602.

Bell:1991:C

[Bel91] William R. Bell. Comment. Journal of Business & Economic Statistics, 9(2):176–177, 1991. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1991.10509841.

Bell:1996:C

[Bel96] William R. Bell. Comment. Journal of Business & Economic Statistics, 14(3):387–388, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524664.

Belluzzo:2004:SAW

[Bel04] Walter Belluzzo, Jr. Semiparametric approaches to welfare evaluations in binary response models. Journal of Business & Economic Statistics, 22(3):322–330, 2004. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000172.

Berkowitz:2001:TDF

[Ber01] Jeremy Berkowitz. Testing density forecasts, with applications to risk management. Journal of Business & Economic Statistics, 19(4):465–474, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
07350010152596718.

Beyer:2001:FSE

[Bey01] Andreas Beyer. A formalization of seasonal encompassing with an application to a German Macromodel. Journal of Business & Economic Statistics, 19(3):315–323, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500101681019963.

Boeschoten:1989:WWP

[BF89] W. C. Boeschoten and M. M. G. Fase. The way we pay with money. Journal of Business & Economic Statistics, 7 (3):319–326, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509741.

Boswijk:2005:EBD

[BF05] H. Peter Boswijk and Philip Hans Franses. On the econometrics of the Bass diffusion model. Journal of Business & Economic Statistics, 23(3):255–268, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000604.

Baardsen:2015:FES

[BF15] Gunnar Bårdsen and Luca Fanelli. Frequentist evaluation of small DSGE models. Journal of Business & Economic Statistics, 33(3):307–322, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.948724.

Bjelland:2011:EEF

[BFHM11] Melissa Bjelland, Bruce Fallick, John Haltiwanger, and Erika McEntarfer. Employer-to-employer flows in the United States: Estimates using linked employer-employee data. Journal of Business & Economic Statistics, 29(4):493–505, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2011.08053.

Bijwaard:2006:MPR

[BFP06] Govert E. Bijwaard, Philip Hans Franses, and Richard Paap. Modeling purchases as repeated events. Journal of Business & Economic Statistics, 24(4):487–502, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000242.

Behrman:1997:DSD

[BFR97] Jere R. Behrman, Andrew Foster, and Mark R. Rosenzweig. Dynamic savings decisions in agricultural environments with incomplete markets. Journal of Business & Economic Statistics, 15(2):282–292, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524706.

Bauwens:1994:EEU

[BFS94] Luc Bauwens, Denzil G. Fiebig, and Mark F. J. Steel. Estimating end-use demand: a Bayesian approach. Journal of Business & Economic Statistics, 12(2):221–231, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10510009.

Bai:2016:SIB

[BFT16] Jushan Bai, Jianqing Fan, and Ruey Tsay. Special issue on big data. Journal of Business & Economic Statistics, 34(4):487–488, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1197681.

Beyhum:2022:NIR

[BFV22] Jad Beyhum, Jean-Pierre Florens, and Ingrid Van Keilegom. Nonparametric instrumental regression with right censored duration outcomes. Journal of Business & Economic Statistics, 40(3):1034– 1045, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1895814.

Backus:1998:AOA

[BFZ98a] David Backus, Silverio Foresi, and Stanley Zin. Arbitrage opportunities in arbitrage-free models of bond pricing. Journal of Business & Economic Statistics, 16(1):13–26, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524731.

Bishop:1998:ITG

[BFZ98b] John A. Bishop, John P. Formby, and Buhong Zheng. Inference tests for Gini-based tax progressivity indexes. Journal of Business & Economic Statistics, 16(3):322–330, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524771.

Buse:1988:BR

[BG88] A. Buse and James E. Gentle. Book reviews. Journal of Business & Economic Statistics, 6(1):141–143, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509647.

Blattberg:1992:EUP

[BG92] Robert C. Blattberg and Edward I. George. Estimation under profit-driven loss functions. Journal of Business & Economic Statistics, 10(4):437–444, 1992. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1992.10509919.

Backus:1993:TRB

[BG93] David K. Backus and Allan W. Gregory. Theoretical relations between risk premiums and conditional variances. Journal of Business & Economic Statistics, 11(2):177–185, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509946.

Bollerslev:1996:PAC

[BG96] Tim Bollerslev and Eric Ghysels. Periodic autoregressive conditional heteroscedasticity. Journal of Business & Economic Statistics, 14(2):139–151, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524640.

Baltagi:2001:ERA

[BG01] Badi H. Baltagi and James M. Griffin. The econometrics of rational addiction. Journal of Business & Economic Statistics, 19(4):449–454, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
07350010152596691.

Bodapati:2005:PFB

[BG05] Anand V. Bodapati and Sachin Gupta. Purchase-frequency bias in random-coefficients brand-choice models. Journal of Business & Economic Statistics, 23(4):473–484, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000569.

Beyhum:2023:FFL

[BG23] Jad Beyhum and Eric Gautier. Factor and factor loading augmented estimators for panel regression with possibly nonstrong factors. Journal of Business & Economic Statistics, 41(1):270– 281, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.2011300.

Bhattacharya:2022:SQM

[BGG22] Jayeeta Bhattacharya, Nathalie Gimenes, and Emmanuel Guerre. Semiparametric quantile models for ascending auctions with asymmetric bidders. Journal of Business & Economic Statistics, 40(3):1020–1033, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1895813.

Banerjee:2004:SMH

[BGKS04] S. Banerjee, A. E. Gelfand, J. R. Knight, and C. F. Sirmans. Spatial modeling of house prices using normalized distance-weighted sums of stationary processes. Journal of Business & Economic Statistics, 22(2):206–213, 2004. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000091.

Brownlees:2020:CDP

[BGL20] Christian Brownlees, Gumundur Stefán Gumundsson, and Gábor Lugosi. Community detection in partial correlation network models. Journal of Business & Economic Statistics, 40(1): 216–226, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1798241.

Basu:2021:RGR

[BGMP21] Ayanendranath Basu, Abhik Ghosh, Nirian Martin, and Leandro Pardo. A robust generalization of the Rao test. Journal of Business & Economic Statistics, 40(2):868–879, 2021.

CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2021.1876711.

Berndt:1993:APP

[BGR93] Ernst R. Berndt, Zvi Griliches, and Joshua G. Rosett. Auditing the producer price index: Micro evidence from prescription pharmaceutical preparations. Journal of Business & Economic Statistics, 11(3):251–264, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509953.

Bianchi:2015:MFS

[BGR15] Daniele Bianchi, Massimo Guidolin, and Francesco Ravazzolo. Macroeconomic factors strike back: a Bayesian changepoint model of time-varying risk exposures and premia in the U.S. cross-section. Journal of Business & Economic Statistics, 35(1):110–129, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1061436.

Boivin:2020:DEC

[BGS20] Jean Boivin, Marc P. Giannoni, and Dalibor Stevanović. Dynamic effects of credit shocks in a data-rich environment. Journal of Business & Economic Statistics, 38(2):272–284, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1497507.

Babii:2022:MLT

[BGS22] Andrii Babii, Eric Ghysels, and Jonas Striaukas. Machine learning time series regressions with an application to nowcasting. Journal of Business & Economic Statistics, 40(3):1094– 1106, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1899933.

Bell:1984:IIS

[BH84a] William R. Bell and Steven C. Hillmer. Issues involved with the seasonal adjustment of economic time series. Journal of Business & Economic Statistics, 2(4):291–320, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1984.10509398. See comment

[Sim85].

Bell:1984:R

[BH84b] William R. Bell and Steven C. Hillmer. Reply. Journal of Business & Economic Statistics, 2(4):343–349, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1984.10509407.

Bowers:1984:KTA

[BH84c] Norman Bowers and Francis W. Horvath. Keeping time: an analysis of errors in the measurement of unemployment duration. Journal of Business & Economic Statistics, 2(2): 140–149, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509382.

Bell:1985:R

[BH85] William R. Bell and Steven C. Hiilmer. Reply. Journal of Business & Economic Statistics, 3(1):95–97, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1985.10509432.

Brunner:1993:HLI

[BH93] Allan D. Brunner and Gregory D. Hess. Are higher levels of inflation less predictable? A state-dependent conditional heteroscedasticity approach. Journal of Business & Economic Statistics, 11(2):187–197, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509947.

Barnett:1994:FFP

[BH94] William A. Barnett and Jeong Ho Hahm. Financial-firm production of monetary services: a generalized symmetric baroett variable-profit-function approach. Journal of Business & Economic Statistics, 12(1):33–46, 1994. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1994.10509989.

Bera:1997:ABC

[BH97] Anil K. Bera and Matthew L. Higgins. ARCH and bilinearity as competing models for nonlinear dependence. Journal of Business & Economic Statistics, 15(1):43–50, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524685.

Bell:2002:IIS

[BH02] William R. Bell and Steven C. Hillmer. Issues involved with the seasonal adjustment of economic time series. Journal of Business & Economic Statistics, 20(1):98–127, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102753410426.

Busetti:2003:ST

[BH03] Fabio Busetti and Andrew Harvey. Seasonality tests. Journal of Business & Economic Statistics, 21(3):420–436, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500103288619061.

Bester:2009:IME

[BH09a] C. Alan Bester and Christian Hansen. Identification of marginal effects in a nonparametric correlated random effects model. Journal of Business & Economic Statistics, 27(2):235–250, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2009.0017.

Bester:2009:PFA

[BH09b] C. Alan Bester and Christian Hansen. A penalty function approach to bias reduction in nonlinear panel models with fixed effects. Journal of Business & Economic Statistics, 27(2):131–148, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1198/jbes.2009.0012.

Bajari:2010:ESM

[BHKN10] Patrick Bajari, Han Hong, John Krainer, and Denis Nekipelov. Estimating static models of strategic interactions. Journal of Business & Economic Statistics, 28(4):469–482, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2009.07264.

Bera:1992:IBA

[BHL92] Anil K. Bera, Matthew L. Higgins, and Sangkyu Lee. Interaction between autocorrelation and conditional heteroscedasticity: a random-coefficient approach. Journal of Business & Economic Statistics, 10(2):133–142, 1992. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1992.10509893.

Benedetto:2007:UWF

[BHLM07] Gary Benedetto, John Haltiwanger, Julia Lane, and Kevin McKinney. Using worker flows to measure firm dynamics. Journal of Business & Economic Statistics, 25(3):299–313, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000620.

Bibinger:2019:ESC

[BHMR19] Markus Bibinger, Nikolaus Hautsch, Peter Malec, and Markus Reiss. Estimating the spot covariation of asset prices-statistical theory and empirical evidence. Journal of Business & Economic Statistics, 37(3):419–435, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1356728.

Bali:2009:PIR

[BHW09] Turan Bali, Massoud Heidari, and Liuren Wu. Predictability of interest rates and interest-rate portfolios. Journal of Business & Economic Statistics, 27(4):517–527, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.06124.

Barassi:2020:CPD

[BHZ20] Marco Barassi, Lajos Horváth, and Yuqian Zhao. Change-point detection in the conditional correlation structure of multivariate volatility models. Journal of Business & Economic Statistics, 38(2):340–349, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1505630.

Belongia:2015:IRM

[BI15] Michael T. Belongia and Peter N. Ireland. Interest rates and money in the measurement of monetary policy. Journal of Business & Economic Statistics, 33(2):255–269, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.946132.

Bertanha:2020:EVF

[BI20] Marinho Bertanha and Guido W. Imbens. External validity in fuzzy regression discontinuity designs. Journal of Business & Economic Statistics, 38(3):593–612, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2018.1546590.

Bierens:2000:NNC

[Bie00] Herman J. Bierens. Nonparametric nonlinear cotrending analysis, with an application to interest and inflation in the United States. Journal of Business & Economic Statistics, 18 (3):323–337, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524874.

Bollerslev:1999:ETV

[BJ99] Tim Bollerslev and Dan Jubinski. Equity trading volume and volatility: Latent information arrivals and common longrun dependencies. Journal of Business & Economic Statistics, 17(1):9–21, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524793.

Brandt:2006:VFR

[BJ06] Michael W. Brandt and Christopher S. Jones. Volatility forecasting with range-based EGARCH models. Journal of Business & Economic Statistics, 24(4):470–486, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000206.

Birchenhall:1999:PUB

[BJOS99] Chris R. Birchenhall, Hans Jessen, Demise R. Osborn, and Paul Simpson. Predicting U.S. business-cycle regimes. Journal of Business & Economic Statistics, 17(3):313–323, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524820.

Biemer:1991:ABD

[BK91] Paul P. Biemer and Sheryl E. Kimes. An application of bootstrapping for determining a decision rule for site location. Journal of Business & Economic Statistics, 9(2): 189–196, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509844.

Bajari:2005:EHD

[BK05] Patrick Bajari and Matthew E. Kahn. Estimating housing demand with an application to explaining racial segregation in cities. Journal of Business & Economic Statistics, 23(1):20–33, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000334.

Baillie:2007:TNN

[BK07] Richard T. Baillie and George Kapetanios. Testing for neglected nonlinearity in long-memory models. Journal of Business & Economic Statistics, 25(4):447–461, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000305.

Bansal:2011:CLR

[BK11] Ravi Bansal and Dana Kiku. Cointegration and long-run asset allocation. Journal of Business & Economic Statistics, 29 (1):161–173, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2010.08062.

Baumeister:2012:RTF

[BK12a] Christiane Baumeister and Lutz Kilian. Real-time forecasts of the real price of oil. Journal of Business & Economic Statistics, 30(2):326–336, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2011.648859.

Boswijk:2012:WFM

[BK12b] H. Peter Boswijk and Franc Klaassen. Why frequency matters for unit root testing in financial time series. Journal of Business & Economic Statistics, 30(3):351–357, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2011.648858.

Baumeister:2015:FRP

[BK15] Christiane Baumeister and Lutz Kilian. Forecasting the real price of oil in a changing world: a forecast combination approach. Journal of Business & Economic Statistics, 33(3):338–351, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2014.949342.

Bauer:2016:RMC

[BK16] Daniel Bauer and Florian Kramer. The risk of a mortality catastrophe. Journal of Business & Economic Statistics, 34(3):391–405, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1040117.

Bayer:2011:NIE

[BKT11] Patrick Bayer, Shakeeb Khan, and Christopher Timmins. Nonparametric identification and estimation in a Roy model with common nonpecuniary returns. Journal of Business & Economic Statistics, 29(2):201–215, 2011. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2010.08083.

Bosma:2019:TCF

[BKW19] Jakob J. Bosma, Michael Koetter, and Michael Wedow. Too connected to fail? Inferring network ties from price co-movements. Journal of Business & Economic Statistics, 37(1):67–80, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1272459.

Biattenberger:1988:R

[BL88a] Gail Biattenberger and Frank Lad. Reply. Journal of Business & Economic Statistics, 6(4):475–477, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509695.

Blattenberger:1988:AOS

[BL88b] Gail Blattenberger and Frank Lad. An application of operationalsubjective statistical methods to rational expectations. Journal of Business & Economic Statistics, 6(4):453–464, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1988.10509690.

Berthelot:1993:IED

[BL93] Jean-Marie Berthelot and Michel Latouche. Improving the efficiency of data collection: a generic respondent follow-up strategy for economic surveys. Journal of Business & Economic Statistics, 11(4):417–424, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509975.

Bauwens:2005:NCM

[BL05] Luc Bauwens and Sébastien Laurent. A new class of multivariate skew densities, with application to generalized autoregressive conditional heteroscedasticity models. Journal of Business & Economic Statistics, 23(3):346–354, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000523.

Blanchard:1987:C

[Bla87] Olivier Jean Blanchard. Comment. Journal of Business & Economic Statistics, 5(4):449–451, 1987. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1987.10509611.

Blackburn:1989:PMI

[Bla89] McKinley L. Blackburn. Poverty measurement: an index related to a Theil measure of inequality. Journal of Business & Economic Statistics, 7(4):475–481, 1989. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1989.10509760.

Bollerslev:2020:GJR

[BLC20] Tim Bollerslev, Jia Li, and Leonardo Salim Saker Chaves. Generalized jump regressions for local moments. Journal of Business & Economic Statistics, 39(4):1015–1025, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1753526.

Bai:2016:EIF

[BLL16] Jushan Bai, Kunpeng Li, and Lina Lu. Estimation and inference of FAVAR models. Journal of Business & Economic Statistics, 34(4):620–641, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1111222.

Bloemen:2019:CLS

[Blo19] Hans G. Bloemen. Collective labor supply, taxes, and intrahousehold allocation: an empirical approach. Journal of Business & Economic Statistics, 37(3):471–483, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1379407.

Banerjee:1992:RST

[BLS92] Anindya Banerjee, Robin L. Lumsdaine, and James H. Stock. Recursive and sequential tests of the unit-root and trend-break hypotheses: Theory and international evidence. Journal of Business & Economic Statistics, 10(3):271–287, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1992.10509905.

Butler:1987:III

[BM87] Richard J. Butler and James B. McDonald. Interdistributional income inequality. Journal of Business & Economic Statistics, 5(1):13–18, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509555.

Bordley:1993:EAA

[BM93] Robert F. Bordley and James B. McDonald. Estimating aggregate automotive income elasticities from the population incomeshare elasticity. Journal of Business & Economic Statistics, 11 (2):209–214, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509949.

Barseghyan:2023:R

[BM23a] Levon Barseghyan and Francesca Molinari. Rejoinder. Journal of Business & Economic Statistics, 41(4):1046–1049, 2023. CO-DEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.

2023.2239870. See [BM23b, CMM23, Gua23, Hon23, Mor23].

Barseghyan:2023:RPT

[BM23b] Levon Barseghyan and Francesca Molinari. Risk preference types, limited consideration, and welfare. Journal of Business & Economic Statistics, 41(4):1011–1029, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2023.

2239949. See discussion [CMM23, Gua23, Hon23, Mor23] and rejoinder [BM23a].

Blundell:1997:SLM

[BMM97] Richard Blundell, Thierry Magnac, and Costas Meghir. Savings and labor-market transitions. Journal of Business & Economic Statistics, 15(2):153–164, 1997. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1997.10524697.

Bia:2020:ACE

[BMM20] Michela Bia, Alessandra Mattei, and Andrea Mercatanti. Assessing causal effects in a longitudinal observational study with “truncated” outcomes due to unemployment and nonignorable missing data. Journal of Business & Economic Statistics, 40(2):718–729, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1862672.

Bakshi:2010:DIJ

[BMP10] Gurdip Bakshi, Dilip Madan, and George Panayotov. Deducing the implications of jump models for the structure of stock market crashes, rallies, jump arrival rates, and extremes. Journal of Business & Economic Statistics, 28(3):380–396, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.06176.

Berg:2004:DIC

[BMY04] Andreas Berg, Renate Meyer, and Jun Yu. Deviance information criterion for comparing stochastic volatility models. Journal of Business & Economic Statistics, 22(1):107–120, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500103288619430.

Brown:2002:GMM

[BN02] Bryan W. Brown and Whitney K. Newey. Generalized method of moments, efficient bootstrapping, and improved inference. Journal of Business & Economic Statistics, 20(4):507–517, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500102288618649.

Bai:2005:TSK

[BN05] Jushan Bai and Serena Ng. Tests for skewness, kurtosis, and normality for time series data. Journal of Business & Economic Statistics, 23(1):49–60, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000271.

Bai:2007:DNP

[BN07] Jushan Bai and Serena Ng. Determining the number of primitive shocks in factor models. Journal of Business & Economic Statistics, 25(1):52–60, 2007. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500106000000413.

Burgette:2012:TRA

[BN12] Lane F. Burgette and Erik V. Nordheim. The trace restriction: an alternative identification strategy for the Bayesian multinomial probit model. Journal of Business & Economic Statistics, 30(3):404–410, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.680416.

Barndorff-Nielsen:2006:C

[BNS06] Ole E. Barndorff-Nielsen and Neil Shephard. Comment. Journal of Business & Economic Statistics, 24(2):179–181, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000099. See

[HL06a, HL06b].

Beran:2001:VSM

[BO01] Jan Beran and Dirk Ocker. Volatility of stock-market indexes — an analysis based on SEMIFAR models. Journal of Business & Economic Statistics, 19(1):103–116, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/07350010152472661.

Bauwens:2016:MDC

[BO16] Luc Bauwens and Edoardo Otranto. Modeling the dependence of conditional correlations on market volatility. Journal of Business & Economic Statistics, 34(2):254–268, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1037882.

Bohara:1991:TRE

[Boh91] Alok K. Bohara. Testing the rational-expectations hypothesis: Further evidence. Journal of Business & Economic Statistics, 9(3):337–340, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509860.

Bonhomme:2020:DTS

[Bon20] Stéphane Bonhomme. Discussion of “Transparency in Structural Research” by Isaiah Andrews, Matthew Gentzkow, and Jesse Shapiro. Journal of Business & Economic Statistics, 38 (4):723–725, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1790377.

Bodnar:2023:OSB

[BOP23] Taras Bodnar, Yarema Okhrin, and Nestor Parolya. Optimal shrinkage-based portfolio selection in high dimensions. Journal of Business & Economic Statistics, 41(1):140–156, 2023. CO-DEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2021.2004897.

Bordley:1985:REC

[Bor85] Robert F. Bordley. Relating elasticities to changes in demand. Journal of Business & Economic Statistics, 3(2): 156–158, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509444.

Bordley:1989:GME

[Bor89] Robert F. Bordley. Generating market elasticity estimates using cross-sectional first-choice and second-choice data. Journal of Business & Economic Statistics, 7(1):141–146, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1989.10509722.

Boswijk:1996:TIC

[Bos96] H. Peter Boswijk. Testing identifiability of cointegrating vectors. Journal of Business & Economic Statistics, 14(2): 153–160, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524641.

Boot:2018:OFM

[BP18] Tom Boot and Andreas Pick. Optimal forecasts from Markov switching models. Journal of Business & Economic Statistics, 36(4):628–642, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1219264.

Barendse:2022:CPA

[BP22] Sander Barendse and Andrew J. Patton. Comparing predictive accuracy in the presence of a loss function shape parameter. Journal of Business & Economic Statistics, 40(3):1057– 1069, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1896527.

Bera:1989:TSD

[BR89] Anil K. Bera and Peter M. Robinson. Tests for serial dependence and other specification analysis in models of markets in disequilibrium. Journal of Business & Economic Statistics, 7 (3):343–352, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509743.

Bollerslev:1995:DNR

[BR95] Tim Bollerslev and Peter E. Rossi. Dan Nelson remembered. Journal of Business & Economic Statistics, 13(4): 361–364, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524611.

Bandi:2006:C

[BR06] Federico M. Bandi and Jeffrey R. Russell. Comment. Journal of Business & Economic Statistics, 24(2):167–173, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000107. See

[HL06a, HL06b].

Balduzzi:2008:MPE

[BR08] Pierluigi Balduzzi and Cesare Robotti. Mimicking portfolios, economic risk premia, and tests of multi-beta models. Journal of Business & Economic Statistics, 26(3):354–368, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500108000000042.

Braun:1990:ECQ

[Bra90] Steven N. Braun. Estimation of current-quarter gross national product by pooling preliminary labor-market data. Journal of Business & Economic Statistics, 8(3):293–304, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1990.10509800.

Bradley:2007:ABR

[Bra07] Ralph Bradley. Analytical bias reduction for small samples in the U.S. consumer price index. Journal of Business & Economic Statistics, 25(3):337–346, 2007. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500106000000639.

Bramoulle:2013:C

[Bra13] Yann Bramoullé. Comment. Journal of Business & Economic Statistics, 31(3):264–266, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.792265. See [GPI13b, GPI13a].

Breitung:2001:RTN

[Bre01] Jörg Breitung. Rank tests for nonlinear cointegration. Journal of Business & Economic Statistics, 19(3):331–340, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500101681019981.

Breunig:2019:TMR

[Bre19] Christoph Breunig. Testing missing at random using instrumental variables. Journal of Business & Economic Statistics, 37(2):223–234, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1302879.

Brown:1988:C

[Bro88] Stephen J. Brown. Comment. Journal of Business & Economic Statistics, 6(1):21–23, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509632. See [CR88b, CR88a].

Bouezmarni:2012:NCB

[BRT12] Taoufik Bouezmarni, Jeroen V. K. Rombouts, and Abderrahim Taamouti. Nonparametric copula-based test for conditional independence with applications to Granger causality. Journal of Business & Economic Statistics, 30(2):275–287, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2011.638831.

Brunner:1992:CAR

[Bru92] Allan D. Brunner. Conditional asymmetries in real GNP: a seminonparametric approach. Journal of Business & Economic Statistics, 10(1):65–72, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509887.

Bauer:2012:CEB

[BRW12] Michael D. Bauer, Glenn D. Rudebusch, and Jing Cynthia Wu. Correcting estimation bias in dynamic term structure models.

Journal of Business & Economic Statistics, 30(3):454–467, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2012.
693855.

Bandi:2013:RVF

[BRY13] Federico M. Bandi, Jeffrey R. Russell, and Chen Yang. Realized volatility forecasting in the presence of time-varying noise. Journal of Business & Economic Statistics, 31(3):331–345, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2013.803866.

Basmann:1987:STC

[BS87a] R. L. Basmann and D. J. Slottje. The sensitivity of the true cost of living to price-induced and income-induced changes in aggregate consumers’ tastes. Journal of Business & Economic Statistics, 5(4):483–498, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509617.

Belongia:1987:IEW

[BS87b] Michael T. Belongia and Richard G. Sheehan. The informational efficiency of weekly money announcements an econometric critique. Journal of Business & Economic Statistics, 5 (3):351–356, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509599.

Brownstone:1989:EEN

[BS89] David Brownstone and Kenneth A. Small. Efficient estimation of nested logit models. Journal of Business & Economic Statistics, 7(1):67–74, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509714.

Bowden:1992:PB

[BS92] Roger J. Bowden and Ah Boon Sim. The privacy bootstrap. Journal of Business & Economic Statistics, 10(3): 337–345, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509909.

Breslaw:1995:MWC

[BS95] Jon A. Breslaw and J. Barry Smith. Measuring welfare changes when quantity is constrained. Journal of Business & Economic Statistics, 13(1):95–103, 1995. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1995.10524582.

Bhattacharya:1998:AMF

[BS98] Uttpal Bhattacharya and Matthew Spiegel. Anatomy of a market failure: NYSE trading suspensions (1974–1988). Journal of Business & Economic Statistics, 16(2):216–226, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524755.

Brezger:2008:MRB

[BS08] Andreas Brezger and Winfried J. Steiner. Monotonic regression based on Bayesian P-splines. Journal of Business & Economic Statistics, 26(1):90–104, 2008. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000223.

Bormann:2020:DSD

[BS20a] Carsten Bormann and Melanie Schienle. Detecting structural differences in tail dependence of financial time series. Journal of Business & Economic Statistics, 38(2):380–392, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1506343.

Borup:2020:SJF

[BS20b] Daniel Borup and Erik Christian Montes Schütte. In search of a job: Forecasting employment growth using Google Trends. Journal of Business & Economic Statistics, 40(1):186–200, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1791133.

Beg:2001:TAI

[BSS01] A. B. M. Rabiul A. Beg, Mervyn J. Silvapulle, and Paramsothy Silvapulle. Tests against inequality constraints when some nuisance parameters are present only under the alternative. Journal of Business & Economic Statistics, 19(2):245–253, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500101316970467.

Bai:2022:TSM

[BSS22] Yuehao Bai, Andres Santos, and Azeem M. Shaikh. A twostep method for testing many moment inequalities. Journal of Business & Economic Statistics, 40(3):1070–1080, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1897016.

Bordignon:1989:OUP

[BT89] Silvano Bordignon and Ugo Trivellato. The optimal use of provisional data in forecasting with dynamic models. Journal of Business & Economic Statistics, 7(2):275–286, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509737.

Burridge:2001:PRB

[BT01] Peter Burridge and A. M. Robert Taylor. On the properties of regression-based tests for seasonal unit roots in the presence of higher-order serial correlation. Journal of Business & Economic Statistics, 19(3):374–379, 2001. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500101681019918.

Busetti:2003:VSS

[BT03] Fabio Busetti and A. M. Robert Taylor. Variance shifts, structural breaks, and stationarity tests. Journal of Business & Economic Statistics, 21(4):510–531, 2003. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500103288619269.

Buhai:2014:TPE

[BT14] I. Sebastian Buhai and Coen N. Teulings. Tenure profiles and efficient separation in a stochastic productivity model. Journal of Business & Economic Statistics, 32(2):245–258, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2013.866568.

Barigozzi:2022:TCT

[BT22] Matteo Barigozzi and Lorenzo Trapani. Testing for common trends in nonstationary large datasets. Journal of Business & Economic Statistics, 40(3):1107–1122, 2022. CODEN ????

ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1901719.

Bansal:2004:RSR

[BTZ04] Ravi Bansal, George Tauchen, and Hao Zhou. Regime shifts, risk premiums in the term structure, and the business cycle. Journal of Business & Economic Statistics, 22(4):396–409, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500104000000398.

Buja:1996:C

[Buj96] Andreas Buja. Comment. Journal of Business & Economic Statistics, 14(1):128–129, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524636.

Burman:1984:C

[Bur84] J. Peter Burman. Comment. Journal of Business & Economic Statistics, 2(4):325–327, 1984. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1984.10509401.

Burnside:1994:HJB

[Bur94] Craig Burnside. Hansen–Jagannathan bounds as classical tests of asset-pricing models. Journal of Business & Economic Statistics, 12(1):57–79, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10509991.

Butz:1985:CFA

[But85] William P. Butz. Comment: The future of administrative records in the census Bureau’s demographic activities. Journal of Business & Economic Statistics, 3(4):393–395, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1985.10509474.

Bernard:1987:PDF

[BV87] Jean-Thomas Bernard and Michael R. Veall. The probability distribution of future demand. Journal of Business & Economic Statistics, 5(3):417–424, 1987. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1987.10509606.

Bunzel:2005:PTF

[BV05] Helle Bunzel and Timothy J. Vogelsang. Powerful trend function tests that are robust to strong serial correlation, with an application to the prebisch-Singer hypothesis. Journal of Business & Economic Statistics, 23(4):381–394, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000631.

Burridge:1984:UCM

[BW84] Peter Burridge and Kenneth F. Wallis. Unobserved-components models for seasonal adjustment filters. Journal of Business & Economic Statistics, 2(4):350–359, 1984. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1984.10509408.

Belsley:1988:C

[BW88] David A. Belsley and Roy E. Welsch. Comment. Journal of Business & Economic Statistics, 6(4):442–447, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509687.

Bai:2015:IBE

[BW15] Jushan Bai and Peng Wang. Identification and Bayesian estimation of dynamic factor models. Journal of Business & Economic Statistics, 33(2):221–240, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.941467.

Berry:2019:BFM

[BW19] Lindsay R. Berry and Mike West. Bayesian forecasting of many count-valued time series. Journal of Business & Economic Statistics, 38(4):872–887, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2019.1604372.

Bonhomme:2022:PAE

[BW22] Stéphane Bonhomme and Martin Weidner. Posterior average effects. Journal of Business & Economic Statistics, 40(4):1849– 1862, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1984928.

Burmeister:1986:EUE

[BWH86] Edwin Burmeister, Kent D. Wall, and James D. Hamilton. Estimation of unobserved expected monthly inflation using Kalman filtering. Journal of Business & Economic Statistics, 4(2):147–160, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509510.

Bykhovskaya:2023:TSA

[Byk23] Anna Bykhovskaya. Time series approach to the evolution of networks: Prediction and estimation. Journal of Business & Economic Statistics, 41(1):170–183, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
2006669.

Bianchi:1999:NAR

[BZ99] Marco Bianchi and Gylfi Zoega. A nonparametrsc analysis of regional unemployment dynamics in Britain. Journal of Business & Economic Statistics, 17(2):205–216, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524811.

Bennett:2013:EDE

[BZ13] Christopher J. Bennett and Ricardas Zitikis. Examining the distributional effects of military service on earnings: a test of initial dominance. Journal of Business & Economic Statistics, 31(1):1–15, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.741053.

Boswijk:2020:ATC

[BZ20] H. Peter Boswijk and Yang Zu. Adaptive testing for cointegration with nonstationary volatility. Journal of Business & Economic Statistics, 40(2):744–755, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1867558.

Cagetti:2003:WAL

[Cag03] Marco Cagetti. Wealth accumulation over the life cycle and precautionary savings. Journal of Business & Economic Statistics, 21(3):339–353, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619007.

Cai:1994:MMS

[Cai94] Jun Cai. A Markov model of switching-regime ARCH. Journal of Business & Economic Statistics, 12(3):309–316, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1994.10524546.

Callan:1991:SPG

[Cal91] Scott J. Callan. The sensitivity of productivity growth measures to alternative structural and behavioral assumptions: an application to electric utilities 1951–1984. Journal of Business & Economic Statistics, 9(2):207–213, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509846.

Campbell:2007:MVP

[Cam07] Sean D. Campbell. Macroeconomic volatility, predictability, and uncertainty in the great moderation. Journal of Business & Economic Statistics, 25(2):191–200, 2007. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500106000000558.

Canova:1992:AAM

[Can92] Fabio Canova. An alternative approach to modeling and forecasting seasonal time series. Journal of Business & Economic Statistics, 10(1):97–108, 1992. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1992.10509890.

Caner:1998:LOS

[Can98] Mehmet Caner. A locally optimal seasonal unit-root test. Journal of Business & Economic Statistics, 16(3):349–356, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1998.10524774.

Canova:2009:C

[Can09] Fabio Canova. Comment. Journal of Business & Economic Statistics, 27(3):311–315, 2009. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.08262. See [KM09b, KM09a].

Cartwright:1984:NUS

[Car84] Phillip A. Cartwright. A note on using state-dependent models with a time-dependent variance. Journal of Business & Economic Statistics, 2(4):410–413, 1984. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1984.10509417.

Carroll:1985:CUA

[Car85] John J. Carroll. Comment: Uses of administrative records: a social security point of view. Journal of Business & Economic Statistics, 3(4):396–397, 1985. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1985.10509475.

Carrasco:2001:BCB

[Car01] Raquel Carrasco. Binary choice with binary endogenous regressors in panel data. Journal of Business & Economic Statistics, 19(4):385–394, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
07350010152596637.

Carstensen:2006:SMD

[Car06] Kai Carstensen. Stock market downswing and the stability of European monetary union money demand. Journal of Business & Economic Statistics, 24(4):395–402, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000369.

Catania:2020:SVM

[Cat20] Leopoldo Catania. A stochastic volatility model with a general leverage specification. Journal of Business & Economic Statistics, 40(2):678–689, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1855187.

Cavicchioli:2022:MSG

[Cav22] Maddalena Cavicchioli. Markov switching Garch models: Higher order moments, kurtosis measures, and volatility evaluation in recessions and pandemic. Journal of Business & Economic Statistics, 40(4):1772–1783, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1974459.

Chatterjee:1987:BR

[CBC^{+}87] Samprit Chatterjee, Peter Bloomfield, Samprit Chatterjee, James B. McDonald, and Donald Richter. Book reviews. Journal of Business & Economic Statistics, 5(1):159–162, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1987.10509572.

Caves:1983:D

[CC83] Douglas W. Caves and Laurits R. Christensen. Discussion. Journal of Business & Economic Statistics, 1(3): 219–220, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509343.

Cheung:1997:FIU

[CC97] Yin-Wong Cheung and Menzie D. Chinn. Further investigation of the uncertain unit root in GNP. Journal of Business & Economic Statistics, 15(1):68–73, 1997. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1997.10524688.

Cattaneo:2014:C

[CC14] Matias D. Cattaneo and Richard K. Crump. Comment. Journal of Business & Economic Statistics, 32(3):324–329, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2014.928220.

See [Mül14a, Mül14b].

Callot:2017:STD

[CCKR17] Laurent Callot, Mehmet Caner, Anders Bredahl Kock, and Juan Andres Riquelme. Sharp threshold detection based on supnorm error rates in high-dimensional models. Journal of Business & Economic Statistics, 35(2):250–264, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1052461.

Carriero:2016:CDV

[CCM16] Andrea Carriero, Todd E. Clark, and Massimiliano Marcellino. Common drifting volatility in large Bayesian VARs. Journal of Business & Economic Statistics, 34(3):375–390, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1040116.

Callot:2019:NRA

[CCÖU19] Laurent Callot, Mehmet Caner, A. Özlem Önder, and Esra Ulasan. A nodewise regression approach to estimating large portfolios. Journal of Business & Economic Statistics, 39 (2):520–531, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1683018.

Canals-Cerda:2013:ATE

[CCP13] José J. Canals-Cerdá and Jason Pearcy. Arriving in time: Estimation of English auctions with a stochastic number of bidders. Journal of Business & Economic Statistics, 31(2):125–135, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2012.747825.

Christoffersen:1998:CLH

[CD98] Peter F. Christoffersen and Francis X. Diebold. Cointegration and long-horizon forecasting. Journal of Business & Economic Statistics, 16(4):450–456, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524784.

Campbell:2009:SRE

[CD09] Sean D. Campbell and Francis X. Diebold. Stock returns and expected business conditions: Half a century of direct evidence. Journal of Business & Economic Statistics, 27(2):266–278, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2009.0025.

Cakmakli:2022:USI

[ÇD22] Cem Çakmaklı and Hamza Demircan. Using survey information for improving the density nowcasting of U.S. GDP. Journal of Business & Economic Statistics, 41(3):667–682, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2058000.

Cowell:2015:GFA

[CDF15] Frank A. Cowell, Russell Davidson, and Emmanuel Flachaire. Goodness of fit: an axiomatic approach. Journal of Business & Economic Statistics, 33(1):54–67, 2015. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2014.922470.

Christiano:1996:SSP

[CdH96] Lawrence J. Christiano and Wouter J. den Haan. Small-sample properties of GMM for business-cycle analysis. Journal of Business & Economic Statistics, 14(3):309–327, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524659.

Christoffersen:2010:VCA

[CDJW10] Peter Christoffersen, Christian Dorion, Kris Jacobs, and Yintian Wang. Volatility components, affine restrictions, and nonnormal innovations. Journal of Business & Economic Statistics, 28(4):483–502, 2010. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.06122.

Carlini:2019:IFC

[CdM19] Federico Carlini and Paolo Santucci de Magistris. On the identification of fractionally cointegrated VAR models with the F(d) condition. Journal of Business & Economic Statistics, 37(1):134–146, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1294077.

Catania:2020:DDM

[CDS20] Leopoldo Catania, Roberto Di Mari, and Paolo Santucci de Magistris. Dynamic discrete mixtures for high-frequency prices. Journal of Business & Economic Statistics, 40(2):559–577, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1840994.

Chen:2013:UIP

[CDY13] Willa W. Chen, Rohit S. Deo, and Yanping Yi. Uniform inference in predictive regression models. Journal of Business & Economic Statistics, 31(4):525–533, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.818008.

Cheung:2005:ERM

[CE05] Yin-Wong Cheung and Ulf G. Erlandsson. Exchange rates and Markov switching dynamics. Journal of Business & Economic Statistics, 23(3):314–320, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000488.

Choo:2008:IPD

[CE08] Eugene Choo and Jean Eid. Interregional price difference in the New Orleans auctions market for slaves. Journal of Business & Economic Statistics, 26(4):486–509, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500107000000421.

Cunningham:2012:SSA

[CEJ^{+}12] Alastair Cunningham, Jana Eklund, Chris Jeffery, George Kapetanios, and Vincent Labhard. A state space approach to extracting the signal from uncertain data. Journal of Business & Economic Statistics, 30(2):173–180, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1198/jbes.2009.08171.

Chen:1999:FSS

[CF99] Rong Chen and Thomas B. Fomby. Forecasting with stable seasonal pattern models with an application to Hawaiian tourism data. Journal of Business & Economic Statistics, 17 (4):497–504, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524839.

Carrasco:2002:SBM

[CF02] Marine Carrasco and Jean-Pierre Florens. Simulation-based method of moments and efficiency. Journal of Business & Economic Statistics, 20(4):482–492, 2002. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500102288618621.

Chen:2015:BTE

[CF15] Xuan Chen and Carlos A. Flores. Bounds on treatment effects in the presence of sample selection and noncompliance: The wage effects of job Corps. Journal of Business & Economic Statistics, 33(4):523–540, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.975229.

Caudill:1995:FEF

[CFG95] Steven B. Caudill, Jon M. Ford, and Daniel M. Gropper. Frontier estimation and firm-specific inefficiency measures in the presence of heteroscedasticity. Journal of Business & Economic Statistics, 13(1):105–111, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524583.

Cai:2019:IPV

[CFLS19] Zongwu Cai, Ying Fang, Ming Lin, and Jia Su. Inferences for a partially varying coefficient model with endogenous regressors. Journal of Business & Economic Statistics, 37(1):158–170, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1294079.

Chirinko:2011:NAE

[CFM11] Robert S. Chirinko, Steven M. Fazzari, and Andrew P. Meyer. A new approach to estimating production function parameters: The elusive capital-labor substitution elasticity. Journal of Business & Economic Statistics, 29(4):587–594, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2011.08119.

Cohen:2005:C

[CFR05] William W. Cohen, Stephen E. Fienberg, and Pradeep Ravikumar. Comments. Journal of Business & Economic Statistics, 23(2):160–162, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000703. See [AV05b, AV05a].

Corradi:2009:IRP

[CFS09] Valentina Corradi, Andres Fernandez, and Norman R. Swanson. Information in the revision process of real-time datasets. Journal of Business & Economic Statistics, 27(4):455–467, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2009.07209.

Christoffersen:2000:IRA

[CG00] Peter F. Christoffersen and Lorenzo Giorgianni. Interest-rate arbitrage in currency baskets: Forecasting weights and measuring risk. Journal of Business & Economic Statistics, 18 (2):242–253, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524866.

Chotikapanich:2002:ELC

[CG02] Duangkamon Chotikapanich and William E. Griffiths. Estimating Lorenz curves using a Dirichlet distribution. Journal of Business & Economic Statistics, 20(2):290–295, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102317352029.

Clements:2008:MFM

[CG08] Michael P. Clements and Ana Beatriz Galvão. Macroeconomic forecasting with mixed-frequency data. Journal of Business & Economic Statistics, 26(4):546–554, 2008. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500108000000015.

Clements:2012:IRT

[CG12] Michael P. Clements and Ana Beatriz Galvão. Improving real-time estimates of output and inflation gaps with multiplevintage models. Journal of Business & Economic Statistics, 30(4):554–562, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.707588.

Clements:2017:PED

[CG17] Michael P. Clements and Ana Beatriz Galvão. Predicting early data revisions to U.S. GDP and the effects of releases on equity markets. Journal of Business & Economic Statistics, 35(3):389–406, 2017. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1076726.

Chen:2019:SGB

[CG19] Wilson Ye Chen and Richard H. Gerlach. Semiparametric GARCH via Bayesian model averaging. Journal of Business & Economic Statistics, 39(2):437–452, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2019.1668796.

Chudik:2022:EIR

[CG22] Alexander Chudik and Georgios Georgiadis. Estimation of impulse response functions when shocks are observed at a higher frequency than outcome variables. Journal of Business & Economic Statistics, 40(3):965–979, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1889567.

Campello:2019:TSH

[CGJ19] Murillo Campello, Antonio F. Galvao, and Ted Juhl. Testing for slope heterogeneity bias in panel data models. Journal of Business & Economic Statistics, 37(4):749–760, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1421545.

Conrad:1997:PST

[CGK97] Jennifer Conrad, Mustafa N. Gultekin, and Gautam Kaul. Profitability of short-term contrarian strategies: Implications for market efficiency. Journal of Business & Economic Statistics, 15 (3):379–386, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524715.

Chen:2013:EPL

[CGL13] Jia Chen, Jiti Gao, and Degui Li. Estimation in partially linear single-index panel data models with fixed effects. Journal of Business & Economic Statistics, 31(3):315–330, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2013.775093.

Camacho:2020:NAD

[CGL20] Maximo Camacho, María Dolores Gadea, and Ana Gómez Loscos. A new approach to dating the reference cycle. Journal of Business & Economic Statistics, 40(1):66–81, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1773834.

Chen:2018:NEF

[CGLS18] Xiangjin B. Chen, Jiti Gao, Degui Li, and Param Silvapulle. Nonparametric estimation and forecasting for time-varying coefficient realized volatility models. Journal of Business & Economic Statistics, 36(1):88–100, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1138118.

Cameron:2011:RIM

[CGM11] A. Colin Cameron, Jonah B. Gelbach, and Douglas L. Miller. Robust inference with multiway clustering. Journal of Business & Economic Statistics, 29(2):238–249, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2010.07136.

Coroneo:2016:UMF

[CGM16] Laura Coroneo, Domenico Giannone, and Michele Modugno. Unspanned macroeconomic factors in the yield curve. Journal of Business & Economic Statistics, 34(3):472–485, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1052456.

Chen:2020:FLR

[CGQ20] Cheng Chen, Shaojun Guo, and Xinghao Qiao. Functional linear regression: Dependence and error contamination. Journal of Business & Economic Statistics, 40(1):444–457, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1832503.

Chotikapanich:2007:ECN

[CGR07] Duangkamon Chotikapanich, William E. Griffiths, and D. S. Prasada Rao. Estimating and combining national income distributions using limited data. Journal of Business & Economic Statistics, 25(1):97–109, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000224.

Creal:2014:MBC

[CGT14] Drew D. Creal, Robert B. Gramacy, and Ruey S. Tsay. Marketbased credit ratings. Journal of Business & Economic Statistics, 32(3):430–444, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.902763.

Cai:2017:NCB

[CGT17] Biqing Cai, Jiti Gao, and Dag Tjøstheim. A new class of bivariate threshold cointegration models. Journal of Business & Economic Statistics, 35(2):288–305, 2017. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1062385.

Cheng:2019:BBE

[CGZ19] Tingting Cheng, Jiti Gao, and Xibin Zhang. Bayesian bandwidth estimation in nonparametric time-varying coefficient models. Journal of Business & Economic Statistics, 37(1):1–12, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1255216.

Canova:1995:SPC

[CH95] Fabio Canova and Bruce E. Hansen. Are seasonal patterns constant over time? A test for seasonal stability. Journal of Business & Economic Statistics, 13(3):237–252, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524598.

Caner:2014:SCN

[CH14] Mehmet Caner and Xu Han. Selecting the correct number of factors in approximate factor models: The large panel case with group bridge estimators. Journal of Business & Economic Statistics, 32(3):359–374, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.880349.

Cheng:2016:C

[CH16] Xu Cheng and Bruce E. Hansen. Comment. Journal of Business & Economic Statistics, 34(3):345–347, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2016.1189338. See [CR16a,

CR16b].

Chalfant:1987:GFA

[Cha87] James A. Chalfant. A globally flexible, almost ideal demand system. Journal of Business & Economic Statistics, 5 (2):233–242, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509581.

Chamberlain:1992:CSM

[Cha92] Gary Chamberlain. Comment: Sequential moment restrictions in panel data. Journal of Business & Economic Statistics, 10(1):20–26, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509881.

Chatfield:1993:CIF

[Cha93a] Chris Chatfield. Calculating interval forecasts. Journal of Business & Economic Statistics, 11(2):121–135, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509938.

Chatfield:1993:R

[Cha93b] Chris Chatfield. Reply. Journal of Business & Economic Statistics, 11(2):143–144, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509942.

Chan:2015:SVM

[Cha15] Joshua C. C. Chan. The stochastic volatility in mean model with time-varying parameters: an application to inflation modeling. Journal of Business & Economic Statistics, 35(1):17–28, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1052459.

Chan:2020:LBV

[Cha20a] Joshua C. C. Chan. Large Bayesian VARs: a flexible Kronecker error covariance structure. Journal of Business & Economic Statistics, 38(1):68–79, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1451336.

Chan:2020:MSA

[Cha20b] Kin Wai Chan. Mean-structure and autocorrelation consistent covariance matrix estimation. Journal of Business & Economic Statistics, 40(1):201–215, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2020.1796397.

Chan:2022:LHT

[Cha22] Joshua C. C. Chan. Large hybrid time-varying parameter VARs. Journal of Business & Economic Statistics, 41(3):890– 905, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2080683.

Cheung:1993:LMF

[Che93] Yin-Wong Cheung. Long memory in foreign-exchange rates. Journal of Business & Economic Statistics, 11(1):93–101, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1993.10509935.

Chen:2007:MBC

[Che07a] Yi-Ting Chen. Moment-based copula tests for financial returns. Journal of Business & Economic Statistics, 25(4): 377–397, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500107000000115.

Chernov:2007:RRP

[Che07b] Mikhail Chernov. On the role of risk premia in volatility forecasting. Journal of Business & Economic Statistics, 25 (4):411–426, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000350.

Chen:2012:BSU

[Che12] Baoline Chen. A balanced system of U.S. industry accounts and distribution of the aggregate statistical discrepancy by industry. Journal of Business & Economic Statistics, 30(2):202–211, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2012.
669667.

Chen:2018:UAE

[Che18] Yi-Ting Chen. A unified approach to estimating and testing income distributions with grouped data. Journal of Business & Economic Statistics, 36(3):438–455, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2016.1194762.

Cheung:2020:LMF

[Che20] Ying Lun Cheung. Long memory factor model: On estimation of factor memories. Journal of Business & Economic Statistics, 40(2):756–769, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1867559.

Chen:2022:CPC

[Che22] Mingjing Chen. Circularly projected common factors for grouped data. Journal of Business & Economic Statistics, 41(2):636– 649, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2051520.

Chimerine:1985:SSM

[Chi85] Lawrence Chimerine. A supply-side miracle? Journal of Business & Economic Statistics, 3(2):101–103, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1985.10509436.

Conway:1988:ESA

[CHL88] Roger Conway, James Hrubovcak, and Michael LeBlanc. Estimating the structure of agricultural investment: a stochasticcoefficients approach. Journal of Business & Economic Statistics, 6(2):231–240, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509657.

Caner:2018:AEN

[CHL18] Mehmet Caner, Xu Han, and Yoonseok Lee. Adaptive elastic net GMM estimation with many invalid moment conditions: Simultaneous model and moment selection. Journal of Business & Economic Statistics, 36(1):24–46, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1129344.

Czajka:1992:PAD

[CHLR92] John L. Czajka, Sharon M. Hirabayashi, Roderick J. A. Little, and Donald B. Rubin. Projecting from advance data using propensity modeling: an application to income and tax statistics. Journal of Business & Economic Statistics, 10 (2):117–131, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509892.

Cheung:1997:CPC

[CHN97] Yin-Wong Cheung, Jia He, and Lilian K. Ng. Common predictable components in regional stock markets. Journal of Business & Economic Statistics, 15(1):35–42, 1997. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1997.10524684.

Cui:2022:RAH

[CHNY22] Guowei Cui, Kazuhiko Hayakawa, Shuichi Nagata, and Takashi Yamagata. A robust approach to heteroscedasticity, error serial correlation and slope heterogeneity in linear models with interactive effects for large panel data. Journal of Business & Economic Statistics, 41(3):862–875, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2077349.

Christiano:1990:LQA

[Chr90a] Lawrence J. Christiano. Linear-quadratic approximation and value-function iteration: a comparison. Journal of Business & Economic Statistics, 8(1):99–113, 1990. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1990.10509780.

Christiano:1990:SSG

[Chr90b] Lawrence J. Christiano. Solving the stochastic growth model by linear-quadratic approximation and by value-function iteration. Journal of Business & Economic Statistics, 8(1):23–26, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/07350015.1990.10509768.

Christiano:1992:SBG

[Chr92] Lawrence J. Christiano. Searching for a break in GNP. Journal of Business & Economic Statistics, 10(3):237–250, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1992.10509903.

Christiano:1996:SSS

[Chr96] Lawrence J. Christiano. Special section on small-sample properties of generalized method of moments (GMM): Associate Editor’s introduction. Journal of Business & Economic Statistics, 14(3):261, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524655.

Christiano:2007:C

[Chr07] Lawrence J. Christiano. Comment. Journal of Business & Economic Statistics, 25(2):143–151, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500107000000061. See [DSSW07a,

DSSW07b].

Card:2004:MMC

[CHSS04] David Card, Andrew K. G. Hildreth, and Lara D. Shore-Sheppard. The measurement of medicaid coverage in the SIPP. Journal of Business & Economic Statistics, 22(4):410–420, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500104000000208.

Carvalho:2007:NCC

[CHT07] Vasco Carvalho, Andrew Harvey, and Thomas Trimbur. A note on common cycles, common trends, and convergence. Journal of Business & Economic Statistics, 25(1):12–20, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000431.

Chan:2017:GEI

[CHT17] Kung-Sik Chan, Bruce E. Hansen, and Allan Timmermann. Guest Editors’ introduction: Regime switching and threshold models. Journal of Business & Economic Statistics, 35(2):159–161, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1236521.

Chen:2020:SFM

[CHW20] Yi-Ting Chen, Yu-Chin Hsu, and Hung-Jen Wang. A stochastic frontier model with endogenous treatment status and mediator. Journal of Business & Economic Statistics, 38(2):243–256, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1497504.

Clements:1987:MIS

[CI87] Kenneth W. Clements and H. Y. Izan. The measurement of inflation: a stochastic approach. Journal of Business & Economic Statistics, 5(3):339–350, 1987. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1987.10509598.

Chamberlain:2003:NAB

[CI03] Gary Chamberlain and Guido W. Imbens. Nonparametric applications of Bayesian inference. Journal of Business & Economic Statistics, 21(1):12–18, 2003. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102288618711.

Chan:2017:TEG

[CILY17] Ngai Hang Chan, Ching-Kang Ing, Yuanbo Li, and Chun Yip Yau. Threshold estimation via group orthogonal greedy algorithm. Journal of Business & Economic Statistics, 35(2):334–345, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1064820.

Chen:2023:CIC

[CJJ^{+}23] Rong Chen, Yuanyuan Ji, Guolin Jiang, Han Xiao, Ruoqing Xie, and Pingfang Zhu. Composite index construction with expert opinion. Journal of Business & Economic Statistics, 41(1):67– 79, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.2000418.

Canto:1986:REK

[CJW86] Victor A. Canto, Douglas H. Joines, and Robert I. Webb. The revenue effects of the Kennedy and Reagan tax cuts: Some time series estimates. Journal of Business & Economic Statistics, 4(3):281–288, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509525.

Chen:2018:TCM

[CJZZ18] Tao Chen, Yuanyuan Ji, Yahong Zhou, and Pingfang Zhu. Testing conditional mean independence under symmetry. Journal of Business & Economic Statistics, 36(4):615–627, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2016.1219263.

Clements:2003:BCA

[CK03] Michael P. Clements and Hans-Martin Krolzig. Business cycle asymmetries. Journal of Business & Economic Statistics, 21(1):196–211, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618892.

Choi:2008:RNT

[CK08] Hwan-Sik Choi and Nicholas M. Kiefer. Robust nonnested testing and the demand for money. Journal of Business & Economic Statistics, 26(1):9–17, 2008. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000179.

Chalak:2019:MEP

[CK19] Karim Chalak and Daniel Kim. Measurement error without the proxy exclusion restriction. Journal of Business & Economic Statistics, 39(1):200–216, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2019.1617156.

Chan:2022:PAD

[CK22] Marc K. Chan and Simon S. Kwok. The PCDID approach: Difference-in-differences when trends are potentially unparallel and stochastic. Journal of Business & Economic Statistics, 40 (3):1216–1233, 2022. CODEN ???? ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2021.1914636.

Creal:2011:DMH

[CKL11] Drew Creal, Siem Jan Koopman, and André Lucas. A dynamic multivariate heavy-tailed model for time-varying volatilities and correlations. Journal of Business & Economic Statistics, 29(4):552–563, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2011.10070.

Chan:2012:TVD

[CKLGS12] Joshua C. C. Chan, Gary Koop, Roberto Leon-Gonzalez, and Rodney W. Strachan. Time varying dimension models. Journal of Business & Economic Statistics, 30(3):358–367, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2012.663258.

Chiang:2022:MCR

[CKMS22] Harold D. Chiang, Kengo Kato, Yukun Ma, and Yuya Sasaki. Multiway cluster robust double/debiased machine learning. Journal of Business & Economic Statistics, 40(3):1046–1056, 2022.

CODEN ???? ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1895815.

Chan:2013:NMT

[CKP13] Joshua C. C. Chan, Gary Koop, and Simon M. Potter. A new model of trend inflation. Journal of Business & Economic Statistics, 31(1):94–106, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.741549.

Callen:1985:FER

[CKY85] Jeffrey L. Callen, Clarence C. Y. Kwan, and Patrick C. Y. Yip. Foreign-exchange rate dynamics: an empirical study using maximum entropy spectral analysis. Journal of Business & Economic Statistics, 3(2):149–155, 1985. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1985.10509443.

Cartwright:1987:TAE

[CL87] Phillip A. Cartwright and Cheng F. Lee. Time aggregation and the estimation of the market model: Empirical evidence. Journal of Business & Economic Statistics, 5(1): 131–143, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509568.

Cheung:1993:FCA

[CL93] Yin-Wong Cheung and Kon S. Lai. A fractional cointegration analysis of purchasing power parity. Journal of Business & Economic Statistics, 11(1):103–112, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509936.

Cheung:1995:LOC

[CL95] Yin-Wong Cheung and Kon S. Lai. Lag order and critical values of the augmented Dickey–Fuller test. Journal of Business & Economic Statistics, 13(3):277–280, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524601.

Chen:2017:AFA

[CL17] Ying Chen and Bo Li. An adaptive functional autoregressive forecast model to predict electricity price curves. Journal of Business & Economic Statistics, 35(3):371–388, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1092976.

Clayton:1994:CBE

[Cla94] Howard R. Clayton. A combined bound for errors in auditing based on Hoeffding’s inequality and the bootstrap. Journal of Business & Economic Statistics, 12(4):437–448, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1994.10524566.

Clark:1996:SSP

[Cla96] Todd E. Clark. Small-sample properties of estimators of nonlinear models of covariance structure. Journal of Business & Economic Statistics, 14(3):367–373, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524662.

Clark:2011:RTD

[Cla11] Todd E. Clark. Real-time density forecasts from Bayesian vector autoregressions with stochastic volatility. Journal of Business & Economic Statistics, 29(3):327–341, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2010.09248.

Clements:1985:SNZ

[Cle85] R. T. Clements. Savings in new Zealand during inflationary times: Measurement, determinants, and implications. Journal of Business & Economic Statistics, 3(3):188–208, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1985.10509451.

Cleveland:1998:C

[Cle98] William P. Cleveland. Comment. Journal of Business & Economic Statistics, 16(2):153–155, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524744.

Clements:2014:FUE

[Cle14] Michael P. Clements. Forecast uncertainty- Ex Ante and Ex Post: U.S. inflation and output growth. Journal of Business & Economic Statistics, 32(2):206–216, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2013.859618.

Clements:2017:AMU

[Cle17] Michael P. Clements. Assessing macro uncertainty in real-time when data are subject to revision. Journal of Business & Economic Statistics, 35(3):420–433, 2017. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2015.1081596.

Christofides:2003:RTS

[CLLM03] Louis N. Christofides, Qi Li, Zhenjuan Liu, and Insik Min. Recent two-stage sample selection procedures with an application to the gender wage gap. Journal of Business & Economic Statistics, 21(3):396–405, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619043.

Chen:2020:NMD

[CLQ20] Xuerong Chen, Denis Heng-Yan Leung, and Jing Qin. Nonignorable missing data, single index propensity score and profile synthetic distribution function. Journal of Business & Economic Statistics, 40(2):705–717, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1860065.

Christensen:2014:DCB

[CLR14] Jens H. E. Christensen, Jose A. Lopez, and Glenn D. Rudebusch. Do central bank liquidity facilities affect interbank lending rates? Journal of Business & Economic Statistics, 32(1):136–151, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.858631.

Calvi:2022:LMM

[CLT22] Rossella Calvi, Arthur Lewbel, and Denni Tommasi. LATE with missing or mismeasured treatment. Journal of Business & Economic Statistics, 40(4):1701–1717, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1970573.

Chan:1988:ICR

[CM88] M. W. Luke Chan and Dean C. Mountain. The interactive and causal relationships involving precious metal price movements an analysis of the gold and Silver markets. Journal of Business & Economic Statistics, 6(1):69–77, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509638.

Campbell:1990:PIC

[CM90] John Y. Campbell and N. Gregory Mankiw. Permanent income, current income, and consumption. Journal of Business & Economic Statistics, 8(3):265–279, 1990. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1990.10509798.

Colledge:1993:QMD

[CM93] Michael Colledge and Mary March. Quality management: Development of a framework for a statistical agency. Journal of Business & Economic Statistics, 11(2):157–165, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509944.

Cushing:1996:PSM

[CM96] Matthew J. Cushing and Mary G. McGarvey. The persistence of shocks to macroeconomic time series: Some evidence from economic theory. Journal of Business & Economic Statistics, 14(2):179–187, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524644.

Chan:2002:CJD

[CM02] Wing H. Chan and John M. Maheu. Conditional jump dynamics in stock market returns. Journal of Business & Economic Statistics, 20(3):377–389, 2002. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102288618513.

Clark:2009:TEP

[CM09] Todd E. Clark and Michael W. McCracken. Tests of equal predictive ability with real-time data. Journal of Business & Economic Statistics, 27(4):441–454, 2009. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.07204.

Clark:2012:RCC

[CM12] Todd E. Clark and Michael W. McCracken. Reality checks and comparisons of nested predictive models. Journal of Business & Economic Statistics, 30(1):53–66, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1198/jbes.2011.10278.

Cornea-Madeira:2019:BHU

[CMHM19] Adriana Cornea-Madeira, Cars Hommes, and Domenico Massaro. Behavioral heterogeneity in U.S. inflation dynamics. Journal of Business & Economic Statistics, 37(2):288–300, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1321548.

Camba-Mendez:2003:TRR

[CMKSW03] Gonzalo Camba-Mendez, George Kapetanios, Richard J. Smith, and Martin R. Weale. Tests of rank in reduced rank regression models. Journal of Business & Economic Statistics, 21 (1):145–155, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618847.

Cattaneo:2023:CDH

[CMM23] Matias D. Cattaneo, Xinwei Ma, and Yusufcan Masatlioglu. Context-dependent heterogeneous preferences: a comment on barseghyan and molinari (2023). Journal of Business & Economic Statistics, 41(4):1030–1034, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2023.2216740. See [BM23b].

Camponovo:2020:ELH

[CMO20] Lorenzo Camponovo, Yukitoshi Matsushita, and Taisuke Otsu. Empirical likelihood for high frequency data. Journal of Business & Economic Statistics, 38(3):621–632, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2018.1549051.

Christopoulos:2019:DET

[CMT19] Dimitris Christopoulos, Peter McAdam, and Elias Tzavalis. Dealing with endogeneity in threshold models using copulas. Journal of Business & Economic Statistics, 39(1):166–178, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1647213.

Carr:2023:RTM

[CMW23] Michael D. Carr, Robert A. Moffitt, and Emily E. Wiemers. Reconciling trends in male earnings volatility: Evidence from the SIPP survey and administrative data. Journal of Business & Economic Statistics, 41(1):26–32, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2126845.

Cavaliere:2020:BNA

[CNR20] Giuseppe Cavaliere, Heino Bohn Nielsen, and Anders Rahbek. Bootstrapping noncausal autoregressions: With applications to explosive bubble modeling. Journal of Business & Economic Statistics, 38(1):55–67, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1448830.

Cavaliere:2020:AIH

[CNT20] Giuseppe Cavaliere, Morten Ørregaard Nielsen, and A. M. Robert Taylor. Adaptive inference in heteroscedastic fractional time series models. Journal of Business & Economic Statistics, 40(1):50–65, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1773275.

Cox:1999:CSD

[CO99] James C. Cox and Ronald L. Oaxaca. Can supply and demand parameters be recovered from data generated by market institutions? Journal of Business & Economic Statistics, 17(3):285–297, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524818.

Cederburg:2019:URR

[CO19] Scott Cederburg and Michael S. O’Doherty. Understanding the risk-return relation: The aggregate wealth proxy actually matters. Journal of Business & Economic Statistics, 37(4):721–735, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1419140.

Cochrane:1989:RLE

[Coc89] John H. Cochrane. The return of the liquidity effect: a study of the short-run relation between money growth and interest rates. Journal of Business & Economic Statistics, 7 (1):75–83, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509715.

Coleman:1990:SSG

[Col90] Wilbur John Coleman II. Solving the stochastic growth model by policy-function iteration. Journal of Business & Economic Statistics, 8(1):27–29, 1990. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1990.10509769.

Collins:1991:PTB

[Col91] Sean Collins. Prediction techniques for Box–Cox regression models. Journal of Business & Economic Statistics, 9(3): 267–277, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509852.

Coleman:1993:SND

[Col93] Wilbur John Coleman. Solving nonlinear dynamic models on parallel computers. Journal of Business & Economic Statistics, 11(3):325–330, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509960.

Considine:1990:SCV

[Con90] Timothy J. Considine. Symmetry constraints and variable returns to scale in logit models. Journal of Business & Economic Statistics, 8(3):347–353, 1990. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1990.10509805.

Cook:1988:C

[Coo88] R. Dennis Cook. Comment. Journal of Business & Economic Statistics, 6(4):429–432, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509685.

Cooper:1998:MRU

[Coo98] Suzanne J. Cooper. Multiple regimes in U.S. output fluctuations. Journal of Business & Economic Statistics, 16 (1):92–100, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524738.

Chintagunta:1998:EID

[CP98] Pradeep K. Chintagunta and Alok R. Prasad. An empirical investigation of the “dynamic McFadden” model of purchase timing and brand choice: Implications for market structure. Journal of Business & Economic Statistics, 16(1):2–12, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524730.

Chauvet:2008:CRT

[CP08] Marcelle Chauvet and Jeremy Piger. A comparison of the realtime performance of business cycle dating methods. Journal of Business & Economic Statistics, 26(1):42–49, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500107000000296.

Crossley:2010:CSS

[CP10] Thomas F. Crossley and Krishna Pendakur. The common-scaling social cost-of-living index. Journal of Business & Economic Statistics, 28(4):523–538, 2010. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.06139.

Clinet:2019:DSH

[CP19] Simon Clinet and Yoann Potiron. Disentangling sources of high frequency market microstructure noise. Journal of Business & Economic Statistics, 39(1):18–39, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2019.1617158.

Chao:2015:CCM

[CPDH15] Shih-Kang Chao, Katharina Proksch, Holger Dette, and Wolfgang Karl Härdle. Confidence corridors for multivariate generalized quantile regression. Journal of Business & Economic Statistics, 35(1):70–85, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1054493.

Cho:2018:PKS

[CPP18] Jin Seo Cho, Myung-Ho Park, and Peter C. B. Phillips. Practical Kolmogorov–Smirnov testing by minimum distance applied to measure top income shares in Korea. Journal of Business & Economic Statistics, 36(3):523–537, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2016.1200983.

Conway:1983:RRF

[CR83] Delores A. Conway and Harry V. Roberts. Reverse regression, fairness, and employment discrimination. Journal of Business & Economic Statistics, 1(1):75–85, 1983. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07350015.1983.10509326. See comments [FG84,

Mil84] and rejoinder [CR84].

Conway:1984:RCR

[CR84] Delores A. Conway and Harry V. Roberts. Rejoinder to comments on “Reverse Regression, Fairness, and Employment Discrimination”. Journal of Business & Economic Statistics, 2 (2):126–139, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509381. See [FG84, Mil84, CR83].

Conway:1988:R

[CR88a] Delores A. Conway and Marc R. Reinganum. Reply. Journal of Business & Economic Statistics, 6(1):24–28, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1988.10509633.

See [CR88b, fC88, Job88, Sta88a, Bro88].

Conway:1988:SFS

[CR88b] Delores A. Conway and Marc R. Reinganum. Stable factors in security returns: Identification using cross-validation. Journal of Business & Economic Statistics, 6(1):1–15, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509628. See com-

ments [fC88, Job88, Sta88a, Bro88] and reply [CR88a].

Cecchetti:2001:SEU

[CR01] Stephen G. Cecchetti and Robert W. Rich. Structural estimates of the U.S. sacrifice ratio. Journal of Business & Economic Statistics, 19(4):416–427, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
07350010152596664.

Corsi:2012:DTV

[CR12] Fulvio Corsi and Roberto Renò. Discrete-time volatility forecasting with persistent leverage effect and the link with continuoustime volatility modeling. Journal of Business & Economic Statistics, 30(3):368–380, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.663261.

Carrasco:2016:SIF

[CR16a] Marine Carrasco and Barbara Rossi. In-sample inference and forecasting in misspecified factor models. Journal of Business & Economic Statistics, 34(3):313–338, 2016. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2016.1186029. See comments

[Sto16, Gia16, CH16, Swa16] and rejoinder [CR16b].

Carrasco:2016:RSI

[CR16b] Marine Carrasco and Barbara Rossi. Rejoinder: In-sample inference and forecasting in misspecified factor models. Journal of Business & Economic Statistics, 34(3):353–356, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.

1191500. See [CR16a].

Cramer:1986:VTC

[Cra86] J. S. Cramer. The volume of transactions and the circulation of money in the United States, 1950–1979. Journal of Business & Economic Statistics, 4(2):225–232, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509517.

Creal:2017:CNG

[Cre17] Drew D. Creal. A class of non-Gaussian state space models with exact likelihood inference. Journal of Business & Economic Statistics, 35(4):585–597, 2017. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1092977.

Chatterjee:1985:BR

[CRL^{+}85] Sangit Chatterjee, James B. Ramsey, Robert Litterman, Samprit Chatterjee, and Roman Frydman. Book reviews. Journal of Business & Economic Statistics, 3(4):410–414, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1985.10509481.

Croushore:2012:C

[Cro12] Dean Croushore. Comment. Journal of Business & Economic Statistics, 30(1):17–20, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.634340. See [PT12a, PT12b].

Chatterjee:1986:BR

[CRTO86] Samprit Chatterjee, M. R. Rao, George Thomas, and Alan J. Oppenheim. Book reviews. Journal of Business & Economic Statistics, 4(4):491–493, 1986. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1986.10509548.

Clapp:2019:RAC

[CRZ19] John M. Clapp, Stephen L. Ross, and Tingyu Zhou. Retail agglomeration and competition externalities: Evidence from openings and closings of multiline department stores in the U.S. Journal of Business & Economic Statistics, 37(1):81–96, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1272460.

Chavas:1986:SAD

[CS86] Jean-Paul Chavas and Kathleen Segerson. Singularity and autoregressive disturbances in linear logit models. Journal of Business & Economic Statistics, 4(2):161–169, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509511.

Coppejans:2005:KEA

[CS05] Mark Coppejans and Holger Sieg. Kernel estimation of average derivatives and differences. Journal of Business & Economic Statistics, 23(2):211–225, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000497.

Chao:2009:C

[CS09] John C. Chao and Norman R. Swanson. Comment. Journal of Business & Economic Statistics, 27(3):316–318, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1198/jbes.2009.08264.

See [KM09b, KM09a].

Conrad:2020:TOM

[CS20] Christian Conrad and Melanie Schienle. Testing for an omitted multiplicative long-term component in GARCH models. Journal of Business & Economic Statistics, 38(2):229–242, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1482759.

Cecchetti:1994:VRT

[CsL94] Stephen G. Cecchetti and Pok sang Lam. Variance-ratio tests: Small-sample properties with an application to international output data. Journal of Business & Economic Statistics, 12 (2):177–186, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10510006.

Casarin:2018:BMS

[CST18] Roberto Casarin, Domenico Sartore, and Marco Tronzano. A Bayesian Markov-switching correlation model for contagion analysis on exchange rate markets. Journal of Business & Economic Statistics, 36(1):101–114, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1137757.

Chang:2022:CHM

[CSZ22] Jinyuan Chang, Zhentao Shi, and Jia Zhang. Culling the herd of moments with penalized empirical likelihood. Journal of Business & Economic Statistics, 41(3):791–805, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2071903.

Chen:2017:RCE

[CSZZ17] Songnian Chen, Jichun Si, Hanghui Zhang, and Yahong Zhou. Root-N consistent estimation of a panel data binary response model with unknown correlated random effects. Journal of Business & Economic Statistics, 35(4):559–571, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1130635.

Chen:1990:RLS

[CT90] Chung Chen and George C. Tiao. Random level-shift time series models, ARIMA approximations, and level-shift detection. Journal of Business & Economic Statistics, 8(1):83–97, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1990.10509779.

Cameron:1993:TIP

[CT93] A. Colin Cameron and Pravin K. Trivedi. Tests of independence in parametric models with applications and illustrations. Journal of Business & Economic Statistics, 11 (1):29–43, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509931.

Carrington:1997:MSS

[CT97] William J. Carrington and Kenneth R. Troske. On measuring segregation in samples with small units. Journal of Business & Economic Statistics, 15(4):402–409, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524718.

Chung:2001:TTZ

[CT01] Chae-Shick Chung and George Tauchen. Testing target-zone models using efficient method of moments. Journal of Business & Economic Statistics, 19(3):255–277, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500101681019891.

Capistran:2009:FCE

[CT09] Carlos Capistrán and Allan Timmermann. Forecast combination with entry and exit of experts. Journal of Business & Economic Statistics, 27(4):428–440, 2009. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.07211.

Candelon:2016:NTG

[CT16] Bertrand Candelon and Sessi Tokpavi. A nonparametric test for Granger causality in distribution with application to financial contagion. Journal of Business & Economic Statistics, 34(2):240–253, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1026774.

Chua:2018:BAM

[CT18] Chew Lian Chua and Sarantis Tsiaplias. A Bayesian approach to modeling time-varying cointegration and cointegrating rank. Journal of Business & Economic Statistics, 36(2):267–277, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1166117.

Chiang:2022:ELU

[CT22] Harold D. Chiang and Bing Yang Tan. Empirical likelihood and uniform convergence rates for dyadic kernel density estimation. Journal of Business & Economic Statistics, 41(3):906– 914, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2080684.

Cui:2020:LEI

[Cui20] Wenhao Cui. Laplace estimator of integrated volatility when sampling times are endogenous. Journal of Business & Economic Statistics, 40(2):651–663, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1855185.

Cowell:2006:DDT

[CVF06] Frank A. Cowell and Maria-Pia Victoria-Feser. Distributional dominance with trimmed data. Journal of Business & Economic Statistics, 24(3):291–300, 2006. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500105000000207.

Clemen:1986:CEF

[CW86a] Robert T. Clemen and Robert L. Winkler. Combining economic forecasts. Journal of Business & Economic Statistics, 4(1):39–46, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509492.

Cupingood:1986:SAT

[CW86b] Leonard A. Cupingood and William W. S. Wei. Seasonal adjustment of time series using one-sided filters. Journal of Business & Economic Statistics, 4(4):473–484, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509546.

Chu:1992:DTC

[CW92] Chia-Shang James Chu and Halbert White. A direct test for changing trend. Journal of Business & Economic Statistics, 10(3):289–299, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509906.

Cameron:1996:SMC

[CW96] A. Colin Cameron and Frank A. G. Windmeijer. R-squared measures for count data regression models with applications to health-care utilization. Journal of Business & Economic Statistics, 14(2):209–220, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524648.

Chib:2001:MCM

[CW01] Siddhartha Chib and Rainer Winkelmann. Markov chain Monte Carlo analysis of correlated count data. Journal of Business & Economic Statistics, 19(4):428–435, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/07350010152596673.

Collard-Wexler:2008:C

[CW08] Allan Collard-Wexler. Comment. Journal of Business & Economic Statistics, 26(3):303–307, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500108000000150. See [ALT08a,

ALT08b].

Chen:2020:DSF

[CWW20] Likai Chen, Weining Wang, and Wei Biao Wu. Dynamic semiparametric factor model with structural breaks. Journal of Business & Economic Statistics, 39(3):757–771, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1730857.

Caner:2014:AEN

[CZ14] Mehmet Caner and Hao Helen Zhang. Adaptive elastic net for generalized methods of moments. Journal of Business & Economic Statistics, 32(1):30–47, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.836104.

Cui:2018:MLC

[CZ18] Qiurong Cui and Zhengjun Zhang. Max-linear competing factor models. Journal of Business & Economic Statistics, 36(1):62– 74, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2015.1137761.

Chib:2019:WFR

[CZ19] Siddhartha Chib and Xiaming Zeng. Which factors are risk factors in asset pricing? A model scan framework. Journal of Business & Economic Statistics, 38(4):771–783, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2019.1573684.

Chen:2022:HDE

[CZZ22] Xin Chen, Jia Zhang, and Wang Zhou. High-dimensional elliptical sliced inverse regression in non-Gaussian distributions. Journal of Business & Economic Statistics, 40(3):1204–1215, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1910041.

deAlba:1988:DFB

[dA88] Enrique de Alba. Disaggregation and forecasting: a Bayesian analysis. Journal of Business & Economic Statistics, 6(2): 197–206, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509654.

Dagum:1987:MEA

[Dag87] Camiio Dagum. Measuring the economic affluence between populations of income receivers. Journal of Business & Economic Statistics, 5(1):5–12, 1987. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1987.10509554.

Danaher:1992:MCM

[Dan92] Peter J. Danaher. A Markov-chain model for multivariate magazine-exposure distributions. Journal of Business & Economic Statistics, 10(4):401–407, 1992. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1992.10509915.

Danielsson:1994:C

[Dan94] Jon Danielsson. Comment. Journal of Business & Economic Statistics, 12(4):393–395, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10524555.

Das:1991:EFC

[Das91] Sanghamitra Das. Estimation of fuel coefficients of cement production: a fixed-effects approach to nonlinear regression. Journal of Business & Economic Statistics, 9(4): 469–474, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509874.

Das:2008:SAG

[Das08] Mitali Das. Semiparametric analysis with grouped dependent variables and application to physicians’ provision of charity care. Journal of Business & Economic Statistics, 26 (4):446–459, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500108000000079.

Davidson:2004:MMP

[Dav04] James Davidson. Moment and memory properties of linear conditional heteroscedasticity models, and a new model. Journal of Business & Economic Statistics, 22(1):16–29, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500103288619359.

Dagum:2008:HSR

[DB08] Estela Bee Dagum and Silvia Bianconcini. The Henderson smoother in reproducing kernel Hilbert space. Journal of Business & Economic Statistics, 26(4):536–545, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500107000000322.

Dirick:2019:MEF

[DBCB19] Lore Dirick, Tony Bellotti, Gerda Claeskens, and Bart Baesens. Macro-economic factors in credit risk calculations: Including timevarying covariates in mixture cure models. Journal of Business & Economic Statistics, 37(1):40–53, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2016.1260471.

deBrouwer:1998:MIA

[dBE98] Gordon de Brouwer and Neil R. Ericsson. Modeling inflation in Australia. Journal of Business & Economic Statistics, 16 (4):433–449, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524783.

denButter:1990:SAU

[dBM90] F. A. G. den Butter and T. J. Mourik. Seasonal adjustment using structural time series models: an application and a comparison with the census X-11 method. Journal of Business & Economic Statistics, 8(4):385–394, 1990. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1990.10509809.

Dixon:1984:RCE

[DBS84] Bruce L. Dixon, Marvin T. Batte, and Steven T. Sonka. Random coefficients estimation of average total product costs for multiproduct firms. Journal of Business & Economic Statistics, 2(4):360–366, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509409.

Dhrymes:1985:BR

[DC85] Phoebus J. Dhrymes and Samprit Chatterjee. Book reviews. Journal of Business & Economic Statistics, 3(2): 174–178, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509448.

Delaney:1986:UBC

[DC86] Nancy Jo Delaney and Sangit Chatterjee. Use of the bootstrap and cross-validation in ridge regression. Journal of Business & Economic Statistics, 4(2):255–262, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509520.

Dagum:1986:BR

[DCW^{+}86] Camilo Dagum, Samprit Chatterjee, Berry Wilson, Glenn Heller, and Poduri S. R. S. Rao. Book reviews. Journal of Business & Economic Statistics, 4(3):391–394, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509534.

Delgado:2013:CSD

[DE13] Miguel A. Delgado and Juan Carlos Escanciano. Conditional stochastic dominance testing. Journal of Business & Economic Statistics, 31(1):16–28, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.723556.

DeVos:2019:BCC

[DE19] Ignace De Vos and Gerdie Everaert. Bias-corrected common correlated effects pooled estimation in dynamic panels. Journal of Business & Economic Statistics, 39(1):294–306, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1654879.

Dehejia:2003:WTR

[Deh03] Rajeev H. Dehejia. Was there a Riverside miracle? A hierarchical framework for evaluating programs with grouped data. Journal of Business & Economic Statistics, 21(1):1–11, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500102288618702.

Delong:1986:C

[Del86] David M. Delong. Comment. Journal of Business & Economic Statistics, 4(4):417–418, 1986. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1986.10509538.

Deschenes:2007:EEF

[Des07] Olivier Deschênes. Estimating the effects of family background on the return to schooling. Journal of Business & Economic Statistics, 25(3):265–277, 2007. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500106000000567.

Devereux:2007:IEV

[Dev07] Paul J. Devereux. Improved errors-in-variables estimators for grouped data. Journal of Business & Economic Statistics, 25 (3):278–287, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000189.

DAmico:2011:FSM

[DF11] Stefania D’Amico and Mira Farka. The Fed and the stock market: an identification based on intraday futures data. Journal of Business & Economic Statistics, 29(1):126–137, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.08019.

Duan:2015:DTM

[DF15] Jin-Chuan Duan and Andras Fulop. Density-tempered marginalized sequential Monte Carlo samplers. Journal of Business & Economic Statistics, 33(2):192–202, 2015. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2014.940081.

Diewert:2020:SBM

[DF20] W. Erwin Diewert and Kevin J. Fox. Substitution bias in multilateral methods for CPI construction. Journal of Business & Economic Statistics, 40(1):355–369, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1816176.

Dufour:2019:PTC

[DFK19] Jean-Marie Dufour, Emmanuel Flachaire, and Lynda Khalaf. Permutation tests for comparing inequality measures. Journal of Business & Economic Statistics, 37(3):457–470, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1371027.

Donald:2011:LGR

[DFP11] Stephen G. Donald, Natércia Fortuna, and Vladas Pipiras. Local and global rank tests for multivariate varying-coefficient models. Journal of Business & Economic Statistics, 29(2):295–306, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2010.07303.

Dhyne:2011:LPA

[DFPS11] Emmanuel Dhyne, Catherine Fuss, M. Hashem Pesaran, and Patrick Sevestre. Lumpy price adjustments: a microeconometric analysis. Journal of Business & Economic Statistics, 29 (4):529–540, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2011.09066.

Davis:1996:MSM

[DG96] George C. Davis and Jean Gauger. Measuring substitution in monetary-asset demand systems. Journal of Business & Economic Statistics, 14(2):203–208, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524647.

Duggan:1999:EEC

[DG99] James E. Duggan and Robert Gillingham. The effect of errors in the CPI on social security finances. Journal of Business & Economic Statistics, 17(2):161–169, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524806.

Durham:2002:NTM

[DG02] Garland B. Durham and A. Ronald Gallant. Numerical techniques for maximum likelihood estimation of continuous-time diffusion processes. Journal of Business & Economic Statistics, 20(3):297–338, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618397.

Domenech:2006:EPO

[DG06] Rafael Doménech and Víctor Gómez. Estimating potential output, core inflation, and the NAIRU as latent variables. Journal of Business & Economic Statistics, 24(3):354–365, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500105000000315.

DiTraglia:2020:FEI

[DGJ20] Francis J. DiTraglia and Camilo García-Jimeno. A framework for eliciting, incorporating, and disciplining identification beliefs in linear models. Journal of Business & Economic Statistics, 39 (4):1038–1053, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1753528.

Dong:2020:VCP

[DGP20] Chaohua Dong, Jiti Gao, and Bin Peng. Varying-coefficient panel data models with nonstationarity and partially observed factor structure. Journal of Business & Economic Statistics, 39(3):700–711, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1721294.

Dubin:1988:EEA

[DH88] Jeffrey A. Dubin and Steven E. Henson. An engineering/econometric analysis of seasonal energy demand and conservation in the Pacific northwest. Journal of Business & Economic Statistics, 6(1):121–134, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509644.

Duncan:1989:AQH

[DH89] Greg J. Duncan and Daniel H. Hill. Assessing the quality of household panel data: The case of the panel study of income dynamics. Journal of Business & Economic Statistics, 7(4):441–452, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509756.

Demetrescu:2012:URT

[DH12] Matei Demetrescu and Christoph Hanck. Unit root testing in heteroscedastic panels using the Cauchy estimator. Journal of Business & Economic Statistics, 30(2):256–264, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2011.638839.

Demetrescu:2020:NPS

[DH20] Matei Demetrescu and Benjamin Hillmann. Nonlinear predictability of stock returns? parametric versus nonparametric inference in predictive regressions. Journal of Business & Economic Statistics, 40(1):382–397, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1819821.

Dimitriadis:2022:RQ

[DH22] Timo Dimitriadis and Roxana Halbleib. Realized quantiles. Journal of Business & Economic Statistics, 40(3):1346–1361, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1929249.

deHaan:2014:SDT

[dHK14] Jan de Haan and Frances Krsinich. Scanner data and the treatment of quality change in nonrevisable price indexes. Journal of Business & Economic Statistics, 32(3):341–358, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2014.880059.

Donald:2014:TUA

[DHL14] Stephen G. Donald, Yu-Chin Hsu, and Robert P. Lieli. Testing the unconfoundedness assumption via inverse probability weighted estimators of (L)ATT. Journal of Business & Economic Statistics, 32(3):395–415, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.888290.

denHaan:1990:SSG

[dHM90] Wouter J. den Haan and Albert Marcet. Solving the stochastic growth model by parameterizing expectations. Journal of Business & Economic Statistics, 8(1):31–34, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1990.10509770.

Deuchert:2019:DIE

[DHS19] Eva Deuchert, Martin Huber, and Mark Schelker. Direct and indirect effects based on difference-in-differences with an application to political preferences following the Vietnam draft lottery. Journal of Business & Economic Statistics, 37(4):710–720, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1419139.

DeLosSantos:2017:SLD

[DHW17] Babur De Los Santos, Ali Hortaçsu, and Matthijs R. Wildenbeest. Search with learning for differentiated products: Evidence from e-commerce. Journal of Business & Economic Statistics, 35(4):626–641, 2017. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1123633.

Damrongplasit:2010:DMS

[DHZ10] Kannika Damrongplasit, Cheng Hsiao, and Xueyan Zhao. Decriminalization and marijuana smoking prevalence: Evidence from Australia. Journal of Business & Economic Statistics, 28 (3):344–356, 2010. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.06129.

Dickey:1988:C

[Dic88] D. A. Dickey. Comment. Journal of Business & Economic Statistics, 6(4):410–412, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509681.

Diebold:1988:C

[Die88a] Francis X. Diebold. Comment. Journal of Business & Economic Statistics, 6(4):470–472, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509693.

Diebold:1988:SCC

[Die88b] Francis X. Diebold. Serial correlation and the combination of forecasts. Journal of Business & Economic Statistics, 6 (1):105–111, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509642.

Diebold:2006:C

[Die06] Francis X. Diebold. Comment. Journal of Business & Economic Statistics, 24(2):181–183, 2006. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500106000000143. See [HL06a, HL06b].

Diebold:2015:CPA

[Die15a] Francis X. Diebold. Comparing predictive accuracy, twenty years later: a personal perspective on the use and abuse of Diebold–Mariano tests. Journal of Business & Economic Statistics, 33(1):1, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.983236. See comments [Ino15, Wri15,

Kil15, HT15, Pat15] and rejoinder [Die15b].

Diebold:2015:R

[Die15b] Francis X. Diebold. Rejoinder. Journal of Business & Economic Statistics, 33(1):24, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.983237. See [Die15a].

DiezdeLosRios:2015:NLE

[Die15c] Antonio Diez de Los Rios. A new linear estimator for Gaussian dynamic term structure models. Journal of Business & Economic Statistics, 33(2):282–295, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.948176.

DeJong:1996:BAC

[DIW96] David N. DeJong, Beth Fisher Ingram, and Charles H. Whiteman. A Bayesian approach to calibration. Journal of Business & Economic Statistics, 14(1):1–9, 1996. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1996.10524625.

deJong:1987:RED

[dJ87] Piet de Jong. Rational economic data revisions. Journal of Business & Economic Statistics, 5(4):539–548, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509622.

deJong:2000:TSC

[dJ00] Frank de Jong. Time series and cross-section information in affine term-structure models. Journal of Business & Economic Statistics, 18(3):300–314, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524872.

Daniel:1988:BR

[DJA88] T. E. Daniel, Jeffrey Jarrett, and Donald A. Anderson. Book reviews. Journal of Business & Economic Statistics, 6(2): 283–286, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509662.

Diebold:2000:URT

[DK00] Francis X. Diebold and Lutz Kilian. Unit-root tests are useful for selecting forecasting models. Journal of Business & Economic Statistics, 18(3):265–273, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524869.

Durlauf:2017:ST

[DKT17] Steven N. Durlauf, Andros Kourtellos, and Chih Ming Tan. Status traps. Journal of Business & Economic Statistics, 35 (2):265–287, 2017. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1189339.

Dagum:1984:C

[DL84] Estela Bee Dagum and Normand J. D. Laniel. Comment. Journal of Business & Economic Statistics, 2(4):328–332, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1984.10509402.

Dagum:1987:RTC

[DL87] Estela Bee Dagum and Normand Laniel. Revisions of trend-cycle estimators of moving average seasonal adjustment methods. Journal of Business & Economic Statistics, 5(2):177–189, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1987.10509576.

Duncan:1989:RDM

[DL89] George Duncan and Diane Lambert. The risk of disclosure for microdata. Journal of Business & Economic Statistics, 7 (2):207–217, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509729.

deLeeuw:1990:RUG

[dL90] Frank de Leeuw. The reliability of U.S. gross national product. Journal of Business & Economic Statistics, 8(2): 191–203, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509790.

Dasgupta:1992:CSA

[DL92] Susmita Dasgupta and Kajal Lahiri. A comparative study of alternative methods of quantifying qualitative survey responses using NAPM data. Journal of Business & Economic Statistics, 10(4):391–400, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509914.

deLima:1998:NNS

[dL98] Pedro J. F. de Lima. Nonlinearities and nonstationarities in stock returns. Journal of Business & Economic Statistics, 16 (2):227–236, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524756.

Dennis:2004:C

[DL04] Richard Dennis and Jose A. Lopez. Comment. Journal of Business & Economic Statistics, 22(2):165–169, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000037. See [PSW04a] and

rejoinder [PSW04b].

Debarsy:2020:BMA

[DL20] Nicolas Debarsy and James P. LeSage. Bayesian model averaging for spatial autoregressive models based on convex combinations of different types of connectivity matrices. Journal of Business & Economic Statistics, 40(2):547–558, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1840993.

Dunbar:2019:IRR

[DLP19] Geoffrey R. Dunbar, Arthur Lewbel, and Krishna Pendakur. Identification of random resource shares in collective households without preference similarity restrictions. Journal of Business & Economic Statistics, 39(2):402–421, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2019.1665532.

delaVina:1985:ESM

[dlV85] Lynda Y. de la Viña. Econometrics software for microcomputers. Journal of Business & Economic Statistics, 3 (1):89–91, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509430.

Dagum:1984:BIS

[DM84] Estela Bee Dagum and Marietta Morry. Basic issues on the seasonal adjustment of the Canadian consumer price index. Journal of Business & Economic Statistics, 2(3): 250–259, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509392.

DeCoster:1991:NMD

[DM91] Gregory P. DeCoster and Douglas W. Mitchell. Nonlinear monetary dynamics. Journal of Business & Economic Statistics, 9(4):455–461, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509872.

DeCoster:1994:R

[DM94a] Gregory P. DeCoster and Douglas W. Mitchell. Reply. Journal of Business & Economic Statistics, 12(1):136–137, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1994.10509998.

Durland:1994:DDT

[DM94b] J. Michael Durland and Thomas H. McCurdy. Durationdependent transitions in a Markov model of U.S. GNP growth. Journal of Business & Economic Statistics, 12(3):279–288, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/07350015.1994.10524543.

Diebold:1995:CPA

[DM95a] Francis X. Diebold and Roberto S. Mariano. Comparing predictive accuracy. Journal of Business & Economic Statistics, 13(3):253–263, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524599.

Dorsey:1995:GAE

[DM95b] Robert E. Dorsey and Walter J. Mayer. Genetic algorithms for estimation problems with multiple optima, nondifferentiability, and other irregular features. Journal of Business & Economic Statistics, 13(1):53–66, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524579.

DeGooijer:1999:LRR

[DM99] Jan G. De Gooijer and Ian B. MacNeill. Lagged regression residuals and serial-correlation tests. Journal of Business & Economic Statistics, 17(2):236–247, 1999. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1999.10524813.

DeAlba:2001:FAS

[DM01] Enrique De Alba and Manuel Mendoza. Forecasting an accumulated series based on partial accumulation. Journal of Business & Economic Statistics, 19(1):95–102, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/07350010152472652.

Diebold:2002:CPA

[DM02] Francis X. Diebold and Robert S. Mariano. Comparing predictive accuracy. Journal of Business & Economic Statistics, 20(1):134–144, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102753410444.

Davidson:2010:WBT

[DM10] Russell Davidson and James G. MacKinnon. Wild bootstrap tests for IV regression. Journal of Business & Economic Statistics, 28(1):128–144, 2010. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.07221.

Duan:2016:DCL

[DM16] Jin-Chuan Duan and Weimin Miao. Default correlations and large-portfolio credit analysis. Journal of Business & Economic Statistics, 34(4):536–546, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1087855.

Duan:1983:CAM

[DMMN83] Naihua Duan, Willard G. Manning, Carl N. Morris, and Joseph P. Newhouse. A comparison of alternative models for the demand for medical care. Journal of Business & Economic Statistics, 1(2):115–126, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/

07350015.1983.10509330. See corrigendum [Ano84b].

Duan:1984:CBS

[DMMN84] Naihua Duan, Willard G. Manning, Carl N. Morris, and Joseph P. Newhouse. Choosing between the sample-selection model and the multi-part model. Journal of Business & Economic Statistics, 2(3):283–289, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509396.

DeLuca:2019:CUS

[DMP19] Giuseppe De Luca, Jan R. Magnus, and Franco Peracchi. Comments on “Unobservable Selection and Coefficient Stability: Theory and Evidence” and “Poorly Measured Confounders are More Useful on the Left Than on the Right”. Journal of Business & Economic Statistics, 37(2):217–222, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2019.1575743. See [Ost19,

PPS19].

Denton:1999:ATC

[DMS99] Frank T. Denton, Dean C. Mountain, and Byron G. Spencer. Age, trend, and cohort effects in a macro model of Canadian expenditure patterns. Journal of Business & Economic Statistics, 17(4):430–443, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524831.

Deb:2006:PIS

[DMT06] Partha Deb, Murat K. Munkin, and Pravin K. Trivedi. Private insurance, selection, and health care use. Journal of Business & Economic Statistics, 24(4):403–415, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000323.

deNew:2023:SRB

[dNS23] Sonja C. de New and Stefanie Schurer. Survey response behavior as a proxy for unobserved ability: Theory and evidence. Journal of Business & Economic Statistics, 41(1):197– 212, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.2008404.

Drost:1998:ETM

[DNW98] Feike C. Drost, Theo E. Nijman, and Bas J. M. Werker. Estimation and testing in models containing both jumps and conditional heteroscedasticity. Journal of Business & Economic Statistics, 16(2):237–243, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524757.

DaleN:1995:VES

[DO95] Jorgen DaléN and Esbjörn Ohlsson. Variance estimation in the Swedish consumer price index. Journal of Business & Economic Statistics, 13(3):347–356, 1995. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1995.10524608.

Dole:1999:CCS

[Dol99] David Dole. CoSmo: a constrained scatterplot smoother for estimating convex, monotonic transformations. Journal of Business & Economic Statistics, 17(4):444–455, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524832.

Dong:2019:RDD

[Don19] Yingying Dong. Regression discontinuity designs with sample selection. Journal of Business & Economic Statistics, 37 (1):171–186, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1302880.

Donoho:2022:DCF

[Don22] David Donoho. Data come first: Discussion of “Co-citation and Co-authorship Networks of Statisticians”. Journal of Business & Economic Statistics, 40(2):491, 2022. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2022.

2055356. See [JJKL21].

Dostie:2005:JTR

[Dos05] Benoit Dostie. Job turnover and the returns to seniority. Journal of Business & Economic Statistics, 23(2):192–199, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500104000000299.

Dickey:1987:DOD

[DP87] David A. Dickey and Sastry G. Pantula. Determining the order of differencing in autoregressive processes. Journal of Business & Economic Statistics, 5(4):455–461, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509614.

Dickey:2002:DOD

[DP02] David A. Dickey and Sastry G. Pantula. Determining the order of differencing in autoregressive processes. Journal of Business & Economic Statistics, 20(1):18–24, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102753410363.

Donaldson:2006:IFC

[DP06] David Donaldson and Krishna Pendakur. The identification of fixed costs from consumer behavior. Journal of Business & Economic Statistics, 24(3):255–265, 2006. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500106000000035.

Durham:2013:BSV

[DP13] Garland Durham and Yang-Ho Park. Beyond stochastic volatility and jumps in returns and volatility. Journal of Business & Economic Statistics, 31(1):107–121, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2013.747800.

DeWinne:2019:IFP

[DP19] Jasmien De Winne and Gert Peersman. The impact of food prices on conflict revisited. Journal of Business & Economic Statistics, 39(2):547–560, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1684301.

Dufour:2022:PMM

[DP22] Jean-Marie Dufour and Denis Pelletier. Practical methods for modeling weak VARMA processes: Identification, estimation and specification with a macroeconomic application. Journal of Business & Economic Statistics, 40(3):1140–1152, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1904960.

Dehejia:2019:LGE

[DPES19] Rajeev Dehejia, Cristian Pop-Eleches, and Cyrus Samii. From local to global: External validity in a fertility natural experiment. Journal of Business & Economic Statistics, 39 (1):217–243, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1639407.

Davison:2022:TRI

[DPS22] Anthony C. Davison, Simone A. Padoan, and Gilles Stupfler. Tail risk inference via expectiles in heavy-tailed time series. Journal of Business & Economic Statistics, 41(3):876–889, 2022. CO-DEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2022.2078332.

Jong:2011:DCR

[dRJH11] de Robert Jong and Ana María Herrera. Dynamic censored regression and the open market desk reaction function. Journal of Business & Economic Statistics, 29(2):228–237, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2010.07181.

Dolado:2020:LTJ

[DRV20] Juan J. Dolado, Heiko Rachinger, and Carlos Velasco. LM tests for joint breaks in the dynamics and level of a long-memory time series. Journal of Business & Economic Statistics, 40(2):629–650, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1855184.

Diebold:1990:PDB

[DS90] Francis X. Diebold and Steven A. Sharpe. Post-deregulation bankdeposit-rate pricing: The multivariate dynamics. Journal of Business & Economic Statistics, 8(3):281–291, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1990.10509799.

Demos:1998:EAC

[DS98] Antonis Demos and Enrique Sentana. An EM algorithm for conditionally heteroscedastic factor models. Journal of Business & Economic Statistics, 16(3):357–361, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524775.

Dufour:2013:FAV

[DS13] Jean-Marie Dufour and Dalibor Stevanović. Factor-augmented VARMA models with macroeconomic applications. Journal of Business & Economic Statistics, 31(4):491–506, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2013.818005.

DelNegro:2007:FNK

[DSSW07a] Marco Del Negro, Frank Schorfheide, Frank Smets, and Rafael Wouters. On the fit of new Keynesian models. Journal of Business & Economic Statistics, 25(2):123–143, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500107000000016. See

comments [Chr07, Gal07, Sim07, Fau07, Kil07] and rejoinder [DSSW07b].

DelNegro:2007:R

[DSSW07b] Marco Del Negro, Frank Schorfheide, Frank Smets, and Rafael Wouters. Rejoinder. Journal of Business & Economic Statistics, 25(2):159–162, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/

073500107000000070. See [DSSW07a, Chr07, Gal07, Sim07, Fau07, Kil07].

Driver:2006:CBT

[DTU06] Ciaran Driver, Paul Temple, and Giovanni Urga. Contrasts between types of assets in fixed investment equations as a way of testing real options theory. Journal of Business & Economic Statistics, 24(4):432–443, 2006. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500106000000062.

Dumitru:2012:IJF

[DU12] Ana-Maria Dumitru and Giovanni Urga. Identifying jumps in financial assets: a comparison between nonparametric jump tests. Journal of Business & Economic Statistics, 30(2):242–255, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2012.
663250.

Dua:1988:MFI

[Dua88] Pami Dua. Multiperiod forecasts of interest rates. Journal of Business & Economic Statistics, 6(3):381–384, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509675.

Dueker:1997:MSG

[Due97] Michael J. Dueker. Markov switching in GARCH processes and mean-reverting stock-market volatility. Journal of Business & Economic Statistics, 15(1):26–34, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524683.

Dueker:1999:CHQ

[Due99] Michael Dueker. Conditional heteroscedasticity in qualitative response models of time series: a Gibbs-sampling approach to the bank prime rate. Journal of Business & Economic Statistics, 17(4):466–472, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524834.

Dueker:2005:DFQ

[Due05] Michael Dueker. Dynamic forecasts of qualitative variables. Journal of Business & Economic Statistics, 23(1):96–104, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500104000000613.

Dufour:2009:C

[Duf09] Jean-Marie Dufour. Comment. Journal of Business & Economic Statistics, 27(3):318–321, 2009. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.08283. See [KM09b, KM09a].

Dunn:1983:D

[Dun83] Douglas M. Dunn. Discussion. Journal of Business & Economic Statistics, 1(1):5, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509315.

Durevall:1998:DCI

[Dur98] Dick Durevall. The dynamics of chronic inflation in Brazil, 1968–1985. Journal of Business & Economic Statistics, 16 (4):423–432, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524782.

Dustmann:2004:ASF

[DV04] Christian Dustmann and Arthur Van Soest. An analysis of speaking fluency of immigrants using ordered response models with classification errors. Journal of Business & Economic Statistics, 22(3):312–321, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000163.

Derrick:1985:EPA

[DW85] Frederick W. Derrick and John D. Wolken. The effects of price aggregation bias in systems of demand equations. Journal of Business & Economic Statistics, 3(4):325–331, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1985.10509464.

Diewert:1988:NQS

[DW88] W. E. Diewert and T. J. Wales. Normalized quadratic systems of consumer demand functions. Journal of Business & Economic Statistics, 6(3):303–312, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509667.

DeJong:1993:EMA

[DW93] David N. DeJong and Charles H. Whiteman. Estimating moving average parameters: Classical pileups and Bayesian posteriors. Journal of Business & Economic Statistics, 11 (3):311–317, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509958.

Drost:2004:SDM

[DW04] Feike C. Drost and Bas J. M. Werker. Semiparametric duration models. Journal of Business & Economic Statistics, 22(1):40–50, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619395.

Dette:2020:PLS

[DW20] Holger Dette and Weichi Wu. Prediction in locally stationary time series. Journal of Business & Economic Statistics, 40(1):370–381, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1819296.

Deltas:2004:SFC

[DZ04] George Deltas and Eleftherios Zacharias. Sampling frequency and the comparison between matched-model and hedonic regression price indexes. Journal of Business & Economic Statistics, 22(1):94–106, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619421.

Easton:1988:C

[Eas88] George S. Easton. Comment. Journal of Business & Economic Statistics, 6(4):432–441, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509686.

Engle:2006:TVD

[EC06] Robert Engle and Riccardo Colacito. Testing and valuing dynamic correlations for asset allocation. Journal of Business & Economic Statistics, 24(2):238–253, 2006. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500106000000017.

Edwards:2008:HRP

[Edw08] Ryan D. Edwards. Health risk and portfolio choice. Journal of Business & Economic Statistics, 26(4):472–485, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500107000000287.

Enders:1998:URT

[EG98] Walter Enders and C. W. J. Granger. Unit-root tests and asymmetric adjustment with an example using the term structure of interest rates. Journal of Business & Economic Statistics, 16(3):304–311, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524769.

Engle:1991:SAM

[EGR91] Robert F. Engle and Gloriá Gonzalez-Rivera. Semiparametric ARCH models. Journal of Business & Economic Statistics, 9(4):345–359, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509863.

Eichenbaum:1990:EMI

[EH90] Martin Eichenbaum and Lars Peter Hansen. Estimating models with intertemporal substitution using aggregate time series data. Journal of Business & Economic Statistics, 8 (1):53–69, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509777.

Escanciano:2019:MAM

[EH19] Juan Carlos Escanciano and Javier Hualde. Measuring asset market linkages: Nonlinear dependence and tail risk. Journal of Business & Economic Statistics, 39(2):453–465, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1668797.

Einmahl:2023:EVE

[EH23] John H. J. Einmahl and Yi He. Extreme value estimation for heterogeneous data. Journal of Business & Economic Statistics, 41(1):255–269, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.2008408.

Ericsson:1998:ECE

[EHM98] Neil R. Ericsson, David F. Hendry, and Grayham E. Mizon. Exogeneity, cointegration, and economic policy analysis. Journal of Business & Economic Statistics, 16(4): 370–387, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524778.

Elliott:2005:OPT

[EJP05] Graham Elliott, Michael Jansson, and Elena Pesavento. Optimal power for testing potential cointegrating vectors with known parameters for nonstationarity. Journal of Business & Economic Statistics, 23(1):34–48, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000307.

Engle:1993:R

[EK93a] Robert F. Engle and Sharon Kozicki. Reply. Journal of Business & Economic Statistics, 11(4):393–395, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509972.

Engle:1993:TCF

[EK93b] Robert F. Engle and Sharon Kozicki. Testing for common features. Journal of Business & Economic Statistics, 11 (4):369–380, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509966.

Engle:2012:DE

[EK12] Robert Engle and Bryan Kelly. Dynamic equicorrelation. Journal of Business & Economic Statistics, 30(2):212–228, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2011.652048.

Ehrman:1987:SST

[EKM87] Chaim Meyer Ehrman, Abba Krieger, and Klaus J. Miescke. Subset selection toward optimizing the best performance at a second stage. Journal of Business & Economic Statistics, 5 (2):295–303, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509589.

Elliott:2000:ERC

[Ell00] Graham Elliott. Estimating restricted cointegrating vectors. Journal of Business & Economic Statistics, 18(1):91–99, 2000. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.2000.10524850.

Elyakime:1997:ABE

[ELLV97] Bernard Elyakime, Jean-Jacques Laffont, Patrice Loisel, and Quang Vuong. Auctioning and bargaining: an econometric study of timber auctions with secret reservation prices. Journal of Business & Economic Statistics, 15(2):209–220, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524701.

Engle:2019:LDC

[ELW19] Robert F. Engle, Olivier Ledoit, and Michael Wolf. Large dynamic covariance matrices. Journal of Business & Economic Statistics, 37(2):363–375, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1345683.

Escanciano:2013:AST

[ELZ13] Juan Carlos Escanciano, Ignacio N. Lobato, and Lin Zhu. Automatic specification testing for vector autoregressions and multivariate nonlinear time series models. Journal of Business & Economic Statistics, 31(4):426–437, 2013. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2013.803973.

Engle:2004:C

[EM04] Robert F. Engle and Simone Manganelli. CAViaR. Journal of Business & Economic Statistics, 22(4):367–381, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500104000000370.

Engelberg:2009:CPP

[EMW09] Joseph Engelberg, Charles F. Manski, and Jared Williams. Comparing the point predictions and subjective probability distributions of professional forecasters. Journal of Business & Economic Statistics, 27(1):30–41, 2009. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.0003.

Elyasiani:1994:ASV

[EN94] Elyas Elyasiani and Alireza Nasseh. The appropriate scale variable in the U.S. money demand: an application of nonnested tests of consumption versus income measures. Journal of Business & Economic Statistics, 12(1):47–55, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10509990.

Engle:1994:C

[Eng94] Robert F. Engle. Comment. Journal of Business & Economic Statistics, 12(4):395–396, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10524556.

Engle:2002:DCC

[Eng02] Robert Engle. Dynamic conditional correlation. Journal of Business & Economic Statistics, 20(3):339–350, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102288618487.

Escanciano:2010:BPV

[EO10] J. Carlos Escanciano and Jose Olmo. Backtesting parametric value-at-risk with estimation risk. Journal of Business & Economic Statistics, 28(1):36–51, 2010. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/jbes.2009.07063.

Escanciano:2015:SER

[EPFK15] Juan Carlos Escanciano, Juan Carlos Pardo-Fernández, and Ingrid Van Keilegom. Semiparametric estimation of riskreturn relationships. Journal of Business & Economic Statistics, 35(1):40–52, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1052879.

Eraker:2001:MAD

[Era01] Bjørn Eraker. MCMC analysis of diffusion models with application to finance. Journal of Business & Economic Statistics, 19(2):177–191, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500101316970403.

Ergemen:2019:SEP

[Erg19] Yunus Emre Ergemen. System estimation of panel data models under long-range dependence. Journal of Business & Economic Statistics, 37(1):13–26, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1255217.

Ericsson:1993:C

[Eri93] Neil R. Ericsson. Comment. Journal of Business & Economic Statistics, 11(4):380–383, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509967.

Ericsson:1998:SSE

[Eri98] Neil R. Ericsson. Special section on exogeneity, cointegration, and economic policy analysis: Associate Editor’s introduction. Journal of Business & Economic Statistics, 16 (4):369, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524777.

Ermini:1991:RTA

[Erm91] Luigi Ermini. Reinterpreting a temporally aggregated consumption CAP model. Journal of Business & Economic Statistics, 9(3):325–328, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509858.

Easton:1988:CRM

[ERT88] George Easton, Harry V. Roberts, and George C. Tiao. Conference report making statistics more effective in schools of business. Journal of Business & Economic Statistics, 6(2): 247–260, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509659.

Engle:1993:CVI

[ES93] Robert F. Engle and Raul Susmel. Common volatility in international equity markets. Journal of Business & Economic Statistics, 11(2):167–176, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509945.

Enders:2001:CTA

[ES01a] Walter Enders and Pierre L. Siklos. Cointegration and threshold adjustment. Journal of Business & Economic Statistics, 19(2):166–176, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500101316970395.

Erdem:2001:TCD

[ES01b] Tülin Erdem and Baohong Sun. Testing for choice dynamics in panel data. Journal of Business & Economic Statistics, 19(2):142–152, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500101316970377.

Estrella:1998:NMF

[Est98] Arturo Estrella. A new measure of fit for equations with dichotomous dependent variables. Journal of Business & Economic Statistics, 16(2):198–205, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524753.

Einmahl:2020:TMR

[EYZ20] John H. J. Einmahl, Fan Yang, and Chen Zhou. Testing the multivariate regular variation model. Journal of Business & Economic Statistics, 39(4):907–919, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2020.1737533.

Fase:1984:C

[Fas84] Martin M. G. Fase. Comment. Journal of Business & Economic Statistics, 2(4):333–334, 1984. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1984.10509403.

Faust:2007:C

[Fau07] Jon Faust. Comment. Journal of Business & Economic Statistics, 25(2):154–156, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/

073500107000000043. See [DSSW07a, DSSW07b].

Foster:1992:DMI

[FB92] Kenneth A. Foster and Oscar R. Burt. A dynamic model of investment in the U.S. beef-cattle industry. Journal of Business & Economic Statistics, 10(4):419–426, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1992.10509917.

Fong:1997:AUB

[FB97] Duncan K. H. Fong and Gary E. Bolton. Analyzing ultimatum bargaining: a Bayesian approach to the comparison of two potency curves under shape constraints. Journal of Business & Economic Statistics, 15(3):335–344, 1997. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1997.10524711.

Frost:1999:AGE

[FB99] Denise Frost and Roger Bowden. An asymmetry generator for error-coorection mechanisms, with application to bank mortgagerate dynamics. Journal of Business & Economic Statistics, 17 (2):253–263, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524815.

Feng:2006:LLT

[FBB06] Shuaizhang Feng, Richard V. Burkhauser, and J. S. Butler. Levels and long-term trends in earnings inequality. Journal of Business & Economic Statistics, 24(1):57–62, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500105000000144.

Chen:1988:C

[fC88] Nai fu Chen. Comment. Journal of Business & Economic Statistics, 6(1):16, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509629. See [CR88b, CR88a].

Francke:2000:ECH

[Fdv00] M. K. Francke and A. F. de vos. Efficient computation of hierarchical trends. Journal of Business & Economic Statistics, 18(1):51–57, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524847.

Frasso:2022:DSP

[FE22] Gianluca Frasso and Paul H. C. Eilers. Direct semi-parametric estimation of the state price density implied in option prices. Journal of Business & Economic Statistics, 40(3):1179–1190, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1906686.

Ferrall:1997:UIE

[Fer97] Christopher Ferrall. Unemployment insurance eligibility and the scholl-to-work transition in Canada and the United States. Journal of Business & Economic Statistics, 15(2):115–129, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1997.10524695.

Feve:2018:ECR

[FFK18] Frédérique Fève, Jean-Pierre Florens, and Ingrid Van Keilegom. Estimation of conditional ranks and tests of exogeneity in nonparametric nonseparable models. Journal of Business & Economic Statistics, 36(2):334–345, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1166120.

Flores:2013:PIL

[FFL13] Carlos A. Flores and Alfonso Flores-Lagunes. Partial identification of local average treatment effects with an invalid instrument. Journal of Business & Economic Statistics, 31(4):534–545, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2013.
822760.

Fan:2016:IGI

[FFX16] Jianqing Fan, Alex Furger, and Dacheng Xiu. Incorporating global industrial classification standard into portfolio allocation: a simple factor-based large covariance matrix estimator with high-frequency data. Journal of Business & Economic Statistics, 34(4):489–503, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1052458.

Ferber:1984:WKF

[FG84] Marianne A. Ferber and Carole A. Green. What kind of fairness is fair? A comment on Conway and Roberts. Journal of Business & Economic Statistics, 2(2):111–113, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1984.10509376. See [CR83,

Mil84, CR84].

Freidlin:2000:CTD

[FG00] Boris Freidlin and Joseph L. Gastwirth. Changepoint tests designed for the analysis of hiring data arising in employment discrimination cases. Journal of Business & Economic Statistics, 18(3):315–322, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524873.

Fosten:2020:TNM

[FG20] Jack Fosten and Daniel Gutknecht. Testing nowcast monotonicity with estimated factors. Journal of Business & Economic Statistics, 38(1):107–123, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1458623.

Fernandes:2019:SQR

[FGH19] Marcelo Fernandes, Emmanuel Guerre, and Eduardo Horta. Smoothing quantile regressions. Journal of Business & Economic Statistics, 39(1):338–357, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2019.1660177.

Farbmacher:2020:IVT

[FGK20] Helmut Farbmacher, Raphael Guber, and Sven Klaassen. Instrument validity tests with causal forests. Journal of Business & Economic Statistics, 40(2):605–614, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1847122.

Fernandez:1990:SUT

[FH90] F. Javier Fernández and Andrew C. Harvey. Seemingly unrelated time series equations and a test for homogeneity. Journal of Business & Economic Statistics, 8(1):71–81, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1990.10509778.

Franses:1994:EAO

[FH94] Philip Hans Franses and Niels Haldrup. The effects of additive outliers on tests for unit roots and cointegration. Journal of Business & Economic Statistics, 12(4):471–478, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10524569.

Fridman:1998:MLA

[FH98] Moshe Fridman and Lawrence Harris. A maximum likelihood approach for non-Gaussian stochastic volatility models. Journal of Business & Economic Statistics, 16(3): 284–291, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524767.

Frolich:2019:ICR

[FH19] Markus Frölich and Martin Huber. Including covariates in the regression discontinuity design. Journal of Business & Economic Statistics, 37(4):736–748, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1421544.

Frijters:2011:IMP

[FHDS11] Paul Frijters, John P. Haisken-DeNew, and Michael A. Shields. The increasingly mixed proportional hazard model: an application to socioeconomic status, health shocks, and mortality. Journal of Business & Economic Statistics, 29(2):271–281, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2010.08082.

Farrell:2000:DLR

[FHLW00] Lisa Farrell, Roger Hartley, Gauthier Lanot, and Ian Walker. The demand for lotto: The role of conscious selection. Journal of Business & Economic Statistics, 18(2):228–241, 2000. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.2000.10524865.

Fan:2020:ECA

[FHLZ20] Qingliang Fan, Yu-Chin Hsu, Robert P. Lieli, and Yichong Zhang. Estimation of conditional average treatment effects with highdimensional data. Journal of Business & Economic Statistics, 40(1):313–327, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1811102.

Fan:2018:SIB

[FHWZ18] Yan Fan, Wolfgang Karl Härdle, Weining Wang, and Lixing Zhu. Single-index-based CoVaR with very high-dimensional covariates. Journal of Business & Economic Statistics, 36(2):212– 226, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2016.1180990.

Fan:2016:WDV

[FID16] Jianqing Fan, Michael B. Imerman, and Wei Dai. What does the volatility risk premium say about liquidity provision and demand for hedging tail risk? Journal of Business & Economic Statistics, 34(4):519–535, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1152968.

Filardo:1994:BCP

[Fil94] Andrew J. Filardo. Business-cycle phases and their transitional dynamics. Journal of Business & Economic Statistics, 12(3):299–308, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10524545.

Fan:2023:OCB

[FIL^{+}23] Jianqing Fan, Kosuke Imai, Inbeom Lee, Han Liu, Yang Ning, and Xiaolin Yang. Optimal covariate balancing conditions in propensity score estimation. Journal of Business & Economic Statistics, 41(1):97–110, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.2002159.

Findley:1983:C

[Fin83] David F. Findley. Comment. Journal of Business & Economic Statistics, 1(4):309–311, 1983. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1983.10509357.

Findley:1996:C

[Fin96] David F. Findley. Comment. Journal of Business & Economic Statistics, 14(3):389–393, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524666.

Fisher:1992:MDV

[Fis92] Douglas Fisher. Money-demand variability: a demand-systems approach. Journal of Business & Economic Statistics, 10 (2):143–151, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509894.

Fredriksson:2008:DTA

[FJ08] Peter Fredriksson and Per Johansson. Dynamic treatment assignment. Journal of Business & Economic Statistics, 26 (4):435–445, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500108000000033.

Fackler:1986:AVT

[FK86] James S. Fackler and Sandra C. Krieger. An application of vector time series techniques to macroeconomic forecasting. Journal of Business & Economic Statistics, 4(1):71–80, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1986.10509495.

Forbes:1999:BAT

[FKK99] Catherine S. Forbes, Guyonne R. J. Kalb, and Paul Kofhian. Bayesian arbitrage threshold analysis. Journal of Business & Economic Statistics, 17(3):364–372, 1999. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1999.10524825.

Fong:1997:JVR

[FKO97] Wai Mun Fong, Seng Kee Koh, and Sam Ouliaris. Joint varianceratio tests of the martingale hypothesis for exchange rates. Journal of Business & Economic Statistics, 15(1):51–59, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1997.10524686.

Freedman:1983:NIR

[FL83] David Freedman and David Lane. A nonstochastic interpretation of reported significance levels. Journal of Business & Economic Statistics, 1(4):292–298, 1983. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1983.10509354.

Faust:1997:WDL

[FL97] Jon Faust and Eric M. Leeper. When do long-run identifying restrictions give reliable results? Journal of Business & Economic Statistics, 15(3):345–353, 1997. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1997.10524712.

Franses:1998:ODC

[FL98] Philip Hans Franses and André Lucas. Outlier detection in cointegration analysis. Journal of Business & Economic Statistics, 16(4):459–468, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524785.

Flinn:1997:EWD

[Fli97] Christopher J. Flinn. Equilibrium wage and dismissal processes. Journal of Business & Economic Statistics, 15(2): 221–236, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524702.

Feir:2016:WIF

[FLM16] Donna Feir, Thomas Lemieux, and Vadim Marmer. Weak identification in fuzzy regression discontinuity designs. Journal of Business & Economic Statistics, 34(2):185–196, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1024836.

Forastiere:2019:EET

[FLM^{+}19] Laura Forastiere, Patrizia Lattarulo, Marco Mariani, Fabrizia Mealli, and Laura Razzolini. Exploring encouragement, treatment, and spillover effects using principal stratification, with application to a field experiment on teens’ museum attendance. Journal of Business & Economic Statistics, 39 (1):244–258, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1647843.

Feng:2020:INS

[FLM20] Long Feng, Binghui Liu, and Yanyuan Ma. An inverse norm sign test of location parameter for high-dimensional data. Journal of Business & Economic Statistics, 39(3):807–815, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1736084.

Fougere:2007:HCP

[FLS07] Denis Fougère, Hervé Le Bihan, and Patrick Sevestre. Heterogeneity in consumer price stickiness. Journal of Business & Economic Statistics, 25(3):247–264, 2007. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000214.

Fernandez:2004:BAI

[FLSVP04] Carmen Fernández, Carmelo J. León, Mark F. J. Steel, and Francisco José Vázquez-Polo. Bayesian analysis of interval data contingent valuation models and pricing policies. Journal of Business & Economic Statistics, 22(4):431–442, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000415.

Fang:2019:MAP

[FLTS19] Fang Fang, Wei Lan, Jingjing Tong, and Jun Shao. Model averaging for prediction with fragmentary data. Journal of Business & Economic Statistics, 37(3):517–527, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1383263.

Fan:1996:SES

[FLW96] Yanqin Fan, Qi Li, and Alfons Weersink. Semiparametric estimation of stochastic production frontier models. Journal of Business & Economic Statistics, 14(4):460–468, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524675.

Feng:2020:MAN

[FLYZ20] Yang Feng, Qingfeng Liu, Qingsong Yao, and Guoqing Zhao. Model averaging for nonlinear regression models. Journal of Business & Economic Statistics, 40(2):785–798, 2020. CO-DEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1870477.

Fiebig:1990:SAD

[FM90] Denzil G. Fiebig and Esfandiar Maasoumi. Specification analysis in dynamic models. Journal of Business & Economic Statistics, 8(4):443–451, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509815.

Feenberg:1997:IAN

[FM97] Daniel Feenberg and Jeffrey A. Miron. Improving the accessibility of the NBER’s historical data. Journal of Business & Economic Statistics, 15(3):293–299, 1997. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1997.10524707.

Frolich:2013:UQT

[FM13] Markus Frölich and Blaise Melly. Unconditional quantile treatment effects under endogeneity. Journal of Business & Economic Statistics, 31(3):346–357, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.803869.

Findley:1998:NCM

[FMB^{+}98a] David F. Findley, Brian C. Monsell, William R. Bell, Mark C. Otto, and Bor-Chung Chen. New capabilities and methods of the X-12-ARIMA seasonal-adjustment program. Journal of Business & Economic Statistics, 16(2):127–152, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524743.

Findley:1998:R

[FMB^{+}98b] David F. Findley, Brian C. Monsell, William R. Bell, Mark C. Otto, and Bor-Chung Chen. Reply. Journal of Business & Economic Statistics, 16(2):169–177, 1998. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1998.10524750.

Farrell:1997:EBS

[FMT97] Patrick J. Farrell, Brenda MacGibbon, and Thomas J. Tomberlin. Empirical Bayes small-area estimation using logistic regression models and summary statistics. Journal of Business & Economic Statistics, 15(1):101–108, 1997. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1997.10524692.

Fry:2010:NCT

[FMT10] Renée Fry, Vance L. Martin, and Chrismin Tang. A new class of tests of contagion with applications. Journal of Business & Economic Statistics, 28(3):423–437, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2010.06060.

Francis:2005:MPM

[FO05] Neville Francis and Michael T. Owyang. Monetary policy in a Markov-switching vector error-correction model. Journal of Business & Economic Statistics, 23(3):305–313, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000325.

Fomby:1986:CFA

[Fom86] Thomas B. Fomby. A comparison of forecasting accuracies of alternative regional production index methodologies. Journal of Business & Economic Statistics, 4(2):177–186, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509513.

Favero:2020:IRP

[FOTY20] Carlo A. Favero, Fulvio Ortu, Andrea Tamoni, and Haoxi Yang. Implications of return predictability for consumption dynamics and asset pricing. Journal of Business & Economic Statistics, 38(3):527–541, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1527702.

Freedman:1984:BEM

[FP84] David A. Freedman and Stephen C. Peters. Bootstrapping an econometric model: Some empirical results. Journal of Business & Economic Statistics, 2(2):150–158, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1984.10509383.

Fafchamps:1997:PSC

[FP97] Marcel Fafchamps and John Pender. Precautionary saving, credit constraints, and irreversible investment: Theory and evidence from serniarid India. Journal of Business & Economic Statistics, 15(2):180–194, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524699.

Fiorentini:2001:ONA

[FP01] Gabriele Fiorentini and Christophe Planas. Overcoming nonadmissibility in ARIMA-model-based signal extraction. Journal of Business & Economic Statistics, 19(4):455–464, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/07350010152596709.

Fong:2002:RUB

[FPAB02] Duncan K. H. Fong, Scott E. Pammer, Steven F. Arnold, and Gary E. Bolton. Reanalyzing ultimatum bargaining-comparing nondecreasing curves without shape constraints. Journal of Business & Economic Statistics, 20(3):423–430, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102288618540.

Fan:2014:QML

[FQX14a] Jianqing Fan, Lei Qi, and Dacheng Xiu. Quasi-maximum likelihood estimation of GARCH models with heavy-tailed likelihoods. Journal of Business & Economic Statistics, 32(2):178– 191, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.

2013.840239. See comments [And14, FS14, FZ14, Yao14b, LZ14] and rejoinder [FQX14b].

Fan:2014:R

[FQX14b] Jianqing Fan, Lei Qi, and Dacheng Xiu. Rejoinder. Journal of Business & Economic Statistics, 32(2):204–205, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2014.898448.

See [FQX14a].

Franke:2014:C

[Fra14] Juergen Franke. Comment. Journal of Business & Economic Statistics, 32(2):171–172, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.903652. See [GPI13b, GPI13a].

Frandsen:2019:TCP

[Fra19] Brigham R. Frandsen. Testing censoring point independence. Journal of Business & Economic Statistics, 37(3):496–505, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1383261.

Freedman:1985:MVM

[Fre85] D. A. Freedman. The mean versus the median: a case study in 4-r act litigation. Journal of Business & Economic Statistics, 3(1):1–13, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509421.

Freedman:1986:CSN

[Fre86a] David A. Freedman. A case study in nonresponse: Plaintiff vs. California state board of equalization. Journal of Business & Economic Statistics, 4(1):123–124, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509501.

Freedman:1986:MWP

[Fre86b] David A. Freedman. A model for water pricing. Journal of Business & Economic Statistics, 4(1):131–133, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509505.

Freedman:1986:R

[Fre86c] David A. Freedman. Reply. Journal of Business & Economic Statistics, 4(1):126–127, 1986. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1986.10509503.

Frees:1988:ECW

[Fre88] Edward W. Frees. Estimating the cost of a warranty. Journal of Business & Economic Statistics, 6(1):79–86, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1988.10509639.

Freedman:1989:C

[Fre89] D. A. Freedman. Comment. Journal of Business & Economic Statistics, 7(1):31–32, 1989. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1989.10509709.

Frees:1992:FSS

[Fre92] Edward W. Frees. Forecasting state-to-state migration rates. Journal of Business & Economic Statistics, 10(2):153–167, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/07350015.1992.10509895.

Fries:2022:CMN

[Fri22] Sébastien Fries. Conditional moments of noncausal alphastable processes and the prediction of bubble crash odds. Journal of Business & Economic Statistics, 40(4):1596–1616, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1953508.

Freedman:1983:EPM

[FRS83a] David Freedman, Thomas Rothenberg, and Richard Sutch. On energy policy models. Journal of Business & Economic Statistics, 1(1):24–32, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/

07350015.1983.10509318. See comments [Hog83, Smi83] and rejoinder [FRS83b].

Freedman:1983:R

[FRS83b] David Freedman, Thomas Rothenberg, and Richard Sutch. Rejoinder. Journal of Business & Economic Statistics, 1(1):36, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/07350015.1983.10509321.

See [FRS83a, Hog83, Smi83].

Fichtenbaum:1988:TBM

[FS88] Rudy Fichtenbaum and Hushang Shahidi. Truncation bias and the measurement of income inequality. Journal of Business & Economic Statistics, 6(3):335–337, 1988. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1988.10509670.

Fomby:1991:ASR

[FS91] Thomas B. Fomby and Subarna K. Samanta. Application of Stein rules to combination forecasting. Journal of Business & Economic Statistics, 9(4):391–407, 1991. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1991.10509867.

French:1993:CPV

[FS93] Mark W. French and Daniel E. Sichel. Cyclical patterns in the variance of economic activity. Journal of Business & Economic Statistics, 11(1):113–119, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509937.

Fleissig:1997:DAI

[FS97] Adrian R. Fleissig and James L. Swofford. Dynamic asymptotically ideal models and finite approximation. Journal of Business & Economic Statistics, 15(4):482–492, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524726.

Fiorentini:2014:C

[FS14] Gabriele Fiorentini and Enrique Sentana. Comment. Journal of Business & Economic Statistics, 32(2):193–198, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2013.878661.

See [FQX14a, FQX14b].

Florens:2019:GPB

[FS19] Jean-Pierre Florens and Anna Simoni. Gaussian processes and Bayesian moment estimation. Journal of Business & Economic Statistics, 39(2):482–492, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2019.1668799.

Ferrara:2022:WGD

[FS22] Laurent Ferrara and Anna Simoni. When are Google Data useful to Nowcast GDP? An approach via preselection and shrinkage. Journal of Business & Economic Statistics, 41(4):1188– 1202, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2116025.

Francq:2023:VEW

[FS23] Christian Francq and Genaro Sucarrat. Volatility estimation when the zero-process is nonstationary. Journal of Business & Economic Statistics, 41(1):53–66, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1999821.

Fiorentini:2003:MLE

[FSC03] Gabriele Fiorentini, Enrique Sentana, and Giorgio Calzolari. Maximum likelihood estimation and inference in multivariate conditionally heteroscedastic dynamic regression models with student t innovations. Journal of Business & Economic Statistics, 21(4):532–546, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619232.

Frohwirth-Schnatter:2008:MBC

[FSK08] Sylvia Fröhwirth-Schnatter and Sylvia Kaufmann. Model-based clustering of multiple time series. Journal of Business & Economic Statistics, 26(1):78–89, 2008. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000106.

Fruhwirth-Schnatter:2004:BAH

[FSTO04] Sylvia Frühwirth-Schnatter, Regina Tüchler, and Thomas Otter. Bayesian analysis of the heterogeneity model. Journal of Business & Economic Statistics, 22(1):2–15, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500103288619331.

Feunou:2012:SVM

[FT12] Bruno Feunou and Roméo Tédongap. A stochastic volatility model with conditional skewness. Journal of Business & Economic Statistics, 30(4):576–591, 2012. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2012.715958.

Fuhrer:1992:ICE

[Fuh92] Jeffrey C. Fuhrer. Inferring changes in expectation behavior over time: an application of nonlinear time-varying-parameters estimation. Journal of Business & Economic Statistics, 10 (2):169–177, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509896.

Fuller:1983:D

[Ful83] Wayne A. Fuller. Discussion. Journal of Business & Economic Statistics, 1(1):6, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509316.

Foster:1995:CST

[FV95] F. Douglas Foster and S. Viswanathan. Can speculative trading explain the volume-volatility relation? Journal of Business & Economic Statistics, 13(4):379–396, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524613.

Fleissig:2003:NPB

[FW03] Adrian R. Fleissig and Gerald A. Whitney. A new PC-based test for Varian’s weak separability conditions. Journal of Business & Economic Statistics, 21(1):133–144, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102288618838.

Fleissig:2005:TSV

[FW05] Adrian R. Fleissig and Gerald A. Whitney. Testing for the significance of violations of Afriat’s inequalities. Journal of Business & Economic Statistics, 23(3):355–362, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000253.

Faust:2009:CGR

[FW09] Jon Faust and Jonathan H. Wright. Comparing greenbook and reduced form forecasts using a large realtime dataset. Journal of Business & Economic Statistics, 27(4):468–479, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2009.07214.

Fang:2020:REA

[FXMFY20] Yan Fang, Lan Xue, Carlos Martins-Filho, and Lijian Yang. Robust estimation of additive boundaries with quantile regression and shape constraints. Journal of Business & Economic Statistics, 40(2):615–628, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1847123.

Francq:2013:EML

[FZ13] Christian Francq and Jean-Michel Zakoïan. Estimating the marginal law of a time series with applications to heavytailed distributions. Journal of Business & Economic Statistics, 31(4):412–425, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.801776.

Francq:2014:C

[FZ14] Christian Francq and Jean-Michel Zakoïan. Comment. Journal of Business & Economic Statistics, 32(2):198–201, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2013.879829.

See [FQX14a, FQX14b].

Fan:2018:NAI

[FZ18] Qingliang Fan and Wei Zhong. Nonparametric additive instrumental variable estimator: a group shrinkage estimation perspective. Journal of Business & Economic Statistics, 36(3):388–399, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1180991.

Galeano:2010:GMD

[GA10] Pedro Galeano and M. Concepción Ausín. The Gaussian mixture dynamic conditional correlation model: Parameter estimation, value at risk calculation, and portfolio selection. Journal of Business & Economic Statistics, 28(4):559–571, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.07238.

Gagnon:1990:SSG

[Gag90] Joseph E. Gagnon. Solving the stochastic growth model by deterministic extended path. Journal of Business & Economic Statistics, 8(1):35–36, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509771.

Gallant:1993:RMT

[Gal93] A. Ronald Gallant. Remarks on my term at JBES. Journal of Business & Economic Statistics, 11(4):426, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509978.

Gallant:2007:C

[Gal07] A. Ronald Gallant. Comment. Journal of Business & Economic Statistics, 25(2):151–152, 2007. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000034. See [DSSW07a, DSSW07b].

Garen:1989:JMQ

[Gar89] John E. Garen. Job-match quality as an error component and the wage-tenure profiler a comparison and test of alternative estimators. Journal of Business & Economic Statistics, 7(2):245–252, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509733.

Gastwirth:1985:SCC

[Gas85] Joseph L. Gastwirth. Short communications: Comment on “Measurement of Economic Distance Between Blacks and Whites” by H. D. Vinod. Journal of Business & Economic Statistics, 3(4):405–407, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509479. See [Vin85a, Vin85b].

Gyimah-Brempong:1987:EFS

[GB87] Kwabena Gyimah-Brempong. Elasticity of factor substitution in police agencies evidence from Florida. Journal of Business & Economic Statistics, 5(2):257–265, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509584.

Gbur:1989:CAE

[Gbu89] Edward Gbur. A comparison of alternative estimators for revised monthly import statistics. Journal of Business & Economic Statistics, 7(1):21–25, 1989. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1989.10509706.

Gerlach:2011:BTV

[GCC11] Richard H. Gerlach, Cathy W. S. Chen, and Nancy Y. C. Chan. Bayesian time-varying quantile forecasting for value-at-risk in financial markets. Journal of Business & Economic Statistics, 29(4):481–492, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2010.08203.

Gracia-Diez:1989:CCL

[GD89] Mercedes Gracia-Diez. Compositional changes of the labor force and the increase of the unemployment rate: an estimate for the United States. Journal of Business & Economic Statistics, 7(2):237–243, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509732.

Gardes:2005:PPP

[GDG^{+}05] François Gardes, Greg J. Duncan, Patrice Gaubert, Marc Gurgand, and Christophe Starzec. Panel and pseudo-panel estimation of cross-sectional and time series elasticities of food consumption. Journal of Business & Economic Statistics, 23 (2):242–253, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000587.

Graham:2016:EED

[GdXPE16] Bryan S. Graham, Cristine Campos de Xavier Pinto, and Daniel Egel. Efficient estimation of data combination models by the method of auxiliary-to-study tilting (AST). Journal of Business & Economic Statistics, 34(2):288–301, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1038544.

Gurmu:2012:FBC

[GE12] Shiferaw Gurmu and John Elder. Flexible bivariate count data regression models. Journal of Business & Economic Statistics, 30(2):265–274, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2011.638816.

Gertler:1988:LVM

[Ger88] Paul J. Gertler. A latent-variable model of quality determination. Journal of Business & Economic Statistics, 6 (1):97–104, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509641.

Geweke:1988:C

[Gew88a] John Geweke. Comment. Journal of Business & Economic Statistics, 6(4):465–466, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509691.

Geweke:1988:SCB

[Gew88b] John Geweke. The secular and cyclical behavior of real GDP in 19 OECD countries, 1957–1983. Journal of Business & Economic Statistics, 6(4):479–486, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509696.

Geweke:1991:ER

[Gew91] John Geweke. Editor’s report. Journal of Business & Economic Statistics, 9(3):343, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509862.

Geweke:1992:ER

[Gew92] John Geweke. Editor’s report. Journal of Business & Economic Statistics, 10(4):375, 1992. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1992.10509912.

Geweke:1993:RMT

[Gew93] John Geweke. Remarks on my term at JBES. Journal of Business & Economic Statistics, 11(4):427, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509979.

Geweke:1994:C

[Gew94] John Geweke. Comment. Journal of Business & Economic Statistics, 12(4):397–399, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10524557.

Geweke:1998:C

[Gew98] John Geweke. Comment. Journal of Business & Economic Statistics, 16(3):269–271, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524762.

Garcia-Ferrer:1987:ACA

[GFdH87] Antonio Garcia-Ferrer and Juan del Hoyo. Analysis of the car accident indexes in Spain: a multiple time series approach. Journal of Business & Economic Statistics, 5 (1):27–38, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509557.

Garcia-Ferrer:1987:MFU

[GFHPZ87] A. Garcia-Ferrer, R. A. Highfield, F. Palm, and A. Zellner. Macroeconomic forecasting using pooled international data. Journal of Business & Economic Statistics, 5(1):53–67, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1987.10509560.

Gillingham:1990:IFT

[GG90] Robert Gillingham and John S. Greenlees. Indexing the Federal tax system: a cost-of-living approach. Journal of Business & Economic Statistics, 8(4):465–473, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1990.10509818.

Gonzalo:1995:ECL

[GG95] Jesus Gonzalo and Clive Granger. Estimation of common longmemory components in cointegrated systems. Journal of Business & Economic Statistics, 13(1):27–35, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524576.

Gelfand:1998:STM

[GGKS98] Alan E. Gelfand, Sujit K. Ghosh, John R. Knight, and C. F. Sirmans. Spatio-temporal modeling of residential sales data. Journal of Business & Economic Statistics, 16(3): 312–321, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524770.

Ghysels:1996:SAL

[GGS96a] Eric Ghysels, Clive W. J. Granger, and Pierre L. Siklos. Is seasonal adjustment a linear or nonlinear data-filtering process? Journal of Business & Economic Statistics, 14(3): 374–386, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524663.

Ghysels:1996:R

[GGS96b] Eric Ghysels, Clive W. J. Granger, and Pierre L. Siklos. Reply. Journal of Business & Economic Statistics, 14(3): 396–397, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524668.

Guan:2014:VNB

[GGW14] Guoyu Guan, Jianhua Guo, and Hansheng Wang. Varying naïve Bayes models with applications to classification of Chinese text documents. Journal of Business & Economic Statistics, 32(3):445–456, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.903086.

Gillis:1991:ESO

[GH91] Kurt D. Gillis and Jesse S. Hixson. Efficacy of statistical outlier analysis for monitoring quality of care. Journal of Business & Economic Statistics, 9(3):241–252, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509850.

Gruber:1995:LME

[GH95] Jonathan Gruber and Maria Hanratty. The labor-marker effects of introducing national health insurance: Evidence from Canada. Journal of Business & Economic Statistics, 13(2): 163–173, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524590.

Ghysels:2002:EIT

[GH02a] Eric Ghysels and Alastair Hall. Editors’ introduction to Twentieth anniversary commemorative issue of the Journal of Business and Economic Statistics. Journal of Business & Economic Statistics, 20(1):1–4, 2002. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102753410345.

Ghysels:2002:EIJ

[GH02b] Eric Ghysels and Alastair R. Hall. Editors’ introduction to JBES Twentieth anniversary issue on the generalized method of moments. Journal of Business & Economic Statistics, 20(4):441, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618568.

Ghysels:2002:ER

[GH02c] Eric Ghysels and Alastair R. Hall. Editors’ report 2001. Journal of Business & Economic Statistics, 20(4):545, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102288618676.

Ghysels:2003:ER

[GH03] Eric Ghysels and Alastair R. Hall. Editors’ report 2002. Journal of Business & Economic Statistics, 21(4):583, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/0735001032886191313.

Ghysels:2004:EA

[GH04a] Eric Ghysels and Alastair Hall. Editorial announcement. Journal of Business & Economic Statistics, 22(1):1, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500103288619377.

Ghysels:2004:ER

[GH04b] Eric Ghysels and Alastair R. Hall. Editors’ report 2003. Journal of Business & Economic Statistics, 22(4):488, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000479.

Garlick:2022:QEE

[GH22] Robert Garlick and Joshua Hyman. Quasi-experimental evaluation of alternative sample selection corrections. Journal of Business & Economic Statistics, 40(3):950–964, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1889566.

Georgiev:2019:BST

[GHLT19] Iliyan Georgiev, David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. A bootstrap stationarity test for predictive regression invalidity. Journal of Business & Economic Statistics, 37(3):528–541, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1385467.

Gustafsson:2009:LTK

[GHNS09] J. Gustafsson, M. Hagmann, J. P. Nielsen, and O. Scaillet. Local transformation kernel density estimation of loss distributions. Journal of Business & Economic Statistics, 27(2):161–175, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2009.0011.

Goncalves:2022:BTS

[GHPS22] Sílvia Gonçalves, Ulrich Hounyo, Andrew J. Patton, and Kevin Sheppard. Bootstrapping two-stage quasi-maximum likelihood estimators of time series models. Journal of Business & Economic Statistics, 41(3):683–694, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2058949.

Green:1987:SEE

[GHR87] Richard Green, William Hahn, and David Rocke. Standard errors for elasticities: a comparison of bootstrap and asymptotic standard errors. Journal of Business & Economic Statistics, 5(1):145–149, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509569.

Giacomini:2009:IDM

[GHS09] Enzo Giacomini, Wolfgang Härdle, and Vladimir Spokoiny. Inhomogeneous dependence modeling with time-varying copulae. Journal of Business & Economic Statistics, 27(2):224–234, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2009.0016.

Ghysels:1987:SEP

[Ghy87] Eric Ghysels. Seasonal extraction in the presence of feedback. Journal of Business & Economic Statistics, 5(2):

191–194, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509577.

Ghysels:1990:URT

[Ghy90] Eric Ghysels. Unit-root tests and the statistical pitfalls of seasonal adjustment: The case of U.S. postwar real gross national product. Journal of Business & Economic Statistics, 8 (2):145–152, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509785.

Grosskopf:1992:MED

[GHY92] Shawna Grosskopf, Kathy Hayes, and Suthathip Yaisawarng. Measuring economies of diversification: a frontier approach. Journal of Business & Economic Statistics, 10(4):453–459, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1992.10509921.

Ghysels:1994:PSB

[Ghy94] Eric Ghysels. On the periodic structure of the business cycle. Journal of Business & Economic Statistics, 12(3): 289–298, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10524544.

Ghysels:1997:SAO

[Ghy97] Eric Ghysels. Seasonal adjustment and other data transformations. Journal of Business & Economic Statistics, 15 (4):410–418, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524719.

Ghysels:1998:C

[Ghy98] Eric Ghysels. Comment. Journal of Business & Economic Statistics, 16(2):165–167, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524748.

Ghysels:2000:SER

[Ghy00] Eric Ghysels. Some econometric recipes for high-frequency data cooking. Journal of Business & Economic Statistics, 18 (2):154–163, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524858.

Gabaix:2011:RSW

[GI11] Xavier Gabaix and Rustam Ibragimov. Rank-1/2: a simple way to improve the OLS estimation of tail exponents. Journal of Business & Economic Statistics, 29(1):24–39, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.06157.

Gelman:2019:WHO

[GI19] Andrew Gelman and Guido Imbens. Why high-order polynomials should not be used in regression discontinuity designs. Journal of Business & Economic Statistics, 37(3):447–456, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1366909.

Giannone:2016:C

[Gia16] Domenico Giannone. Comment. Journal of Business & Economic Statistics, 34(3):342–344, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1190280. See [CR16a, CR16b].

Ganics:2019:CIB

[GIR19] Gergely Ganics, Atsushi Inoue, and Barbara Rossi. Confidence intervals for bias and size distortion in IV and local projections-IV models. Journal of Business & Economic Statistics, 39 (1):307–324, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1660175.

Graham:2020:IEB

[GIR20] Bryan S. Graham, Guido W. Imbens, and Geert Ridder. Identification and efficiency bounds for the average match function under conditionally exogenous matching. Journal of Business & Economic Statistics, 38(2):303–316, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2018.1497509.

Ghysels:1994:C

[GJ94] Eric Ghysels and Joanna Jasiak. Comment. Journal of Business & Economic Statistics, 12(4):399–401, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10524558.

Gredenhoff:2001:BTL

[GJ01] Mikael Gredenhoff and Tor Jacobson. Bootstrap testing linear restrictions on cointergrating vectors. Journal of Business & Economic Statistics, 19(1):63–72, 2001. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/07350010152472625.

Gourieroux:2022:GCE

[GJ22] Christian Gourieroux and Joann Jasiak. Generalized covariance estimator. Journal of Business & Economic Statistics, 41(4):1315– 1327, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2120486.

Guiso:2002:EAE

[GJP02] Luigi Guiso, Tullio Jappelli, and Luigi Pistaferri. An empirical analysis of earnings and employment risk. Journal of Business & Economic Statistics, 20(2):241–253, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102317351985.

Gersch:1983:PTS

[GK83] Will Gersch and Genshiro Kitagawa. The prediction of time series with trends and seasonalities. Journal of Business & Economic Statistics, 1(3):253–264, 1983. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1983.10509349.

Garber:1992:C

[GK92] Steven Garber and Mark Kamlet. Comment. Journal of Business & Economic Statistics, 10(1):44–48, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1992.10509884.

Gross:1996:TSA

[GK96] John Gross and Dan Kaiser. Two simple algorithms for generating a subset of data consistent with WARP and other binary relations. Journal of Business & Economic Statistics, 14 (2):251–255, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524652.

Groen:2003:LBC

[GK03] Jan J. J. Groen and Frank Kleibergen. Likelihood-based cointegration analysis in panels of vector error-correction models. Journal of Business & Economic Statistics, 21(2):295–318, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500103288618972.

Giacomini:2005:ECC

[GK05] Raffaella Giacomini and Ivana Komunjer. Evaluation and combination of conditional quantile forecasts. Journal of Business & Economic Statistics, 23(4):416–431, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500105000000018.

Giordani:2008:EBI

[GK08] Paolo Giordani and Robert Kohn. Efficient Bayesian inference for multiple change-point and mixture innovation models. Journal of Business & Economic Statistics, 26(1):66–77, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500107000000241.

Galvao:2014:EIL

[GK14] Antonio F. Galvao and Kengo Kato. Estimation and inference for linear panel data models under misspecification when both n and T are large. Journal of Business & Economic Statistics, 32(2):285–309, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.875473.

Gu:2015:UHI

[GK15] Jiaying Gu and Roger Koenker. Unobserved heterogeneity in income dynamics: an empirical Bayes perspective. Journal of Business & Economic Statistics, 35(1):1–16, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1052457.

Gayle:2018:MLE

[GK18] Wayne-Roy Gayle and Natalia Khorunzhina. Micro-level estimation of optimal consumption choice with intertemporal nonseparability in preferences and measurement errors. Journal of Business & Economic Statistics, 36(2):227–238, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2016.1149071.

Gunawan:2020:MMC

[GKK20] David Gunawan, Mohamad A. Khaled, and Robert Kohn. Mixed marginal copula modeling. Journal of Business & Economic Statistics, 38(1):137–147, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1469998.

Garratt:2009:RTP

[GKMV09] Anthony Garratt, Gary Koop, Emi Mise, and Shaun P. Vahey. Real-time prediction with U.K. monetary aggregates in the presence of model uncertainty. Journal of Business & Economic Statistics, 27(4):480–491, 2009. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.07208.

Golan:2000:ECP

[GKP00] Amos Golan, Larry S. Karp, and Jeffrey M. Perloff. Estimating Coke’s and Pepsi’s price and advertising strategies. Journal of Business & Economic Statistics, 18(4): 398–409, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524880.

Gospodinov:2012:FRL

[GKR12] Nikolay Gospodinov, Raymond Kan, and Cesare Robotti. Further results on the limiting distribution of GMM sample moment conditions. Journal of Business & Economic Statistics, 30(4):494–504, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.694743.

Giacomini:2022:NRPa

[GKR22a] Raffaella Giacomini, Toru Kitagawa, and Matthew Read. Narrative restrictions and proxies. Journal of Business & Economic Statistics, 40(4):1415–1425, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2115496. See com-

ments [RR22, Kil22, PM22] and rejoinder [?, GKR22b].

Giacomini:2022:NRPb

[GKR22b] Raffaella Giacomini, Toru Kitagawa, and Matthew Read. Narrative restrictions and proxies: Rejoinder. Journal of Business & Economic Statistics, 40(4):1438–1441, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.

2115710. See [GKR22a, RR22, Kil22, PM22].

Gungor:2013:TLF

[GL13] Sermin Gungor and Richard Luger. Testing linear factor pricing models with large cross sections: a distribution-free approach. Journal of Business & Economic Statistics, 31(1):66–77, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2012.
740435.

Gungor:2016:MTM

[GL16] Sermin Gungor and Richard Luger. Multivariate tests of meanvariance efficiency and spanning with a large number of assets and time-varying covariances. Journal of Business & Economic Statistics, 34(2):161–175, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1019510.

Gorodnichenko:2019:FEV

[GL19] Yuriy Gorodnichenko and Byoungchan Lee. Forecast error variance decompositions with local projections. Journal of Business & Economic Statistics, 38(4):921–933, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2019.1610661.

Guo:2020:HSI

[GL20] Chaohui Guo and Jialiang Li. Homogeneity and structure identification in semiparametric factor models. Journal of Business & Economic Statistics, 40(1):408–422, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1831516.

Gabler:1993:SEB

[GLL93] Siegfried Gabler, François Laisney, and Michael Lechner. Seminonparametric estimation of binary-choice models with an application to labor-force participation. Journal of Business & Economic Statistics, 11(1):61–80, 1993. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1993.10509933.

Guerin:2020:MST

[GLLM20] Pierre Guérin, Danilo Leiva-Leon, and Massimiliano Marcellino. Markov-switching three-pass regression filter. Journal of Business & Economic Statistics, 38(2):285–302, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2018.1497508.

Gaglianone:2011:EVR

[GLLS11] Wagner Piazza Gaglianone, Luiz Renato Lima, Oliver Linton, and Daniel R. Smith. Evaluating value-at-risk models via quantile regression. Journal of Business & Economic Statistics, 29 (1):150–160, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2010.07318.

Guttman:1983:BIM

[GM83] Irwin Guttman and Ulrich Menzefricke. Bayesian inference in multivariate regression with missing observations on the response variables. Journal of Business & Economic Statistics, 1(3):239–248, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509347.

Gritz:1997:MIU

[GM97] R. Mark Gritz and Thomas MaCurdy. Measuring the influence of unemployment insurance on unemployment experiences. Journal of Business & Economic Statistics, 15(2): 130–152, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524696.

Garcia:2006:C

[GM06] René Garcia and Nour Meddahi. Comment. Journal of Business & Economic Statistics, 24(2):184–192, 2006. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500106000000161. See [HL06a, HL06b].

Guerin:2013:MSM

[GM13] Pierre Guérin and Massimiliano Marcellino. Markov-switching MIDAS models. Journal of Business & Economic Statistics, 31(1):45–56, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.727721.

Griffin:2020:BQT

[GM20] Jim E. Griffin and Gelly Mitrodima. A Bayesian quantile time series model for asset returns. Journal of Business & Economic Statistics, 40(1):16–27, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2020.1766470.

Gottlieb:1985:IEI

[GMP85] Daniel Gottlieb, Rafi Melnick, and Sylvia Piterman. Inflationary expectations in Israel: a multiple indicators approach. Journal of Business & Economic Statistics, 3(2): 112–117, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509438.

Gospodinov:2011:SIR

[GMP11] Nikolay Gospodinov, Alex Maynard, and Elena Pesavento. Sensitivity of impulse responses to small low-frequency comovements: Reconciling the evidence on the effects of technology shocks. Journal of Business & Economic Statistics, 29(4):455–467, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2011.10042.

Grossman:1987:ECT

[GMS87] S. J. Grossman, A. Melino, and R. J. Shiller. Estimating the continuous-time consumption-based asset-pricing model. Journal of Business & Economic Statistics, 5(3):315–327, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1987.10509594.

Gospodinov:2015:MDE

[GN15] Nikolay Gospodinov and Serena Ng. Minimum distance estimation of possibly noninvertible moving average models. Journal of Business & Economic Statistics, 33(3):403–417, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2014.955175.

Gastwirth:1986:LST

[GNK86] Joseph L. Gastwirth, Tapan K. Nayak, and Abba M. Krieger. Large sample theory for the bounds on the Gini and related indices of inequality estimated from grouped data. Journal of Business & Economic Statistics, 4(2):269–273, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509522.

Gagne:1998:CFF

[GO98] Robert Gagné and Pierre Ouellette. On the choice of functional forms: Summary of a Monte Carlo experiment. Journal of Business & Economic Statistics, 16(1):118–124, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524741.

Goidin:1985:SAC

[Goi85] Ephraim Goidin. Statistical analysis of coins lost in circulation. Journal of Business & Economic Statistics, 3 (1):36–42, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509424.

Goldberger:1984:RRR

[Gol84] Arthur S. Goldberger. Redirecting reverse regression. Journal of Business & Economic Statistics, 2(2):114–116, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1984.10509377.

Gomez:1999:TEM

[Góm99] Victor Gómez. Three equivalent methods for filtering finite nonstationary time series. Journal of Business & Economic Statistics, 17(1):109–116, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524800.

Gomez:2001:UBF

[Góm01] Víctor Gómez. The use of Butterworth filters for trend and cycle estimation in economic time series. Journal of Business & Economic Statistics, 19(3):365–373, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500101681019909.

Goodwin:1993:BCA

[Goo93] Thomas H. Goodwin. Business-cycle analysis with a Markovswitching model. Journal of Business & Economic Statistics, 11(3):331–339, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509961.

Goodwin:1996:SDF

[Goo96] Barry K. Goodwin. Semiparametric (distribution-free) testing of the expectations hypothesis in a parimutuel gambling market. Journal of Business & Economic Statistics, 14(4): 487–496, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524678.

Gordon:1989:RFA

[Gor89] Daniel V. Gordon. A revenue-function approach to the measurement of output-substitution possibilities in agriculture. Journal of Business & Economic Statistics, 7(4): 483–487, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509761.

Gospodinov:2002:BBI

[Gos02] Nikolay Gospodinov. Bootstrap-based inference in models with a nearly noninvertible moving average component. Journal of Business & Economic Statistics, 20(2):254–268, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102317351994.

Gospodinov:2010:INN

[Gos10] Nikolay Gospodinov. Inference in nearly nonstationary SVAR models with long-run identifying restrictions. Journal of Business & Economic Statistics, 28(1):1–12, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.08116.

Goudreau:1984:AQS

[GOV84] Karen Goudreau, Howard Oberheu, and Denton Vaughan. An assessment of the quality of survey reports of income from the aid to families with dependent children (AFDC) program. Journal of Business & Economic Statistics, 2(2): 179–186, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509386.

Garber:1988:C

[GP88] Steven Garber and Dale J. Poirier. Comment. Journal of Business & Economic Statistics, 6(4):466–469, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509692.

Geweke:1998:PDR

[GP98] John Geweke and Lea Petrella. Prior density-ratio class robustness in econometrics. Journal of Business & Economic Statistics, 16(4):469–478, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524786.

Gonzalo:2012:RSP

[GP12] Jesús Gonzalo and Jean-Yves Pitarakis. Regime-specific predictability in predictive regressions. Journal of Business & Economic Statistics, 30(2):229–241, 2012. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2011.652053.

Gonzalo:2017:IPI

[GP17] Jesùs Gonzalo and Jean-Yves Pitarakis. Inferring the predictability induced by a persistent regressor in a predictive threshold model. Journal of Business & Economic Statistics, 35(2):202–217, 2017. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1164054.

Goncalves:2015:BPI

[GPD15] Sílvia Gonçalves, Benoit Perron, and Antoine Djogbenou. Bootstrap prediction intervals for factor models. Journal of Business & Economic Statistics, 35(1):53–69, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1054492.

Goldsmith-Pinkham:2013:R

[GPI13a] Paul Goldsmith-Pinkham and Guido Imbens. Rejoinder. Journal of Business & Economic Statistics, 31(3):279–281, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2013.792260.

See [GPI13b].

Goldsmith-Pinkham:2013:SNI

[GPI13b] Paul Goldsmith-Pinkham and Guido W. Imbens. Social networks and the identification of peer effects. Journal of Business & Economic Statistics, 31(3):253–264, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2013.801251. See comments

[Bra13, Gra13, Jac13, Man13, Sac13, KT13] and rejoinder [GPI13a].

Guerrero:1998:MIE

[GPP98] Victor M. Guerrero, Daniel Peña, and Pilar Poncela. Measuring intervention effects on multiple time series subjected to linear restrictions: a banking example. Journal of Business & Economic Statistics, 16(4):489–497, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524788.

Groen:2013:RTI

[GPR13] Jan J. J. Groen, Richard Paap, and Francesco Ravazzolo. Realtime inflation forecasting in a changing world. Journal of Business & Economic Statistics, 31(1):29–44, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2012.727718.

Giovannins:1989:ERD

[GR89] Alberto Giovannins and Julio J. Rotemberg. Exchange-rate dynamics with sticky prices: The Deutsche mark, 1974– 1982. Journal of Business & Economic Statistics, 7(2): 169–178, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509725.

Gneiting:2011:CDF

[GR11] Tilmann Gneiting and Roopesh Ranjan. Comparing density forecasts using threshold- and quantile-weighted scoring rules. Journal of Business & Economic Statistics, 29(3):411–422, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2010.08110.

Grishchenko:2012:RHA

[GR12] Olesya V. Grishchenko and Marco Rossi. The role of heterogeneity in asset pricing: The effect of a clustering approach. Journal of Business & Economic Statistics, 30(2):297–311, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2012.670544.

Gianetto:2015:TIC

[GR15] Quentin Giai Gianetto and Hamdi Raïssi. Testing instantaneous causality in presence of nonconstant unconditional covariance. Journal of Business & Economic Statistics, 33(1):46–53, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2014.
920614.

Granger:1984:C

[Gra84] Clive W. J. Granger. Comment. Journal of Business & Economic Statistics, 2(4):335–336, 1984. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1984.10509404.

Granger:1986:C

[Gra86] C. W. J. Granger. Comment. Journal of Business & Economic Statistics, 4(1):16–17, 1986. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1986.10509487.

Gramley:1987:EFO

[Gra87] Lyle E. Gramley. The economic and financial outlook. Journal of Business & Economic Statistics, 5(1):1–4, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509553.

Granger:1993:C

[Gra93] Clive W. J. Granger. Comment. Journal of Business & Economic Statistics, 11(4):384–385, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509968.

Graddy:1997:DFF

[Gra97] Kathryn Graddy. Do fast-food chains price discriminate on the race and income characteristics of an area? Journal of Business & Economic Statistics, 15(4):391–401, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524717.

Granger:1998:C

[Gra98] Clive W. J. Granger. Comment. Journal of Business & Economic Statistics, 16(3):268–269, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524761.

Graham:2013:C

[Gra13] Bryan S. Graham. Comment. Journal of Business & Economic Statistics, 31(3):266–270, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.792261. See [GPI13b, GPI13a].

Gradzewicz:2019:WHA

[Gra19] Michał Gradzewicz. What happens after an investment spikeinvestment events and firm performance. Journal of Business & Economic Statistics, 39(3):636–651, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2019.1708369.

Greene:1984:RRA

[Gre84] William H. Greene. Reverse regression: The algebra of discrimination. Journal of Business & Economic Statistics, 2 (2):117–120, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509378.

Gregory:1989:NTA

[Gre89] Allan W. Gregory. A nonparametric test for autoregressive conditional heteroscedasticity: a Markov-chain approach. Journal of Business & Economic Statistics, 7(1):107–115, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1989.10509718.

Gregory:1994:TCL

[Gre94] Allan W. Gregory. Testing for cointegration in linear quadratic models. Journal of Business & Economic Statistics, 12(3): 347–360, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10524550.

Grillenzoni:1993:APA

[Gri93] Carlo Grillenzoni. ARIMA processes with ARIMA parameters. Journal of Business & Economic Statistics, 11(2): 235–250, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509952.

Gonzalez-Rivera:2013:CRI

[GRL13] Gloria González-Rivera and Wei Lin. Constrained regression for interval-valued data. Journal of Business & Economic Statistics, 31(4):473–490, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.818004.

Grossman:1986:OSD

[Gro86] Jean Baldwin Grossman. Optimal sample designs with preliminary tests of significance. Journal of Business & Economic Statistics, 4(2):171–176, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509512.

Gonzalez-Rivera:2011:ADS

[GRSY11] Gloria González-Rivera, Zeynep Senyuz, and Emre Yoldas. Autocontours: Dynamic specification testing. Journal of Business & Economic Statistics, 29(1):186–200, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2010.08144.

Graham:2022:TCA

[GRTZ22] Bryan S. Graham, Geert Ridder, Petra Thiemann, and Gema Zamarro. Teacher-to-classroom assignment and student achievement. Journal of Business & Economic Statistics, 41(4):1328– 1340, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2126480.

Godfrey:1984:MBS

[GRW84] James Godfrey, Alan Roshwalb, and Roger L. Wright. Modelbased stratification in inventory cost estimation. Journal of Business & Economic Statistics, 2(1):01–09, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1984.10509365.

Gregory:1991:TLR

[GS91a] Allan W. Gregory and Michael J. Sampson. Testing long-run properties of stationary time series. Journal of Business & Economic Statistics, 9(3):287–295, 1991. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1991.10509854.

Gregory:1991:CTI

[GS91b] Allan W. Gregory and Gregor W. Smith. Calibration as testing: Inference in simulated macroeconomic models. Journal of Business & Economic Statistics, 9(3):297–303, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509855.

Guay:2003:IIN

[GS03] Alain Guay and Olivier Scaillet. Indirect inference, nuisance parameter, and threshold moving average models. Journal of Business & Economic Statistics, 21(1):122–132, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102288618829.

Ghysels:2006:C

[GS06a] Eric Ghysels and Arthur Sinko. Comment. Journal of Business & Economic Statistics, 24(2):192–194, 2006. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500106000000080. See [HL06a, HL06b].

Guo:2006:IVS

[GS06b] Hui Guo and Robert Savickas. Idiosyncratic volatility, stock market volatility, and expected stock returns. Journal of Business & Economic Statistics, 24(1):43–56, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500105000000180.

Gourieroux:2010:DPW

[GS10] Christian Gourieroux and Razvan Sufana. Derivative pricing with Wishart multivariate stochastic volatility. Journal of Business & Economic Statistics, 28(3):438–451, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.08105.

Giudici:2016:GNM

[GS16] P. Giudici and A. Spelta. Graphical network models for international financial flows. Journal of Business & Economic Statistics, 34(1):128–138, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1017643.

Gordon:2004:ARS

[GSA04] Stephen Gordon and Pascal St-Amour. Asset returns and statedependent risk preferences. Journal of Business & Economic Statistics, 22(3):241–252, 2004. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000127.

Gordon:1996:BES

[GSC96] Stephen Gordon, Lucie Samson, and Benoît Carmichael. Bayesian estimation of stochastic discount factors. Journal of Business & Economic Statistics, 14(4):412–420, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524670.

Gelman:2003:RMM

[GSC03] Andrew Gelman, Matt Stevens, and Valerie Chan. Regression modeling and meta-analysis for decision making. Journal of Business & Economic Statistics, 21(2):213–225, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500103288618909.

Gurkaynak:2007:MBM

[GSS07] Refet S. Gürkaynak, Brian P. Sack, and Eric T. Swanson. Marketbased measures of monetary policy expectations. Journal of Business & Economic Statistics, 25(2):201–212, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000387.

Gregory:2001:TFC

[GSY01] Allan W. Gregory, Gregor W. Smith, and James Yetman. Testing for forecast consensus. Journal of Business & Economic Statistics, 19(1):34–43, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
07350010152472599.

Gurmu:1996:EZC

[GT96] Shiferaw Gurmu and Pravin K. Trivedi. Excess zeros in count models for recreational trips. Journal of Business & Economic Statistics, 14(4):469–477, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524676.

Gualdani:2023:DRP

[Gua23] Cristina Gualdani. Discussion of “Risk Preference Types, Limited Consideration, and Welfare” by Levon Barseghyan and Francesca Molinari. Journal of Business & Economic Statistics, 41(4):1035– 1038, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2023.2216255. See [BM23b].

Guo:1998:RPV

[Guo98] Dajiang Guo. The risk premium of volatility implicit in currency options. Journal of Business & Economic Statistics, 16 (4):498–507, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524789.

Gurmu:1991:TDO

[Gur91] Shiferaw Gurmu. Tests for detecting overdispersion in the positive Poisson regression model. Journal of Business & Economic Statistics, 9(2):215–222, 1991. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1991.10509847.

George:1985:EDC

[GW85] Edward I. George and William E. Wecker. Estimating damages in a class action litigation. Journal of Business & Economic Statistics, 3(2):132–139, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509441.

Gupta:1988:CEF

[GW88] Sunil Gupta and Peter C. Wilton. Combination of economic forecasts: an odds-matrix approach. Journal of Business & Economic Statistics, 6(3):373–379, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509674.

Ghysels:2009:FPF

[GW09] Eric Ghysels and Jonathan H. Wright. Forecasting professional forecasters. Journal of Business & Economic Statistics, 27(4):504–516, 2009. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.06044.

Ghysels:2014:MID

[GW14] Eric Ghysels and Fangfang Wang. Moment-implied densities: Properties and applications. Journal of Business & Economic Statistics, 32(1):88–111, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.847842.

Greenlees:1984:GTM

[GZ84] John S. Greenlees and Kimberly D. Zieschang. Grouping tests for misspecification: an application to housing demand. Journal of Business & Economic Statistics, 2(2): 159–169, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509384.

Haan:1996:HAU

[Haa96] Wouter J. Dend Haan. Heterogeneity, aggregate uncertainty, and the short-term interest rate. Journal of Business & Economic Statistics, 14(4):399–411, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524669.

Hadri:1999:EDH

[Had99] Kaddour Hadri. Estimation of a doubly heteroscedastic stochastic frontier cost function. Journal of Business & Economic Statistics, 17(3):359–363, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524824.

Haggstrom:1983:LRD

[Hag83] Gus W. Haggstrom. Logistic regression and discriminant analysis by ordinary least squares. Journal of Business & Economic Statistics, 1(3):229–238, 1983. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1983.10509346.

Hall:1990:LMT

[Hal90] Alastair Hall. Lagrange multiplier tests for normality against seminonparametric alternatives. Journal of Business & Economic Statistics, 8(4):417–426, 1990. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1990.10509812.

Haldrup:1994:STD

[Hal94a] Niels Haldrup. Semiparametric tests for double unit roots. Journal of Business & Economic Statistics, 12(1):109–122, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1994.10509994.

Hall:1994:TUR

[Hal94b] Alastair Hall. Testing for a unit root in time series with pretest data-based model selection. Journal of Business & Economic Statistics, 12(4):461–470, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10524568.

Hamilton:1991:C

[Ham91a] James D. Hamilton. Comment. Journal of Business & Economic Statistics, 9(4):388–389, 1991. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1991.10509865.

Hamilton:1991:QBA

[Ham91b] James D. Hamilton. A quasi-Bayesian approach to estimating parameters for mixtures of normal distributions. Journal of Business & Economic Statistics, 9(1):27–39, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509824.

Hamilton:1994:AEI

[Ham94] James D. Hamilton. Associate Editor’s introduction: Changes in regime and the business cycle. Journal of Business & Economic Statistics, 12(3):267, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10524541.

Hammond:2013:QCP

[Ham13] Robert G. Hammond. Quantifying consumer perception of a financially distressed company. Journal of Business & Economic Statistics, 31(4):398–411, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.799998.

Hansen:1986:C

[Han86] Lars Peter Hansen. Comment. Journal of Business & Economic Statistics, 4(4):418–421, 1986. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1986.10509539.

Hansen:1992:TPI

[Han92] Bruce E. Hansen. Tests for parameter instability in regressions with 1(1) processes. Journal of Business & Economic Statistics, 10(3):321–335, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509908.

Hansen:1993:C

[Han93] Bruce E. Hansen. Comment. Journal of Business & Economic Statistics, 11(4):385–386, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509969.

Hansen:1997:AAV

[Han97] Bruce E. Hansen. Approximate asymptotic P values for structural-change tests. Journal of Business & Economic Statistics, 15(1):60–67, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524687.

Hanson:2002:TPI

[Han02] Bruce E. Hanson. Tests for parameter instability in regressions with I(1) processes. Journal of Business & Economic Statistics, 20(1):45–59, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102753410381.

Hansen:2005:TSP

[Han05] Peter Reinhard Hansen. A test for superior predictive ability. Journal of Business & Economic Statistics, 23(4): 365–380, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500105000000063.

Hansen:2017:RKU

[Han17] Bruce E. Hansen. Regression kink with an unknown threshold. Journal of Business & Economic Statistics, 35(2):228–240, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1073595.

Han:2019:SEF

[Han19] Xu Han. Shrinkage estimation of factor models with global and group-specific factors. Journal of Business & Economic Statistics, 39(1):1–17, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1617157.

Hartman:1983:ESR

[Har83] Raymond S. Hartman. The estimation of short-run household electricity demand using pooled aggregate data. Journal of Business & Economic Statistics, 1(2):127–135, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1983.10509331.

Harvey:1985:TCM

[Har85] A. C. Harvey. Trends and cycles in macroeconomic time series. Journal of Business & Economic Statistics, 3(3): 216–227, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509453.

Hartman:1991:MCA

[Har91] Raymond S. Hartman. A Monte Carlo analysis of alternative estimators in models involving selectivity. Journal of Business & Economic Statistics, 9(1):41–49, 1991. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1991.10509825.

Hasbrouck:1987:NFD

[Has87] Joel Hasbrouck. A note on forecaster discord and consensus prediction error. Journal of Business & Economic Statistics, 5(1):151–154, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509570.

Haug:1991:CGB

[Hau91] Alfred A. Haug. Cointegration and government borrowing constraints: Evidence for the United States. Journal of Business & Economic Statistics, 9(1):97–101, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509831.

Haug:1996:BMC

[Hau96] Alfred A. Haug. Blanchard’s model of consumption: an empirical study. Journal of Business & Economic Statistics, 14 (2):169–177, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524643.

Hausman:1999:CTN

[Hau99] Jerry Hausman. Cellular telephone, new products, and the CPI. Journal of Business & Economic Statistics, 17(2): 188–194, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524809.

Hayes:1986:TOT

[Hay86] Kathy J. Hayes. Third-order translog utility functions. Journal of Business & Economic Statistics, 4(3):339–346, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1986.10509530.

Hayes:1989:STC

[Hay89] K. J. Hayes. A specification test for choosing the “right” public-good price. Journal of Business & Economic Statistics, 7(2):267–273, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509736.

Hayashi:1992:C

[Hay92] Fumio Hayashi. Comment. Journal of Business & Economic Statistics, 10(1):15–17, 1992. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1992.10509879.

Hill:1994:SPEa

[HC94a] R. Carter Hill and P. A. Cartwright. The statistical properties of the equity estimator. Journal of Business & Economic Statistics, 12(2):141–147, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10510000.

Hill:1994:SPEb

[HC94b] R. Carter Hill and P. A. Cartwright. The statistical properties of the equity estimator: a rejoinder. Journal of Business & Economic Statistics, 12(2):155, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10510002.

Han:2011:IDM

[HCP11] Chirok Han, Jin Seo Cho, and Peter C. B. Phillips. Infinite density at the median and the typical shape of stock return distributions. Journal of Business & Economic Statistics, 29(2):282–294, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2010.07327.

Heien:1985:TCL

[HD85] Dale Heien and James Dunn. The true cost-of-living index with changing preferences. Journal of Business & Economic Statistics, 3(4):332–335, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509465.

Hoga:2021:TEC

[HD21] Yannick Hoga and Timo Dimitriadis. On testing equal conditional predictive ability under measurement error. Journal of Business & Economic Statistics, 41(2):364–376, 2021. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
2021923.

Heien:1983:SUC

[Hei83] Dale Heien. Seasonality in U.S. consumer demand. Journal of Business & Economic Statistics, 1(4):280–284, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1983.10509352.

Heiler:2022:ECB

[Hei22] Phillip Heiler. Efficient covariate balancing for the local average treatment effect. Journal of Business & Economic Statistics, 40(4):1569–1582, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1946067.

Henzi:2022:CED

[Hen22] Alexander Henzi. Consistent estimation of distribution functions under increasing concave and convex stochastic ordering. Journal of Business & Economic Statistics, 41(4):1203– 1214, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2116026.

Hogue:1986:MGF

[HF86] Carma R. Hogue and Paul O. Flaim. Measuring gross flows in the labor force: an overview of a special conference. Journal of Business & Economic Statistics, 4(1):111–121, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509500.

Hung:1987:REC

[HF87] Hsien-Ming Hung and Wayne A. Fuller. Regression estimation of crop acreages with transformed Landsat data as auxiliary variables. Journal of Business & Economic Statistics, 5 (4):475–482, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509616.

Harvey:1989:R

[HF89a] A. C. Harvey and C. Fernandas. Reply. Journal of Business & Economic Statistics, 7(4):422, 1989. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1989.10509753.

Harvey:1989:TSM

[HF89b] A. C. Harvey and C. Fernandes. Time series models for count or qualitative observations. Journal of Business & Economic Statistics, 7(4):407–417, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509750.

Hill:1997:SIN

[HF97] Robert J. Hill and Kevin J. Fox. Splicing index numbers. Journal of Business & Economic Statistics, 15(3): 387–389, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524716.

Huber:2019:ASB

[HF19] Florian Huber and Martin Feldkircher. Adaptive shrinkage in Bayesian vector autoregressive models. Journal of Business & Economic Statistics, 37(1):27–39, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2016.1256217.

Ho:1998:C

[HfJL98] Hwai-Chung Ho and Chien fu Jeff Lin. Comment. Journal of Business & Economic Statistics, 16(3):272, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524763.

Hajargasht:2012:IID

[HGB^{+}12] Gholamreza Hajargasht, William E. Griffiths, Joseph Brice, D. S. Prasada Rao, and Duangkamon Chotikapanich. Inference for income distributions using grouped data. Journal of Business & Economic Statistics, 30(4):563–575, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2012.707590.

Huang:1983:ECF

[HH83] Kuo S. Huang and Richard C. Haidacher. Estimation of a composite food demand system for the United States. Journal of Business & Economic Statistics, 1(4):285–291, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1983.10509353.

Honore:2004:PSR

[HH04] Bo E. Honoré and Luojia Hu. On the performance of some robust instrumental variables estimators. Journal of Business & Economic Statistics, 22(1):30–39, 2004. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/073500103288619368.

Han:2011:TCI

[HH11a] Lu Han and Seung-Hyun Hong. Testing cost inefficiency under free entry in the real estate brokerage industry. Journal of Business & Economic Statistics, 29(4):564–578, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2011.08314.

Hendry:2011:CDF

[HH11b] David F. Hendry and Kirstin Hubrich. Combining disaggregate forecasts or combining disaggregate information to forecast an aggregate. Journal of Business & Economic Statistics, 29(2):216–227, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.07112.

Hansen:2016:EGM

[HH16] Peter Reinhard Hansen and Zhuo Huang. Exponential GARCH modeling with realized measures of volatility. Journal of Business & Economic Statistics, 34(2):269–287, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1038543.

Hardle:2019:ADH

[HH19a] Wolfgang Karl Härdle and Li-Shan Huang. Analysis of deviance for hypothesis testing in generalized partially linear models. Journal of Business & Economic Statistics, 37(2):322–333, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1330693.

Heufer:2019:HET

[HH19b] Jan Heufer and Per Hjertstrand. Homothetic efficiency: Theory and applications. Journal of Business & Economic Statistics, 37(2):235–247, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1319372.

Hafner:2021:DSD

[HH21] Christian M. Hafner and Helmut Herwartz. Dynamic scoredriven independent component analysis. Journal of Business & Economic Statistics, 41(2):298–308, 2021. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
2013244.

Hahn:2011:AED

[HHK11] Jinyong Hahn, Keisuke Hirano, and Dean Karlan. Adaptive experimental design using the propensity score. Journal of Business & Economic Statistics, 29(1):96–108, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.08161.

Han:2019:SMA

[HHK19] Xiaoyi Han, Chih-Sheng Hsieh, and Stanley I. M. Ko. Spatial modeling approach for dynamic network formation and interactions. Journal of Business & Economic Statistics, 39 (1):120–135, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1639395.

Hauzenberger:2022:FFB

[HHKO22] Niko Hauzenberger, Florian Huber, Gary Koop, and Luca Onorante. Fast and flexible Bayesian inference in time-varying parameter regression models. Journal of Business & Economic Statistics, 40(4):1904–1918, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1990772.

Hahn:2018:BFM

[HHL18] P. Richard Hahn, Jingyu He, and Hedibert Lopes. Bayesian factor model shrinkage for linear IV regression with many instruments. Journal of Business & Economic Statistics, 36(2):278– 287, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2016.1172968.

Han:2022:RSR

[HHL^{+}22] Dongxiao Han, Jian Huang, Yuanyuan Lin, Lei Liu, Lianqiang Qu, and Liuquan Sun. Robust signal recovery for highdimensional linear log-contrast models with compositional covariates. Journal of Business & Economic Statistics, 41(3):957– 967, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2097911.

Hansen:2008:EMI

[HHN08] Christian Hansen, Jerry Hausman, and Whitney Newey. Estimation with many instrumental variables. Journal of Business & Economic Statistics, 26(4):398–422, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500108000000024.

Haldrup:2007:CPC

[HHPS07] Niels Haldrup, Svend Hylleberg, Gabriel Pons, and Andreu Sansó. Common periodic correlation features and the interaction of stocks and flows in daily airport data. Journal of Business & Economic Statistics, 25(1):21–32, 2007. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/073500106000000459.

He:2019:SIR

[HHPS19] Yi He, Yanxi Hou, Liang Peng, and Jiliang Sheng. Statistical inference for a relative risk measure. Journal of Business & Economic Statistics, 37(2):301–311, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1321549.

Holmes:1991:FFF

[HHW91] James M. Holmes, Patricia A. Hutton, and Edward Weber. A functional-form-free test of the research and development/firm size relationship. Journal of Business & Economic Statistics, 9(1):85–90, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509829.

Hansen:1996:FSP

[HHY96] Lars Peter Hansen, John Heaton, and Amir Yaron. Finite-sample properties of some alternative GMM estimators. Journal of Business & Economic Statistics, 14(3):262–280, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524656.

Hess:1997:MCB

[HI97] Gregory D. Hess and Shigeru Iwata. Measuring and comparing business-cycle features. Journal of Business & Economic Statistics, 15(4):432–444, 1997. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1997.10524721.

Hyslop:2001:BCO

[HI01] Dean R. Hyslop and Guido W. Imbens. Bias from classical and other forms of measurement error. Journal of Business & Economic Statistics, 19(4):475–481, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/07350010152596727.

Hillmer:1985:MVM

[Hil85] Steven C. Hillmer. Measures of variability for model-based seasonal adjustment procedures. Journal of Business & Economic Statistics, 3(1):60–68, 1985. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1985.10509427.

Hill:1986:CBA

[Hil86] Bruce M. Hill. Comment: a Bayesian approach to the nonresponse problem, using covariates à la Freedman. Journal of Business & Economic Statistics, 4(1):125–126, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509502.

Hill:1989:C

[Hil89] Bruce M. Hill. Comment. Journal of Business & Economic Statistics, 7(1):29–30, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509708.

Hill:1991:C

[Hil91] Bruce M. Hill. Comment. Journal of Business & Economic Statistics, 9(2):149–150, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509836.

Ham:2000:TFI

[HJ00] John C. Ham and Kris Jacobs. Testing for full insurance using exogenous information. Journal of Business & Economic Statistics, 18(4):387–397, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524879.

Horrace:2022:LSF

[HJL22] William C. Horrace, Hyunseok Jung, and Yoonseok Lee. LASSO for stochastic frontier models with many efficient firms. Journal of Business & Economic Statistics, 41(4):1132–1142, 2022. CO-DEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2022.2110881.

Harbo:1998:AIC

[HJNR98] Ingrid Harbo, Søren Johansen, Bent Nielsen, and Anders Rahbek. Asymptotic inference on cointegrating rank in partial systems. Journal of Business & Economic Statistics, 16 (4):388–399, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524779.

Horrace:2020:NCT

[HJS20] William C. Horrace, Hyunseok Jung, and Shane Sanders. Network competition and team chemistry in the NBA. Journal of Business & Economic Statistics, 40(1):35–49, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1773273.

Hazilla:1986:SEC

[HK86] Michael Hazilla and Raymond J. Kopp. Systematic effects of capital service price definition on perceptions of input substitution. Journal of Business & Economic Statistics, 4(2): 209–224, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509516.

Harvey:1992:DCU

[HK92] Andrew C. Harvey and Siem Jan Koopman. Diagnostic checking of unobserved-components time series models. Journal of Business & Economic Statistics, 10(4):377–389, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1992.10509913.

Herriges:1994:RDE

[HK94] Joseph A. Herriges and Kathleen Kuester King. Residential demand for electricity under inverted block rates: Evidence from a controlled experiment. Journal of Business & Economic Statistics, 12(4):419–430, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10524564.

Hasegawa:2001:BAE

[HK01] Hikaru Hasegawa and Hideo Kozumi. Bayesian analysis on Engel curves estimation with measurement errors and an instrumental variable. Journal of Business & Economic Statistics, 19(3):292–298, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500101681019945.

Han:2014:ATQ

[HK14] Heejoon Han and Dennis Kristensen. Asymptotic theory for the QMLE in GARCH-x models with stationary and nonstationary covariates. Journal of Business & Economic Statistics, 32(3):416–429, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.897954.

Honore:2017:ESN

[HK17] Bo E. Honoré and Michaela Kesina. Estimation of some nonlinear panel data models with both time-varying and time-invariant explanatory variables. Journal of Business & Economic Statistics, 35(4):543–558, 2017. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1123635.

Hafner:2019:ETG

[HK19] Christian M. Hafner and Dimitra Kyriakopoulou. Exponentialtype GARCH models with linear-in-variance risk premium. Journal of Business & Economic Statistics, 39(2):589–603, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1691564.

Huisman:2001:TIE

[HKKP01] Ronald Huisman, Kees G. Koedijk, Clemens J. M. Kool, and Franz Palm. Tail-index estimates in small samples. Journal of Business & Economic Statistics, 19(2):208–216, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500101316970421.

Huber:2020:ISS

[HKO20] Florian Huber, Gary Koop, and Luca Onorante. Inducing sparsity and shrinkage in time-varying parameter models. Journal of Business & Economic Statistics, 39(3):669–683, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1713796.

Harvey:1997:MSA

[HKR97] Andrew Harvey, Siem Jan Koopman, and Marco Riani. The modeling and seasonal adjustment of weekly observations. Journal of Business & Economic Statistics, 15(3): 354–368, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524713.

He:2020:LDF

[HKYZ20] Yong He, Xinbing Kong, Long Yu, and Xinsheng Zhang. Largedimensional factor analysis without moment constraints. Journal of Business & Economic Statistics, 40(1):302–312, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1811101.

Hanssens:1983:LSR

[HL83] Dominique M. Hanssens and Lon-Mu Liu. Lag specification in rational distributed lag structural models. Journal of Business & Economic Statistics, 1(4):316–325, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1983.10509360.

Hart:2004:FCF

[HL04] Chad E. Hart and Sergio H. Lence. Financial constraints and farm investment. Journal of Business & Economic Statistics, 22(1):51–63, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619403.

Hansen:2006:RVM

[HL06a] Peter R. Hansen and Asger Lunde. Realized variance and market microstructure noise. Journal of Business & Economic Statistics, 24(2):127–161, 2006. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/

073500106000000071. See comments [ASMZ06, BR06, ABFN06, BNS06, Die06, GM06, GS06a, Oom06a, PY06] and rejoinder [HL06b].

Hansen:2006:R

[HL06b] Peter R. Hansen and Asger Lunde. Rejoinder. Journal of Business & Economic Statistics, 24(2):208–218, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000170. See

[HL06a, ASMZ06, BR06, ABFN06, BNS06, Die06, GM06, GS06a, Oom06a, PY06].

Han:2016:BAS

[HL16] Xiaoyi Han and Lung-Fei Lee. Bayesian analysis of spatial panel autoregressive models with time-varying endogenous spatial weight matrices, common factors, and random coefficients.

Journal of Business & Economic Statistics, 34(4):642–660, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1167058.

Horowitz:2022:ICO

[HL22] Joel L. Horowitz and Sokbae Lee. Inference in a class of optimization problems: Confidence regions and finite sample bounds on errors in coverage probabilities. Journal of Business & Economic Statistics, 41(3):927–938, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2093883.

Hsu:2020:CTE

[HLL20] Yu-Chin Hsu, Tsung-Chih Lai, and Robert P. Lieli. Counterfactual treatment effects: Estimation and inference. Journal of Business & Economic Statistics, 40(1):240–255, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1800479.

Harris:2005:PST

[HLM05] David Harris, Stephen Leybourne, and Brendan McCabe. Panel stationarity tests for purchasing power parity with cross-sectional dependence. Journal of Business & Economic Statistics, 23 (4):395–409, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500105000000090.

Hurn:2015:EPS

[HLM15] A. S. Hurn, K. A. Lindsay, and A. J. McClelland. Estimating the parameters of stochastic volatility models using option price data. Journal of Business & Economic Statistics, 33(4):579–594, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.981634.

Huber:2016:FSP

[HLM16] Martin Huber, Michael Lechner, and Giovanni Mellace. The finite sample performance of estimators for mediation analysis under sequential conditional independence. Journal of Business & Economic Statistics, 34(1):139–160, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1017644.

Harvey:1998:TFE

[HLN98] David I. Harvey, Stephen J. Leybourne, and Paul Newbold. Tests for forecast encompassing. Journal of Business & Economic Statistics, 16(2):254–259, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524759.

Hurlin:2017:RMI

[HLQS17] Christophe Hurlin, Sébastien Laurent, Rogier Quaedvlieg, and Stephan Smeekes. Risk measure inference. Journal of Business & Economic Statistics, 35(4):499–512, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1127815.

Hay:1987:OLS

[HLR87] Joel W. Hay, Robert Leu, and Paul Rohrer. Ordinary least squares and sample-selection models of health-care demand Monte Carlo comparison. Journal of Business & Economic Statistics, 5(4):499–506, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509618.

Horvath:2022:TSF

[HLR^{+}22] Lajos Horváth, Zhenya Liu, Gregory Rice, Shixuan Wang, and Yaosong Zhan. Testing stability in functional event observations with an application to IPO performance. Journal of Business & Economic Statistics, 41(4):1262–1273, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2118127.

Huang:2014:FSU

[HLW14] Danyang Huang, Runze Li, and Hansheng Wang. Feature screening for ultrahigh dimensional categorical data with applications. Journal of Business & Economic Statistics, 32(2):237–244, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2013.
863158.

Huang:2014:EMG

[HLWY14] Mian Huang, Runze Li, Hansheng Wang, and Weixin Yao. Estimating mixture of Gaussian processes by kernel smoothing. Journal of Business & Economic Statistics, 32(2):259–270, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2013.868084.

Hong:2004:SPD

[HLZ04] Yongmiao Hong, Haitao Li, and Feng Zhao. Out-of-sample performance of discrete-time spot interest rate models. Journal of Business & Economic Statistics, 22(4):457–473, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000433.

Huang:2022:UFS

[HLZ22] Wei Huang, Oliver Linton, and Zheng Zhang. A unified framework for specification tests of continuous treatment effect models. Journal of Business & Economic Statistics, 40(4):1817– 1830, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1981915.

Hall:1989:MCS

[HM89] A. D. Hall and Michael McAleer. A Monte Carlo study of some tests of model adequacy in time series analysis. Journal of Business & Economic Statistics, 7(1):95–106, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509717.

Hendry:2007:CB

[HM07] David F. Hendry and Michael Massmann. Co-breaking. Journal of Business & Economic Statistics, 25(1):33–51, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000422.

Hubrich:2014:C

[HM14] Kirstin Hubrich and Simone Manganelli. Comment. Journal of Business & Economic Statistics, 32(4):506–509, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2014.956874.

See [AGO^{+}14a, AGO^{+}14b].

Hillebrand:2016:NBL

[HM16] Eric Hillebrand and Marcelo C. Medeiros. Nonlinearity, breaks, and long-range dependence in time-series models. Journal of Business & Economic Statistics, 34(1):23–41, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.985828.

Hsiao:1995:BIE

[HMI95] Cheng Hsiao, Dean C. Mountain, and Kathleen Ho Illman. A Bayesian integration of end-use metering and conditionaldemand analysis. Journal of Business & Economic Statistics, 13(3):315–326, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524605.

Hansen:2010:IVE

[HMN10] Christian Hansen, James B. McDonald, and Whitney K. Newey. Instrumental variables estimation with flexible distributions. Journal of Business & Economic Statistics, 28(1):13–25, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2009.06161.

Horvath:2020:NCC

[HMR20] Lajos Horváth, Curtis Miller, and Gregory Rice. A new class of change point test statistics of Rényi type. Journal of Business & Economic Statistics, 38(3):570–579, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2018.1537923.

Hahn:2017:LTN

[HMS17] Jinyong Hahn, Hyungsik Roger Moon, and Connan Snider. LM test of neglected correlated random effects and its application. Journal of Business & Economic Statistics, 35(3):359–370, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1063426.

Horny:2012:JDW

[HMvdB12] Guillaume Horny, Rute Mendes, and Gerard J. van den Berg. Job durations with worker- and firm-specific effects: MCMC estimation with longitudinal employer–employee data. Journal of Business & Economic Statistics, 30(3):468–480, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2012.698142.

Hay:1984:LTE

[HO84] Joel W. Hay and Randall J. Olsen. Let them eat cake: a note on comparing alternative models of the demand for medical care. Journal of Business & Economic Statistics, 2 (3):279–282, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509395.

Hoffman:1991:TSR

[Hof91] Dennis L. Hoffman. Two-step and related estimators in contemporary rational-expectations models: an analysis of smallsample properties. Journal of Business & Economic Statistics, 9(1):51–61, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509826.

Hogan:1983:C

[Hog83] William W. Hogan. Comment. Journal of Business & Economic Statistics, 1(1):33, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/

07350015.1983.10509319. See [FRS83a, FRS83b, Smi83].

Hogg:1988:C

[Hog88] Robert V. Hogg. Comment. Journal of Business & Economic Statistics, 6(4):428, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509684.

Hoga:2019:CIC

[Hog19] Yannick Hoga. Confidence intervals for conditional tail risk measures in ARMA-GARCH models. Journal of Business & Economic Statistics, 37(4):613–624, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1401543.

Hoga:2022:EDB

[Hog22] Yannick Hoga. Extremal dependence-based specification testing of time series. Journal of Business & Economic Statistics, 41(4):1274–1287, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2120483.

Hong:1991:C

[Hon91] Chansik Hong. Comment. Journal of Business & Economic Statistics, 9(2):150–152, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509837.

Hong:2005:C

[Hon05] Han Hong. Comment. Journal of Business & Economic Statistics, 23(2):158–160, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000695. See [AV05b, AV05a].

Hong:2008:C

[Hon08] Han Hong. Comment. Journal of Business & Economic Statistics, 26(3):292–294, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/

073500108000000132. See [ALT08a, ALT08b].

Honka:2023:DRP

[Hon23] Elisabeth Honka. Discussion of “Risk Preference Types, Limited Consideration, and Welfare” by Levon Barseghyan and Francesca Molinari. Journal of Business & Economic Statistics, 41(4):1039– 1041, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2023.2217870. See [BM23b].

Horgan:1998:SSS

[Hor98] Jane M. Horgan. Stabilized sieve sampling: a point-estimator analysis. Journal of Business & Economic Statistics, 16 (1):42–51, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524733.

Hoshino:2018:SSA

[Hos18] Tadao Hoshino. Semiparametric spatial autoregressive models with endogenous regressors: With an application to crime data. Journal of Business & Economic Statistics, 36(1):160–172, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1146145.

Hoshino:2019:TSE

[Hos19] Tadao Hoshino. Two-step estimation of incomplete information social interaction models with sample selection. Journal of Business & Economic Statistics, 37(4):598–612, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1394861.

Hinich:1985:END

[HP85] Melvin J. Hinich and Douglas M. Patterson. Evidence of nonlinearity in daily stock returns. Journal of Business & Economic Statistics, 3(1):69–77, 1985. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1985.10509428.

Hendershott:1989:HSU

[HP89] Patric H. Hendershott and Joe Peek. Household saving in the United States: Measurement and behavior. Journal of Business & Economic Statistics, 7(1):11–19, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509705.

Herrera:2005:DUO

[HP05] Ana María Herrera and Elena Pesavento. The decline in U.S. output volatility. Journal of Business & Economic Statistics, 23(4):462–472, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000596.

Hong:2010:TLC

[HP10] Seung Hyun Hong and Peter C. B. Phillips. Testing linearity in cointegrating relations with an application to purchasing power parity. Journal of Business & Economic Statistics, 28(1):96–114, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1198/jbes.2009.07182.

Harding:2011:EAS

[HP11] Don Harding and Adrian Pagan. An econometric analysis of some models for constructed binary time series. Journal of Business & Economic Statistics, 29(1):86–95, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.08005.

Hautsch:2013:PBE

[HP13] Nikolaus Hautsch and Mark Podolskij. Preaveraging-based estimation of quadratic variation in the presence of noise and jumps: Theory, implementation, and empirical evidence. Journal of Business & Economic Statistics, 31(2):165–183, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2012.754313.

Hayes:1987:DLT

[HPH87] K. Hayes and S. Porter-Hudak. Deadweight loss: Theoretical size relationships and the precision of measurement. Journal of Business & Economic Statistics, 5(1):47–52, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509559.

Hung-pin:2023:PSF

[HpK23] Lai Hung-pin and Subal C. Kumbhakar. Panel stochastic frontier model with endogenous inputs and correlated random components. Journal of Business & Economic Statistics, 41(1):80– 96, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.2001341.

Hoogstrate:2000:PDP

[HPP00] André J. Hoogstrate, Franz C. Palm, and Gerard A. Pfann. Pooling in dynamic panel-data models: an application to forecasting GDP growth rates. Journal of Business & Economic Statistics, 18(3):274–283, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524870.

Ho:1996:CTA

[HPS96] Mun S. Ho, William R. M. Perraudin, and Bent E. Sørensen. A continuous-time arbitrage-pricing model with stochastic volatility and jumps. Journal of Business & Economic Statistics, 14(1):31–43, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524627.

Hamada:1988:RSA

[HPSW88] Robert Hamada, James M. Patell, Richard Staelin, and William E. Wecker. The role of statistics in accounting, marketing, finance, and production. Journal of Business & Economic Statistics, 6(2):261–272, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509660.

He:2021:RAG

[HPZZ21] Yi He, Liang Peng, Dabao Zhang, and Zifeng Zhao. Risk analysis via generalized Pareto distributions. Journal of Business & Economic Statistics, 40(2):852–867, 2021. CODEN RFSCFJ.

ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2021.
1874390.

Hall:1991:ESA

[HR91] Alastair Hall and Robert J. Rossana. Estimating the speed of adjustment in partial adjustment models. Journal of Business & Economic Statistics, 9(4):441–453, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509871.

Harvey:1994:C

[HR94] Andrew C. Harvey and Esther Ruiz. Comment. Journal of Business & Economic Statistics, 12(4):402–403, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10524559.

Hoonhout:2019:NAM

[HR19] Pierre Hoonhout and Geert Ridder. Nonignorable attrition in multi-period panels with refreshment samples. Journal of Business & Economic Statistics, 37(3):377–390, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1345744.

Hordahl:2020:ERP

[HRV20] Peter Hördahl, Eli M. Remolona, and Giorgio Valente. Expectations and risk premia at 8:30 a.m.: Deciphering the responses of bond yields to macroeconomic announcements. Journal of Business & Economic Statistics, 38(1):27–42, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2018.1429278.

Hoogerheide:2012:C

[HRvD12] Lennart Hoogerheide, Francesco Ravazzolo, and Herman K. van Dijk. Comment. Journal of Business & Economic Statistics, 30(1):30–33, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.634348. See [PT12a, PT12b].

Haynes:1985:NMS

[HS85] Stephen E. Haynes and Joe A. Stone. Neglected method of separating demand and supply in time series regression. Journal of Business & Economic Statistics, 3(3):238–243, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1985.10509455.

Haynes:1988:R

[HS88] Stephen E. Haynes and Joe A. Stone. Reply. Journal of Business & Economic Statistics, 6(1):138–140, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509646.

Hughes:1989:PAM

[HS89] James W. Hughes and Edward A. Snyder. Policy analysis of medical malpractice reforms: What can we learn from claims data? Journal of Business & Economic Statistics, 7 (4):423–431, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509754.

Hidiroglou:1993:PAD

[HS93] M. A. Hidiroglou and K. P. Srinath. Problems associated with designing subannual business surveys. Journal of Business & Economic Statistics, 11(4):397–405, 1993. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1993.10509973.

Hansen:1996:EEL

[HS96a] Lars Peter Hansen and Kenneth J. Singleton. Efficient estimation of linear asset-pricing models with moving average errors. Journal of Business & Economic Statistics, 14 (1):53–68, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524629.

Harvey:1996:EAS

[HS96b] Andrew C. Harvey and Neil Shephard. Estimation of an asymmetric stochastic volatility model for asset returns. Journal of Business & Economic Statistics, 14(4):429–434, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524672.

Holmes:1996:NBB

[HS96c] Thomas J. Holmes and James A. Schmitz. Nonresponse bias and business turnover rates: The case of the characteristics of business owners survey. Journal of Business & Economic Statistics, 14(2):231–241, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524650.

Hall:1999:SST

[HS99a] Alastair R. Hall and Amit Sen. Structural stability testing in models estimated by generalized method of moments. Journal of Business & Economic Statistics, 17(3):335–348, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524822.

Hong:1999:NTA

[HS99b] Yongmlao Hong and Ramsey D. Shehadeh. A new test for ARCH effects and its finite-sample performance. Journal of Business & Economic Statistics, 17(1):91–108, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524799.

Hodge:2016:SVT

[HS16] Andrew Hodge and Sriram Shankar. Single-variable threshold effects in ordered response models with an application to estimating the income-happiness gradient. Journal of Business & Economic Statistics, 34(1):42–52, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.991785.

Hansen:2019:DMV

[HS19a] Peter Reinhard Hansen and Matthias Schmidtblaicher. A dynamic model of vaccine compliance: How fake news undermined the Danish HPV vaccine program. Journal of Business & Economic Statistics, 39(1):259–271, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2019.1623045.

Huber:2019:FSI

[HS19b] Martin Huber and Andreas Steinmayr. A framework for separating individual-level treatment effects from spillover effects. Journal of Business & Economic Statistics, 39(2):422–436, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1668795.

Huang:1987:ESU

[HSA87] Cliff J. Huang, Frank A. Sloan, and Killard W. Adamache. Estimation of seemingly unrelated Tobit regressions via the EM algorithm. Journal of Business & Economic Statistics, 5 (3):425–430, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509607.

Hsieh:1989:MHD

[Hsi89] David A. Hsieh. Modeling heteroscedasticity in daily foreignexchange rates. Journal of Business & Economic Statistics, 7 (3):307–317, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509740.

Huber:2020:DIE

[HSS20] Martin Huber, Mark Schelker, and Anthony Strittmatter. Direct and indirect effects based on changes-in-changes. Journal of Business & Economic Statistics, 40(1):432–443, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1831929.

Hsiao:2022:TEP

[HSZ22a] Cheng Hsiao, Zhentao Shi, and Qiankun Zhou. Transformed estimation for panel interactive effects models. Journal of Business & Economic Statistics, 40(4):1831–1848, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1983438.

Huang:2022:DUH

[HSZ22b] Wenxin Huang, Liangjun Su, and Yuan Zhuang. Detecting unobserved heterogeneity in efficient prices via classifierlasso. Journal of Business & Economic Statistics, 41(2):509– 522, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2036613.

Harvey:1983:FET

[HT83a] A. C. Harvey and P. H. J. Todd. Forecasting economic time series with structural and Box–Jenkins models: a case study. Journal of Business & Economic Statistics, 1(4):

299–307, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509355.

Harvey:1983:R

[HT83b] A. C. Harvey and P. H. J. Todd. Response. Journal of Business & Economic Statistics, 1(4):313–315, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1983.10509359.

Hill:2006:SEW

[HT06] Robert J. Hill and Marcel P. Timmer. Standard errors as weights in multilateral price indexes. Journal of Business & Economic Statistics, 24(3):366–377, 2006. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500105000000270.

Hu:2014:PVC

[HT14a] Yu-Pin Hu and Ruey S. Tsay. Principal volatility component analysis. Journal of Business & Economic Statistics, 32(2):153–164, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.

1080/07350015.2013.818006. See comments [Lin14, Yao14a, YL14, AG14, Fra14, HW14, McA14] and rejoinder [HT14b].

Hu:2014:R

[HT14b] Yu-Pin Hu and Ruey S. Tsay. Rejoinder. Journal of Business & Economic Statistics, 32(2):176–177, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.902236. See [HT14a].

Hansen:2015:C

[HT15] Peter Reinhard Hansen and Allan Timmermann. Comment. Journal of Business & Economic Statistics, 33(1):17–21, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2014.983601.

See [AGO^{+}14a, AGO^{+}14b].

Hyslop:2020:EDM

[HT20] Dean R. Hyslop and Wilbur Townsend. Earnings dynamics and measurement error in matched survey and administrative data. Journal of Business & Economic Statistics, 38(2):457–469, 2020.

ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1514308.

Horvath:2022:CDH

[HT22] Lajos Horváth and Lorenzo Trapani. Changepoint detection in heteroscedastic random coefficient autoregressive models. Journal of Business & Economic Statistics, 41(4):1300– 1314, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2120485.

Huber:2015:CPD

[Hub15] Martin Huber. Causal pitfalls in the decomposition of wage gaps. Journal of Business & Economic Statistics, 33(2):179–191, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2014.
937437.

Hanssens:1987:TSS

[HV87] Dominique M. Hanssens and Pierre M. Vanden Abeele. A timeseries study of the formation and predictive performance of EEC production survey expectations. Journal of Business & Economic Statistics, 5(4):507–519, 1987. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1987.10509619.

Hodgson:2003:EEC

[HV03] Douglas J. Hodgson and Keith P. Vorkink. Efficient estimation of conditional asset-pricing models. Journal of Business & Economic Statistics, 21(2):269–283, 2003. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500103288618954.

Hirschey:1984:AMV

[HW84] Mark Hirschey and Dean W. Wichern. Accounting and marketvalue measures of profitability: Consistency, determinants, and uses. Journal of Business & Economic Statistics, 2 (4):375–383, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509411.

Heien:1990:DSE

[HW90]  Dale Heien and Cathy Roheim Wesseils.  Demand sys-

tems estimation with microdata: a censored regression approach. Journal of Business & Economic Statistics, 8(3): 365–371, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509807.

Hassler:1995:LMI

[HW95] Uwe Hassler and Jürgen Wolters. Long memory in inflation rates: International evidence. Journal of Business & Economic Statistics, 13(1):37–45, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524577.

Hansen:2002:GMM

[HW02] Bruce E. Hansen and Kenneth D. West. Generalized method of moments and macroeconomics. Journal of Business & Economic Statistics, 20(4):460–469, 2002. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102288618603.

Houser:2004:HDB

[HW04] Daniel Houser and Joachim Winter. How do behavioral assumptions affect structural inference? Evidence from a laboratory experiment. Journal of Business & Economic Statistics, 22(1):64–79, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619386.

Hurvich:2010:PJT

[HW10] Clifford M. Hurvich and Yi Wang. A pure-jump transactionlevel price model yielding cointegration. Journal of Business & Economic Statistics, 28(4):539–558, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.07116.

Hirano:2011:ER

[HW11] Keisuke Hirano and Jonathan Wright. Editors’ report 2011. Journal of Business & Economic Statistics, 29(4):597, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2011.294er.

Hardle:2014:C

[HW14] Wolfgang Karl Härdle and Weining Wang. Comment. Journal of Business & Economic Statistics, 32(2):173–174, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2014.898585.

See [GPI13b, GPI13a].

Hirshberg:2020:DIA

[HW20a] David A. Hirshberg and Stefan Wager. Debiased inference of average partial effects in single-index models: Comment on Wooldridge and Zhu. Journal of Business & Economic Statistics, 38(1):19–24, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2019.1681277. See [WZ20a, WZ20b].

Horrace:2020:SPP

[HW20b] William C. Horrace and Ian A. Wright. Stationary points for parametric stochastic frontier models. Journal of Business & Economic Statistics, 38(3):516–526, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2018.1526088.

Hylleberg:1996:C

[Hyl96] Svend Hylleberg. Comment. Journal of Business & Economic Statistics, 14(3):388–389, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524665.

Hylleberg:1998:C

[Hyl98] Svend Hylleberg. Comment. Journal of Business & Economic Statistics, 16(2):167–168, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524749.

Huang:2016:SMS

[HYSW16] Danyang Huang, Jun Yin, Tao Shi, and Hansheng Wang. A statistical model for social network labeling. Journal of Business & Economic Statistics, 34(3):368–374, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1039014.

Huang:2022:LCQ

[HZ22] Xiao Huang and Zhaoguo Zhan. Local composite quantile regression for regression discontinuity. Journal of Business & Economic Statistics, 40(4):1863–1875, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1990072.

Ilmakunnas:1990:FPC

[Ilm90] Pekka Ilmakunnas. Forecast pretesting and correction. Journal of Business & Economic Statistics, 8(4):475–480, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1990.10509819.

Idson:1997:IDS

[IM97] Todd Idson and Cynthia Miller. The implications of demographicspecific inflation rates for trends in real educational wage differentials. Journal of Business & Economic Statistics, 15 (4):464–469, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524724.

Ibragimov:2010:SBC

[IM10] Rustam Ibragimov and Ulrich K. Müller. t-statistic based correlation and heterogeneity robust inference. Journal of Business & Economic Statistics, 28(4):453–468, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.08046.

Imbens:2002:GMM

[Imb02] Guido W. Imbens. Generalized method of moments and empirical likelihood. Journal of Business & Economic Statistics, 20(4):493–506, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618630.

Inclan:1993:DMC

[Inc93] Carla Inclán. Detection of multiple changes of variance using posterior odds. Journal of Business & Economic Statistics, 11(3):289–300, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509956.

Ingram:1990:EMA

[Ing90a] Beth Fisher Ingram. Equilibrium modeling of asset prices: Rationality versus rules of thumb. Journal of Business & Economic Statistics, 8(1):115–125, 1990. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1990.10509781.

Ingram:1990:SSG

[Ing90b] Beth Fisher Ingram. Solving the stochastic growth model by backsolving with an expanded shock space. Journal of Business & Economic Statistics, 8(1):37–38, 1990. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1990.10509772.

Inoue:2015:C

[Ino15] Atsushi Inoue. Comment. Journal of Business & Economic Statistics, 33(1):9–11, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.969428. See [Die15a, Die15b].

Iacone:2021:STO

[INT21] Fabrizio Iacone, Morten Ørregaard Nielsen, and A. M. Robert Taylor. Semiparametric tests for the order of integration in the possible presence of level breaks. Journal of Business & Economic Statistics, 40(2):880–896, 2021. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2021.
1876712.

Ioannatos:1995:CRE

[Ioa95] Petros E. Ioannatos. Censored regression estimation under unobserved heterogeneity: a stochastic parameter approach. Journal of Business & Economic Statistics, 13(3): 327–335, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524606.

Israel:1990:UCE

[IOS90] Ronen Israel, Aharon R. Ofer, and Daniel R. Siegel. The use of changes in equity value as a measure of the information content of announcements of changes in financial policy. Journal of Business & Economic Statistics, 8(2):209–216, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1990.10509792.

Iglesias:2012:AUE

[IP12] Emma M. Iglesias and Garry D. A. Phillips. Almost unbiased estimation in simultaneous equation models with strong and/or weak instruments. Journal of Business & Economic Statistics, 30(4):505–520, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.715959.

Inoue:2005:RPT

[IR05] Atsushi Inoue and Barbara Rossi. Recursive predictability tests for real-time data. Journal of Business & Economic Statistics, 23(3):336–345, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000668.

Iacopini:2022:PSR

[IRR22] Matteo Iacopini, Francesco Ravazzolo, and Luca Rossini. Proper scoring rules for evaluating density forecasts with asymmetric loss functions. Journal of Business & Economic Statistics, 41(2):482– 496, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2035229.

Ignatieva:2015:EAA

[IRS15] Katja Ignatieva, Paulo Rodrigues, and Norman Seeger. Empirical analysis of affine versus nonaffine variance specifications in jump-diffusion models for equity indices. Journal of Business & Economic Statistics, 33(1):68–75, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.922471.

Isaki:1990:SAE

[Isa90] Cary T. Isaki. Small-area estimation of economic statistics. Journal of Business & Economic Statistics, 8(4):435–441, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1990.10509814.

Ishihara:2022:PDQ

[Ish22] Takuya Ishihara. Panel data quantile regression for treatment effect models. Journal of Business & Economic Statistics, 41(3):

720–736, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2061495.

Iskrev:2020:SIM

[Isk20] Nikolay Iskrev. On the sources of information in the moment structure of dynamic macroeconomic models. Journal of Business & Economic Statistics, 40(1):272–284, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1803079.

Imbens:1995:ECC

[IV95] Guido Imbens and Wilbert Van Der Klaauw. Evaluating the cost of conscription in the Netherlands. Journal of Business & Economic Statistics, 13(2):207–215, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524595.

Iversen:2001:SDR

[Ive01] Edwin S. Iversen, Jr. Spatially disaggregated real estate indices. Journal of Business & Economic Statistics, 19(3): 341–357, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500101681019990.

Iwata:1991:NRR

[Iwa91] Shigeru Iwata. A note on reverse regression, collinearity, and employment discrimination. Journal of Business & Economic Statistics, 9(3):341–342, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509861.

Jabine:1985:R

[Jab85] Thomas B. Jabine. Reply. Journal of Business & Economic Statistics, 3(4):402–404, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509478.

Jacobs:2000:ACP

[Jac00] Kris Jacobs. Aggregate consumption and the predictability of asset returns. Journal of Business & Economic Statistics, 18(1):58–76, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524848.

Jackson:2013:C

[Jac13] Matthew O. Jackson. Comment. Journal of Business & Economic Statistics, 31(3):270–273, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.794095. See [GPI13b, GPI13a].

Jaeger:1997:RON

[Jae97] David A. Jaeger. Reconciling the old and new Census Bureau education questions: Recommendations for researchers. Journal of Business & Economic Statistics, 15(3):300–309, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1997.10524708.

Jain:1989:SAP

[Jai89] Raj K. Jain. The seasonal adjustment procedures for the consumer price indexes: Some empirical results. Journal of Business & Economic Statistics, 7(4):461–469, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509758.

Jain:2019:PIP

[Jai19] Monica Jain. Perceived inflation persistence. Journal of Business & Economic Statistics, 37(1):110–120, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1281814.

Janson:1988:CRT

[Jan88a] Marius A. Janson. Combining robust and traditional least squares methods: a critical evaluation. Journal of Business & Economic Statistics, 6(4):415–427, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509683.

Janson:1988:R

[Jan88b] Marius A. Janson. Reply. Journal of Business & Economic Statistics, 6(4):450–451, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509689.

Jarque:1987:SSA

[Jar87] Carlos M. Jarque. Sample splitting and applied econometric modeling. Journal of Business & Economic Statistics, 5 (2):267–274, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509585.

Jefferson:2008:EAC

[Jef08] Philip N. Jefferson. Educational attainment and the cyclical sensitivity of employment. Journal of Business & Economic Statistics, 26(4):526–535, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500108000000060.

Ju:2019:NPE

[JGL19] Gaosheng Ju, Li Gan, and Qi Li. Nonparametric panel estimation of labor supply. Journal of Business & Economic Statistics, 37(2):260–274, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1321546.

Jaeger:2020:CTE

[JJK20] David A. Jaeger, Theodore J. Joyce, and Robert Kaestner. A cautionary tale of evaluating identifying assumptions: Did reality TV really cause a decline in teenage childbearing? Journal of Business & Economic Statistics, 38(2):317–326, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2018.1497510.

Ji:2021:CCC

[JJKL21] Pengsheng Ji, Jiashun Jin, Zheng Tracy Ke, and Wanshan Li. Co-citation and co-authorship networks of statisticians. Journal of Business & Economic Statistics, 40(2):469–485, 2021. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2021.1978469. See discussion [WF22, Don22,

MLZ22, ZK22, LC22] and rejoinder [JJKL22].

Ji:2022:RCC

[JJKL22] Pengsheng Ji, Jiashun Jin, Zheng Tracy Ke, and Wanshan Li. Rejoinder: “Co-citation and Co-authorship Networks of Statisticians”. Journal of Business & Economic Statistics, 40(2):499–504, 2022. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2022.2055358. See [JJKL21].

Jondeau:2018:MCA

[JJR18] Eric Jondeau, Emmanuel Jurczenko, and Michael Rockinger. Moment component analysis: an illustration with international stock markets. Journal of Business & Economic Statistics, 36(4):576–598, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1216851.

Jiang:2002:ECT

[JK02] George J. Jiang and John L. Knight. Estimation of continuoustime processes via the empirical characteristic function. Journal of Business & Economic Statistics, 20(2):198–212, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500102317351958.

Jentsch:2022:AVB

[JL22] Carsten Jentsch and Kurt G. Lunsford. Asymptotically valid bootstrap inference for proxy SVARs. Journal of Business & Economic Statistics, 40(4):1876–1891, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1990770.

Jondeau:2004:AGM

[JLG04] Eric Jondeau, Hervé Le Bihan, and Clémentine Gallès. Assessing generalized method-of-moments estimates of the Federal Reserve reaction function. Journal of Business & Economic Statistics, 22(2):225–239, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000118.

Jing:2014:EIV

[JLK14] Bing-Yi Jing, Zhi Liu, and Xin-Bing Kong. On the estimation of integrated volatility with jumps and microstructure noise. Journal of Business & Economic Statistics, 32(3):457–467, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2014.906350.

Jiang:2023:TMT

[JLP23] Lei Jiang, Weimin Liu, and Liang Peng. Test for market timing using daily fund returns. Journal of Business & Economic Statistics, 41(1):184–196, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.2006670.

Jung:2011:DFM

[JLR11] Robert C. Jung, Roman Liesenfeld, and Jean-François Richard. Dynamic factor models for multivariate count data: an application to stock-market trading activity. Journal of Business & Economic Statistics, 29(1):73–85, 2011. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.08212.

Jun:2016:TEU

[JLS16] Sung Jae Jun, Yoonseok Lee, and Youngki Shin. Treatment effects with unobserved heterogeneity: a set identification approach. Journal of Business & Economic Statistics, 34(2):302–311, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1044008.

Jeong:1991:MET

[JM91] Jinook Jeong and G. S. Maddala. Measurement errors and tests for rationality. Journal of Business & Economic Statistics, 9 (4):431–439, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509870.

Jun:2016:TDM

[JMP16] Duk B. Jun, Jihwan Moon, and Sungho Park. Temporal disaggregation: Methods, information loss, and diagnostics. Journal of Business & Economic Statistics, 34(1):53–61, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2014.995797.

James:2018:SVM

[JMZ18] Lancelot F. James, Gernot Müller, and Zhiyuan Zhang. Stochastic volatility models based on OU-gamma time change: Theory and estimation. Journal of Business & Economic Statistics, 36(1):75–87, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1133427.

Jobson:1988:C

[Job88] J. D. Jobson. Comment. Journal of Business & Economic Statistics, 6(1):16–20, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509630. See [CR88b, CR88a].

Jochmans:2018:SAN

[Joc18] Koen Jochmans. Semiparametric analysis of network formation. Journal of Business & Economic Statistics, 36(4):705–713, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1286242.

Johnson:1987:DIO

[Joh87] Wesley Johnson. The detection of influential observations for allocation, separation, and the determination of probabilities in a Bayesian framework. Journal of Business & Economic Statistics, 5(3):369–381, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509601.

Johnson:2002:VMT

[Joh02] Timothy C. Johnson. Volatility, momentum, and time-varying skewness in foreign exchange returns. Journal of Business & Economic Statistics, 20(3):390–411, 2002. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500102288618522.

Johansen:2004:C

[Joh04] Søren Johansen. Comment. Journal of Business & Economic Statistics, 22(2):169–172, 2004. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/043500104000000046. See [PSW04a] and rejoinder

[PSW04b].

Jorda:1999:RTA

[Jor99] Oscar Jordá. Random-time aggregation in partial adjustment models. Journal of Business & Economic Statistics, 17 (3):382–395, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524827.

Jacquier:1994:BAS

[JPR94a] Eric Jacquier, Nicholas G. Polson, and Peter E. Rossi. Bayesian analysis of stochastic volatility models. Journal of Business & Economic Statistics, 12(4):371–389, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10524553.

Jacquier:1994:R

[JPR94b] Eric Jacquier, Nicholas G. Polson, and Peter E. Rossi. Reply. Journal of Business & Economic Statistics, 12(4): 413–417, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10524563.

Jacquier:2002:BAS

[JPR02] Eric Jacquier, Nicholas G. Polson, and Peter E. Rossi. Bayesian analysis of stochastic volatility models. Journal of Business & Economic Statistics, 20(1):69–87, 2002. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500102753410408.

Jain:1990:EPD

[JR90] Dipak C. Jain and Ram C. Rao. Effect of price on the demand for durables: Modeling, estimation, and findings. Journal of Business & Economic Statistics, 8(2):163–170, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1990.10509787.

Jalan:2003:EBI

[JR03] Jyotsna Jalan and Martin Ravallion. Estimating the benefit incidence of an antipoverty program by propensity-score matching. Journal of Business & Economic Statistics, 21 (1):19–30, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618720.

Juodis:2022:IPP

[JR22] Arturas Juodis and Simon Reese. The incidental parameters problem in testing for remaining cross-section correlation. Journal of Business & Economic Statistics, 40(3):1191–1203, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1906687.

Jabine:1985:GSU

[JS85] Thomas B. Jabine and Fritz Scheuren. Goals for statistical uses of administrative records: The next 10 years. Journal of Business & Economic Statistics, 3(4):380–391, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1985.10509472.

Jorgenson:1987:ACB

[JS87] Dale W. Jorgenson and Daniel T. Slesnick. Aggregate consumer behavior and household equivalence scales. Journal of Business & Economic Statistics, 5(2):219–232, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509580.

Jaditz:1998:SFP

[JS98] Ted Jaditz and Chera L. Sayers. Out-of-sample forecast performance as a test for nonlinearity in time series. Journal of Business & Economic Statistics, 16(1):110–117, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524740.

Jorgenson:1999:IGP

[JS99] Dale W. Jorgenson and Daniel T. Slesnick. Indexing government programs for changes in the cost of living. Journal of Business & Economic Statistics, 17(2):170–181, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524807.

Juarez:2010:MBC

[JS10] Miguel A. Juárez and Mark F. J. Steel. Model-based clustering of non-Gaussian panel data based on skew-t distributions. Journal of Business & Economic Statistics, 28(1):52–66, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2009.07145.

Jo:2019:MUT

[JS19] Soojin Jo and Rodrigo Sekkel. Macroeconomic uncertainty through the lens of professional forecasters. Journal of Business & Economic Statistics, 37(3):436–446, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1356729.

Juodis:2020:LEF

[JS20] Arturas Juodis and Vasilis Sarafidis. A linear estimator for factor-augmented fixed-T panels with endogenous regressors. Journal of Business & Economic Statistics, 40(1):1–15, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1766469.

Jorgenson:1988:TSB

[JSS88] Dale W. Jorgenson, Daniel T. Slesnick, and Thomas M. Stoker. Two-stage budgeting and exact aggregation. Journal of Business & Economic Statistics, 6(3):313–325, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509668.

Jacobs:2020:CGM

[JSSvN20] Jan P. A. M. Jacobs, Samad Sarferaz, Jan-Egbert Sturm, and Simon van Norden. Can GDP measurement be further improved? Data revision and reconciliation. Journal of Business & Economic Statistics, 40(1):423–431, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2020.1831928.

Jagannathan:2002:GMM

[JSW02] Ravi Jagannathan, Georgios Skoulakis, and Zhenyu Wang. Generalized methods of moments. Journal of Business & Economic Statistics, 20(4):470–481, 2002. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102288618612.

Jaggia:1995:CTO

[JT95] Sanjiv Jaggia and Satish Thosar. Contested tender offers: an estimate of the hazard function. Journal of Business & Economic Statistics, 13(1):113–119, 1995. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1995.10524584.

Juhl:2014:NTP

[Juh14] Ted Juhl. A nonparametric test of the predictive regression model. Journal of Business & Economic Statistics, 32(3):387– 394, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2014.887013.

Juodis:2018:PPD

[Juo18] Arturas Juodis. Pseudo panel data models with cohort interactive effects. Journal of Business & Economic Statistics, 36(1):47–61, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1137759.

Juselius:1998:SVD

[Jus98] Katarina Juselius. A structured VAR for Denmark under changing monetary regimes. Journal of Business & Economic Statistics, 16(4):400–411, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524780.

Jain:1994:RCL

[JVC94] Dipak C. Jain, Naufel J. Vilcassim, and Pradeep K. Chintagunta. A random-coefficients logit brand-choice model applied to panel data. Journal of Business & Economic Statistics, 12 (3):317–328, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10524547.

Jorgenson:1999:SSC

[JW99] Dale W. Jorgenson and Mark W. Watson. Special section on consumer price research: Introduction. Journal of Business & Economic Statistics, 17(2):137–140, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524803.

Jiang:2021:NCS

[JY21] Rong Jiang and Keming Yu. No-crossing single-index quantile regression curve estimation. Journal of Business & Economic Statistics, 41(2):309–320, 2021. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.2013245.

Kakwani:1984:RDC

[Kak84a] Nanak Kakwani. The relative deprivation curve and its applications. Journal of Business & Economic Statistics, 2 (4):384–394, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509412. See comment [App85].

Kakwani:1984:R

[Kak84b] Nanak Kakwani. Reply. Journal of Business & Economic Statistics, 2(4):400–405, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509415.

Kakwani:1985:R

[Kak85] Nanak Kakwani. Reply. Journal of Business & Economic Statistics, 3(2):171–173, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509447.

Kalnina:2022:INH

[Kal22] Ilze Kalnina. Inference for nonparametric high-frequency estimators with an application to time variation in betas. Journal of Business & Economic Statistics, 41(2):538–549, 2022. CO-DEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2022.2040520.

Kang:1985:EDG

[Kan85] Heejoon Kang. The effects of detrending in Granger causality tests. Journal of Business & Economic Statistics, 3 (4):344–349, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509467.

Kang:1986:UAF

[Kan86] Heejoon Kang. Univariate ARIMA forecasts of defined variables. Journal of Business & Economic Statistics, 4(1):81–86, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/07350015.1986.10509496.

Kapetanios:2010:TPD

[Kap10] George Kapetanios. A testing procedure for determining the number of factors in approximate factor models with large datasets. Journal of Business & Economic Statistics, 28(3):397–409, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2009.07239.

Karolyi:1995:MGM

[Kar95] G. Andrew Karolyi. A multivariate GARCH model of international transmissions of stock returns and volatility: The case of the United States and Canada. Journal of Business & Economic Statistics, 13(1):11–25, 1995. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1995.10524575.

Kasahara:2009:TIT

[Kas09] Hiroyuki Kasahara. Temporary increases in tariffs and investment: The Chilean experience. Journal of Business & Economic Statistics, 27(1):113–127, 2009. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.0009.

Kashaev:2022:IEM

[Kas22] Nail Kashaev. Identification and estimation of multinomial choice models with latent special covariates. Journal of Business & Economic Statistics, 41(3):695–707, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2060987.

Koopman:2004:SSM

[KB04] Siem Jan Koopman and Charles S. Bos. State space models with a common stochastic variance. Journal of Business & Economic Statistics, 22(3):346–357, 2004. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000190.

Koenker:2010:MMQ

[KB10] Roger Koenker and Gilbert W. Bassett Jr. March madness, quantile regression bracketology, and the Hayek hypothesis. Journal of Business & Economic Statistics, 28(1):26–35, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2009.07093.

Kim:1995:MDP

[KBR95] Byung-Do Kim, Robert C. Blattberg, and Peter E. Rossi. Modeling the distribution of price sensitivity and implications for optimal retail pricing. Journal of Business & Economic Statistics, 13(3):291–303, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524603.

Kim:1989:STM

[KBZ89] Moshe Kim and Uri Ben-Zion. The structure of technology in a multioutput branch banking firm. Journal of Business & Economic Statistics, 7(4):489–496, 1989. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1989.10509762.

Karson:1985:EME

[KC85] Marvin J. Karson and David C. Cheng. Estimation of multiperiod expected rates of return when investment relatives are lognormally distributed. Journal of Business & Economic Statistics, 3(2):140–148, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509442.

Kokic:1997:MPP

[KCBB97] Philip Kokic, Ray Chambers, Jens Breckling, and Stewe Beare. A measure of production performance. Journal of Business & Economic Statistics, 15(4):445–451, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524722.

Kolesar:2015:IIM

[KCF^{+}15] Michal Kolesár, Raj Chetty, John Friedman, Edward Glaeser, and Guido W. Imbens. Identification and inference with many invalid instruments. Journal of Business & Economic Statistics, 33(4):474–484, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.978175.

Kruger:2017:UET

[KCR17] Fabian Krüger, Todd E. Clark, and Francesco Ravazzolo. Using entropic tilting to combine BVAR forecasts with external nowcasts. Journal of Business & Economic Statistics, 35(3):470–485, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1087856.

Kilian:2000:RBT

[KD00] Lutz Kilian and Ufuk Demiroglu. Residual-based tests for normality in autoregressions: Asymptotic theory and simulation evidence. Journal of Business & Economic Statistics, 18(1):40–50, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524846.

Kerr:2008:CSB

[KD08] Sougata Kerr and Lucia Dunn. Consumer search behavior in the changing credit card market. Journal of Business & Economic Statistics, 26(3):345–353, 2008. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000133.

Kappe:2020:STF

[KDM20] Eelco Kappe, Wayne S. DeSarbo, and Marcelo C. Medeiros. A smooth transition finite mixture model for accommodating unobserved heterogeneity. Journal of Business & Economic Statistics, 38(3):580–592, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1543126.

Keane:1992:NIM

[Kea92] Michael P. Keane. A note on identification in the multinomial probit model. Journal of Business & Economic Statistics, 10(2):193–200, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509898.

Keane:1997:MHS

[Kea97] Michael P. Keane. Modeling heterogeneity and state dependence in consumer choice behavior. Journal of Business & Economic Statistics, 15(3):310–327, 1997. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1997.10524709.

Kennedy:1995:RTE

[Ken95] Fetter E. Kennedy. Randomization tests in econometrics. Journal of Business & Economic Statistics, 13(1):85–94, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1995.10524581.

Kenny:2014:C

[Ken14] G. Kenny. Comment. Journal of Business & Economic Statistics, 32(4):500–504, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.956636. See [AGO^{+}14a, AGO^{+}14b].

Kersten:1985:NAC

[Ker85] H. M. P. Kersten. Nonresponse assessment of a consumer price index. Journal of Business & Economic Statistics, 3 (4):336–343, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509466.

Keweloh:2020:GMM

[Kew20] Sascha Alexander Keweloh. A generalized method of moments estimator for structural vector autoregressions based on higher moments. Journal of Business & Economic Statistics, 39(3):772–782, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1730858.

Koebel:2003:ITC

[KFL03] Bertrand Koebel, Martin Falk, and François Laisney. Imposing and testing curvature conditions on a Box–Cox cost function. Journal of Business & Economic Statistics, 21 (2):319–335, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288618981.

Kalli:2015:FMD

[KG15] Maria Kalli and Jim Griffin. Flexible modeling of dependence in volatility processes. Journal of Business & Economic Statistics, 33(1):102–113, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.925457.

Kumbhakar:1991:GPF

[KGM91] Subal C. Kumbhakar, Soumendra Ghosh, and J. Thomas McGuckin. A generalized production frontier approach for estimating determinants of inefficiency in U.S. dairy farms. Journal of Business & Economic Statistics, 9(3):279–286, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1991.10509853.

Khan:2008:C

[Kha08] Shakeeb Khan. Comment. Journal of Business & Economic Statistics, 26(3):294–295, 2008. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500108000000097. See [ALT08a, ALT08b].

Kuan:2005:UCM

[KHT05] Chung-Ming Kuan, Yu-Lieh Huang, and Ruey S. Tsay. An unobserved-component model with switching permanent and transitory innovations. Journal of Business & Economic Statistics, 23(4):443–454, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500105000000054.

Kiefer:2010:DEE

[Kie10] Nicholas M. Kiefer. Default estimation and expert information. Journal of Business & Economic Statistics, 28(2):320–328, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2009.07236.

Kiefer:2014:C

[Kie14] Nicholas M. Kiefer. Comment. Journal of Business & Economic Statistics, 32(3):322–323, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.926816. See [Mül14a, Mül14b].

Kilian:2007:C

[Kil07] Lutz Kilian. Comment. Journal of Business & Economic Statistics, 25(2):156–159, 2007. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000025. See [DSSW07a, DSSW07b].

Kilian:2015:C

[Kil15] Lutz Kilian. Comment. Journal of Business & Economic Statistics, 33(1):13–17, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.969430.

Kilian:2022:CGK

[Kil22] Lutz Kilian. Comment on Giacomini, Kitagawa, and Read’s “Narrative Restrictions and Proxies”. Journal of Business & Economic Statistics, 40(4):1429–1433, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.

2102022. See [GKR22a, GKR22b].

Kim:1990:ERI

[Kim90] Yoonbai Kim. Exchange rates and import prices in the United States: a varying-parameter estimation of exchange-rate pass-through. Journal of Business & Economic Statistics, 8 (3):305–315, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509801.

Kim:1993:UCT

[Kim93] Chang-Jin Kim. Unobserved-component time series models with Markov-switching heteroscedasticity: Changes in regime and the link between inflation rates and inflation uncertainty. Journal of Business & Economic Statistics, 11(3): 341–349, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509962.

Kim:2001:BAB

[Kim01] Jae H. Kim. Bootstrap-after-bootstrap prediction intervals for autoregressive models. Journal of Business & Economic Statistics, 19(1):117–128, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
07350010152472670.

Kim:2005:CRP

[Kim05] Jaebeom Kim. Convergence rates to purchasing power parity for traded and nontraded goods. Journal of Business & Economic Statistics, 23(1):76–86, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000226.

Kim:2022:RID

[Kim22] Min Seong Kim. Robust inference for diffusion-index forecasts with cross-sectionally dependent data. Journal of Business & Economic Statistics, 40(3):1153–1167, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1906258.

Keeler:1985:RSW

[KJAG85] James P. Keeler, William L. James, and Mohamed Abdel-Ghany. The relative size of windfall income and the permanent income hypothesis. Journal of Business & Economic Statistics, 3(3):209–215, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509452.

Kessels:2009:EAC

[KJGV09] Roselinde Kessels, Bradley Jones, Peter Goos, and Martina Vandebroek. An efficient algorithm for constructing Bayesian optimal choice designs. Journal of Business & Economic Statistics, 27(2):279–291, 2009. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.0026.

Koedijk:1992:TEE

[KK92] Kees G. Koedijk and Clemens J. M. Kool. Tail estimates of East European exchange rates. Journal of Business & Economic Statistics, 10(1):83–96, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509889.

Kishor:2012:VEF

[KK12] N. Kundan Kishor and Evan F. Koenig. VAR estimation and forecasting when data are subject to revision. Journal of Business & Economic Statistics, 30(2):181–190, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1198/jbes.2010.08169.

Kim:2015:BIR

[KK15] Chang-Jin Kim and Jaeho Kim. Bayesian inference in regimeswitching ARMA models with absorbing states: The dynamics of the ex-ante real interest rate under regime shifts. Journal of Business & Economic Statistics, 33(4):566–578, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.979995.

Kiefer:1994:MPE

[KKB94] Nicholas M. Kiefer, Thomas J. Kelly, and Kenneth Burdett. Menu pricing: an experimental approach. Journal of Business & Economic Statistics, 12(3):329–337, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10524548.

Kaul:2022:SSC

[KKPS22] Ashok Kaul, Stefan Klößner, Gregor Pfeifer, and Manuel Schieler. Standard synthetic control methods: The case of using all preintervention outcomes together with covariates. Journal of Business & Economic Statistics, 40(3):1362–1376, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1930012.

Klaassen:2022:TMH

[KKS22] Sven Klaassen, Jannis Kueck, and Martin Spindler. Transformation models in high dimensions. Journal of Business & Economic Statistics, 40(3):1168–1178, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1906259.

Kanoh:1990:MEM

[KL90] Satoru Kanoh and Zhi Dong Li. A method of exploring the mechanism of inflationary expectations based on qualitative survey data. Journal of Business & Economic Statistics, 8(4):395–403, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509810.

Koopman:2008:NGP

[KL08] Siem Jan Koopman and André Lucas. A non-Gaussian panel time series model for estimating and decomposing default risk. Journal of Business & Economic Statistics, 26(4):510–525, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500108000000051.

Klaassen:2005:LSE

[Kla05] Franc Klaassen. Long swings in exchange rates. Journal of Business & Economic Statistics, 23(1):87–95, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000505.

Koop:2010:DPR

[KLGS10] Gary Koop, Roberto Leon-Gonzalez, and Rodney W. Strachan. Dynamic probabilities of restrictions in state space models: an application to the Phillips curve. Journal of Business & Economic Statistics, 28(3):370–379, 2010. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.07335.

Kling:2001:IIV

[Kli01] Jeffrey R. Kling. Interpreting instrumental variables estimates of the returns to schooling. Journal of Business & Economic Statistics, 19(3):358–364, 2001. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500101681020006.

Kline:2016:IDC

[Kli16] Brendan Kline. Identification of the direction of a causal effect by instrumental variables. Journal of Business & Economic Statistics, 34(2):176–184, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1021925.

Klinenberg:2022:SCT

[Kli22] Danny Klinenberg. Synthetic control with time varying coefficients a state space approach with Bayesian shrinkage. Journal of Business & Economic Statistics, 41(4):1065–1076, 2022. CO-DEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2022.2102025.

Kahn-Lang:2020:PPD

[KLL20] Ariella Kahn-Lang and Kevin Lang. The promise and pitfalls of differences-in-differences: Reflections on 16 and Pregnant and other applications. Journal of Business & Economic Statistics, 38(3):613–620, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1546591.

Kiefer:1985:HLS

[KLN85] Nicholas M. Kiefer, Shelly J. Lundberg, and George R. Neumann. How long is a spell of unemployment? Illusions and biases in the use of CPS data. Journal of Business & Economic Statistics, 3(2):118–128, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509439.

Kapetanios:2008:FUB

[KLP08] George Kapetanios, Vincent Labhard, and Simon Price. Forecasting using Bayesian and information-theoretic model averaging. Journal of Business & Economic Statistics, 26 (1):33–41, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500107000000232.

Koopman:2012:DFM

[KLS12] Siem Jan Koopman, André Lucas, and Bernd Schwaab. Dynamic factor models with macro, frailty, and industry effects for U.S. default counts: The credit crisis of 2008. Journal of Business & Economic Statistics, 30(4):521–532, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2012.700859.

Koopman:2015:NAI

[KLS15] Siem Jan Koopman, André Lucas, and Marcel Scharth. Numerically accelerated importance sampling for nonlinear non-Gaussian state-space models. Journal of Business & Economic Statistics, 33(1):114–127, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.925807.

Ke:2020:HDD

[KLZ20] Yuan Ke, Heng Lian, and Wenyang Zhang. High-dimensional dynamic covariance matrices with homogeneous structure. Journal of Business & Economic Statistics, 40(1):96–110, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1779079.

Kormendi:1991:C

[KM91] Roger C. Kormendi and Philip Meguire. Comment. Journal of Business & Economic Statistics, 9(2):152–159, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1991.10509838.

Klier:2008:CAS

[KM08] Thomas Klier and Daniel P. McMillen. Clustering of auto supplier plants in the United States. Journal of Business & Economic Statistics, 26(4):460–471, 2008. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000188.

Kleibergen:2009:R

[KM09a] Frank Kleibergen and Sophocles Mavroeidis. Rejoinder. Journal of Business & Economic Statistics, 27(3):331–339, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1198/jbes.2009.08350.

See [KM09b].

Kleibergen:2009:WIR

[KM09b] Frank Kleibergen and Sophocles Mavroeidis. Weak instrument robust tests in GMM and the new Keynesian Phillips curve. Journal of Business & Economic Statistics, 27(3):293–311, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1198/jbes.2009.08280.

See comments [Can09, CS09, Duf09, Mik09, Wri09, Yog09, ZC09] and rejoinder [KM09a].

Kim:2005:MPE

[KMF05] Jin Gyo Kim, Ulrich Menzefricke, and Fred M. Feinberg. Modeling parametric evolution in a random utility framework. Journal of Business & Economic Statistics, 23(3):282–294, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500104000000550.

Kelly:2021:TS

[KMM21] Bryan Kelly, Asaf Manela, and Alan Moreira. Text selection. Journal of Business & Economic Statistics, 39(4):859–879, 2021. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2021.

1947843. See discussion [Pel21, XCK21, Sin21].

Koop:2022:REM

[KMMP22] Gary Koop, Stuart McIntyre, James Mitchell, and Aubrey Poon. Reconciled estimates of monthly GDP in the United States. Journal of Business & Economic Statistics, 41(2):563– 577, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2044336.

Kim:2005:SBE

[KMN05] Chang-Jin Kim, James C. Morley, and Charles R. Nelson. The structural break in the equity premium. Journal of Business & Economic Statistics, 23(2):181–191, 2005. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000352.

Kowal:2019:FAS

[KMR19] Daniel R. Kowal, David S. Matteson, and David Ruppert. Functional autoregression for sparsely sampled data. Journal of Business & Economic Statistics, 37(1):97–109, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1279058.

Koopman:2010:ATS

[KMV10] Siem Jan Koopman, Max I. P. Mallee, and Michel Van der Wel. Analyzing the term structure of interest rates using the dynamic Nelson–Siegel model with time-varying parameters. Journal of Business & Economic Statistics, 28(3):329–343, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2009.07295.

Kim:1989:TVP

[KN89] Chang-Jin Kim and Charles R. Nelson. The time-varying-parameter model for modeling changing conditional variance: The case of the Lucas hypothesis. Journal of Business & Economic Statistics, 7(4):433–440, 1989. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1989.10509755.

Kim:2004:LVU

[KNP04] Chang-Jin Kim, Charles R. Nelson, and Jeremy Piger. The lessvolatile U.S. economy. Journal of Business & Economic Statistics, 22(1):80–93, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619412.

Knuppel:2009:TBC

[Knü09] Malte Knüppel. Testing business cycle asymmetries based on autoregressions with a Markov-switching intercept. Journal of Business & Economic Statistics, 27(4):544–552, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.06117.

Knuppel:2015:ECM

[Knü15] Malte Knüppel. Evaluating the calibration of multi-step-ahead density forecasts using raw moments. Journal of Business & Economic Statistics, 33(2):270–281, 2015. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2014.948175.

Karavias:2022:SBI

[KNW22] Yiannis Karavias, Paresh Kumar Narayan, and Joakim Westerlund. Structural breaks in interactive effects panels and the stock market reaction to COVID-19. Journal of Business & Economic Statistics, 41(3):653–666, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2053690.

Kociecki:2010:PIR

[Koc10] Andrzej Kociecki. A prior for impulse responses in Bayesian structural VAR models. Journal of Business & Economic Statistics, 28(1):115–127, 2010. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.07278.

Koenker:1988:C

[Koe88] Roger Koenker. Comment. Journal of Business & Economic Statistics, 6(4):447–449, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509688.

Koebel:2006:ELD

[Koe06] Bertrand Koebel. Exports and labor demand. Journal of Business & Economic Statistics, 24(1):91–103, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500105000000234.

Kohler:1983:BPE

[Koh83a] Daniel F. Kohler. The bias in price elasticity estimates under homothetic separability: Implications for analysis of peak-load electricity pricing. Journal of Business & Economic Statistics, 1(3):202–210, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509340.

Kohler:1983:R

[Koh83b] Daniel F. Kohler. Response. Journal of Business & Economic Statistics, 1(3):226–228, 1983. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1983.10509345.

Kokoski:1987:PMC

[Kok87] Mary F. Kokoski. Problems in the measurement of consumer cost-of-living indexes. Journal of Business & Economic Statistics, 5(1):39–46, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509558.

Kolokolov:2020:EJA

[Kol20] Aleksey Kolokolov. Estimating jump activity using multipower variation. Journal of Business & Economic Statistics, 40(1):128–140, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1784745.

Kongsted:1998:MPQ

[Kon98] Hans Christian Kongsted. Modeling price and quantity relations for Danish manufacturing exports. Journal of Business & Economic Statistics, 16(1):81–91, 1998. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1998.10524737.

Koop:1991:IPR

[Koo91] Gary Koop. Intertemporal properties of real output: a Bayesian analysis. Journal of Business & Economic Statistics, 9(3):253–265, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509851.

Korniotis:2010:EPM

[Kor10] George M. Korniotis. Estimating panel models with internal and external habit formation. Journal of Business & Economic Statistics, 28(1):145–158, 2010. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.08041.

Korobilis:2019:HDM

[Kor19] Dimitris Korobilis. High-dimensional macroeconomic forecasting using message passing algorithms. Journal of Business & Economic Statistics, 39(2):493–504, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2019.1677472.

Koop:1994:BEA

[KOS94a] Gary Koop, Jacek Osiewalski, and Mark F. J. Steel. Bayesian efficiency analysis with a flexible form: The AIM cost function. Journal of Business & Economic Statistics, 12(3): 339–346, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10524549.

Koop:1994:PPL

[KOS94b] Gary Koop, Jacek Osiewalski, and Mark F. J. Steel. Posterior properties of long-run impulse responses. Journal of Business & Economic Statistics, 12(4):489–492, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10524571.

Koop:2000:MSO

[KOS00] Gary Koop, Jacek Osiewalski, and Mark F. J. Steel. Modeling the sources of output growth in a panel of countries. Journal of Business & Economic Statistics, 18(3):284–299, 2000. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.2000.10524871.

Kott:1984:STA

[Kot84] Phillip S. Kott. A superpopulation theory approach to the design of price index estimators with small sampling biases. Journal of Business & Economic Statistics, 2(1):83–90, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1984.10509372.

Kott:1985:NMB

[Kot85] Phillip S. Kott. A note on model-based stratification. Journal of Business & Economic Statistics, 3(3):284–286, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1985.10509458.

Kott:1987:NPS

[Kot87] Phillip S. Kott. Nonresponse in a periodic sample survey. Journal of Business & Economic Statistics, 5(2):287–293, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1987.10509588.

Koulayev:2013:SDP

[Kou13] Sergei Koulayev. Search with Dirichlet priors: Estimation and implications for consumer demand. Journal of Business & Economic Statistics, 31(2):226–239, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.764696.

Kowalski:2016:CQI

[Kow16] Amanda Kowalski. Censored quantile instrumental variable estimates of the price elasticity of expenditure on medical care.

Journal of Business & Economic Statistics, 34(1):107–117, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1004072.

Kozbur:2020:IAS

[Koz20] Damian Kozbur. Inference in additively separable models with a high-dimensional set of conditioning variables. Journal of Business & Economic Statistics, 39(4):984–1000, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1753524.

Klemm:1987:BR

[KP87] Rebecca Klemm and Adrian Pagan. Book reviews. Journal of Business & Economic Statistics, 5(4):549–550, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1987.10509623.

Koreisha:1995:CBD

[KP95] Sergio G. Koreisha and Tarmo Pukkila. A comparison between different order-determination criteria for identification of ARIMA models. Journal of Business & Economic Statistics, 13(1):127–131, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524586.

Koop:1999:DAU

[KP99] Gary Koop and Simon M. Potter. Dynamic asymmetries in U.S. unemployment. Journal of Business & Economic Statistics, 17(3):298–312, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524819.

Koop:2003:BAE

[KP03] Gary Koop and Simon M. Potter. Bayesian analysis of endogenous delay threshold models. Journal of Business & Economic Statistics, 21(1):93–103, 2003. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102288618801.

Kejriwal:2010:TMS

[KP10] Mohitosh Kejriwal and Pierre Perron. Testing for multiple structural changes in cointegrated regression models. Journal of Business & Economic Statistics, 28(4):503–522, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.07220.

Kreider:2011:IEO

[KP11] Brent Kreider and John V. Pepper. Identification of expected outcomes in a data error mixing model with multiplicative mean independence. Journal of Business & Economic Statistics, 29(1):49–60, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.07223.

Kazak:2022:BPP

[KP22] Ekaterina Kazak and Winfried Pohlmeier. Bagged pretested portfolio selection. Journal of Business & Economic Statistics, 41 (4):1116–1131, 2022. CODEN ???? ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2022.2110880.

Koo:2000:MRC

[KPS00] Jahyeong Koo, Keith R. Phillips, and Fiona D. Sigalla. Measuring regional cost of living. Journal of Business & Economic Statistics, 18(1):127–136, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524853.

Koop:2013:IBD

[KPS13] Gary Koop, M. Hashem Pesaran, and Ron P. Smith. On identification of Bayesian DSGE models. Journal of Business & Economic Statistics, 31(3):300–314, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.773905.

Kemp:2020:DVM

[KPS20] Gordon C. R. Kemp, Paulo M. D. C. Parente, and J. M. C. Santos Silva. Dynamic vector mode regression. Journal of Business & Economic Statistics, 38(3):647–661, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2018.1562935.

Koch:1988:ECD

[KR88] Paul D. Koch and Robert H. Rasche. An examination of the commerce department leading-indicator approach. Journal of Business & Economic Statistics, 6(2):167–187, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509652.

Koch:1990:DRB

[KR90] Paul D. Koch and Jeffrey A. Rosensweig. The dynamic relationship between the dollar and components of U.S. trade. Journal of Business & Economic Statistics, 8(3):355–364, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1990.10509806.

Keane:1992:EPD

[KR92a] Michael P. Keane and David E. Runkle. On the estimation of panel-data models with serial correlation when instruments are not strictly exogenous. Journal of Business & Economic Statistics, 10(1):1–9, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509877.

Keane:1992:R

[KR92b] Michael P. Keane and David E. Runkle. Reply. Journal of Business & Economic Statistics, 10(1):26–29, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1992.10509882.

Krishnamurthi:1994:SPEa

[KR94a] Lakshman Krishnamurthi and Arvind Rangaswamy. The statistical properties of the equity estimator: a reply. Journal of Business & Economic Statistics, 12(2):149–153, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10510001.

Krishnamurthi:1994:SPEb

[KR94b] Lakshman Krishnamurthi and Arvind Rangaswamy. The statistical properties of the equity estimator: Reply to rejoinder. Journal of Business & Economic Statistics, 12 (2):155, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10510003.

Krasker:1986:TSB

[Kra86] William S. Krasker. Two-stage bounded-influence estimators for simultaneous-equations models. Journal of Business & Economic Statistics, 4(4):437–444, 1986. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1986.10509543.

Krauth:2007:PSE

[Kra07] Brian V. Krauth. Peer and selection effects on youth smoking in California. Journal of Business & Economic Statistics, 25(3):288–298, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000396.

Kristensen:2017:DIS

[Kri17] Johannes Tang Kristensen. Diffusion indexes with sparse loadings. Journal of Business & Economic Statistics, 35(3):434–451, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1084308.

Krohn:1983:MEP

[Kro83] Gregory A. Krohn. Measuring the experience-productivity relationship: The case of major league baseball. Journal of Business & Economic Statistics, 1(4):273–279, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1983.10509351.

Kim:2014:MIM

[KRW^{+}14] Hang J. Kim, Jerome P. Reiter, Quanli Wang, Lawrence H. Cox, and Alan F. Karr. Multiple imputation of missing or faulty values under linear constraints. Journal of Business & Economic Statistics, 32(3):375–386, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.885435.

Koop:1994:DTA

[KS94] Gary Koop and Mark F. J. Steel. A decision-theoretic analysis of the unit-root hypothesis using mixtures of elliptical models. Journal of Business & Economic Statistics, 12 (1):95–107, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10509993.

Koschat:1996:IGM

[KS96a] Martin A. Koschat and Deborah F. Swayne. Interactive graphical methods in the analysis of customer panel data. Journal of Business & Economic Statistics, 14(1):113–126, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524634.

Koschat:1996:R

[KS96b] Martin A. Koschat and Deborah F. Swayne. Reply. Journal of Business & Economic Statistics, 14(1):130–132, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1996.10524637.

Kiefer:1998:BAP

[KS98] Nicholas M. Kiefer and Mark F. J. Steel. Bayesian analysis of the prototypal search model. Journal of Business & Economic Statistics, 16(2):178–186, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524751.

Kim:2008:MCB

[KS08] Yangseon Kim and Peter Schmidt. Marginal comparisons with the best and the efficiency measurement problem. Journal of Business & Economic Statistics, 26(2):253–260, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500107000000331.

Kim:2017:TJT

[KS17] Young-Joo Kim and Myung Hwan Seo. Is there a jump in the transition? Journal of Business & Economic Statistics, 35(2):241–249, 2017. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1164055.

Kashaev:2020:DSC

[KS20] Nail Kashaev and Bruno Salcedo. Discerning solution concepts for discrete games. Journal of Business & Economic Statistics, 39(4):1001–1014, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1753525.

Kvanli:1998:CCI

[KSD98] Alan H. Kvanli, Yaung Kaung Shen, and Lih Yuan Deng. Construction of confidence intervals for the mean of a population containing many zero values. Journal of Business & Economic Statistics, 16(3):362–368, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524776.

Kennickell:1997:RRH

[KSM97] Arthur B. Kennickell and Martha Starr-McCluer. Retrospective reporting of household wealth: Evidence from the 1983–1989 survey of consumer finances. Journal of Business & Economic Statistics, 15(4):452–463, 1997. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1997.10524723.

Koulovatianos:2009:NHT

[KSS09] Christos Koulovatianos, Carsten Schrder, and Ulrich Schmidt. Nonmarket household time and the cost of children. Journal of Business & Economic Statistics, 27(1):42–51, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2009.0004.

Khan:2011:HTM

[KST11] Shakeeb Khan, Youngki Shin, and Elie Tamer. Heteroscedastic transformation models with covariate dependent censoring. Journal of Business & Economic Statistics, 29(1):40–48, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2009.07227.

Kneip:2016:THN

[KSW16] Alois Kneip, Léopold Simar, and Paul W. Wilson. Testing hypotheses in nonparametric models of production. Journal of Business & Economic Statistics, 34(3):435–456, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1049747.

Kim:2022:OGI

[KSW22] Donggyu Kim, Minseok Shin, and Yazhen Wang. Overnight GARCH–Itô volatility models. Journal of Business & Economic Statistics, 41(4):1215–1227, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2116027.

Kim:1995:EER

[KT95]  Taeil Kim and Lowell J. Taylor.  The employment ef-

fect in retail trade of California’s 1988 minimum wage increase. Journal of Business & Economic Statistics, 13(2): 175–182, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524591.

Kline:2013:C

[KT13] Brendan Kline and Elie Tamer. Comment. Journal of Business & Economic Statistics, 31(3):276–279, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2013.792264. See [GPI13b,

GPI13a].

Khan:2018:DSE

[KT18] S. Khan and E. Tamer. Discussion of “Simple Estimators for Invertible Index Models” by H. Ahn, H. Ichimura, J. Powell, and P. Ruud. Journal of Business & Economic Statistics, 36(1):11–15, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1392312. See [AIPR18b, AIPR18a].

Kitagawa:2019:EMT

[KT19] Toru Kitagawa and Aleksey Tetenov. Equality-minded treatment choice. Journal of Business & Economic Statistics, 39 (2):561–574, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1688664.

Tse:2012:EHF

[kTY12] Yiu kuen Tse and Thomas Tao Yang. Estimation of highfrequency volatility: an autoregressive conditional duration approach. Journal of Business & Economic Statistics, 30(4):533– 545, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2012.707582.

Kumbhakar:1987:PFP

[Kum87] Subal C. Kumbhakar. Production frontiers and panel data: an application to U.S. class 1 railroads. Journal of Business & Economic Statistics, 5(2):249–255, 1987. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1987.10509583.

Kumbhakar:1989:ETE

[Kum89] Subal C. Kumbhakar. Estimation of technical efficiency using flexible functional form and panel data. Journal of Business & Economic Statistics, 7(2):253–258, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509734.

Kuttner:1994:EPO

[Kut94] Kenneth N. Kuttner. Estimating potential output as a latent variable. Journal of Business & Economic Statistics, 12 (3):361–368, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10524551.

Kilian:2013:DOP

[KV13] Lutz Kilian and Robert J. Vigfusson. Do oil prices help forecast U.S. real GDP? The role of nonlinearities and asymmetries. Journal of Business & Economic Statistics, 31(1):78–93, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2012.740436.

Kao:1990:TSE

[KW90] Chihwa Kao and Chunchi Wu. Two-step estimation of linear models with ordinal unobserved variables: The case of corporate bonds. Journal of Business & Economic Statistics, 8 (3):317–325, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509802.

Koschat:1992:CTT

[KW92] Martin A. Koschat and Samaradasa Weerahandi. Chow-type tests under heteroscedasticity. Journal of Business & Economic Statistics, 10(2):221–228, 1992. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1992.10509901.

Keane:1997:IJS

[KW97] Michael P. Keane and Kenneth I. Wolpin. Introduction to the jbes special issue on structural estimation in applied microeconomics. Journal of Business & Economic Statistics, 15 (2):111–114, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524694.

Kadane:2004:HME

[KW04a] Joseph B. Kadane and George G. Woodworth. Hierarchical models for employment decisions. Journal of Business & Economic Statistics, 22(2):182–193, 2004. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000073.

Kamakura:2004:EBP

[KW04b] Wagner A. Kamakura and Michel Wedel. An empirical Bayes procedure for improving individual-level estimates and predictions from finite mixtures of multinomial logit models. Journal of Business & Economic Statistics, 22(1):121–125, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500103288619449.

Klimenka:2019:MRM

[KW19] Filip Klimenka and James Lewis Wolter. Multiple regression model averaging and the focused information criterion with an application to portfolio choice. Journal of Business & Economic Statistics, 37(3):506–516, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1383262.

Kaddoura:2022:EPD

[KW22] Yousef Kaddoura and Joakim Westerlund. Estimation of panel data models with random interactive effects and multiple structural breaks when T is fixed. Journal of Business & Economic Statistics, 41(3):778–790, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2067546.

Kalli:2013:MCD

[KWD13] Maria Kalli, Stephen G. Walker, and Paul Damien. Modeling the conditional distribution of daily stock index returns: an alternative Bayesian semiparametric model. Journal of Business & Economic Statistics, 31(4):371–383, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2013.794142.

Wong:1990:RTS

[kWM90] Wing keung Wong and Robert B. Miller. Repeated time series analysis of ARIMA-noise models. Journal of Business & Economic Statistics, 8(2):243–250, 1990. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1990.10509796.

Kobayashi:2021:BAL

[KYK^{+}21] Genya Kobayashi, Yuta Yamauchi, Kazuhiko Kakamu, Yuki Kawakubo, and Shonosuke Sugasawa. Bayesian approach to Lorenz curve using time series grouped data. Journal of Business & Economic Statistics, 40(2):897–912, 2021. CO-DEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2021.1883438.

Kruger:2020:GCF

[KZ20] Fabian Krüger and Johanna F. Ziegel. Generic conditions for forecast dominance. Journal of Business & Economic Statistics, 39(4):972–983, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1741376.

Lillard:1984:TDE

[LA84] Lee A. Lillard and Dennis J. Aigner. Time-of-day electricity consumption response to temperature and the ownership of air conditioning appliances. Journal of Business & Economic Statistics, 2(1):40–53, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509368.

Labadie:1990:SSG

[Lab90] Pamela Labadie. Solving the stochastic growth model by using a recursive mapping based on least squares projection. Journal of Business & Economic Statistics, 8(1):39–40, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1990.10509773.

Lahiri:2012:C

[Lah12] Kajal Lahiri. Comment. Journal of Business & Economic Statistics, 30(1):20–25, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.634342. See [PT12a, PT12b].

Lancaster:1997:ESI

[Lan97] Tony Lancaster. Exact structural inference in optimal jobsearch models. Journal of Business & Economic Statistics, 15 (2):165–179, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524698.

Lavetti:2020:ECW

[Lav20] Kurt Lavetti. The estimation of compensating wage differentials: Lessons from the Deadliest Catch. Journal of Business & Economic Statistics, 38(1):165–182, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2018.1470000.

Lo:2016:OJR

[LB16] Melody Lo and Yong Bao. Are overall journal rankings a good mapping for article quality in specialty fields? Journal of Business & Economic Statistics, 34(1):62–67, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.995798.

Lee:1985:MLM

[LC85] Hau L. Lee and Morris A. Cohen. A multinomial logit model for the spatial distribution of hospital utilization. Journal of Business & Economic Statistics, 3(2):159–168, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1985.10509445.

Loyal:2022:DCC

[LC22] Joshua Daniel Loyal and Yuguo Chen. Discussion of “Cocitation and Co-authorship Networks of Statisticians”. Journal of Business & Economic Statistics, 40(2):497–498, 2022. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2022.2044828. See [JJKL21].

Lan:2022:IHM

[LCZT22] Wei Lan, Xuerong Chen, Tao Zou, and Chih-Ling Tsai. Imputations for high missing rate data in covariates via semi-supervised learning approach. Journal of Business & Economic Statistics, 40(3):1282–1290, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1922120.

Li:1996:PTP

[LD96] David T. Li and Jeffrey H. Dorfman. Predicting turning points through the integration of multiple models. Journal of Business & Economic Statistics, 14(4):421–428, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524671.

Lan:2019:FAM

[LD19] Wei Lan and Lilun Du. A factor-adjusted multiple testing procedure with application to mutual fund selection. Journal of Business & Economic Statistics, 37(1):147–157, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1294078.

Luginbuhl:1999:BAU

[LdV99] Rob Luginbuhl and Aart de Vos. Bayesian analysis of an unobserved-component time series model of GDP with Markovswitching and time-varying growths. Journal of Business & Economic Statistics, 17(4):456–465, 1999. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1999.10524833.

Lechner:1995:SST

[Lec95] Michael Lechner. Some specification tests for probit models estimated on panel data. Journal of Business & Economic Statistics, 13(4):475–488, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524621.

Lechner:1999:EEE

[Lec99] Michael Lechner. Earnings and employment effects of continuous off-the-job training in East Germany after unification. Journal of Business & Economic Statistics, 17 (1):74–90, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524798.

Lechner:2009:SCM

[Lec09] Michael Lechner. Sequential causal models for the evaluation of labor market programs. Journal of Business & Economic Statistics, 27(1):71–83, 2009. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.0006.

Lee:2018:CVE

[Lee18] Seojeong Lee. A consistent variance estimator for 2SLS when instruments identify different LATEs. Journal of Business & Economic Statistics, 36(3):400–410, 2018. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2016.1186555.

LeSage:1991:ADL

[LeS91] James P. LeSage. Analysis and development of leading indicators using a Bayesian turning-points approach. Journal of Business & Economic Statistics, 9(3):305–316, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509856.

LeSage:1992:CTV

[LeS92] James P. LeSage. A comparison of time-varying parameter and multiprocess mixture models in the case of money-supply announcements. Journal of Business & Economic Statistics, 10 (2):201–211, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509899.

Lewbel:2012:UHI

[Lew12] Arthur Lewbel. Using heteroscedasticity to identify and estimate mismeasured and endogenous regressor models. Journal of Business & Economic Statistics, 30(1):67–80, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2012.643126.

Leyes:1985:ID

[Ley85] John Leyes. Introduction to discussants. Journal of Business & Economic Statistics, 3(4):392, 1985. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1985.10509473.

Lan:2018:CME

[LFWT18] Wei Lan, Zheng Fang, Hansheng Wang, and Chih-Ling Tsai. Covariance matrix estimation via network structure. Journal of Business & Economic Statistics, 36(2):359–369, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2016.1173558.

Lichtenberg:1989:EMO

[LG89] Frank R. Lichtenberg and Zvi Griliches. Errors of measurement in output deflators. Journal of Business & Economic Statistics, 7(1):1–9, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509704.

Lu:2020:NET

[LG20] Lu Lu and Sujit K. Ghosh. Nonparametric estimation and testing for positive quadrant dependent bivariate copula. Journal of Business & Economic Statistics, 40(2):664–677, 2020.

CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1855186.

Liang:2022:SSA

[LGG22] Xuan Liang, Jiti Gao, and Xiaodong Gong. Semiparametric spatial autoregressive panel data model with fixed effects and timevarying coefficients. Journal of Business & Economic Statistics, 40(4):1784–1802, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1979564.

Li:2016:MMV

[LGLY16] Weiming Li, Jing Gao, Kunpeng Li, and Qiwei Yao. Modeling multivariate volatilities via latent common factors. Journal of Business & Economic Statistics, 34(4):564–573, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1092975.

Leybourne:1999:MST

[LH99] S. J. Leybourne and B. P. M. HcCabe. Modified stationarity tests with data-dependent model-selection rules. Journal of Business & Economic Statistics, 17(2):264–270, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524816.

Li:2002:SSC

[LHLF02] Qi Li, Cliff J. Huang, Dong Li, and Tsu-Tan Fu. Semiparametric smooth coefficient models. Journal of Business & Economic Statistics, 20(3):412–422, 2002. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102288618531.

Lin:2020:TSS

[LHM20] Wei Lin, Jianhua Z. Huang, and Tucker McElroy. Time series seasonal adjustment using regularized singular value decomposition. Journal of Business & Economic Statistics, 38 (3):487–501, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1515081.

Li:2022:SRM

[LHYZ22] Gaorong Li, Lei Huang, Jin Yang, and Wenyang Zhang. A synthetic regression model for large portfolio allocation. Journal of Business & Economic Statistics, 40(4):1665–1677, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1961787.

Li:1999:TSP

[Li99] Kai Li. Testing symmetry and proportionality in PPP: a paneldata approach. Journal of Business & Economic Statistics, 17 (4):409–418, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524829.

Li:2005:WCR

[Li05] Yuming Li. The wealth-consumption ratio and the consumptionhabit ratio. Journal of Business & Economic Statistics, 23 (2):226–241, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000361.

Li:2011:SBA

[Li11] Junye Li. Sequential Bayesian analysis of time-changed infinite activity derivatives pricing models. Journal of Business & Economic Statistics, 29(4):468–480, 2011. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2010.08310.

Li:2015:SSP

[Li15] Jiahan Li. Sparse and stable portfolio selection with parameter uncertainty. Journal of Business & Economic Statistics, 33(3):381–392, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.954708.

Lian:2012:SEA

[Lia12] Heng Lian. Semiparametric estimation of additive quantile regression models by two-fold penalty. Journal of Business & Economic Statistics, 30(3):337–350, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.693851.

Liesenfeld:1998:DBM

[Lie98] Roman Liesenfeld. Dynamic BivarSate mixture models: Modeling the behavior of prices and trading volume. Journal of Business & Economic Statistics, 16(1):101–109, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524739.

Lin:1986:ALP

[Lin86] Winston T. Lin. Analysis of lumber and pulpwood production in a partial adjustment model with dynamic and variable speeds of adjustment. Journal of Business & Economic Statistics, 4(3):305–316, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509527.

Lin:1997:IRF

[Lin97] Wen-Ling Lin. Impulse response function for conditional volatility in GARCH models. Journal of Business & Economic Statistics, 15(1):15–25, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524682.

Ling:2014:Ca

[Lin14] Shiqing Ling. Comment. Journal of Business & Economic Statistics, 32(2):165, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.887016. See [HT14a, HT14b].

Litterman:1983:RWM

[Lit83] Robert B. Litterman. A random walk, Markov model for the distribution of time series. Journal of Business & Economic Statistics, 1(2):169–173, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509336.

Litterman:1986:C

[Lit86a] Robert B. Litterman. Comment. Journal of Business & Economic Statistics, 4(1):17–19, 1986. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1986.10509488.

Litterman:1986:FBV

[Lit86b] Robert B. Litterman. Forecasting with Bayesian vector autoregressions — five years of experience. Journal of Business & Economic Statistics, 4(1):25–38, 1986. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1986.10509491.

Litterman:1986:SAE

[Lit86c] Robert B. Litterman. A statistical approach to economic forecasting. Journal of Business & Economic Statistics, 4 (1):1–4, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509485.

Little:1988:MDA

[Lit88a] Roderick J. A. Little. Missing-data adjustments in large surveys. Journal of Business & Economic Statistics, 6(3): 287–296, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509663.

Little:1988:R

[Lit88b] Roderick J. A. Little. Reply. Journal of Business & Economic Statistics, 6(3):300–301, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509666.

Liu:2014:IEE

[Liu14] Xiaodong Liu. Identification and efficient estimation of simultaneous equations network models. Journal of Business & Economic Statistics, 32(4):516–536, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.907093.

Liu:2021:DFP

[Liu21] Laura Liu. Density forecasts in panel data models: a semiparametric Bayesian perspective. Journal of Business & Economic Statistics, 41(2):349–363, 2021. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.2021922.

Liu:2018:EHP

[LJ18] Guangying Liu and Bing-Yi Jing. On estimation of Hurst parameter under noisy observations. Journal of Business & Economic Statistics, 36(3):483–492, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1191503.

Lockwood:1988:RME

[LK88] Larry J. Lockwood and K. Rao Kadiyala. Risk measurement for event-dependent security returns. Journal of Business & Economic Statistics, 6(1):43–49, 1988. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1988.10509635.

Lee:1993:LMM

[LK93] John H. H. Lee and Maxwell L. King. A locally most mean powerful based score test for ARCH and GARCH regression disturbances. Journal of Business & Economic Statistics, 11 (1):17–27, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509930.

Li:2022:HDI

[LKFL22] Daoji Li, Yinfei Kong, Yingying Fan, and Jinchi Lv. Highdimensional interaction detection with false sign rate control. Journal of Business & Economic Statistics, 40(3):1234–1245, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1917419.

Lamoureux:1990:PVS

[LL90] Christopher G. Lamoureux and William D. Lastrapes. Persistence in variance, structural change, and the GARCH model. Journal of Business & Economic Statistics, 8(2): 225–234, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509794.

Lamoureux:1994:ETV

[LL94] Christopher G. Lamoureux and William D. Lastrapes. Endogenous trading volume and momentum in stock-return volatility. Journal of Business & Economic Statistics, 12 (2):253–260, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10510012.

Larsson:2007:IPC

[LL07] Rolf Larsson and Johan Lyhagen. Inference in panel cointegration models with long panels. Journal of Business & Economic Statistics, 25(4):473–483, 2007. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500106000000549.

Lanne:2010:SVA

[LL10] Markku Lanne and Helmut Lütkepohl. Structural vector autoregressions with nonnormal residuals. Journal of Business & Economic Statistics, 28(1):159–168, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.06003.

Lanne:2019:GEN

[LL19] Markku Lanne and Jani Luoto. GMM estimation of non-Gaussian structural vector autoregression. Journal of Business & Economic Statistics, 39(1):69–81, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2019.1629940.

Liao:2022:VCM

[LLCL22] Yujie Liao, Jingyuan Liu, Donna L. Coffman, and Runze Li. Varying coefficient mediation model and application to analysis of behavioral economics data. Journal of Business & Economic Statistics, 40(4):1759–1771, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1971089.

Li:2022:DMS

[LLF22] Yu-Ning Li, Degui Li, and Piotr Fryzlewicz. Detection of multiple structural breaks in large covariance matrices. Journal of Business & Economic Statistics, 41(3):846–861, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2076686.

Li:2011:STH

[LLL11] Muyi Li, Guodong Li, and Wai Keung Li. Score tests for hyperbolic GARCH models. Journal of Business & Economic Statistics, 29(4):579–586, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2011.10024.

Lin:2019:RAI

[LLL19] Huazhen Lin, Wei Liu, and Wei Lan. Regression analysis with individual-specific patterns of missing covariates. Journal of Business & Economic Statistics, 39(1):179–188, 2019.

ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1635486.

Li:2020:NQR

[LLL20] Degui Li, Qi Li, and Zheng Li. Nonparametric quantile regression estimation with mixed discrete and continuous data. Journal of Business & Economic Statistics, 39(3):741–756, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1730856.

Lanne:2022:ISV

[LLL22] Markku Lanne, Keyan Liu, and Jani Luoto. Identifying structural vector autoregression via leptokurtic economic shocks. Journal of Business & Economic Statistics, 41(4):1341–1351, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2134872.

Lahiri:1989:EIU

[LLN89] Kajal Lahiri, R. Hamilton Lankford, and Richard P. Numrich. The estimation and interpretation of urban density gradients. Journal of Business & Economic Statistics, 7 (2):227–235, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509731.

Li:2015:UTP

[LLP15] Chenxue Li, Deyuan Li, and Liang Peng. Uniform test for predictive regression with AR errors. Journal of Business & Economic Statistics, 35(1):29–39, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1052460.

Lee:2020:WKP

[LLPZ20] Lung-Fei Lee, Xiaodong Liu, Eleonora Patacchini, and Yves Zenou. Who is the key player? A network analysis of juvenile delinquency. Journal of Business & Economic Statistics, 39(3):849–857, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1737082.

Li:2013:OBS

[LLR13] Qi Li, Juan Lin, and Jeffrey S. Racine. Optimal bandwidth selection for nonparametric conditional distribution and quantile functions. Journal of Business & Economic Statistics, 31(1):57–65, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.738955.

Lazarus:2018:HIRa

[LLSW18a] Eben Lazarus, Daniel J. Lewis, James H. Stock, and Mark W. Watson. HAR inference: Recommendations for practice. Journal of Business & Economic Statistics, 36(4):541–559, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.

1506926. See discussion [Wes18a, Mül18].

Lazarus:2018:HIRb

[LLSW18b] Eben Lazarus, Daniel J. Lewis, James H. Stock, and Mark W. Watson. HAR inference: Recommendations for practice rejoinder. Journal of Business & Economic Statistics, 36(4):574–575, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1513251.

Lan:2015:TDL

[LLT^{+}15] Wei Lan, Ronghua Luo, Chih-Ling Tsai, Hansheng Wang, and Yunhong Yang. Testing the diagonality of a large covariance matrix in a regression setting. Journal of Business & Economic Statistics, 33(1):76–86, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.923317.

Lau:1990:DSR

[LLW90] Amy Hing-Ling Lau, Hon-Shiang Lau, and John R. Wingender. The distribution of stock returns: New evidence against the stable model. Journal of Business & Economic Statistics, 8(2):217–223, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509793.

Li:2020:STI

[LLY20] Rui Li, Chenlei Leng, and Jinhong You. Semiparametric tail index regression. Journal of Business & Economic Statistics, 40(1):82–95, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1775616.

Li:2016:TDA

[LLZ16] Dong Li, Shiqing Ling, and Rongmao Zhang. On a threshold double autoregressive model. Journal of Business & Economic Statistics, 34(1):68–80, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.1001028.

Layton:1989:LIS

[LM89] Allan P. Layton and Geoffrey H. Moore. Leading indicators for the service sector. Journal of Business & Economic Statistics, 7(3):379–386, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509746.

Lesage:1990:UBT

[LM90] James P. Lesage and Michael Magura. Using Bayesian techniques for data pooling in regional payroll forecasting. Journal of Business & Economic Statistics, 8(1):127–135, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1990.10509782.

LeSage:1992:MMA

[LM92] James P. LeSage and Michael Magura. A mixture-model approach to combining forecasts. Journal of Business & Economic Statistics, 10(4):445–452, 1992. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1992.10509920.

Leybourne:1994:CTU

[LM94a] S. J. Leybourne and B. P. M. McCabe. A consistent test for a unit root. Journal of Business & Economic Statistics, 12 (2):157–166, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10510004.

Lim:1994:STI

[LM94b] G. C. Lim and Vance L. Martin. A spectral-temporal index with an application to U.S. interest rates. Journal of Business & Economic Statistics, 12(1):81–93, 1994. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1994.10509992.

LaCour:2000:MEA

[LM00] Lisbeth La Cour and Ronald MacDonald. Modeling the ECU against the U.S. dollar: a structural monetary interpretation. Journal of Business & Economic Statistics, 18 (4):436–450, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524883.

Loaiza-Maya:2022:FVB

[LMN22a] Rubén Loaiza-Maya and Didier Nibbering. Fast variational Bayes methods for multinomial probit models. Journal of Business & Economic Statistics, 41(4):1352–1363, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2139267.

Loaiza-Maya:2022:SBE

[LMN22b] Rubén Loaiza-Maya and Didier Nibbering. Scalable Bayesian estimation in the multinomial probit model. Journal of Business & Economic Statistics, 40(4):1678–1690, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1961788.

Leon:2009:PPS

[LMS09] Ángel León, Javier Mencía, and Enrique Sentana. Parametric properties of semi-nonparametric distributions, with applications to option valuation. Journal of Business & Economic Statistics, 27(2):176–192, 2009. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.0013.

Loaiza-Maya:2020:RTM

[LMS20] Rubén Loaiza-Maya and Michael Stanley Smith. Real-time macroeconomic forecasting with a heteroscedastic inversion copula. Journal of Business & Economic Statistics, 38(2):470–486, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1514309.

Leybourne:1996:CET

[LMT96] S. J. Leybourne, B. P. M. McCabe, and A. R. Tremayne. Can economic time series be differenced to stationarity? Journal of Business & Economic Statistics, 14(4):435–446, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524673.

Lahaye:2020:RJV

[LN20] Jérôme Lahaye and Christopher Neely. The role of jumps in volatility spillovers in foreign exchange markets: Meteor shower and heat waves revisited. Journal of Business & Economic Statistics, 38(2):410–427, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1512865.

Lieli:2010:OBP

[LNB10] Robert P. Lieli and Augusto Nieto-Barthaburu. Optimal binary prediction for group decision making. Journal of Business & Economic Statistics, 28(2):308–319, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.06120.

Lewbel:2010:ER

[LNHW10] Arthur Lewbel, Serena Ng, Keisuke Hirano, and Jonathan Wright. Editors’ report 2009. Journal of Business & Economic Statistics, 28(4):574, 2010. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2010.284er.

Loddo:2011:SMS

[LNS11] Antonello Loddo, Shawn Ni, and Dongchu Sun. Selection of multivariate stochastic volatility models via Bayesian stochastic search. Journal of Business & Economic Statistics, 29(3):342–355, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2010.08197.

Lobato:2003:TNA

[Lob03] Ignacio N. Lobato. Testing for nonlinear autoregression. Journal of Business & Economic Statistics, 21(1):164–173, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500102288618865.

Los:1985:MPI

[Los85] Cornelis A. Los. Measurement problems of inflation disaggregation. Journal of Business & Economic Statistics, 3 (3):244–253, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509456.

Loudermilk:2007:EFD

[Lou07] Margaret S. Loudermilk. Estimation of fractional dependent variables in dynamic panel data models with an application to firm dividend policy. Journal of Business & Economic Statistics, 25(4):462–472, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500107000000098.

Lutkepohl:1996:SEF

[LP96] Helmut Lütkepohl and D. S. Poskitt. Specification of echelonform VARMA models. Journal of Business & Economic Statistics, 14(1):69–79, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524630.

Linton:2003:SRP

[LP03] Oliver Linton and Benoit Perron. The shape of the risk premium. Journal of Business & Economic Statistics, 21 (3):354–367, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619052.

Lavergne:2012:OAA

[LP12] Pascal Lavergne and Valentin Patilea. One for all and all for one: Regression checks with many regressors. Journal of Business & Economic Statistics, 30(1):41–52, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1198/jbes.2011.07152.

Lyssiotou:1999:PHR

[LPS99] Panayiota Lyssiotou, Panos Pashardes, and Thanasis Stengos. Preference heterogeneity and the rank of demand systems. Journal of Business & Economic Statistics, 17(2): 248–252, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524814.

Lehrer:2022:MTD

[LPS22] Steven F. Lehrer, R. Vincent Pohl, and Kyungchul Song. Multiple testing and the distributional effects of accountability incentives in education. Journal of Business & Economic Statistics, 40(4):1552–1568, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1941055.

Li:2003:SEO

[LPV03] Tong Li, Isabelle Perrigne, and Quang Vuong. Semiparametric estimation of the optimal reserve price in first-price auctions. Journal of Business & Economic Statistics, 21(1):53–64, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500102288618757.

Lian:2019:HPS

[LQZ19] Heng Lian, Xinghao Qiao, and Wenyang Zhang. Homogeneity pursuit in single index models based panel data analysis. Journal of Business & Economic Statistics, 39(2):386–401, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1665531.

Liesenfeld:2003:EDB

[LR03] Roman Liesenfeld and Jean-François Richard. Estimation of dynamic bivariate mixture models. Journal of Business & Economic Statistics, 21(4):570–576, 2003. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500103288619287.

Li:2008:NEC

[LR08] Qi Li and Jeffrey S. Racine. Nonparametric estimation of conditional CDF and quantile functions with mixed categorical and continuous data. Journal of Business & Economic Statistics, 26(4):423–434, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500107000000250.

Luukkonen:1999:TVN

[LRS99] Ritva Luukkonen, Antti Ripatti, and Pentti Saikkonen. Testing for a valid normalization of cointegrating vectors in vector autoregressive processes. Journal of Business & Economic Statistics, 17(2):195–204, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524810.

Li:2009:EEA

[LRW09] Qi Li, Jeffrey S. Racine, and Jeffrey M. Wooldridge. Efficient estimation of average treatment effects with mixed categorical and continuous data. Journal of Business & Economic Statistics, 27(2):206–223, 2009. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.0015.

Lee:1987:BEP

[LS87] Jack C. Lee and Darius J. Sabavala. Bayesian estimation and prediction for the beta-binomial model. Journal of Business & Economic Statistics, 5(3):357–367, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509600.

Lee:1994:PRG

[LS94] Ray-Shine Lee and Nsrvikar Singh. Patterns in residential gas and electricity consumption: an econometric analysis. Journal of Business & Economic Statistics, 12(2):233–241, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10510010.

Lobato:1998:RSL

[LS98a] I. N. Lobato and N. E. Savin. Real and spurious long-memory properties of stock-market data. Journal of Business & Economic Statistics, 16(3):261–268, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524760.

Lobato:1998:R

[LS98b] I. N. Lobato and N. E. Savin. Reply. Journal of Business & Economic Statistics, 16(3):280–283, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524766.

Lanne:2002:TAS

[LS02] Markku Lanne and Pentti Saikkonen. Threshold autoregressions for strongly autocorrelated time series. Journal of Business & Economic Statistics, 20(2):282–289, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102317352010.

Lanne:2007:MGO

[LS07] Markku Lanne and Pentti Saikkonen. A multivariate generalized orthogonal factor GARCH model. Journal of Business & Economic Statistics, 25(1):61–75, 2007. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/073500106000000404.

Liao:2015:SBD

[LS15] Yin Liao and John Stachurski. Simulation-based density estimation for time series using covariate data. Journal of Business & Economic Statistics, 33(4):595–606, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.982247.

Lee:2016:MST

[LS16] Yoonseok Lee and Donggyun Shin. Measuring social tension from income class segregation. Journal of Business & Economic Statistics, 34(3):457–471, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1051624.

Lam:2020:ESS

[LS20a] Clifford Lam and Pedro C. L. Souza. Estimation and selection of spatial weight matrix in a spatial lag model. Journal of Business & Economic Statistics, 38(3):693–710, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2019.1569526.

Lee:2020:VMD

[LS20b] Chung Eun Lee and Xiaofeng Shao. Volatility martingale difference divergence matrix and its application to dimension reduction for multivariate volatility. Journal of Business & Economic Statistics, 38(1):80–92, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1458621.

Li:2020:LAH

[LS20c] Mengheng Li and Marcel Scharth. Leverage, asymmetry, and heavy tails in the high-dimensional factor stochastic volatility model. Journal of Business & Economic Statistics, 40 (1):285–301, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1806853.

Lutkepohl:2022:HPV

[LS22] Helmut Lütkepohl and Thore Schlaak. Heteroscedastic proxy vector autoregressions. Journal of Business & Economic Statistics, 40(3):1268–1281, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1920962.

Li:2022:RCM

[LSCX22] Danning Li, Arun Srinivasan, Qian Chen, and Lingzhou Xue. Robust covariance matrix estimation for high-dimensional compositional data with application to sales data analysis. Journal of Business & Economic Statistics, 41(4):1090–1100, 2022. CO-DEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2022.2106990.

Lunde:2016:EAV

[LSS16] Asger Lunde, Neil Shephard, and Kevin Sheppard. Econometric analysis of vast covariance matrices using composite realized kernels and their application to portfolio choice. Journal of Business & Economic Statistics, 34(4):504–518, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1064432.

Lucas:2019:BBM

[LSS19] André Lucas, Julia Schaumburg, and Bernd Schwaab. Bank business models at zero interest rates. Journal of Business & Economic Statistics, 37(3):542–555, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2017.1386567.

Long:2011:EFD

[LSU11] Xiangdong Long, Liangjun Su, and Aman Ullah. Estimation and forecasting of dynamic conditional covariance: a semiparametric multivariate model. Journal of Business & Economic Statistics, 29(1):109–125, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.07057.

Li:2015:CAI

[LSX15] Yan Li, Liangjun Su, and Yuewu Xu. A combined approach to the inference of conditional factor models. Journal of Business & Economic Statistics, 33(2):203–220, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.940082.

Lucas:2014:CEA

[LSZ14] André Lucas, Bernd Schwaab, and Xin Zhang. Conditional Euro area sovereign default risk. Journal of Business & Economic Statistics, 32(2):271–284, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.873540.

llmakunnas:1985:TLH

[lT85] Pekka llmakunnas and Hiroki Tsurumi. Testing the Lucas hypothesis on output-inflation trade-offs. Journal of Business & Economic Statistics, 3(1):43–53, 1985. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1985.10509425.

Li:2000:SAM

[LT00] David X. Li and H. J. Turtle. Semiparametric ARCH models: an estimating function approach. Journal of Business & Economic Statistics, 18(2):174–186, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524860.

Lunde:2004:DDS

[LT04] Asger Lunde and Allan Timmermann. Duration dependence in stock prices. Journal of Business & Economic Statistics, 22(3):253–273, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000136.

Li:2012:MHP

[LT12] Qian Li and Pravin K. Trivedi. Medicare health plan choices of the elderly: a choice-with-screening model. Journal of Business & Economic Statistics, 30(1):81–93, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2011.08191.

Lacal:2019:ETN

[LT19] Virginia Lacal and Dag Tjøstheim. Estimating and testing nonlinear local dependence between two time series. Journal of Business & Economic Statistics, 37(4):648–660, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1407777.

Liu:2019:CMD

[LTMZ19] Philip Liu, Konstantinos Theodoridis, Haroon Mumtaz, and Francesco Zanetti. Changing macroeconomic dynamics at the zero lower bound. Journal of Business & Economic Statistics, 37(3):391–404, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1350186.

Li:2019:RTJ

[LTTL19] Jia Li, Viktor Todorov, George Tauchen, and Huidi Lin. Rank tests at jump events. Journal of Business & Economic Statistics, 37(2):312–321, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1328362.

Lee:2015:FEP

[LTU15] Tae-Hwy Lee, Yundong Tu, and Aman Ullah. Forecasting equity premium: Global historical average versus local historical average and constraints. Journal of Business & Economic Statistics, 33(3):393–402, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.955174.

Lundbergh:2003:TVS

[LTvD03] Stefan Lundbergh, Timo Teräsvirta, and Dick van Dijk. Timevarying smooth transition autoregressive models. Journal of Business & Economic Statistics, 21(1):104–121, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102288618810.

Ling:2019:TSC

[LTY19] Shiqing Ling, Ruey S. Tsay, and Yaxing Yang. Testing serial correlation and ARCH effect of high-dimensional timeseries data. Journal of Business & Economic Statistics, 39 (1):136–147, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1647844.

Li:2022:CPD

[LTY^{+}22] Shaobo Li, Shaonan Tian, Yan Yu, Xiaorui Zhu, and Heng Lian. Corporate probability of default: a single-index hazard model approach. Journal of Business & Economic Statistics, 41(4):1288– 1299, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2120484.

Lu:2015:CSC

[Lu15] Xun Lu. A covariate selection criterion for estimation of treatment effects. Journal of Business & Economic Statistics, 33(4):506–522, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.982755.

Lu:2022:SSP

[Lu22] Lina Lu. Simultaneous spatial panel data models with common shocks. Journal of Business & Economic Statistics, 41(2):608– 623, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2046007.

Lucas:2000:NOE

[Luc00] André Lucas. A note on optimal estimation from a riskmanagement perspective under possibly misspecified tail behavior. Journal of Business & Economic Statistics, 18 (1):31–39, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524845.

Lumsdaine:1995:FSP

[Lum95] Robin L. Lumsdaine. Finite-sample properties of the maximum likelihood estimator in GARCH(1,1) and IGARCH(1,1) models: a Monte Carlo investigation. Journal of Business & Economic Statistics, 13(1):1–10, 1995. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1995.10524574.

Luo:2002:CDM

[Luo02] Guo Ying Luo. Collective decision-making and heterogeneity in tastes. Journal of Business & Economic Statistics, 20 (2):213–226, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102317351967.

Lusardi:1996:PIC

[Lus96] Annamaria Lusardi. Permanent income, current income, and consumption: Evidence from two panel data sets. Journal of Business & Economic Statistics, 14(1):81–90, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524631.

Lutkepohl:1984:FCA

[Lüt84] Helmut Lütkepohl. Forecasting contemporaneously aggregated vector ARMA processes. Journal of Business & Economic Statistics, 2(3):201–214, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509388.

Lutkepohl:1986:FVA

[Lüt86] Helmut Lütkepohl. Forecasting vector ARMA processes with systematically missing observations. Journal of Business & Economic Statistics, 4(3):375–390, 1986. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1986.10509533.

Lutkepohl:1989:PTS

[Lüt89] Helmut Lütkepohl. Prediction tests for structural stability of multiple time series. Journal of Business & Economic Statistics, 7(1):129–135, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509720.

Lux:2008:MSM

[Lux08] Thomas Lux. The Markov-switching multifractal model of asset returns. Journal of Business & Economic Statistics, 26 (2):194–210, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500107000000403.

Lobato:2000:LMS

[LV00] Ignacio N. Lobato and Carlos Velasco. Long memory in stockmarket trading volume. Journal of Business & Economic Statistics, 18(4):410–427, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524881.

Lasak:2015:FCR

[LV15] Katarzyna Lasak and Carlos Velasco. Fractional cointegration rank estimation. Journal of Business & Economic Statistics, 33(2):241–254, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.945589.

Lillard:1997:UHS

[LW97] Lee A. Lillard and Yoram Weiss. Uncertain health and survival: Effects on end-of-life consumption. Journal of Business & Economic Statistics, 15(2):254–268, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524704.

Lee:2009:RUC

[LW09] Sokbae Lee and Ralf A. Wilke. Reform of unemployment compensation in Germany: a nonparametric bounds analysis using register data. Journal of Business & Economic Statistics, 27(2):193–205, 2009. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.0014.

Lin:2015:STC

[LW15] Juan Lin and Ximing Wu. Smooth tests of copula specifications. Journal of Business & Economic Statistics, 33(1):128–143, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2014.
932696.

Luo:2018:IQB

[LW18] Yao Luo and Yuanyuan Wan. Integrated-quantile-based estimation for first-price auction models. Journal of Business & Economic Statistics, 36(1):173–180, 2018. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2016.1166119.

Li:2019:EQE

[LW19a] Deyuan Li and Huixia Judy Wang. Extreme quantile estimation for autoregressive models. Journal of Business & Economic Statistics, 37(4):661–670, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1408469.

Lin:2019:NES

[LW19b] Haizhen Lin and Matthijs R. Wildenbeest. Nonparametric estimation of search costs for differentiated products: Evidence from Medigap. Journal of Business & Economic Statistics, 38 (4):754–770, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1573683.

Lin:2022:PSI

[LXHT22] Yuanyuan Lin, Jinhan Xie, Ruijian Han, and Niansheng Tang. Post-selection inference of high-dimensional logistic regression under case-control design. Journal of Business & Economic Statistics, 41(2):624–635, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2050245.

Lui:2022:GBC

[LXY22] Yiu Lim Lui, Weilin Xiao, and Jun Yu. The grid bootstrap for continuous time models. Journal of Business & Economic Statistics, 40(3):1390–1402, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1930014.

Lahiri:2018:CBR

[LY18] Kajal Lahiri and Liu Yang. Confidence bands for ROC curves with serially dependent data. Journal of Business & Economic Statistics, 36(1):115–130, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1073593.

Lee:2022:QEG

[LYY22] Lung-Fei Lee, Chao Yang, and Jihai Yu. QML and efficient GMM estimation of spatial autoregressive models with dominant (popular) units. Journal of Business & Economic Statistics, 41(2): 550–562, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2041424.

Liu:2022:QBI

[LYZ22] Xiaobin Liu, Thomas Tao Yang, and Yichong Zhang. Quasi-Bayesian inference for production frontiers. Journal of Business & Economic Statistics, 40(3):1334–1345, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1927745.

Lahiri:1987:MFU

[LZ87] Kajal Lahiri and Mark Zaporowski. More flexible use of survey data on expectations in macroeconomic models. Journal of Business & Economic Statistics, 5(1):69–76, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509561.

Levine:1995:BAH

[LZ95] Phillip B. Levine and David J. Zimmerman. The benefit of additional high-school math and science classes for young men and women. Journal of Business & Economic Statistics, 13 (2):137–149, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524588.

Ling:2014:Cb

[LZ14] Shiqing Ling and Ke Zhu. Comment. Journal of Business & Economic Statistics, 32(2):202–203, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.907059. See [FQX14a,

FQX14b].

Li:2018:VRP

[LZ18] Junye Li and Gabriele Zinna. The variance risk premium: Components, term structures, and stock return predictability. Journal of Business & Economic Statistics, 36(3):411–425, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1191502.

Luo:2019:MUS

[LZ19] Wei Luo and Yeying Zhu. Matching using sufficient dimension reduction for causal inference. Journal of Business & Economic Statistics, 38(4):888–900, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2019.1609974.

Luo:2023:TED

[LZGL23] Donghang Luo, Ke Zhu, Huan Gong, and Dong Li. Testing error distribution by kernelized Stein discrepancy in multivariate time series models. Journal of Business & Economic Statistics, 41(1): 111–125, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.2002160.

Li:2017:MDA

[LZLL17] Guodong Li, Qianqian Zhu, Zhao Liu, and Wai Keung Li. On mixture double autoregressive time series models. Journal of Business & Economic Statistics, 35(2):306–317, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1102735.

Li:2022:NPD

[LZW^{+}22] Jialu Li, Wan Zhang, Peiyao Wang, Qizhai Li, Kai Zhang, and Yufeng Liu. Nonparametric prediction distribution from resolution-wise regression with heterogeneous data. Journal of Business & Economic Statistics, 41(4):1157–1172, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2115498.

Liu:2020:MBT

[LZZ20] Mengya Liu, Fukang Zhu, and Ke Zhu. Multifrequencyband tests for white noise under heteroscedasticity. Journal of Business & Economic Statistics, 40(2):799–814, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1870478.

McKinney:2023:MEV

[MA23] Kevin L. McKinney and John M. Abowd. Male earnings volatility in LEHD before, during, and after the great recession. Journal of Business & Economic Statistics, 41(1):33– 39, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2126479.

Moffitt:2023:RTU

[MAB^{+}23] Robert Moffitt, John Abowd, Christopher Bollinger, Michael Carr, Charles Hokayem, Kevin McKinney, Emily Wiemers, Sisi Zhang, and James Ziliak. Reconciling trends in U.S. male earnings volatility: Results from survey and administrative data. Journal of Business & Economic Statistics, 41(1):1– 11, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2102020.

McShane:2008:CMB

[MABF08] Blake McShane, Moshe Adrian, Eric T. Bradlow, and Peter S. Fader. Count models based on Weibull interarrival times. Journal of Business & Economic Statistics, 26(3):369–378, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500107000000278.

MaCurdy:1992:C

[MaC92] Thomas MaCurdy. Comment. Journal of Business & Economic Statistics, 10(1):17–19, 1992. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1992.10509880.

MacKinnon:1994:AAD

[Mac94] James G. MacKinnon. Approximate asymptotic distribution functions for unit-root and cointegration tests. Journal of Business & Economic Statistics, 12(2):167–176, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10510005.

Magnac:2008:C

[Mag08] Thierry Magnac. Comment. Journal of Business & Economic Statistics, 26(3):295–297, 2008. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500108000000141. See [ALT08a, ALT08b].

Malhotra:1983:CPV

[Mal83] Naresh K. Malhotra. A comparison of the predictive validity of procedures for analyzing binary data. Journal of Business & Economic Statistics, 1(4):326–336, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1983.10509361.

Manganelli:2009:FJ

[Man09] Simone Manganelli. Forecasting with judgment. Journal of Business & Economic Statistics, 27(4):553–563, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.08052.

Manski:2013:C

[Man13] Charles F. Manski. Comment. Journal of Business & Economic Statistics, 31(3):273–275, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.792262. See [GPI13b, GPI13a].

Manzan:2015:FDE

[Man15] Sebastiano Manzan. Forecasting the distribution of economic variables in a data-rich environment. Journal of Business & Economic Statistics, 33(1):144–164, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.937436.

Maravall:1983:ANT

[Mar83] Agustin Maravall. An application of nonlinear time series forecasting. Journal of Business & Economic Statistics, 1 (1):66–74, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509325.

Maravall:1984:C

[Mar84] Agustin Maravall. Comment. Journal of Business & Economic Statistics, 2(4):337–339, 1984. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1984.10509405.

Maravall:1985:STS

[Mar85] Agustin Maravall. On structural time series models and the characterization of components. Journal of Business & Economic Statistics, 3(4):350–355, 1985. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1985.10509468.

Maravall:1987:MMS

[Mar87] Agustín Maravall. Minimum mean squared error estimation of the noise in unobserved component models. Journal of Business & Economic Statistics, 5(1):115–120, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509566.

Maravall:1998:C

[Mar98] Agustín Maravall. Comment. Journal of Business & Economic Statistics, 16(2):155–160, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524745.

Marcellino:1999:SCT

[Mar99] Massimiliano Marcellino. Some consequences of temporal aggregation in empirical analysis. Journal of Business & Economic Statistics, 17(1):129–136, 1999. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1999.10524802.

Mayer:1986:CLP

[May86] Lawrence S. Mayer. On cross-lagged panel models with serially correlated errors. Journal of Business & Economic Statistics, 4(3):347–357, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509531.

Michelson:1984:RRE

[MB84] Stephan Michelson and Gail Blattenberger. Reverse regression and employment discrimination. Journal of Business & Economic Statistics, 2(2):121–122, 1984. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1984.10509379.

McElroy:1988:APT

[MB88]  Marjorie B. McElroy and Edwin Burmeister.  Arbitrage

pricing theory as a restricted nonlinear multivariate regression model iterated nonlinear seemingly unrelated regression estimates. Journal of Business & Economic Statistics, 6 (1):29–42, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509634.

Moral-Benito:2013:LBE

[MB13] Enrique Moral-Benito. Likelihood-based estimation of dynamic panels with predetermined regressors. Journal of Business & Economic Statistics, 31(4):451–472, 2013. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2013.818003.

Morry:1998:C

[MC98] Marietta Morry and Norma Chhab. Comment. Journal of Business & Economic Statistics, 16(2):161–163, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524746.

McAleer:2014:C

[McA14] Michael McAleer. Comment. Journal of Business & Economic Statistics, 32(2):174–175, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.898584. See [GPI13b, GPI13a].

McCall:1995:IUI

[McC95] Brian P. McCall. The impact of unemployment insurance benefit levels on recipiency. Journal of Business & Economic Statistics, 13(2):189–198, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524593.

McCulloch:1997:MTT

[McC97] J. Huston McCulloch. Measuring tail thickness to estimate the stable index α: a critique. Journal of Business & Economic Statistics, 15(1):74–81, 1997. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1997.10524689.

McCloskey:2019:AUT

[McC19] Adam McCloskey. Asymptotically uniform tests after consistent model selection in the linear regression model. Journal of Business & Economic Statistics, 38(4):810–825, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1592754.

McDonald:1987:NML

[McD87a] John McDonald. A new model for learning curves, DARM. Journal of Business & Economic Statistics, 5(3):329–335, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1987.10509595.

McDonald:1987:R

[McD87b] John McDonald. Reply. Journal of Business & Economic Statistics, 5(3):338, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509597.

McElroy:2017:MSA

[McE17] Tucker McElroy. Multivariate seasonal adjustment, economic identities, and seasonal taxonomy. Journal of Business & Economic Statistics, 35(4):611–625, 2017. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2015.1123159.

McGrattan:1990:SSG

[McG90] Ellen R. McGrattan. Solving the stochastic growth model by linear-quadratic approximation. Journal of Business & Economic Statistics, 8(1):41–44, 1990. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1990.10509774.

McGarvey:1991:NPC

[McG91] Mary G. McGarvey. The neutrality properties of competing relative price models: Tests using linear feedback. Journal of Business & Economic Statistics, 9(1):15–25, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509823.

McGuckin:1995:EME

[McG95] Robert H. McGuckin. Establishment microdata for economic research and policy analysis: Looking beyond the aggregates. Journal of Business & Economic Statistics, 13 (1):121–126, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524585.

McKenzie:1984:CSA

[McK84] Sandra K. McKenzie. Concurrent seasonal adjustment with census X-11. Journal of Business & Economic Statistics, 2 (3):235–249, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509391.

McKenzie:2018:CBO

[McK18] David McKenzie. Can business owners form accurate counterfactuals? Eliciting treatment and control beliefs about their outcomes in the alternative treatment status. Journal of Business & Economic Statistics, 36(4):714–722, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1305276.

McNees:1986:FAA

[McN86a] Stephen K. McNees. Forecasting accuracy of alternative techniques: a comparison of U.S. macroeconomic forecasts. Journal of Business & Economic Statistics, 4(1):5–15, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509486.

McNees:1986:R

[McN86b] Stephen K. McNees. Reply. Journal of Business & Economic Statistics, 4(1):23, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509490.

Moeltner:2004:CBU

[ME04] Klaus Moeltner and Jeffrey Englin. Choice behavior under timevariant quality. Journal of Business & Economic Statistics, 22 (2):214–224, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000109.

Megdal:1987:EPL

[Meg87] Sharon Bernstein Megdal. The econometrics of piecewise-linear budget constraints. Journal of Business & Economic Statistics, 5(2):243–248, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509582.

Melnick:1990:DMA

[Mel90] Rafi Melnick. The demand for money in Argentina 1978–1987: Before and after the Austral program. Journal of Business & Economic Statistics, 8(4):427–434, 1990. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1990.10509813.

Melser:2005:HRT

[Mel05] Daniel Melser. The hedonic regression time-dummy method and the monotonicity axioms. Journal of Business & Economic Statistics, 23(4):485–492, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000578.

Melser:2018:SDP

[Mel18] Daniel Melser. Scanner data price indexes: Addressing some unresolved issues. Journal of Business & Economic Statistics, 36(3):516–522, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1218339.

Mele:2022:SMH

[Mel22] Angelo Mele. A structural model of homophily and clustering in social networks. Journal of Business & Economic Statistics, 40(3):1377–1389, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1930013.

Merz:1999:TSE

[Mer99] Monika Merz. Time series evidence of unemployment flows: The sample period matters. Journal of Business & Economic Statistics, 17(3):324–334, 1999. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1999.10524821.

Metin:1998:RBI

[Met98] Kivilcim Metin. The relationship between inflation and the budget deficit in Turkey. Journal of Business & Economic Statistics, 16(4):412–422, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524781.

Meyer:1988:CEM

[Mey88] Bruce D. Meyer. Classification-error models and labor-market dynamics. Journal of Business & Economic Statistics, 6 (3):385–390, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509676.

Meyer:1995:NQE

[Mey95] Breed D. Meyer. Natural and quasi-experiments in economics. Journal of Business & Economic Statistics, 13(2): 151–161, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524589.

Menezes-Filho:2005:CSC

[MF05] Naercio Menezes-Filho. Is the consumer sector competitive in the U.K.? A test using household-level demand elasticities and firmlevel price equations. Journal of Business & Economic Statistics, 23(3):295–304, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000514.

Meese:1984:CAU

[MG84] Richard Meese and John Geweke. A comparison of autoregressive univariate forecasting procedures for macroeconomic time series. Journal of Business & Economic Statistics, 2 (3):191–200, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509387.

McCloskey:2017:PER

[MH17] Adam McCloskey and Jonathan B. Hill. Parameter estimation robust to low-frequency contamination. Journal of Business & Economic Statistics, 35(4):598–610, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1093948.

Miyazaki:2019:DTS

[MHB19] Kei Miyazaki, Takahiro Hoshino, and Ulf Böckenholt. Dynamic two stage modeling for category-level and brand-level purchases using potential outcome approach with Bayes inference. Journal of Business & Economic Statistics, 39(3):622–635, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1702047.

Mele:2022:SEL

[MHCP22] Angelo Mele, Lingxin Hao, Joshua Cape, and Carey E. Priebe. Spectral estimation of large stochastic blockmodels with discrete nodal covariates. Journal of Business & Economic Statistics, 41(4):1364–1376, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2139709.

Mikusheva:2009:C

[Mik09] Anna Mikusheva. Comment. Journal of Business & Economic Statistics, 27(3):322–323, 2009. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.08263. See [KM09b, KM09a].

Miller:1984:SOS

[Mil84] John J. Miller. Some observations, a suggestion, and some comments on the Conway–Roberts article. Journal of Business & Economic Statistics, 2(2):123–125, 1984. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1984.10509380. See [CR83, FG84,

CR84].

Miller:1985:BR

[Mil85] Robert B. Miller. Book reviews. Journal of Business & Economic Statistics, 3(1):98–99, 1985. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1985.10509433.

Milhoj:1987:CVM

[Mil87] Anders Milhøj. A conditional variance model for daily deviations of an exchange rate. Journal of Business & Economic Statistics, 5(1):99–103, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509564.

Mishkin:1995:NRT

[Mis95] Frederic S. Mishkin. Nonstationarity of regressors and tests on real-interest-rate behavior. Journal of Business & Economic Statistics, 13(1):47–51, 1995. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1995.10524578.

Mittnik:1990:MFU

[Mit90] Stefan Mittnik. Macroeconomic forecasting using pooled international data. Journal of Business & Economic Statistics, 8 (2):205–208, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509791.

Miyazaki:1986:ELP

[MJY86] Shigetaka Miyazaki, George Judge, and Thomas Yancey. Estimation of location parameters under nonnormal errors and quadratic loss. Journal of Business & Economic Statistics, 4 (2):263–268, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509521.

Ma:2019:MCE

[MJY19] Jun Ma, Hugo Jales, and Zhengfei Yu. Minimum contrast empirical likelihood inference of discontinuity in density. Journal of Business & Economic Statistics, 38(4):934–950, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1617155.

Magdalinos:2001:SAE

[MK01] Michael Magdalinos and Helen Kandilorou. Specification analysis in equations with stochastic regressors. Journal of Business & Economic Statistics, 19(2):226–232, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500101316970449.

Menkveld:2007:MAC

[MKL07] Albert J. Menkveld, Siem Jan Koopman, and André Lucas. Modeling around-the-clock price discovery for cross-listed stocks using state space methods. Journal of Business & Economic Statistics, 25(2):213–225, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000594.

McDonald:1988:ANH

[ML88] Bill McDonald and Cheng-Few Lee. An analysis of nonlinearities, heteroscedasticity, and functional form in the market model. Journal of Business & Economic Statistics, 6 (4):505–509, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509699.

McHugh:1990:ETC

[ML90] Richard McHugh and Julia Lane. Embodied technological change and tests of the internal-adjustment-cost hypothesis. Journal of Business & Economic Statistics, 8(4):

459–464, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509817.

Morgan:1987:MAR

[MlM87] Alison Morgan and leuan Morgan. Measurement of abnormal returns from small firms. Journal of Business & Economic Statistics, 5(1):121–129, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509567.

Millimet:2019:PIE

[MLR19] Daniel L. Millimet, Hao Li, and Punarjit Roychowdhury. Partial identification of economic mobility: With an application to the United States. Journal of Business & Economic Statistics, 38(4):732–753, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1569527.

Ma:2020:TAC

[MLST20] Shujie Ma, Wei Lan, Liangjun Su, and Chih-Ling Tsai. Testing alphas in conditional time-varying factor models with highdimensional assets. Journal of Business & Economic Statistics, 38(1):214–227, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1482758.

MacDonald:2022:DCC

[MLZ22] Peter W. MacDonald, Elizaveta Levina, and Ji Zhu. Discussion of “Co-citation and Co-authorship Networks of Statisticians” by Pengsheng Ji, Jiashun Jin, Zheng Tracy Ke, and Wanshan Li. Journal of Business & Economic Statistics, 40(2):492–493, 2022. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2022.2041423. See [JJKL21].

Mayfield:1992:DDR

[MM92] E. Scott Mayfield and Bruce Mizrach. On determining the dimension of real-time stock-price data. Journal of Business & Economic Statistics, 10(3):367–374, 1992. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1992.10509911.

Maheu:2000:IBB

[MM00] John M. Maheu and Thomas H. McCurdy. Identifying bull and bear markets in stock returns. Journal of Business & Economic Statistics, 18(1):100–112, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524851.

Maliar:2003:PEA

[MM03] Lilia Maliar and Serguei Maliar. Parameterized expectations algorithm and the moving bounds. Journal of Business & Economic Statistics, 21(1):88–92, 2003. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102288618793.

Maheu:2009:HUH

[MM09] John M. Maheu and Thomas H. McCurdy. How useful are historical data for forecasting the long-run equity return distribution? Journal of Business & Economic Statistics, 27(1):95–112, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2009.0008.

Manski:2010:RPE

[MM10] Charles F. Manski and Francesca Molinari. Rounding probabilistic expectations in surveys. Journal of Business & Economic Statistics, 28(2):219–231, 2010. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.08098.

Mumtaz:2019:EIG

[MM19] Haroon Mumtaz and Alberto Musso. The evolving impact of global, region-specific, and country-specific uncertainty. Journal of Business & Economic Statistics, 39(2):466–481, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1668798.

Masini:2020:CAI

[MM20] Ricardo Masini and Marcelo C. Medeiros. Counterfactual analysis and inference with nonstationary data. Journal of Business & Economic Statistics, 40(1):227–239, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1799814.

Majumdar:2006:GSR

[MMG^{+}06] Anandamayee Majumdar, Henry J. Munneke, Alan E. Gelfand, Sudipto Banerjee, and C. F. Sirmans. Gradients in spatial response surfaces with application to urban land values. Journal of Business & Economic Statistics, 24(1):77–90, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500105000000162.

McCracken:2022:BCF

[MMO22] Michael W. McCracken, Joseph T. McGillicuddy, and Michael T. Owyang. Binary conditional forecasts. Journal of Business & Economic Statistics, 40(3):1246–1258, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1920960.

Maheu:2012:CBB

[MMS12] John M. Maheu, Thomas H. McCurdy, and Yong Song. Components of bull and bear markets: Bull corrections and bear rallies. Journal of Business & Economic Statistics, 30(3):391–403, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2012.
680412.

Maasoumi:1988:MMW

[MN88] Esfandiar Maasoumi and Gerald Nickelsburg. Multivariate measures of well-being and an analysis of inequality in the Michigan data. Journal of Business & Economic Statistics, 6 (3):327–334, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509669.

Masulis:1995:ODS

[MN95a] Ronald W. Masulis and Victor K. Ng. Overnight and daytime stock-return dynamites on the London Stock Exchange: The impacts of “Big Bang” and the 1987 stockmarket crash. Journal of Business & Economic Statistics, 13 (4):365–378, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524612.

Miller:1995:UAP

[MN95b] John P. Miller and Paul Newbold. Uncertainty about the persistence of economic shocks. Journal of Business & Economic Statistics, 13(4):435–440, 1995. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1995.10524617.

McCracken:2016:FMM

[MN16] Michael W. McCracken and Serena Ng. FRED-MD: a monthly database for macroeconomic research. Journal of Business & Economic Statistics, 34(4):574–589, 2016. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2015.1086655.

Munro:2020:LDA

[MN20] Evan Munro and Serena Ng. Latent Dirichlet analysis of categorical survey responses. Journal of Business & Economic Statistics, 40(1):256–271, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1802285.

Murnane:1985:CPP

[MNO85] Richard J. Murnane, Stuart Newstead, and Randall J. Olsen. Comparing public and private schools: The puzzling role of selectivity bias. Journal of Business & Economic Statistics, 3(1):23–35, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509423.

MacKinnon:2019:WBA

[MNW19] James G. MacKinnon, Morten Ørregaard Nielsen, and Matthew D. Webb. Wild bootstrap and asymptotic inference with multiway clustering. Journal of Business & Economic Statistics, 39 (2):505–519, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1677473.

Mathiowetz:1988:WMR

[MO88] Nancy A. Mathiowetz and Greg J. Ouncan. Out of work, out of mind: Response errors in retrospective reports of unemployment. Journal of Business & Economic Statistics, 6 (2):221–229, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509656.

Moffitt:1986:EPL

[Mof86] Robert Moffitt. The econometrics of piecewise-linear budget constraints. Journal of Business & Economic Statistics, 4 (3):317–328, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509528.

Muirhead:1986:BR

[MOJ^{+}86] Robb J. Muirhead, Alan J. Oppenheim, Keith N. Johnson, William Greene, and Halina Frydman. Book reviews. Journal of Business & Economic Statistics, 4(1):135–141, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1986.10509506.

Molinari:2010:MT

[Mol10] Francesca Molinari. Missing treatments. Journal of Business & Economic Statistics, 28(1):82–95, 2010. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/jbes.2009.07161.

Montgomery:1988:C

[Mon88] David B. Montgomery. Comment. Journal of Business & Economic Statistics, 6(2):163–164, 1988. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1988.10509650.

Montalvo:1997:GEC

[Mon97] Jose G. Montalvo. GMM estimation of count-panel-data models with fixed effects and predetermined instruments. Journal of Business & Economic Statistics, 15(1):82–89, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524690.

Morrison:1985:PDC

[Mor85] Catherine J. Morrison. Primal and dual capacity utilization: an application to productivity measurement in the U.S. automobile industry. Journal of Business & Economic Statistics, 3(4):312–324, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509463.

Mork:1987:ABF

[Mor87] Knut Anton Mork. Ain’t behavin’: Forecast errors and measurement errors in early GNP estimates. Journal of Business & Economic Statistics, 5(2):165–175, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509575.

Morrison:1992:MUJ

[Mor92] Catherine J. Morrison. Markups in U.S. and Japanese manufacturing: a short-run econometric analysis. Journal of Business & Economic Statistics, 10(1):51–63, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1992.10509886.

Morrison:1993:ICA

[Mor93] Catherine J. Morrison. Investment in capital assets and economic performance: The U.S. chemicals and primary-metals industries in transition. Journal of Business & Economic Statistics, 11(1):45–60, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509932.

Mortimer:2023:DLB

[Mor23] Julie Holland Mortimer. Discussion of Levon Barseghyan and Francesca Molinari’s “Risk Preference Types, Limited Consideration, and Welfare”. Journal of Business & Economic Statistics, 41(4):1042–1045, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2023.2223592. See [BM23b].

Moschini:1991:TPC

[Mos91] Giancarlo Moschini. Testing for preference change in consumer demand: an indirectly separable, semiparametric model. Journal of Business & Economic Statistics, 9(1):111–117, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1991.10509833.

Moschini:1999:ILC

[Mos99] Giancarlo Moschini. Imposing local curvature conditions in flexible demand systems. Journal of Business & Economic Statistics, 17(4):487–490, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524837.

Matsushita:2022:EDR

[MOT22] Yukitoshi Matsushita, Taisuke Otsu, and Keisuke Takahata. Estimating density ratio of marginals to joint: Applications to causal inference. Journal of Business & Economic Statistics, 41(2):467– 481, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2035228.

Moulton:1987:DGE

[Mou87] Brent R. Moulton. Diagnostics for group effects in regression analysis. Journal of Business & Economic Statistics, 5 (2):275–282, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509586.

McGuckin:2007:MTU

[MOZ07] Robert H. McGuckin, Ataman Ozyildirim, and Victor Zarnowitz. A more timely and useful index of leading indicators. Journal of Business & Economic Statistics, 25(1):110–120, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000279.

Maravall:1983:PDE

[MP83] Agustin Maravall and David A. Pierce. Preliminary-data error and monetary aggregate targeting. Journal of Business & Economic Statistics, 1(3):179–186, 1983. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1983.10509337.

Murray:2005:DPH

[MP05] Christian J. Murray and David H. Papell. Do panels help solve the purchasing power parity puzzle? Journal of Business & Economic Statistics, 23(4):410–415, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500105000000072.

Matsumura:1991:CPT

[MPTK91] Ella Mae Matsumura, Robert Plante, Kam-Wah Tsui, and P. Kannan. Comparative performance of two muitinomial-based methods for obtaining lower bounds on the total overstatement error in accounting populations. Journal of Business & Economic Statistics, 9(4):423–429, 1991. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1991.10509869.

Marcellino:2016:STG

[MPV16] Massimiliano Marcellino, Mario Porqueddu, and Fabrizio Venditti. Short-term GDP forecasting with a mixed-frequency dynamic factor model with stochastic volatility. Journal of Business & Economic Statistics, 34(1):118–127, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1006773.

Monache:2020:PDR

[MPV20] Davide Delle Monache, Ivan Petrella, and Fabrizio Venditti. Price dividend ratio and long-run stock returns: a score-driven state space model. Journal of Business & Economic Statistics, 39 (4):1054–1065, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1763805.

Ma:2022:CQR

[MQZ22] Huijuan Ma, Jing Qin, and Yong Zhou. From conditional quantile regression to marginal quantile estimation with applications to missing data and causal inference. Journal of Business & Economic Statistics, 41(4):1377–1390, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2140158.

Miller:1991:QSL

[MR91a] Preston J. Miller and William T. Roberds. The quantitative significance of the Lucas critique. Journal of Business & Economic Statistics, 9(4):361–387, 1991. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1991.10509864.

Miller:1991:R

[MR91b] Preston J. Miller and William T. Roberds. Reply. Journal of Business & Economic Statistics, 9(4):389, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509866.

Molinari:2008:C

[MR08] Francesca Molinari and Adam M. Rosen. Comment. Journal of Business & Economic Statistics, 26(3):297–302, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500108000000088. See

[ALT08a, ALT08b].

Meijer:2012:MEE

[MRW12] Erik Meijer, Susann Rohwedder, and Tom Wansbeek. Measurement error in earnings data: Using a mixture model approach to combine survey and register data. Journal of Business & Economic Statistics, 30(2):191–201, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1198/jbes.2011.08166.

Mitchell:2019:BPM

[MRW19] James Mitchell, Donald Robertson, and Stephen Wright. R^{2} bounds for predictive models: What univariate properties tell us about multivariate predictability. Journal of Business & Economic Statistics, 37(4):681–695, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2017.1415909.

Marshall:1992:PPI

[MRZ92a] Robert C. Marshall, Jean-Francois Richard, and Gary A. Zarkin. Posterior probabilities of the independence axiom with nonexperimental data (or buckle up and fan out). Journal of Business & Economic Statistics, 10(1):31–44, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1992.10509883.

Marshall:1992:R

[MRZ92b] Robert C. Marshall, Jean-Francois Richard, and Gary A. Zarkin. Reply. Journal of Business & Economic Statistics, 10(1):48–49, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509885.

Morrison:1988:GNM

[MS88a] Donald G. Morrison and David C. Schmittlein. Generalizing the NBD model for customer purchases: What are the implications and is it worth the effort? Journal of Business & Economic Statistics, 6(2):145–159, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509648.

Morrison:1988:R

[MS88b] Donald G. Morrison and David C. Schmittlein. Reply. Journal of Business & Economic Statistics, 6(2):165–166, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1988.10509651.

Mork:1989:TLC

[MS89] Knut Anton Mork and V. Kerry Smith. Testing the life-cycle hypothesis with a Norwegian household panel. Journal of Business & Economic Statistics, 7(3):287–296, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509738.

Morrison:1996:PIP

[MS96] Catherine J. Morrison and Amy Ellen Schwartz. Public infrastructure, private input demand, and economic performance in new England manufacturing. Journal of Business & Economic Statistics, 14(1):91–101, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524632.

Miller:1997:MCH

[MS97] Robert A. Miller and Holger Sieg. A microeconometric comparison of household behavior between countries. Journal of Business & Economic Statistics, 15(2):237–253, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524703.

Moulton:1999:OEU

[MS99] Brent R. Moulton and Kenneth J. Stewart. An overview of experimental U.S. consumer price indexes. Journal of Business & Economic Statistics, 17(2):141–151, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524804.

Moriarity:2003:NRS

[MS03] Chris Moriarity and Fritz Scheuren. A note on Rubin’s statistical matching using file concatenation with adjusted weights and multiple imputations. Journal of Business & Economic Statistics, 21(1):65–73, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618766.

Mencia:2018:VRE

[MS18] Javier Mencía and Enrique Sentana. Volatility-related exchange traded assets: an econometric investigation. Journal of Business & Economic Statistics, 36(4):599–614, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2016.1216852.

Mishra:2010:SET

[MSU10] Santosh Mishra, Liangjun Su, and Aman Ullah. Semiparametric estimator of time series conditional variance. Journal of Business & Economic Statistics, 28(2):256–274, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.08118.

Machak:1983:APT

[MSW83] Joseph A. Machak, W. Allen Spivey, and William J. Wrobleski. Analyzing permanent and transient influences in multiple time series models. Journal of Business & Economic Statistics, 1(1):57–65, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509324.

Machak:1985:FTV

[MSW85] Joseph A. Machak, W. Allen Spivey, and William J. Wrobleski. A framework for time varying parameter regression modeling. Journal of Business & Economic Statistics, 3 (2):104–111, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509437.

Murphy:1985:EIT

[MT85] Kevin M. Murphy and Robert H. Topel. Estimation and inference in two-step econometric models. Journal of Business & Economic Statistics, 3(4):370–379, 1985. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1985.10509471.

Morgan:1996:IOC

[MT96] Michael S. Morgan and Minakshi Trivedi. Inferring the order of the choice process using consumer purchase histories. Journal of Business & Economic Statistics, 14(2): 221–229, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524649.

Morgan:1999:LMC

[MT99] I. G. Morgan and R. G. Trevor. Limit moves as censored observations of equilibrium futures price in GARCH processes. Journal of Business & Economic Statistics, 17 (4):397–408, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524828.

Murphy:2002:EIT

[MT02] Kevin M. Murphy and Robert H. Topel. Estimation and inference in two-step econometric models. Journal of Business & Economic Statistics, 20(1):88–97, 2002. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102753410417.

McMillen:2003:AT

[MT03] Daniel P. McMillen and Paul Thorsnes. The aroma of Tacoma. Journal of Business & Economic Statistics, 21 (2):237–246, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288618927.

Meitz:2006:EMA

[MT06] Mika Meitz and Timo Teräsvirta. Evaluating models of autoregressive conditional duration. Journal of Business & Economic Statistics, 24(1):104–124, 2006. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500105000000081.

Millimet:2009:SPS

[MT09] Daniel L. Millimet and Rusty Tchernis. On the specification of propensity scores, with applications to the analysis of trade policies. Journal of Business & Economic Statistics, 27(3):397–415, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1198/jbes.2009.06045.

Mumtaz:2018:CTU

[MT18] Haroon Mumtaz and Konstantinos Theodoridis. The changing transmission of uncertainty shocks in the U.S. Journal of Business & Economic Statistics, 36(2):239–252, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2016.1147357.

Maddala:1997:ESR

[MTLJ97] G. S. Maddala, Robert P. Trost, Hongyi Li, and Frederick Joutz. Estimation of short-run and long-run elasticities of energy demand from panel data using shrinkage estimators. Journal of Business & Economic Statistics, 15(1):90–100, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524691.

Muller:2014:HCS

[Mül14a] Ulrich K. Müller. HAC corrections for strongly autocorrelated time series. Journal of Business & Economic Statistics, 32(3):311–322, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.931238. See comments [Kie14, CC14,

Sun14, Vog14] and rejoinder [Mül14b].

Muller:2014:R

[Mül14b] Ulrich K. Müller. Rejoinder. Journal of Business & Economic Statistics, 32(3):338–340, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.931769. See [Mül14a].

Muller:2018:CHI

[Mül18] Ulrich K. Müller. Comment on “HAR Inference: Recommendations for Practice” by E. Lazarus, D. J. Lewis, J. H. Stock and M. W. Watson. Journal of Business & Economic Statistics, 36(4):563–564, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1497502. See [LLSW18a, Wes18a].

Maeshiro:1988:PED

[MV88] Asatoshi Maeshiro and Shapoor Vali. Pitfalls in the estimation of a differenced model. Journal of Business & Economic Statistics, 6(4):511–515, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509700.

Magee:1991:SRP

[MV91] Lonnie Magee and Michael R. Veall. Selecting regressors for prediction using Press and White t statistics. Journal of Business & Economic Statistics, 9(1):91–96, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509830.

McCracken:2018:AIP

[MV18] Michael W. McCracken and Giorgio Valente. Asymptotic inference for performance fees and the predictability of asset returns. Journal of Business & Economic Statistics, 36(3):426–437, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1215317.

Medeiros:2019:FID

[MVVZ19] Marcelo C. Medeiros, Gabriel F. R. Vasconcelos, Álvaro Veiga, and Eduardo Zilberman. Forecasting inflation in a data-rich environment: The benefits of machine learning methods. Journal of Business & Economic Statistics, 39(1):98–119, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1637745.

Muller:2022:SCR

[MW22] Ulrich K. Müller and Mark W. Watson. Spatial correlation robust inference in linear regression and panel models. Journal of Business & Economic Statistics, 41(4):1050–1064, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2127737.

Milunovich:2018:SES

[MY18] George Milunovich and Minxian Yang. Simultaneous equation systems with heteroscedasticity: Identification, estimation, and stock price elasticities. Journal of Business & Economic Statistics, 36(2):288–308, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1149072.

Moffitt:2023:ETM

[MZ23] Robert Moffitt and Sisi Zhang. Estimating trends in male earnings volatility with the panel study of income dynamics. Journal of Business & Economic Statistics, 41(1):20–25, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2102024.

Nelson:1992:ICU

[NC92] Daniel B. Nelson and Charles Q. Cao. Inequality constraints in the univariate GARCH model. Journal of Business & Economic Statistics, 10(2):229–235, 1992. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1992.10509902.

Neftci:1993:SAS

[Nef93] Salih N. Neftci. Statistical analysis of shapes in macroeconomic time series: Is there a business cycle? Journal of Business & Economic Statistics, 11(2):215–224, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509950.

Nelson:1994:C

[Nel94] Daniel B. Nelson. Comment. Journal of Business & Economic Statistics, 12(4):403–406, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10524560.

Nevo:2003:UWA

[Nev03] Aviv Nevo. Using weights to adjust for sample selection when auxiliary information is available. Journal of Business & Economic Statistics, 21(1):43–52, 2003. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102288618748.

Noh:2015:SCQ

[NEV15] Hohsuk Noh, Anouar El Ghouch, and Ingrid Van Keilegom. Semiparametric conditional quantile estimation through copula-based multivariate models. Journal of Business & Economic Statistics, 33(2):167–178, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.926171.

Newbold:1983:C

[New83] Paul Newbold. Comment. Journal of Business & Economic Statistics, 1(4):311–312, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509358.

Nayak:1989:UDA

[NG89] Tapan K. Nayak and Joseph L. Gastwirth. The use of diversity analysis to assess the relative influence of factors affecting the income distribution. Journal of Business & Economic Statistics, 7(4):453–460, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509757.

Ng:2006:TCS

[Ng06] Serena Ng. Testing cross-section correlation in panel data using spacings. Journal of Business & Economic Statistics, 24(1):12–23, 2006. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500105000000171.

Ng:2008:STN

[Ng08] Serena Ng. A simple test for nonstationarity in mixed panels. Journal of Business & Economic Statistics, 26(1): 113–127, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000675.

Nicol:1989:TTE

[Nic89] Christopher J. Nicol. Testing a theory of exact aggregation. Journal of Business & Economic Statistics, 7(2): 259–265, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509735.

Nielsen:2004:ORB

[Nie04] Morten Ørregaard Nielsen. Optimal residual-based tests for fractional cointegration and exchange rate dynamics. Journal of Business & Economic Statistics, 22(3):331–345, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500104000000181.

Nielsen:2007:LWA

[Nie07] Morten Ørregaard Nielsen. Local Whittle analysis of stationary fractional cointegration and the implied-realized volatility relation. Journal of Business & Economic Statistics, 25 (4):427–446, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000314.

Nelson:1984:PUT

[NK84] Charles R. Nelson and Heejoon Kang. Pitfalls in the use of time as an explanatory variable in regression. Journal of Business & Economic Statistics, 2(1):73–82, 1984. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1984.10509371.

Ng:2008:ER

[NL08] Serena Ng and Arthur Lewbel. Editors’ report 2007. Journal of Business & Economic Statistics, 26(4):557, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500108000000187.

Ng:2009:ER

[NL09] Serena Ng and Arthur Lewbel. Editors’ report 2008. Journal of Business & Economic Statistics, 27(4):566, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.274er.

Nakamura:1987:BR

[NNH87] Alice Nakamura, Masao Nakamura, and Marilyn Hart. Book reviews. Journal of Business & Economic Statistics, 5(2): 309–311, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509591.

Nakamura:1988:BR

[NNW88] Alice Nakamura, Masao Nakamura, and Frank A. Wolak. Book reviews. Journal of Business & Economic Statistics, 6 (4):517–520, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509701.

Norwood:1985:CAS

[Nor85] Janet L. Norwood. Comment: Administrative statistics: a BLS perspective. Journal of Business & Economic Statistics, 3(4):398–400, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509476.

Normandin:1994:PSE

[Nor94] Michel Normandin. Precautionary saving: an explanation for excess sensitivity of consumption. Journal of Business & Economic Statistics, 12(2):205–219, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10510008.

Nordhaus:1999:BCA

[Nor99] William D. Nordhaus. Beyond the CPI: an augmented costof-living index. Journal of Business & Economic Statistics, 17 (2):182–187, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524808.

Nelson:1985:NFR

[NP85] Charles R. Nelson and Stephen C. Peck. The NERC fan: a retrospective analysis of the NERC summary forecasts. Journal of Business & Economic Statistics, 3(3):179–187, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1985.10509450.

Nijman:1986:CUA

[NP86] T. E. Nijman and F. C. Palm. The construction and use of approximations for missing quarterly observations: a modelbased approach. Journal of Business & Economic Statistics, 4(1):47–58, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509493.

Nijman:1990:PAG

[NP90] Theo E. Nijman and Franz C. Palm. Predictive accuracy gain from disaggregate sampling in ARIMA models. Journal of Business & Economic Statistics, 8(4):405–415, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1990.10509811.

Nandram:1997:BAA

[NP97] Balgobin Nandram and Joseph D. Petruccelli. A Bayesian analysis of autoregressive time series panel data. Journal of Business & Economic Statistics, 15(3):328–334, 1997. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1997.10524710.

Nicoletti:2011:EIP

[NPF11] Cheti Nicoletti, Franco Peracchi, and Francesca Foliano. Estimating income poverty in the presence of missing data and measurement error. Journal of Business & Economic Statistics, 29(1):61–72, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2010.07185.

Nijman:1993:PFF

[NPW93] Theo E. Nijman, Franz C. Palm, and Christian C. P. Wolff. Premia in forward foreign exchange as unobserved components: a note. Journal of Business & Economic Statistics, 11 (3):361–365, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509964.

Nelson:2001:MRS

[NPZ01] Charles R. Nelson, Jeremy Piger, and Eric Zivot. Markov regime switching and unit-root tests. Journal of Business & Economic Statistics, 19(4):404–415, 2001. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/07350010152596655.

Neely:2001:RAV

[NRW01] Christopher J. Neely, Amlan Roy, and Charles H. Whiteman. Risk aversion versus intertemporal substitution. Journal of Business & Economic Statistics, 19(4):395–403, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/07350010152596646.

Nankervis:1996:LPB

[NS96] John C. Nankervis and N. E. Savin. The level and power of the bootstrap t test in the AR(1) model with trend. Journal of Business & Economic Statistics, 14(2):161–168, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524642.

Ni:2005:BEV

[NS05] Shawn Ni and Dongchu Sun. Bayesian estimates for vector autoregressive models. Journal of Business & Economic Statistics, 23(1):105–117, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000622.

Nankervis:2010:TSC

[NS10] John C. Nankervis and N. E. Savin. Testing for serial correlation: Generalized Andrews–Ploberger tests. Journal of Business & Economic Statistics, 28(2):246–255, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.08115.

Narasimham:1988:PAU

[NSR88] Gorti V. L. Narasimham, P. A. V. B. Swamy, and R. C. Reed. Productivity analysis of U.S. manufacturing using a stochasticcoefficients production function. Journal of Business & Economic Statistics, 6(3):339–349, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509671.

Ni:2007:IBE

[NSS07] Shawn Ni, Dongchu Sun, and Xiaoqian Sun. Intrinsic Bayesian estimation of vector autoregression impulse responses. Journal of Business & Economic Statistics, 25(2):163–176, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000378.

Nguyen:2022:SRS

[NTGK22] Trong-Nghia Nguyen, Minh-Ngoc Tran, David Gunawan, and Robert Kohn. A statistical recurrent stochastic volatility model for stock markets. Journal of Business & Economic Statistics, 41(2):414–428, 2022. CODEN ???? ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2022.2028631.

Nelson:1994:QAP

[NTP94] Randy A. Nelson, Tim L. Tanguay, and Christopher D. Patterson. A quality-adjusted price index for personal computers. Journal of Business & Economic Statistics, 12 (1):23–31, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10509988.

Nolte:2012:LSI

[NV12] Ingmar Nolte and Valeri Voev. Least squares inference on integrated volatility and the relationship between efficient prices and noise. Journal of Business & Economic Statistics, 30(1):94– 108, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/10473289.
2011.637876.

Neumark:1991:CWI

[NW91] David Neumark and William L. Wascher. Can we improve upon preliminary estimates of payroll employment growth? Journal of Business & Economic Statistics, 9(2):197–205, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1991.10509845.

Neumark:1995:MWE

[NW95] David Neumark and William Wascher. Minimum wage effects on employment and school enrollment. Journal of Business & Economic Statistics, 13(2):199–206, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524594.

Nakajima:2013:BAL

[NW13] Jouchi Nakajima and Mike West. Bayesian analysis of latent threshold dynamic models. Journal of Business & Economic Statistics, 31(2):151–164, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.747847.

Naik:2010:MIB

[NWK10] Prasad A. Naik, Michel Wedel, and Wagner Kamakura. Multiindex binary response analysis of large data sets. Journal of Business & Economic Statistics, 28(1):67–81, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.07170.

Nolde:2020:CEA

[NZ20] Natalia Nolde and Jinyuan Zhang. Conditional extremes in asymmetric financial markets. Journal of Business & Economic Statistics, 38(1):201–213, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1476248.

Ooms:1997:PCB

[OF97] Marius Ooms and Philip Hans Franses. On periodic correlations between estimated seasonal and nonseasonal components in German and U.S. unemployment. Journal of Business & Economic Statistics, 15(4):470–481, 1997. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1997.10524725.

Ohta:1986:APQ

[OG86] Makoto Ohta and Zvi Griliches. Automobile prices and quality: Did the gasoline price increases change consumer tastes in the U.S.? Journal of Business & Economic Statistics, 4 (2):187–198, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509514.

Ozabaci:2014:ANR

[OHS14] Deniz Ozabaci, Daniel J. Henderson, and Liangjun Su. Additive nonparametric regression in the presence of endogenous regressors. Journal of Business & Economic Statistics, 32(4):555–575, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.917590.

Ohta:1987:GCH

[Oht87] Makoto Ohta. Gasoline cost and hedonic price indexes of U.S. used cars for 1970–1983: a note. Journal of Business & Economic Statistics, 5(4):521–528, 1987. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1987.10509620.

Opschoor:2018:NHM

[OJLD18] Anne Opschoor, Pawel Janus, André Lucas, and Dick Van Dijk. New HEAVY models for fat-tailed realized covariances and returns. Journal of Business & Economic Statistics, 36(4):643–657, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1245622.

Okumura:2011:NEL

[Oku11] Tsunao Okumura. Nonparametric estimation of labor supply and demand factors. Journal of Business & Economic Statistics, 29(1):174–185, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2010.08068.

Opschoor:2020:CFM

[OLBvD20] Anne Opschoor, André Lucas, István Barra, and Dick van Dijk. Closed-form multi-factor copula models with observation-driven dynamic factor loadings. Journal of Business & Economic Statistics, 39(4):1066–1079, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2020.1763806.

Oliver:1987:BFS

[Oli87] Robert M. Oliver. Bayesian forecasting with stable seasonal patterns. Journal of Business & Economic Statistics, 5(1):77–85, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509562.

Oller:1986:NES

[Öll86] Lars-Erik Öller. A note on exponentially smoothed seasonal differences. Journal of Business & Economic Statistics, 4 (4):485–489, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509547.

Olmstead:2009:RFV

[Olm09] Sheila M. Olmstead. Reduced-form versus structural models of water demand under nonlinear prices. Journal of Business & Economic Statistics, 27(1):84–94, 2009. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/jbes.2009.0007.

Oomen:2006:C

[Oom06a] Roel C. A. Oomen. Comment. Journal of Business & Economic Statistics, 24(2):195–202, 2006. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500106000000125. See [HL06a, HL06b].

Oomen:2006:PRV

[Oom06b] Roel C. A. Oomen. Properties of realized variance under alternative sampling schemes. Journal of Business & Economic Statistics, 24(2):219–237, 2006. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500106000000044.

Olea:2013:RTW

[OP13] José Luis Montiel Olea and Carolin Pflueger. A robust test for weak instruments. Journal of Business & Economic Statistics, 31(3):358–369, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/00401706.2013.806694.

Oh:2015:MDH

[OP15] Dong Hwan Oh and Andrew J. Patton. Modeling dependence in high dimensions with factor copulas. Journal of Business & Economic Statistics, 35(1):139–154, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1062384.

Oh:2018:TVS

[OP18] Dong Hwan Oh and Andrew J. Patton. Time-varying systemic risk: Evidence from a dynamic copula model of CDS spreads. Journal of Business & Economic Statistics, 36(2):181–195, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1177535.

Ord:1993:C

[Ord93] Keith Ord. Comment. Journal of Business & Economic Statistics, 11(2):138–139, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509940.

Ohanissian:2008:TSL

[ORT08] Arek Ohanissian, Jeffrey R. Russell, and Ruey S. Tsay. True or spurious long memory? A new test. Journal of Business & Economic Statistics, 26(2):161–175, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500107000000340.

Osborn:1989:PPA

[OS89] Denise R. Osborn and Jeremy P. Smith. The performance of periodic autoregressive models in forecasting seasonal U.K. consumption. Journal of Business & Economic Statistics, 7 (1):117–127, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509719.

Oparina:2020:ASW

[OS20] Ekaterina Oparina and Sorawoot Srisuma. Analyzing subjective well-being data with misclassification. Journal of Business & Economic Statistics, 40(2):730–743, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1865169.

Ostry:1984:OEO

[Ost84] Sylvia Ostry. OECD economic outlook: How we got here and where we’re headed. Journal of Business & Economic Statistics, 2(2):105–109, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509374.

Oster:2019:USC

[Ost19] Emily Oster. Unobservable selection and coefficient stability: Theory and evidence. Journal of Business & Economic Statistics, 37(2):187–204, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1227711. See comments [DMP19].

Otneim:2021:LGP

[OT21] Håkon Otneim and Dag Tjøstheim. The locally Gaussian partial correlation. Journal of Business & Economic Statistics, 40(2): 924–936, 2021. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/07350015.2021.1886107.

Otter:1990:CCM

[Ott90] Pieter W. Otter. Canonical correlation in multivariate time series analysis with an application to one-year-ahead and multiyearahead macroeconomic forecasting. Journal of Business & Economic Statistics, 8(4):453–457, 1990. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1990.10509816. See comments

[ZH92].

Otter:1992:R

[Ott92] Pieter W. Otter. Reply. Journal of Business & Economic Statistics, 10(4):471–472, 1992. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1992.10509925.

Oum:1991:NCU

[OTZ91] Tae H. Oum, Michael W. Tretheway, and Yimin Zhang. A note on capacity utilization and measurement of scale economies. Journal of Business & Economic Statistics, 9 (1):119–123, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509834.

Ottenwaelter:1984:EAB

[OV84] B. Ottenwaelter and Q. H. Vuong. An empirical analysis of backlog, inventory, production, and price adjustments: an application of recursive systems of log-linear models. Journal of Business & Economic Statistics, 2(3):224–234, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1984.10509390.

Otsu:2013:EID

[OXM13] Taisuke Otsu, Ke-Li Xu, and Yukitoshi Matsushita. Estimation and inference of discontinuity in density. Journal of Business & Economic Statistics, 31(4):507–524, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2013.818007.

Paass:1988:DRD

[Paa88] Gerhard Paass. Disclosure risk and disclosure avoidance for microdata. Journal of Business & Economic Statistics, 6 (4):487–500, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509697.

Parzen:1985:BR

[PAC85] Emanuel Parzen, Alan Agresti, and Samprit Chatterjee. Book reviews. Journal of Business & Economic Statistics, 3(3): 289–291, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509460.

Pakos:2011:EII

[Pak11] Michal Pakos. Estimating intertemporal and intratemporal substitutions when both income and substitution effects are present: The role of durable goods. Journal of Business & Economic Statistics, 29(3):439–454, 2011. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.07046.

Pantula:1991:ADU

[Pan91] Sastry G. Pantula. Asymptotic distributions of unit-root tests when the process is nearly stationary. Journal of Business & Economic Statistics, 9(1):63–71, 1991. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1991.10509827.

Parks:1983:D

[Par83] Richard W. Parks. Discussion. Journal of Business & Economic Statistics, 1(3):221–225, 1983. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1983.10509344.

Patterson:1986:SAW

[Pat86] Douglas M. Patterson. The speed of adjustment of warrant prices to changes in stock prices. Journal of Business & Economic Statistics, 4(2):233–241, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509518.

Patton:2015:C

[Pat15] Andrew J. Patton. Comment. Journal of Business & Economic Statistics, 33(1):22–24, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.977445. See [AGO^{+}14a, AGO^{+}14b].

Patton:2019:CPM

[Pat19] Andrew J. Patton. Comparing possibly misspecified forecasts. Journal of Business & Economic Statistics, 38(4):796–809, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1585256.

Pfeffermann:1991:SNE

[PB91] Danny Pfeffermann and Charles H. Barnard. Some new estimators for small-area means with application to the assessment of farmland values. Journal of Business & Economic Statistics, 9(1):73–84, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509828.

Phipps:1995:RES

[PBC95] Polly A. Phipps, Shall J. Butani, and Young I. Chun. Research on establishment-survey questionnaire design. Journal of Business & Economic Statistics, 13(3):337–346, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524607.

Peters:2006:TCS

[PdV06] Remco T. Peters and Robin G. de Vilder. Testing the continuous semimartingale hypothesis for the S&P 500. Journal of Business & Economic Statistics, 24(4):444–454, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000341.

Pelger:2021:DTS

[Pel21] Markus Pelger. Discussion of “Text Selection” by Bryan Kelly, Asaf Manela, and Alan Moreira. Journal of Business & Economic Statistics, 39(4):880–882, 2021. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2021.1948420. See [KMM21].

Pena:1990:IOT

[Peñ90] Daniel Peña. Influential observations in time series. Journal of Business & Economic Statistics, 8(2):235–241, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1990.10509795.

Perron:1990:TUR

[Per90] Pierre Perron. Testing for a unit root in a time series with a changing mean. Journal of Business & Economic Statistics, 8(2):153–162, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509786.

Peters:1984:SNB

[PF84] S. C. Peters and D. A. Freedman. Some notes on the bootstrap in regression problems. Journal of Business & Economic Statistics, 2(4):406–409, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509416.

Pereda-Fernandez:2019:CBR

[PF19] Santiago Pereda-Fernández. Copula-based random effects models for clustered data. Journal of Business & Economic Statistics, 39(2):575–588, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1688665.

Pammer:2000:FPN

[PFA00] Scott E. Pammer, Duncan K. H. Fong, and Steven F. Arnold. Forecasting the penetration of a new product — a Bayesian approach. Journal of Business & Economic Statistics, 18 (4):428–435, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524882.

Pfeffermann:1991:ESA

[Pfe91a] Danny Pfeffermann. Estimation and seasonal adjustment of population means using data from repeated surveys. Journal of Business & Economic Statistics, 9(2):163–175, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509840.

Pfeffermann:1991:R

[Pfe91b] Danny Pfeffermann. Reply. Journal of Business & Economic Statistics, 9(2):177, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509842.

Pfeffermann:1998:EAS

[PFS98] Danny Pfeffermann, Moshe Feder, and David Signorelli. Estimation of autocorrelations of survey errors with application to trend estimation in small areas. Journal of Business & Economic Statistics, 16(3):339–348, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524773.

Pace:1993:TPI

[PG93] R. Kelley Pace and Otisy W. Gilley. Translating prior information across specifications to improve predictive accuracy. Journal of Business & Economic Statistics, 11(3): 301–309, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509957.

Philipov:2006:MSV

[PG06] Alexander Philipov and Mark E. Glickman. Multivariate stochastic volatility via Wishart processes. Journal of Business & Economic Statistics, 24(3):313–328, 2006. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500105000000306.

Pierce:1984:SAW

[PGC84] David A. Pierce, Michael R. Grupe, and William P. Cleveland. Seasonal adjustment of the weekly monetary aggregates: a modelbased approach. Journal of Business & Economic Statistics, 2(3):260–270, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509393.

Pantula:1994:CUR

[PGFF94] Sastry G. Pantula, Graciela Gonzalez-Farias, and Wayne A. Fuller. A comparison of unit-root test criteria. Journal of Business & Economic Statistics, 12(4):449–459, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10524567.

Phillips:1987:CFI

[Phi87] Robert F. Phillips. Composite forecasting: an integrated approach and optimality reconsidered. Journal of Business & Economic Statistics, 5(3):389–395, 1987. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1987.10509603.

Phillips:2010:IFG

[Phi10] Robert F. Phillips. Iterated feasible generalized least-squares estimation of augmented dynamic panel data models. Journal of Business & Economic Statistics, 28(3):410–422, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2009.08106.

Pei:2020:NBC

[PHPY20] Youquan Pei, Tao Huang, Heng Peng, and Jinhong You. Networkbased clustering for varying coefficient panel data models. Journal of Business & Economic Statistics, 40(2):578–594, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1841648.

Picci:2001:ELS

[Pic01] Lucio Picci. Explaining long- and short-run interactions in time series data. Journal of Business & Economic Statistics, 19(1):85–94, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
07350010152472643.

Pierce:1983:SAM

[Pie83] David A. Pierce. Seasonal adjustment of the monetary aggregates: Summary of the Federal Reserve’s Committee Report. Journal of Business & Economic Statistics, 1(1):37–42, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1983.10509322.

Pierce:1984:CSA

[Pie84] David A. Pierce. Comment: Seasonal adjustment-models, assumptions, and criteria. Journal of Business & Economic Statistics, 2(4):340–342, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509406.

Pischke:1995:MEE

[Pis95] Jörn-Steffen Pischke. Measurement error and earnings dynamics: Some estimates from the PSID validation study. Journal of Business & Economic Statistics, 13(3):305–314, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524604.

Phillips:2014:TMH

[PJ14] Peter C. B. Phillips and Sainan Jin. Testing the martingale hypothesis. Journal of Business & Economic Statistics, 32(4):537–554, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.908780.

Pesaran:2020:DQS

[PJ20] M. Hashem Pesaran and Ida Johnsson. Double-question survey measures for the analysis of financial bubbles and crashes. Journal of Business & Economic Statistics, 38(2):428–442, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1513845.

Pei:2022:LPO

[PLCW22] Zhuan Pei, David S. Lee, David Card, and Andrea Weber. Local polynomial order in regression discontinuity designs. Journal of Business & Economic Statistics, 40(3):1259– 1267, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1920961.

Potiron:2020:LPE

[PM20] Yoann Potiron and Per Mykland. Local parametric estimation in high frequency data. Journal of Business & Economic Statistics, 38(3):679–692, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2019.1566731.

Plagborg-Moller:2022:DNR

[PM22] Mikkel Plagborg-Møller. Discussion of “Narrative Restrictions and Proxies” by Raffaella Giacomini, Toru Kitagawa, and Matthew Read. Journal of Business & Economic Statistics, 40 (4):1434–1437, 2022. CODEN ???? ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2022.2096042. See [GKR22a,

GKR22b].

Poirier:1989:RB

[Poi89] Dale J. Poirier. A report from the battlefront. Journal of Business & Economic Statistics, 7(1):137–139, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509721.

Poirier:1991:BVN

[Poi91a] Dale J. Poirier. A Bayesian view of nominal money and real output through a new classical macroeconomic window. Journal of Business & Economic Statistics, 9(2):125–148, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509835.

Poirier:1991:RMW

[Poi91b] Dale J. Poirier. Reply: Is my window broken? Journal of Business & Economic Statistics, 9(2):159–161, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509839.

Poirier:1992:RB

[Poi92] Dale J. Poirier. A return to the battlefront. Journal of Business & Economic Statistics, 10(4):473–474, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1992.10509926.

Popkin:1984:BCV

[Pop84] Joel Popkin. The business cycle at various stages of process. Journal of Business & Economic Statistics, 2(3): 215–223, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509389.

Porter:2002:ECM

[Por02] Jack Porter. Effciency of covariance matrix estimators for maximum likelihood estimation. Journal of Business & Economic Statistics, 20(3):431–440, 2002. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102288618559.

Porter:2018:CSE

[Por18] Jack Porter. Comment on “Simple Estimators for Invertible Index Models”. Journal of Business & Economic Statistics, 36(1):16–17, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1396989. See [AIPR18b, AIPR18a].

Poskitt:2000:SCD

[Pos00] D. S. Poskitt. Strongly consistent determination of cointegrating rank via canonical correlations. Journal of Business & Economic Statistics, 18(1):77–90, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524849.

Powell:2022:SCE

[Pow22] David Powell. Synthetic control estimation beyond comparative case studies: Does the minimum wage reduce employment? Journal of Business & Economic Statistics, 40(3):1302– 1314, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1927743.

Pesaran:2011:FCA

[PP11] M. Hashem Pesaran and Andreas Pick. Forecast combination across estimation Windows. Journal of Business & Economic Statistics, 29(2):307–318, 2011. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2010.09018.

Pesaran:1994:CBD

[PPL94] M. Hashem Pesaran, Richard G. Pierse, and Kevin C. Lee. Choice between disaggregate and aggregate specifications estimated by instrumental variables methods. Journal of Business & Economic Statistics, 12(1):11–21, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10509987.

Pastorello:2003:IRE

[PPR03] Sergio Pastorello, Valentin Patilea, and Eric Renault. Iterative and recursive estimation in structural nonadaptive models. Journal of Business & Economic Statistics, 21(4):449–509, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500103288619124.

Pei:2019:PMC

[PPS19] Zhuan Pei, Jörn-Steffen Pischke, and Hannes Schwandt. Poorly measured confounders are more useful on the left than on the right. Journal of Business & Economic Statistics, 37(2):205–216, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.

1462710. See comments [DMP19].

Penm:1992:UBA

[PPT92] Jack H. W. Penm, Jammie H. Penm, and R. D. Terrell. Using the bootstrap as an aid in choosing the approximate representation for vector time series. Journal of Business & Economic Statistics, 10(2):213–219, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509900.

Perron:2010:LML

[PQ10] Pierre Perron and Zhongjun Qu. Long-memory and level shifts in the volatility of stock market return indices. Journal of Business & Economic Statistics, 28(2):275–290, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.06171.

Peck:1987:VIA

[PR87] Stephen C. Peck and Richard G. Richels. The value of information to the acidic deposition debates. Journal of Business & Economic Statistics, 5(2):205–217, 1987. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1987.10509579.

Partridge:1998:GBV

[PR98] Mark D. Partridge and Dan S. Rickman. Generalizing the Bayesian vector autoregression approach for regional interindustry employment forecasting. Journal of Business & Economic Statistics, 16(1):62–72, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524735.

Pena:1984:RMB

[PRC84] Daniel Peña and Javier Ruiz-Castillo. Robust methods of building regression models — an application to the housing sector. Journal of Business & Economic Statistics, 2 (1):10–20, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509366.

Pena:1998:EFE

[PRC98] Daniel Peña and Javier Ruiz-Castillo. The estimation of food expenditures from household budget data in the presence of bulk purchases. Journal of Business & Economic Statistics, 16(3):292–303, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524768.

Press:1995:SAT

[Pre95] S. James Press. Sample-audit tax assessment for businesses: What’s fair? Journal of Business & Economic Statistics, 13 (3):357–359, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524609.

Planas:2008:BAO

[PRF08] Christophe Planas, Alessandro Rossi, and Gabriele Fiorentini. Bayesian analysis of the output gap. Journal of Business & Economic Statistics, 26(1):18–32, 2008. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/073500106000000576.

Pan:2022:NDQ

[PRG^{+}22] Rui Pan, Tunan Ren, Baishan Guo, Feng Li, Guodong Li, and Hansheng Wang. A note on distributed quantile regression by pilot sampling and one-step updating. Journal of Business & Economic Statistics, 40(4):1691–1700, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1961789.

Pesaran:1999:AER

[PRM99] M. Hashem Pesaran and Francisco J. Ruge-Murcia. Analysis of exchange-rate target zones using a limited-dependent rationalexpectations model with jumps. Journal of Business & Economic Statistics, 17(1):50–66, 1999. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1999.10524796.

Prodan:2008:PPD

[Pro08] Ruxandra Prodan. Potential pitfalls in determining multiple structural changes with an application to purchasing power parity. Journal of Business & Economic Statistics, 26 (1):50–65, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500107000000304.

Prowse:2012:MED

[Pro12] Victoria Prowse. Modeling employment dynamics with state dependence and unobserved heterogeneity. Journal of Business & Economic Statistics, 30(3):411–431, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2012.697851.

Penaranda:2022:NST

[PRPS22] Francisco Peñaranda, Juan M. Rodríguez-Poo, and Stefan Sperlich. Nonparametric specification testing of conditional asset pricing models. Journal of Business & Economic Statistics, 40 (4):1455–1469, 2022. CODEN ???? ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2021.1933500.

Pastorello:2000:SIR

[PRT00] Sergio Pastorello, Eric Renault, and Nizar Touzi. Statistical inference for random-variance option pricing. Journal of Business & Economic Statistics, 18(3):358–367, 2000. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.2000.10524876.

Palumbo:2006:RBR

[PRW06] Michael Palumbo, Jeremy Rudd, and Karl Whelan. On the relationships between real consumption, income, and wealth. Journal of Business & Economic Statistics, 24(1):1–11, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500105000000225.

Perraudin:1992:CCC

[PS92] William R. M. Perraudin and Bent E. Sorensen. The creditconstrained consumer: an empirical study of demand and supply in the loan market. Journal of Business & Economic Statistics, 10(2):179–192, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509897.

Petersen:1996:WSI

[PS96] Bruce Petersen and Steven Strongin. Why are some industries more cyclical than others? Journal of Business & Economic Statistics, 14(2):189–198, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524645.

Parker:2016:IUC

[PS16] Jason Parker and Donggyu Sul. Identification of unknown common factors: Leaders and followers. Journal of Business & Economic Statistics, 34(2):227–239, 2016. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2015.1026439.

Psaradakis:2016:UBT

[Psa16] Zacharias Psaradakis. Using the bootstrap to test for symmetry under unknown dependence. Journal of Business & Economic Statistics, 34(3):406–415, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1043368.

Panagiotelis:2014:AAM

[PSD14] Anastasios Panagiotelis, Michael S. Smith, and Peter J. Danaher. From Amazon to Apple: Modeling online retail sales, purchase incidence, and visit behavior. Journal of Business & Economic Statistics, 32(1):14–29, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.835729.

Pakel:2020:FVD

[PSSE20] Cavit Pakel, Neil Shephard, Kevin Sheppard, and Robert F. Engle. Fitting vast dimensional time-varying covariance models. Journal of Business & Economic Statistics, 39(3):652–668, 2020.

ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1713795.

Paap:2009:DLI

[PSvD09] Richard Paap, Rene Segers, and Dick van Dijk. Do leading indicators lead peaks more than troughs? Journal of Business & Economic Statistics, 27(4):528–543, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.07061.

Pesaran:2004:MRI

[PSW04a] M. Hashem Pesaran, Til Schuermann, and Scott M. Weiner. Modeling regional interdependencies using a global error-correcting macroeconometric model. Journal of Business & Economic Statistics, 22(2):129–162, 2004. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.

1198/073500104000000019. See comments [Bal04, DL04, Joh04, Wal04] and rejoinder [PSW04b].

Pesaran:2004:R

[PSW04b] M. Hashem Pesaran, Til Schuermann, and Scott M. Weiner. Rejoinder. Journal of Business & Economic Statistics, 22 (2):175–181, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/

073500104000000064. See [PSW04a, Bal04, DL04, Joh04, Wal04].

Pesaran:1992:SNT

[PT92] M. Hashem Pesaran and Allan Timmermann. A simple nonparametric test of predictive performance. Journal of Business & Economic Statistics, 10(4):461–465, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1992.10509922.

Polson:2000:BPS

[PT00] Nicholas G. Polson and Bernard V. Tew. Bayesian portfolio selection: an empirical analysis of the S&P 500 Index 1970–1996. Journal of Business & Economic Statistics, 18 (2):164–173, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524859.

Poirier:2003:PDO

[PT03] Dale J. Poirier and Justin L. Tobias. On the predictive distributions of outcome gains in the presence of an unidentified parameter. Journal of Business & Economic Statistics, 21 (2):258–268, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288618945.

Patton:2011:POG

[PT11] Andrew J. Patton and Allan Timmermann. Predictability of output growth and inflation: a multi-horizon survey approach. Journal of Business & Economic Statistics, 29(3):397–410, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2010.08347.

Patton:2012:FRT

[PT12a] Andrew J. Patton and Allan Timmermann. Forecast rationality tests based on multi-horizon bounds. Journal of Business & Economic Statistics, 30(1):1–17, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2012.634337. See comments

[Cro12, Lah12, Ros12, HRvD12, Wes12] and rejoinder [PT12b].

Patton:2012:R

[PT12b] Andrew J. Patton and Allan Timmermann. Rejoinder. Journal of Business & Economic Statistics, 30(1):36–40, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2012.634354.

Pettenuzzo:2017:FMV

[PT17] Davide Pettenuzzo and Allan Timmermann. Forecasting macroeconomic variables under model instability. Journal of Business & Economic Statistics, 35(2):183–201, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1051183.

Pustejovsky:2018:SSM

[PT18] James E. Pustejovsky and Elizabeth Tipton. Small-sample methods for cluster-robust variance estimation and hypothesis testing in fixed effects models. Journal of Business & Economic Statistics, 36(4):672–683, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1247004. See corrigendum [PT23].

Pustejovsky:2023:CSS

[PT23] James E. Pustejovsky and Elizabeth Tipton. Corrigendum: Small sample methods for cluster-robust variance estimation and hypothesis testing in fixed effects models. Journal of Business & Economic Statistics, 41(2):650–652, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2023.

2174123. See [PT18].

Poirier:1986:DTN

[PTZ86] Dale J. Poirier, Mario D. Tello, and Stanley E. Zin. A diagnostic test for normality within the power exponential family. Journal of Business & Economic Statistics, 4(3): 359–373, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509532.

Perron:1992:NLS

[PV92a] Pierre Perron and Timothy J. Vogelsang. Nonstationarity and level shifts with an application to purchasing power parity. Journal of Business & Economic Statistics, 10(3): 301–320, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509907.

Perron:1992:TUR

[PV92b] Pierre Perron and Timothy J. Vogelsang. Testing for a unit root in a time series with a changing mean: Corrections and extensions. Journal of Business & Economic Statistics, 10 (4):467–470, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509923.

Paap:2003:BEM

[PV03] Richard Paap and Herman K. Van Dijk. Bayes estimates of Markov trends in possibly cointegrated series. Journal of Business & Economic Statistics, 21(4):547–563, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500103288619296.

Psaradakis:2022:UTA

[PV22] Zacharias Psaradakis and Marián Vávra. Using triples to assess symmetry under weak dependence. Journal of Business & Economic Statistics, 40(4):1538–1551, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1939037.

Parker:2006:SCC

[PvP06] Simon C. Parker and C. Mirjam van Praag. Schooling, capital constraints, and entrepreneurial performance. Journal of Business & Economic Statistics, 24(4):416–431, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000215.

Pendakur:2010:GCG

[PW10] Krishna Pendakur and Simon Woodcock. Glass ceilings or glass doors? Wage disparity within and between firms. Journal of Business & Economic Statistics, 28(1):181–189, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.08124.

Patton:2022:TUH

[PW22] Andrew J. Patton and Brian M. Weller. Testing for unobserved heterogeneity via k-means clustering. Journal of Business & Economic Statistics, 41(3):737–751, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2061983.

Pan:2021:PPD

[PX21] Zhewen Pan and Jianhui Xie. ℓ_{1}-Penalized pairwise difference estimation for a high-dimensional censored regression model. Journal of Business & Economic Statistics, 41(2):283– 297, 2021. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.2013243.

Pelger:2022:ISP

[PX22a] Markus Pelger and Ruoxuan Xiong. Interpretable sparse proximate factors for large dimensions. Journal of Business & Economic Statistics, 40(4):1642–1664, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1961786.

Pelger:2022:SVF

[PX22b] Markus Pelger and Ruoxuan Xiong. State-varying factor models of large dimensions. Journal of Business & Economic Statistics, 40(3):1315–1333, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1927744.

Phillips:2006:C

[PY06] Peter C. B. Phillips and Jun Yu. Comment. Journal of Business & Economic Statistics, 24(2):202–208, 2006. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500106000000116. See [HL06a, HL06b].

Perron:2009:TST

[PY09] Pierre Perron and Tomoyoshi Yabu. Testing for shifts in trend with an integrated or stationary noise component. Journal of Business & Economic Statistics, 27(3):369–396, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2009.07268.

Poskitt:2017:VAM

[PY17] D. S. Poskitt and Wenying Yao. Vector autoregressions and macroeconomic modeling: an error taxonomy. Journal of Business & Economic Statistics, 35(3):407–419, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1077139.

Perron:2019:TCF

[PY19] Pierre Perron and Yohei Yamamoto. Testing for changes in forecasting performance. Journal of Business & Economic Statistics, 39(1):148–165, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1641410.

Patel:1990:SPT

[PZ90] Jayendu Patel and Richard Zeckhauser. Shared price trends: Evidence from U.S. cities and OECD countries. Journal of Business & Economic Statistics, 8(2):179–189, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1990.10509789.

Poirier:2019:EMM

[PZ19] Alexandre Poirier and Nicolas L. Ziebarth. Estimation of models with multiple-valued explanatory variables. Journal of Business & Economic Statistics, 37(4):586–597, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1391694.

Pan:2020:SEC

[PZZ20] Zhewen Pan, Xianbo Zhou, and Yahong Zhou. Semiparametric estimation of a censored regression model subject to nonparametric sample selection. Journal of Business & Economic Statistics, 40(1):141–151, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1784746.

Qi:1999:NPS

[Qi99] Min Qi. Nonlinear predictability of stock returns using financial and economic variables. Journal of Business & Economic Statistics, 17(4):419–429, 1999. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1999.10524830.

Qian:2019:ICS

[Qia19] Hang Qian. Inequality constrained state-space models. Journal of Business & Economic Statistics, 37(2):350–362, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1340300.

Qian:2022:BIC

[Qia22] Hang Qian. Bayesian inference in common microeconometric models with massive datasets by double marginalized subsampling. Journal of Business & Economic Statistics, 40(4):1484– 1497, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1933502.

Qu:2011:TAS

[Qu11] Zhongjun Qu. A test against spurious long memory. Journal of Business & Economic Statistics, 29(3):423–438, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2010.09153.

Quan:1988:PSS

[Qua88] Nguyen T. Quan. The prediction sum of squares as a general measure for regression diagnostics. Journal of Business & Economic Statistics, 6(4):501–504, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509698.

Quah:1993:C

[Qua93] Danny Quah. Comment. Journal of Business & Economic Statistics, 11(4):387–390, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509970.

Quaedvlieg:2019:MHF

[Qua19] Rogier Quaedvlieg. Multi-horizon forecast comparison. Journal of Business & Economic Statistics, 39(1):40–53, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1620074.

Quintos:1995:SDP

[Qui95] Carmela E. Quintos. Sustainability of the deficit process with structural shifts. Journal of Business & Economic Statistics, 13(4):409–417, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524615.

Quast:2020:RRT

[QW20] Josefine Quast and Maik H. Wolters. Reliable real-time output gap estimates based on a modified Hamilton filter. Journal of Business & Economic Statistics, 40(1):152–168, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1784747.

Qu:2019:UIQ

[QY19] Zhongjun Qu and Jungmo Yoon. Uniform inference on quantile effects under sharp regression discontinuity designs. Journal of Business & Economic Statistics, 37(4):625–647, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1407323.

Qin:2015:EAI

[QZL15] Jing Qin, Biao Zhang, and Denis H. Y. Leung. Efficient augmented inverse probability weighted estimation in missing data problems. Journal of Business & Economic Statistics, 35(1):86–97, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1058266.

Racine:1997:CST

[Rac97] Jeff Racine. Consistent significance testing for nonparametric regression. Journal of Business & Economic Statistics, 15 (3):369–378, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524714.

Racine:2001:NPS

[Rac01] Jeffrey Racine. On the nonlinear predictability of stock returns using financial and economic variables. Journal of Business & Economic Statistics, 19(3):380–382, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500101681019927.

Radner:1983:AES

[Rad83] Daniel B. Radner. Adjusted estimates of the size distribution of family money income. Journal of Business & Economic Statistics, 1(2):136–146, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509332.

Rosenberg:1999:HBM

[RAL99] Marjorie A. Rosenberg, Richard W. Andrews, and Peter J. Lenk. A hierarchical Bayesian model for predicting the rate of nonacceptable in-patient hospital utilization. Journal of Business & Economic Statistics, 17(1):1–8, 1999. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1999.10524792.

Ramajo:1994:CRF

[Ram94] Julián Ramajo. Curvature restrictions on flexible functional forms: an application of the minflex Laurent almost ideal demand system to the pattern of Spanish demand, 1954– 1987. Journal of Business & Economic Statistics, 12(4):

431–436, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10524565.

Randolph:1988:HDA

[Ran88] William C. Randolph. Housing depreciation and aging bias in the consumer price index. Journal of Business & Economic Statistics, 6(3):359–371, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509673.

Rathelot:2012:MSW

[Rat12] Roland Rathelot. Measuring segregation when units are small: a parametric approach. Journal of Business & Economic Statistics, 30(4):546–553, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.707586.

Raviv:2006:NEP

[Rav06] Yaron Raviv. New evidence on price anomalies in sequential auctions. Journal of Business & Economic Statistics, 24 (3):301–312, 2006. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000026.

Rayner:1990:BVP

[Ray90] Robert K. Rayner. Bootstrapping p values and power in the first-order autoregression: a Monte Carlo investigation. Journal of Business & Economic Statistics, 8(2):251–263, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1990.10509797.

Reichert:1983:ECI

[RCW83] Alan K. Reichert, Chien-Ching Cho, and George M. Wagner. An examination of the conceptual issues involved in developing credit-scoring models. Journal of Business & Economic Statistics, 1(2):101–114, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509329.

Rotemberg:1995:MOP

[RDP95] Julio J. Rotemberg, John C. Driscoll, and James M. Poterba. Money, output, and prices: Evidence from a new monetary aggregate. Journal of Business & Economic Statistics, 13 (1):67–83, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524580.

Russell:2005:DSC

[RE05] Jeffrey R. Russell and Robert F. Engle. A discrete-state continuous-time model of financial transactions prices and times. Journal of Business & Economic Statistics, 23(2): 166–180, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000541.

Rangel:2012:FSG

[RE12] José Gonzalo Rangel and Robert F. Engle. The factor-spline-GARCH model for high and low frequency correlations. Journal of Business & Economic Statistics, 30(1):109–124, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2012.643132.

Reinsdorf:1999:USD

[Rei99] Marshall B. Reinsdorf. Using scanner data to construct CP1 basic component indexes. Journal of Business & Economic Statistics, 17(2):152–160, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524805.

Rendon:2007:DWE

[Ren07] Sílvio Rendon. Does wealth explain black-white differences in early employment careers? Journal of Business & Economic Statistics, 25(4):484–500, 2007. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000124.

Roy:2001:EAT

[RF01] Anindya Roy and Wayne A. Fuller. Estimation for autoregressive time series with a root near 1. Journal of Business & Economic Statistics, 19(4):482–493, 2001. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/07350010152596736.

Robbins:2015:CCM

[RF15] Michael W. Robbins and Thomas J. Fisher. Cross-correlation matrices for tests of independence and causality between two multivariate time series. Journal of Business & Economic Statistics, 33(4):459–473, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.962699.

Ruseckaite:2020:FMA

[RFG20] Aiste Ruseckaite, Dennis Fok, and Peter Goos. Flexible mixtureamount models using multivariate Gaussian processes. Journal of Business & Economic Statistics, 38(2):257–271, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1497506.

Roosen:2004:TML

[RH04] Jutta Roosen and David A. Hennessy. Testing for the monotone likelihood ratio assumption. Journal of Business & Economic Statistics, 22(3):358–366, 2004. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000235.

Rossi:2012:SFT

[RI12] Barbara Rossi and Atsushi Inoue. Out-of-sample forecast tests robust to the choice of window size. Journal of Business & Economic Statistics, 30(3):432–453, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.693850.

Richardson:1993:TCS

[Ric93] Matthew Richardson. Temporary components of stock prices: a Skeptic’s view. Journal of Business & Economic Statistics, 11(2):199–207, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509948.

Rilstone:1991:SMC

[Ril91] Paul Rilstone. Some Monte Carlo evidence on the relative efficiency of parametric and semiparametric EGLS estimators. Journal of Business & Economic Statistics, 9(2): 179–187, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509843.

Ritchey:1986:ACS

[Rit86] Robert J. Ritchey. An application of the chi-squared goodnessof-fit test to discrete common stock returns. Journal of Business & Economic Statistics, 4(2):243–254, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509519.

Reh:2022:PGM

[RKL22] Laura Reh, Fabian Krüger, and Roman Liesenfeld. Predicting the global minimum variance portfolio. Journal of Business & Economic Statistics, 41(2):440–452, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2035226.

Ryscavage:1984:C

[RL84] Paul Ryscavage and Enrique Lamas. Comment. Journal of Business & Economic Statistics, 2(4):398–399, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1984.10509414.

Racine:2022:OMA

[RLYZ22] Jeffrey S. Racine, Qi Li, Dalei Yu, and Li Zheng. Optimal model averaging of mixed-data kernel-weighted spline regressions. Journal of Business & Economic Statistics, 41(4):1251– 1261, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2118126.

Rose:1988:STS

[RMS88] Elizabeth L. Rose, Joseph A. Machak, and W. Allen Spivey. A survey of the teaching of statistics in M.B.A. programs. Journal of Business & Economic Statistics, 6(2):273–282, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1988.10509661.

Robinson:1998:C

[Rob98] P. M. Robinson. Comment. Journal of Business & Economic Statistics, 16(3):276–279, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524765.

Rodgers:1984:ESM

[Rod84] Willard L. Rodgers. An evaluation of statistical matching. Journal of Business & Economic Statistics, 2(1):91–102, 1984. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1984.10509373.

Rodrigues:2016:MEP

[Rod16] João D. F. Rodrigues. Maximum-entropy prior uncertainty and correlation of statistical economic data. Journal of Business & Economic Statistics, 34(3):357–367, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1038545.

Rogers:1987:UTP

[Rog87] Robert P. Rogers. Unobservable transactions price and the measurement of a supply and demand model for the American steel industry. Journal of Business & Economic Statistics, 5(3):407–415, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509605.

Rosen:1984:C

[Ros84] Sherwin Rosen. Comment. Journal of Business & Economic Statistics, 2(4):395–397, 1984. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1984.10509413.

Rossi:1986:C

[Ros86] Peter E. Rossi. Comment. Journal of Business & Economic Statistics, 4(4):421–422, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509540.

Rossana:1998:SIP

[Ros98] Robert J. Rossana. Structural instability and the productionsmoothing model of inventories. Journal of Business & Economic Statistics, 16(2):206–215, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524754.

Rossi:2005:CIH

[Ros05] Barbara Rossi. Confidence intervals for half-life deviations from purchasing power parity. Journal of Business & Economic Statistics, 23(4):432–442, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500105000000027.

Roskelley:2008:CRR

[Ros08] Kenneth D. Roskelley. Cromwell’s rule and the role of the prior in the economic metric. Journal of Business & Economic Statistics, 26(2):227–236, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500107000000160.

Rossi:2012:C

[Ros12] Barbara Rossi. Comment. Journal of Business & Economic Statistics, 30(1):25–29, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.634343. See [PT12a, PT12b].

Rossi:2014:C

[Ros14] Barbara Rossi. Comment. Journal of Business & Economic Statistics, 32(4):510–514, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.956875. See [AGO^{+}14a, AGO^{+}14b].

Rothe:2015:DCE

[Rot15] Christoph Rothe. Decomposing the composition effect: The role of covariates in determining between-group differences in economic outcomes. Journal of Business & Economic Statistics, 33(3):323–337, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.948959.

Rouse:1995:DDE

[Rou95] Cecilia Elena Rouse. Democratization or diversion? The effect of community colleges on educational attainment. Journal of Business & Economic Statistics, 13(2):217–224, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524596.

Ramsey:1994:CNM

[RR94] James B. Ramsey and Philip Rothman. Comment on “nonlinear monetary dynamics” by DeCoster and Mitchell. Journal of Business & Economic Statistics, 12(1):135–136, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10509997.

Rothwell:1997:OLN

[RR97] Geoffrey Rothwell and John Rust. On the optimal lifetime of nuclear powe plants. Journal of Business & Economic Statistics, 15(2):195–208, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524700.

Rubio-Ramirez:2022:CNR

[RR22] Juan Rubio-Ramírez. Comments on “Narrative Restrictions and Proxies” by Giacomini, Kitagawa, and Read. Journal of Business & Economic Statistics, 40(4):1426–1428, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.

2102021. See [GKR22a, GKR22b].

Rao:1992:CSE

[RS92] D. S. Prasada Rao and E. A. Selvanathan. Computation of standard errors for Geary–Khamis parities and international prices: a stochastic approach. Journal of Business & Economic Statistics, 10(1):109–115, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509891.

Raj:1994:TBA

[RS94a] Baldev Raj and Daniel J. Slottje. The trend behavior of alternative income inequality measures in the United States from 1947– 1990 and the structural break. Journal of Business & Economic Statistics, 12(4):479–487, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10524570.

Ryu:1994:CSV

[RS94b] Hang K. Ryu and Daniel J. Slottje. Coordinate space versus index space representations as estimation methods: an application to how macro activity affects the U.S. income distribution. Journal of Business & Economic Statistics, 12 (2):243–251, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10510011.

Rossana:1995:TAE

[RS95] Robert J. Rossana and John J. Seater. Temporal aggregation and economic time series. Journal of Business & Economic Statistics, 13(4):441–451, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524618.

Rungsuriyawiboon:2007:DEE

[RS07] Supawat Rungsuriyawiboon and Spiro E. Stefanou. Dynamic efficiency estimation. Journal of Business & Economic Statistics, 25(2):226–238, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000288.

Rigobon:2009:BCR

[RS09] Roberto Rigobon and Thomas M. Stoker. Bias from censored regressors. Journal of Business & Economic Statistics, 27(3):340–353, 2009. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.06119.

Rho:2015:ITS

[RS15] Yeonwoo Rho and Xiaofeng Shao. Inference for time series regression models with weakly dependent and heteroscedastic errors. Journal of Business & Economic Statistics, 33(3):444–457, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2014.
962698.

Ross:2022:MSI

[RS22] Stephen L. Ross and Zhentao Shi. Measuring social interaction effects when instruments are weak. Journal of Business & Economic Statistics, 40(3):995–1006, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1895811.

Robertson:2018:URI

[RSW18] Donald Robertson, Vasilis Sarafidis, and Joakim Westerlund. Unit root inference in generally trending and cross-correlated fixed-T panels. Journal of Business & Economic Statistics, 36 (3):493–504, 2018. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1191501.

Ray:2000:LRD

[RT00] Bonnie K. Ray and Ruey S. Tsay. Long-range dependence in daily stock volatilities. Journal of Business & Economic Statistics, 18(2):254–262, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524867.

Robertson:2001:IFF

[RT01] John C. Robertson and Ellis W. Tallman. Improving federal-funds rate forecasts in VAR models used for policy analysis. Journal of Business & Economic Statistics, 19(3):324–330, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500101681019972.

Rinnergschwentner:2012:MSV

[RTW12] Wolfgang Rinnergschwentner, Gottfried Tappeiner, and Janette Walde. Multivariate stochastic volatility via Wishart processes: a comment. Journal of Business & Economic Statistics, 30(1):164, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.634358.

Rockinger:2001:TVP

[RU01] Michael Rockinger and Giovanni Urga. A time varying parameter model to test for predictability and integration in the stock markets of transition economies. Journal of Business & Economic Statistics, 19(1):73–84, 2001. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/07350010152472634.

Rubin:1986:SMU

[Rub86] Donald B. Rubin. Statistical matching using file concatenation with adjusted weights and multiple imputations. Journal of Business & Economic Statistics, 4(1):87–94, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509497.

Runkle:1987:R

[Run87a] David E. Runkle. Reply. Journal of Business & Economic Statistics, 5(4):454, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509613.

Runkle:1987:VAR

[Run87b] David E. Runkle. Vector autoregressions and reality. Journal of Business & Economic Statistics, 5(4):437–442, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1987.10509609.

Runkle:2002:VAR

[Run02] David E. Runkle. Vector autoregressions and reality. Journal of Business & Economic Statistics, 20(1):128–133, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500102753410435.

Russell:2019:SBF

[Rus19] Thomas M. Russell. Sharp bounds on functionals of the joint distribution in the analysis of treatment effects. Journal of Business & Economic Statistics, 39(2):532–546, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1684300.

Racine:2019:SNM

[RV19] Jeffrey S. Racine and Ingrid Van Keilegom. A smooth nonparametric, multivariate, mixed-data location-scale test. Journal of Business & Economic Statistics, 38(4):784–795, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1574227.

Rhodes:1986:EAC

[RW86a] George F. Rhodes and M. Daniel Westbrook. Econometric analysis of costing system components in rail rate regulation. Journal of Business & Economic Statistics, 4(3): 289–303, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509526.

Robins:1986:SAL

[RW86b] Philip K. Robins and Richard W. West. Sample attrition and labor supply response in experimental panel data. Journal of Business & Economic Statistics, 4(3):329–338, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509529.

Roley:1996:SIR

[RW96] V. Vance Roley and Simon M. Wheatley. Shifts in the interestrate response to money announcements: What can we say about when they occur? Journal of Business & Economic Statistics, 14(1):135–138, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524639.

Ryan:1998:SMI

[RW98] David L. Ryan and Terence J. Wales. A simple method for imposing local curvature in some flexible consumer-demand systems. Journal of Business & Economic Statistics, 16(3):

331–338, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524772.

Rudebusch:2009:FRP

[RW09] Glenn D. Rudebusch and John C. Williams. Forecasting recessions: The puzzle of the enduring power of the yield curve. Journal of Business & Economic Statistics, 27(4):492–503, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1198/jbes.2009.07213.

Ruf:2022:HLR

[RW22] Johannes Ruf and Weiguan Wang. Hedging with linear regressions and neural networks. Journal of Business & Economic Statistics, 40(4):1442–1454, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1931241.

Sabavala:1988:C

[Sab88] Darius J. Sabavala. Comment. Journal of Business & Economic Statistics, 6(2):161–162, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509649.

Sacerdote:2013:C

[Sac13] Bruce Sacerdote. Comment. Journal of Business & Economic Statistics, 31(3):275, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2013.792263. See [GPI13b, GPI13a].

Salemi:1995:RPF

[Sal95] Michael K. Salemi. Revealed preference of the Federal Reserve: Using inverse-control theory to interpret the policy equation of a vector autoregression. Journal of Business & Economic Statistics, 13(4):419–433, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524616.

Sande:1988:C

[San88] I. G. Sande. Comment. Journal of Business & Economic Statistics, 6(3):296–297, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509664.

SantAnna:2017:TUR

[San17] Pedro H. C. Sant’Anna. Testing for uncorrelated residuals in dynamic count models with an application to corporate bankruptcy. Journal of Business & Economic Statistics, 35(3):349–358, 2017. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1102732.

SantAnna:2020:NTT

[San20a] Pedro H. C. Sant’Anna. Nonparametric tests for treatment effect heterogeneity with duration outcomes. Journal of Business & Economic Statistics, 39(3):816–832, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1737080.

Santos:2020:IDS

[San20b] Carlos Daniel Santos. Identifying demand shocks from production data. Journal of Business & Economic Statistics, 38(1):93–106, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1458622.

Schmidt:1992:C

[SAW92] Peter Schmidt, Seung C. Ahn, and Donald Wyhowski. Comment. Journal of Business & Economic Statistics, 10(1):10–14, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/07350015.1992.10509878.

Stefanski:1988:BR

[SB88a] L. A. Stefanski and Bruce L. Bowerman. Book reviews. Journal of Business & Economic Statistics, 6(3):399–400, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1988.10509678.

Swamy:1988:C

[SB88b] P. A. V. B. Swamy and J. R. Barth. Comment. Journal of Business & Economic Statistics, 6(4):473–474, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509694.

Steinhorst:1989:AIF

[SB89a] R. Kirk Steinhorst and C. Randall Byers. An alternative interpretation of Freedman’s nonresponse case study. Journal of Business & Economic Statistics, 7(1):27–28, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509707.

Steinhorst:1989:R

[SB89b] R. Kirk Steinhorst and C. Randall Byers. Reply. Journal of Business & Economic Statistics, 7(1):33, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509710.

Sun:2019:INB

[SB19] Zhenting Sun and Brendan K. Beare. Improved nonparametric bootstrap tests of Lorenz dominance. Journal of Business & Economic Statistics, 39(1):189–199, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2019.1647214.

Stuttgen:2018:SRM

[SBK18] Peter Stüttgen, Peter Boatwright, and Joseph B. Kadane. Stockouts and restocking: Monitoring the retailer from the supplier’s perspective. Journal of Business & Economic Statistics, 36(3):471–482, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1200982.

Seetharaman:2003:PHM

[SC03] P. B. Seetharaman and Pradeep K. Chintagunta. The proportional hazard model for purchase timing. Journal of Business & Economic Statistics, 21(3):368–382, 2003. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500103288619025.

Su:2017:TTD

[SC17] Fei Su and Kung-Sik Chan. Testing for threshold diffusion. Journal of Business & Economic Statistics, 35(2):218–227, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1073594.

Schmittlein:1983:SSP

[Sch83] David C. Schmittlein. Some sampling properties of a model for income distribution. Journal of Business & Economic Statistics, 1(2):147–153, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509333.

Schafer:1987:MED

[Sch87] Daniel W. Schafer. Measurement-error diagnostics and the sex discrimination problem. Journal of Business & Economic Statistics, 5(4):529–537, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509621.

Scheuren:1988:C

[Sch88] Fritz Scheuren. Comment. Journal of Business & Economic Statistics, 6(3):298–299, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509665.

Schwert:1989:TUR

[Sch89] G. William Schwert. Tests for unit roots: a Monte Carlo investigation. Journal of Business & Economic Statistics, 7 (2):147–159, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509723.

Schwert:2002:TUR

[Sch02] G. William Schwert. Tests for unit roots. Journal of Business & Economic Statistics, 20(1):5–17, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102753410354.

Scotti:2014:C

[Sco14] Chiara Scotti. Comment. Journal of Business & Economic Statistics, 32(4):504–506, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.956873. See [AGO^{+}14a, AGO^{+}14b].

Shen:2020:MTI

[SCS20] Zhouyu Shen, Yu Chen, and Ruxin Shi. Modeling tail index with autoregressive conditional Pareto model. Journal of Business & Economic Statistics, 40(1):458–466, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1832504.

Sen:1987:STS

[SD87] D. L. Sen and D. A. Dickey. Symmetric test for second differencing in univariate time series. Journal of Business & Economic Statistics, 5(4):463–473, 1987. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1987.10509615.

Seater:1993:WTT

[Sea93] John J. Seater. World temperature-trend uncertainties and their implications for economic policy. Journal of Business & Economic Statistics, 11(3):265–277, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509954.

Selvanathan:1989:NSA

[Sel89] E. A. Selvanathan. A note on the stochastic approach to index numbers. Journal of Business & Economic Statistics, 7 (4):471–474, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509759.

Sen:2003:URT

[Sen03] Amit Sen. On unit-root tests when the alternative is a trend-break stationary process. Journal of Business & Economic Statistics, 21(1):174–184, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618874.

Seo:2020:RTE

[Seo20] Juwon Seo. Randomization tests for equality in dependence structure. Journal of Business & Economic Statistics, 39(4): 1026–1037, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1753527.

Serletis:1988:ERB

[Ser88a] Apostolos Serletis. The empirical relationship between money, prices, and income revisited. Journal of Business & Economic Statistics, 6(3):351–358, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509672.

Serletis:1988:TFF

[Ser88b] Apostolos Serletis. Translog flexible functional forms and substitutability of monetary assets. Journal of Business & Economic Statistics, 6(1):59–67, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509637.

Serletis:1995:RWB

[Ser95] Apostolos Serletis. Random walks, breaking trend functions, and the chaotic structure of the velocity of money. Journal of Business & Economic Statistics, 13(4):453–465, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524619.

Salinas:1987:MPM

[SH87] Teresita S. Salinas and Steven C. Hillmer. MulticolIinearity problems in modeling time series with trading-day variation. Journal of Business & Economic Statistics, 5(3): 431–436, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509608.

Silver:2005:FMI

[SH05] Mick Silver and Saeed Heravi. A failure in the measurement of inflation. Journal of Business & Economic Statistics, 23 (3):269–281, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000343.

Silver:2007:DBH

[SH07] Mick Silver and Saeed Heravi. The difference between hedonic imputation indexes and time dummy hedonic indexes. Journal of Business & Economic Statistics, 25(2):239–246, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000486.

Shea:1992:BEH

[She92] Gary S. Shea. Benchmarking the expectations hypothesis of the interest-rate term structure: an analysis of cointegration vectors. Journal of Business & Economic Statistics, 10 (3):347–366, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509910.

Shea:1993:IOA

[She93] John Shea. The input-output approach to instrument selection. Journal of Business & Economic Statistics, 11(2): 145–155, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509943.

Shen:2019:EID

[She19] Shu Shen. Estimation and inference of distributional partial effects: Theory and application. Journal of Business & Economic Statistics, 37(1):54–66, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1272458.

Shively:2000:SCS

[Shi00] Philip A. Shively. Stationary components in stock prices: an exact pointwise most powerful invariant test. Journal of Business & Economic Statistics, 18(4):489–496, 2000. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.2000.10524887.

Shore:2011:ITI

[Sho11] Stephen H. Shore. The intergenerational transmission of income volatility: Is riskiness inherited? Journal of Business & Economic Statistics, 29(3):372–381, 2011. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2011.08091.

Shrestha:1989:EMI

[Shr89] Keshab Shrestha. Empirical measurement of an inflation index: a multiple-indicators distributed-lag approach. Journal of Business & Economic Statistics, 7(2):219–225, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509730.

Sichel:1994:ITP

[Sic94] Daniel E. Sichel. Inventories and the three phases of the business cycle. Journal of Business & Economic Statistics, 12 (3):269–277, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10524542.

Silver:1986:TRU

[Sil86] J. Lew Silver. Two results useful for implementing Litterman’s procedure for interpolating a time series. Journal of Business & Economic Statistics, 4(1):129–130, 1986. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1986.10509504.

Silva:2001:IDE

[Sil01] J. M. C. Santos Silva. Influence diagnostics and estimation algorithms for Powell’s SCLS. Journal of Business & Economic Statistics, 19(1):55–62, 2001. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/07350010152472616.

Sims:1985:CII

[Sim85] Christopher A. Sims. Comment on “Issues Involved With the Seasonal Adjustment of Economic Time Series” by William R. Bell and Steven C. Hiilmer. Journal of Business & Economic Statistics, 3(1):92–94, 1985. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1985.10509431. See [BH84a].

Sims:1987:C

[Sim87] Christopher A. Sims. Comment. Journal of Business & Economic Statistics, 5(4):443–449, 1987. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1987.10509610.

Sims:1990:SSG

[Sim90] Christopher A. Sims. Solving the stochastic growth model by backsolving with a particular nonlinear form for the decision rule. Journal of Business & Economic Statistics, 8 (1):45–47, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509775.

Sims:2007:C

[Sim07] Christopher A. Sims. Comment. Journal of Business & Economic Statistics, 25(2):152–154, 2007. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000052. See [DSSW07a, DSSW07b].

Sinha:2021:DTS

[Sin21] Nitish Ranjan Sinha. A discussion of “Text Selection”. Journal of Business & Economic Statistics, 39(4):888–891, 2021. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2021.

1961785. See [KMM21].

Sizova:2013:LHR

[Siz13] Natalia Sizova. Long-horizon return regressions with historical volatility and other long-memory variables. Journal of Business & Economic Statistics, 31(4):546–559, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2013.827985.

Sanches:2018:MDE

[SJS18] Fabio Sanches, Daniel Silva Junior, and Sorawoot Srisuma. Minimum distance estimation of search costs using price distribution. Journal of Business & Economic Statistics, 36(4):658–671, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1247003.

Silvapulle:1991:TMA

[SK91] Paramsothy Silvapulle and Maxwell L. King. Testing moving average against autoregressive disturbances in the Li nearregression model. Journal of Business & Economic Statistics, 9(3):329–335, 1991. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1991.10509859.

Shephard:1994:C

[SK94] Neil Shephard and Sangjoon Kim. Comment. Journal of Business & Economic Statistics, 12(4):406–410, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10524561.

Shin:1999:STD

[SK99] Dong Wan Shin and Hyun Jung Kim. Semiparametric tests for double unit roots based on symmetric estimators. Journal of Business & Economic Statistics, 17(1):67–73, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524797.

Smith:2020:BIR

[SK20] Michael Stanley Smith and Nadja Klein. Bayesian inference for regression copulas. Journal of Business & Economic Statistics, 39(3):712–728, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1721295.

So:1999:BUR

[SL99] Mike K. P. So and W. K. Ll. Bayesian unit-root testing in stochastic volatility models. Journal of Business & Economic Statistics, 17(4):491–496, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524838.

Saikkonen:2000:TCR

[SL00] Pentti Saikkonen and Helmut Lütkepohl. Testing for the cointegrating rank of a VAR process with structural shifts. Journal of Business & Economic Statistics, 18(4):451–464, 2000. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.2000.10524884.

Shin:2001:TAP

[SL01] Dong Wan Shin and Oesook Lee. Tests for asymmetry in possibly nonstationary time series data. Journal of Business & Economic Statistics, 19(2):233–244, 2001. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500101316970458.

Sun:2011:DDB

[SL11] Yiguo Sun and Qi Li. Data-driven bandwidth selection for nonstationary semiparametric models. Journal of Business & Economic Statistics, 29(4):541–551, 2011. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2011.09159.

Seng:2020:SEM

[SL20] Loraine Seng and Jialiang Li. Structural equation model averaging: Methodology and application. Journal of Business & Economic Statistics, 40(2):815–828, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1870479.

Slesnick:1998:ODR

[Sle98] Daniel T. Slesnick. Are our data relevant to the theory? The case of aggregate consumption. Journal of Business & Economic Statistics, 16(1):52–61, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524734.

So:1998:SVM

[SLL98] Mike Ec. P. So, K. Lam, and W. K. Li. A stochastic volatility model with Markov switching. Journal of Business & Economic Statistics, 16(2):244–253, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524758.

Slottje:1987:RPC

[Slo87] D. J. Slottje. Relative price changes and inequality in the size distribution of various components of income. Journal of Business & Economic Statistics, 5(1):19–26, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509556.

Swamy:1983:FRZ

[SM83] P. A. V. B. Swamy and J. S. Mehta. Further results on Zellner’s minimum expected loss and full information maximum likelihood estimators for undersized samples. Journal of Business & Economic Statistics, 1(2):154–162, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1983.10509334.

Smith:1983:C

[Smi83] Wray Smith. Comment. Journal of Business & Economic Statistics, 1(1):34–35, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/

07350015.1983.10509320. See [FRS83a, FRS83b, Hog83].

Smith:2000:MST

[Smi00] Michael Smith. Modeling and short-term forecasting of new south Wales electricity system load. Journal of Business & Economic Statistics, 18(4):465–478, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524885.

Smith:2002:MSS

[Smi02] Daniel R. Smith. Markov-switching and stochastic volatility diffusion models of short-term interest rates. Journal of Business & Economic Statistics, 20(2):183–197, 2002. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500102317351949.

Smith:2005:LSI

[Smi05] Aaron Smith. Level shifts and the illusion of long memory in economic time series. Journal of Business & Economic Statistics, 23(3):321–335, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000280.

Smith:2022:SBG

[Smi22] Simon C. Smith. Structural breaks in grouped heterogeneity. Journal of Business & Economic Statistics, 41(3):752– 764, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2063132.

Su:2013:LLG

[SMU13] Liangjun Su, Irina Murtazashvili, and Aman Ullah. Local linear GMM estimation of functional coefficient IV models with an application to estimating the rate of return to schooling. Journal of Business & Economic Statistics, 31(2):184–207, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2012.754314.

Snyder:2001:PIA

[SOK01] Ralph D. Snyder, J. Keith Ord, and Anne B. Koehler. Prediction intervals for ARIMA models. Journal of Business & Economic Statistics, 19(2):217–225, 2001. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500101316970430.

Solon:1986:ERG

[Sol86] Gary Solon. Effects of rotation group bias on estimation of unemployment. Journal of Business & Economic Statistics, 4 (1):105–109, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509499.

Song:2010:QAP

[Son10] Minjae Song. The quality adjusted price index in the pure characteristics demand model. Journal of Business & Economic Statistics, 28(1):190–199, 2010. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.08032.

Song:2012:TPA

[Son12] Kyungchul Song. Testing predictive ability and power robustification. Journal of Business & Economic Statistics, 30(2):288–296, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2012.
663245.

Spivey:1986:C

[Spi86] W. Allen Spivey. Comment. Journal of Business & Economic Statistics, 4(1):19–22, 1986. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1986.10509489.

Spletzer:2000:CEB

[Spl00] James R. Spletzer. The contribution of establishment births and deaths to employment growth. Journal of Business & Economic Statistics, 18(1):113–126, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524852.

Su:2017:STS

[SQ17] Liangjun Su and Xi Qu. Specification test for spatial autoregressive models. Journal of Business & Economic Statistics, 35(4):572–584, 2017. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1102734.

Souza:2018:ETC

[SRdCFB18] Igor Viveiros Melo Souza, Valderio Anselmo Reisen, Glaura da Conceição Franco, and Pascal Bondon. The estimation and testing of the cointegration order based on the frequency domain. Journal of Business & Economic Statistics, 36(4):695–704, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1251442.

Schmidt:1984:PFP

[SS84] Peter Schmidt and Robin C. Sickles. Production frontiers and panel data. Journal of Business & Economic Statistics, 2 (4):367–374, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509410.

Smith:2006:SOF

[SS06] J. Q. Smith and António A. F. Santos. Second-order filter distribution approximations for financial time series with extreme outliers. Journal of Business & Economic Statistics, 24 (3):329–337, 2006. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500105000000199.

Su:2013:NTA

[SS13] Liangjun Su and Martin Spindler. Nonparametric testing for asymmetric information. Journal of Business & Economic Statistics, 31(2):208–225, 2013. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.755127.

Schorfheide:2015:RTF

[SS15] Frank Schorfheide and Dongho Song. Real-time forecasting with a mixed-frequency VAR. Journal of Business & Economic Statistics, 33(3):366–380, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.954707.

Sur:2015:MBD

[SSBD15] Pragya Sur, Galit Shmueli, Smarajit Bose, and Paromita Dubey. Modeling bimodal discrete data using Conway–Maxwell–Poisson mixture models. Journal of Business & Economic Statistics, 33(3):352–365, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.949343.

Seaver:1989:IUM

[ST89] Bill L. Seaver and Konstantinos P. Triantis. The implications of using messy data to estimate production-frontier-based technical efficiency measures. Journal of Business & Economic Statistics, 7(1):49–59, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509712.

Scaillet:2010:TSD

[ST10] Olivier Scaillet and Nikolas Topaloglou. Testing for stochastic dominance efficiency. Journal of Business & Economic Statistics, 28(1):169–180, 2010. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2009.06167.

Song:2018:MNG

[ST18] Xiaojun Song and Abderrahim Taamouti. Measuring nonlinear Granger causality in mean. Journal of Business & Economic Statistics, 36(2):321–333, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1166118.

Song:2020:MGC

[ST20] Xiaojun Song and Abderrahim Taamouti. Measuring Granger causality in quantiles. Journal of Business & Economic Statistics, 39(4):937–952, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1739531.

Song:2022:NDU

[ST22a] Dongho Song and Jenny Tang. News-driven uncertainty fluctuations. Journal of Business & Economic Statistics, 41(3):968– 982, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2097912.

Sun:2022:HDM

[ST22b] Baoluo Sun and Zhiqiang Tan. High-dimensional model-assisted inference for local average treatment effects with instrumental variables. Journal of Business & Economic Statistics, 40(4):1732– 1744, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1970575.

Stambaugh:1988:C

[Sta88a] Robert F. Stambaugh. Comment. Journal of Business & Economic Statistics, 6(1):20–21, 1988. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1988.10509631. See [CR88b, CR88a].

Stasny:1988:MNN

[Sta88b] Elizabeth A. Stasny. Modeling nonignorable nonresponse in categorical panel data with an example in estimating gross laborforce flows. Journal of Business & Economic Statistics, 6 (2):207–219, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509655.

Startz:2008:BAM

[Sta08] Richard Startz. Binomial autoregressive moving average models with an application to U.S. recessions. Journal of Business & Economic Statistics, 26(1):1–8, 2008. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/073500107000000151.

Stekler:1987:WFB

[Ste87] H. O. Stekler. Who forecasts better? Journal of Business & Economic Statistics, 5(1):155–158, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509571.

Stock:1988:RFC

[Sto88a] James H. Stock. A reexamination of Friedman’s consumption puzzle. Journal of Business & Economic Statistics, 6 (4):401–407, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509679.

Stock:1988:R

[Sto88b] James H. Stock. Reply. Journal of Business & Economic Statistics, 6(4):413–414, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509682.

Stokes:1990:ESS

[Sto90] Lynne Stokes. Estimating the size of a subdomain: an application in auditing. Journal of Business & Economic Statistics, 8(3):337–346, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509804.

Stock:2016:C

[Sto16] James H. Stock. Comment. Journal of Business & Economic Statistics, 34(3):339–341, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1186030. See [CR16a, CR16b].

Strachan:2003:VBE

[Str03] Rodney W. Strachan. Valid Bayesian estimation of the cointegrating error correction model. Journal of Business & Economic Statistics, 21(1):185–195, 2003. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102288618883.

Stringham:2022:FBR

[Str22] Thomas Stringham. Fast Bayesian record linkage with recordspecific disagreement parameters. Journal of Business & Economic Statistics, 40(4):1509–1522, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1934478.

Su:2009:TCU

[SU09] Liangjun Su and Aman Ullah. Testing conditional uncorrelatedness. Journal of Business & Economic Statistics, 27(1):18–29, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1198/jbes.2009.0002.

Spreng:2022:CVM

[SU22] Lars Spreng and Giovanni Urga. Combining p-values for multivariate predictive ability testing. Journal of Business & Economic Statistics, 41(3):765–777, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2067545.

Sun:2014:C

[Sun14] Yixiao Sun. Comment. Journal of Business & Economic Statistics, 32(3):330–334, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.926817. See [Mül14a, Mül14b].

Sun:2018:C

[Sun18] Yixiao Sun. Comment. Journal of Business & Economic Statistics, 36(4):565–568, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1505628.

Steckel:1988:HCL

[SV88] Joel H. Steckel and Wilfried R. Vanhonacker. A heterogeneous conditional logit model of choice. Journal of Business & Economic Statistics, 6(3):391–398, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509677.

Smith:2016:AFD

[SV16] Michael S. Smith and Shaun P. Vahey. Asymmetric forecast densities for U.S. macroeconomic variables from a Gaussian copula model of cross-sectional and serial dependence. Journal of Business & Economic Statistics, 34(3):416–434, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1044533.

Swanson:2006:SRA

[SvD06] Norman R. Swanson and Dick van Dijk. Are statistical reporting agencies getting it right? Data rationality and business cycle asymmetry. Journal of Business & Economic Statistics, 24(1):24–42, 2006. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500105000000036.

Swofford:1988:CNT

[SW88] James L. Swofford and Gerald A. Whitney. A comparison of nonparametric tests of weak separability for annual and quarterly data on consumption, leisure, and money. Journal of Business & Economic Statistics, 6(2):241–246, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1988.10509658.

Swofford:1990:BEM

[SW90] James L. Swofford and Gerald A. Whitney. Bounding an economic monetary aggregate under nonhomothetic preferences. Journal of Business & Economic Statistics, 8(1): 137–141, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509783.

Srinivasan:1994:UNC

[SW94] T. C. Srinivasan and Russell S. Winer. Using neoclassical consumer-choice theory to produce a market map from purchase data. Journal of Business & Economic Statistics, 12(1):1–9, 1994. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1994.10509986.

Swanson:1995:MSA

[SW95] Norman R. Swanson and Halbert White. A model-selection approach to assessing the information in the term structure using linear models and artificial neural networks. Journal of Business & Economic Statistics, 13(3):265–275, 1995. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1995.10524600.

Stock:1996:ESI

[SW96] James H. Stock and Mark W. Watson. Evidence on structural instability in macroeconomic time series relations. Journal of Business & Economic Statistics, 14(1):11–30, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524626.

Stock:2002:MFU

[SW02] James H. Stock and Mark W. Watson. Macroeconomic forecasting using diffusion indexes. Journal of Business & Economic Statistics, 20(2):147–162, 2002. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500102317351921.

Stock:2012:GSM

[SW12] James H. Stock and Mark W. Watson. Generalized shrinkage methods for forecasting using many predictors. Journal of Business & Economic Statistics, 30(4):481–493, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2012.715956.

Sasaki:2020:FIC

[SW20] Yuya Sasaki and Yulong Wang. Fixed-k inference for conditional extremal quantiles. Journal of Business & Economic Statistics, 40(2):829–837, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1870985.

Sasaki:2021:DTF

[SW21] Yuya Sasaki and Yulong Wang. Diagnostic testing of finite moment conditions for the consistency and root-n asymptotic normality of the GMM and M estimators. Journal of Business & Economic Statistics, 41(2):339–348, 2021. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
2019047.

Simar:2022:NSF

[SW22a] Léopold Simar and Paul W. Wilson. Nonparametric, stochastic frontier models with multiple inputs and outputs. Journal of Business & Economic Statistics, 41(4):1391–1403, 2022. CO-DEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2022.2110882.

Stauskas:2022:TEF

[SW22b] Ovidijus Stauskas and Joakim Westerlund. Tests of equal forecasting accuracy for nested models with estimated CCE factors*. Journal of Business & Economic Statistics, 40(4):1745– 1758, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1970576.

Swait:2003:FCS

[Swa03a] Joffre Swait. Flexible covariance structures for categorical dependent variables through finite mixtures of generalized extreme value models. Journal of Business & Economic Statistics, 21(1):80–87, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618784.

Swartz:2003:BMC

[Swa03b] Tim Swartz. Bayesian modeling and computations in finaloffer arbitration. Journal of Business & Economic Statistics, 21(1):74–79, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618775.

Swanson:2016:C

[Swa16] Norman R. Swanson. Comment. Journal of Business & Economic Statistics, 34(3):348–353, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1186554. See [CR16a, CR16b].

Su:2019:SET

[SWJ19] Liangjun Su, Xia Wang, and Sainan Jin. Sieve estimation of timevarying panel data models with latent structures. Journal of Business & Economic Statistics, 37(2):334–349, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1340299.

Sandstrom:1988:VEG

[SWW88] Arne Sandström, Jan H. Wretman, and Bertil Walden. Variance estimators of the Gini coefficient-probability sampling. Journal of Business & Economic Statistics, 6(1):113–119, 1988. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1988.10509643.

Stock:2002:SWI

[SWY02] James H. Stock, Jonathan H. Wright, and Motohiro Yogo. A survey of weak instruments and weak identification in generalized method of moments. Journal of Business & Economic Statistics, 20(4):518–529, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618658.

Sun:2020:FBE

[SX20] Yucheng Sun and Wen Xu. A factor-based estimation of integrated covariance matrix with noisy high-frequency data. Journal of Business & Economic Statistics, 40(2):770–784, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1868301.

Shin:1994:ILH

[SY94] Richard T. Shin and John S. Ying. Imposing linear homogeneity on Box–Tidwell flexible functional forms. Journal of Business & Economic Statistics, 12(2):261–265, 1994. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1994.10510013.

Sickles:1998:DDL

[SY98] Robim C. Sickles and Abdo Yazbeck. On the dynamics of demand for leisure and the production of health. Journal of Business & Economic Statistics, 16(2):187–197, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524752.

Schechtman:2008:WDR

[SYA08] Edna Schechtman, Shlomo Yitzhaki, and Yevgeny Artsev. Who does not respond in the household expenditure survey. Journal of Business & Economic Statistics, 26(3):329–344, 2008. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/00.

Shin:2020:NAI

[SZ20] Minchul Shin and Molin Zhong. A new approach to identifying the real effects of uncertainty shocks. Journal of Business & Economic Statistics, 38(2):367–379, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1506342.

Trsvedi:1989:RPM

[TA89] P. K. Trsvedi and J. N. Alexander. Reemployment probability and multiple unemployment spells: a partial-likelihood approach. Journal of Business & Economic Statistics, 7 (3):395–401, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509748.

Taber:2020:TTS

[Tab20] Christopher Taber. Thoughts on “Transparency in Structural Research”. Journal of Business & Economic Statistics, 38 (4):726–727, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1796396.

Tahmiscioglu:2001:IVF

[Tah01] A. Kamil Tahmiscioglu. Intertemporal variation in financial constraints on investment. Journal of Business & Economic Statistics, 19(2):153–165, 2001. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500101316970386.

Tamer:2020:DTS

[Tam20] Elie Tamer. Discussion on “Transparency in Structural Research” by I. Andrews, M. Gentkow and J. Shapiro. Journal of Business & Economic Statistics, 38(4):728–730, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1804917.

Tao:2019:ELR

[Tao19] Jing Tao. Empirical likelihood ratio tests of conditional moment restrictions with unknown functions. Journal of Business & Economic Statistics, 39(1):282–293, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2019.1647845.

Tarozzi:2007:CCS

[Tar07] Alessandro Tarozzi. Calculating comparable statistics from incomparable surveys, with an application to poverty in India. Journal of Business & Economic Statistics, 25(3):314–336, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000233.

Tauchen:1986:R

[Tau86a] George Tauchen. Reply. Journal of Business & Economic Statistics, 4(4):423–425, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509541.

Tauchen:1986:SPG

[Tau86b] George Tauchen. Statistical properties of generalized method-of-moments estimators of structural parameters obtained from financial market data. Journal of Business & Economic Statistics, 4(4):397–416, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509537.

Tauchen:1990:SSG

[Tau90] George Tauchen. Solving the stochastic growth model by using quadrature methods and value-function iterations. Journal of Business & Economic Statistics, 8(1):49–51, 1990. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1990.10509776.

Tauchen:1993:ER

[Tau93a] George Tauchen. Editor’s report. Journal of Business & Economic Statistics, 11(4):367, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509965.

Tauchen:1993:JBE

[Tau93b] George Tauchen. The Journal of Business & Economic Statistics: the first 10 years and a look ahead. Journal of Business & Economic Statistics, 11(4):425, 1993. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1993.10509976.

Tauchen:1993:RMT

[Tau93c] George Tauchen. Remarks on my term at JBES. Journal of Business & Economic Statistics, 11(4):428–431, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509980.

Tauchen:1994:ER

[Tau94] George Tauchen. Editor’s report. Journal of Business & Economic Statistics, 12(4):369, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10524552.

Taylor:2002:RBU

[Tay02] A. M. Robert Taylor. Regression-based unit root tests with recursive mean adjustment for seasonal and nonseasonal time series. Journal of Business & Economic Statistics, 20(2): 269–281, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102317352001.

Taylor:2003:RST

[Tay03] A. M. Robert Taylor. Robust stationarity tests in seasonal time series processes. Journal of Business & Economic Statistics, 21(1):156–163, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618856.

Taylor:2019:FVR

[Tay19] James W. Taylor. Forecasting value at risk and expected shortfall using a semiparametric approach based on the asymmetric Laplace distribution. Journal of Business & Economic Statistics, 37(1):121–133, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2017.1281815.

Turnovsky:1984:ESS

[TD84] Michelle H. L. Turnovsky and William A. Donnelly. Energy substitution, separability, and technical progress in the Australian iron and steel industry. Journal of Business & Economic Statistics, 2(1):54–63, 1984. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1984.10509369.

Terza:1985:RFT

[Ter85] Joseph V. Terza. Reduced-form trinomial probit: a quantal response model without a priori restrictions. Journal of Business & Economic Statistics, 3(1):54–59, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1985.10509426.

Taddy:2016:NBA

[TGCD16] Matt Taddy, Matt Gardner, Liyun Chen, and David Draper. A nonparametric Bayesian analysis of heterogeneous treatment effects in digital experimentation. Journal of Business & Economic Statistics, 34(4):661–672, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1172013.

Trimble:1983:EUI

[TH83] John Trimble and Eric Hirst. Energy use in institutional buildings: Estimates from state energy-audit surveys. Journal of Business & Economic Statistics, 1(4):337–347, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1983.10509362.

Trabelsi:1989:BAF

[TH89] Abdelwahed Trabelsi and Steven C. Hillmer. A benchmarking approach to forecast combination. Journal of Business & Economic Statistics, 7(3):353–362, 1989. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1989.10509744.

Thorsrud:2020:WNN

[Tho20] Leif Anders Thorsrud. Words are the new numbers: a newsy coincident index of the business cycle. Journal of Business & Economic Statistics, 38(2):393–409, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2018.1506344.

Timmermann:2001:SBI

[Tim01] Allan Timmermann. Structural breaks, incomplete information, and stock prices. Journal of Business & Economic Statistics, 19(3):299–314, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500101681019954.

Taddy:2010:BNA

[TK10] Matthew A. Taddy and Athanasios Kottas. A Bayesian nonparametric approach to inference for quantile regression. Journal of Business & Economic Statistics, 28(3):357–369, 2010. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2009.07331.

Thompson:1986:SFB

[TM86] Patrick A. Thompson and Robert B. Miller. Sampling the future: a Bayesian approach to forecasting from univariate time series models. Journal of Business & Economic Statistics, 4 (4):427–436, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509542.

Tabord-Meehan:2019:IDD

[TM19] Max Tabord-Meehan. Inference with dyadic data: Asymptotic behavior of the dyadic-robust t-statistic. Journal of Business & Economic Statistics, 37(4):671–680, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1409630.

Trucios:2023:FCC

[TMH^{+}23] Carlos Trucíos, João H. G. Mazzeu, Marc Hallin, Luiz K. Hotta, Pedro L. Valls Pereira, and Mauricio Zevallos. Forecasting conditional covariance matrices in high-dimensional time series: a general dynamic factor approach. Journal of Business & Economic Statistics, 41(1):40–52, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1996380.

Thompson:1992:EEC

[TN92] Patrick A. Thompson and Thomas Noordewier. Estimating the effects of consumer incentive programs on domestic automobile sales. Journal of Business & Economic Statistics, 10 (4):409–417, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509916.

Trivellato:1986:PDE

[TR86] Ugo Trivellato and Enrico Rettore. Preliminary data errors and their impact on the forecast error of simultaneousequations models. Journal of Business & Economic Statistics, 4(4):445–453, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509544.

Trumbo:1993:ASI

[TR93] Bruce E. Trumbo and Dixie Watts Reaves. An author and subject index to the Journal of Business & Economic Statistics, Volumes 1–10 (1983–1992). Journal of Business & Economic Statistics, 11(4):435–436, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509982.

Taylor:2001:TSU

[TS01] A. M. Robert Taylor and Richard J. Smith. Tests of the seasonal unit-root hypothesis against heteroscedastic seasonal integration. Journal of Business & Economic Statistics, 19 (2):192–207, 2001. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500101316970412.

Tsay:1985:MID

[Tsa85] Ruey S. Tsay. Model identification in dynamic regression (distributed lag) models. Journal of Business & Economic Statistics, 3(3):228–237, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509454.

Tsay:1989:PPV

[Tsa89] Ruey S. Tsay. Parsimonious parameterization of vector autoregressive moving average models. Journal of Business & Economic Statistics, 7(3):327–341, 1989. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1989.10509742.

Tsay:1993:C

[Tsa93a] Ruey S. Tsay. Comment. Journal of Business & Economic Statistics, 11(4):390–392, 1993. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1993.10509971.

Tsay:1993:CAF

[Tsa93b] Ruey S. Tsay. Comment: Adaptive forecasting. Journal of Business & Economic Statistics, 11(2):140–142, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509941.

Tsay:1993:TNM

[Tsa93c] Ruey S. Tsay. Testing for noninvertibie models with applications. Journal of Business & Economic Statistics, 11 (2):225–233, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509951.

Tsay:2000:EIP

[Tsa00] Ruey S. Tsay. Editor’s introduction to panel discussion on analysis of high-frequency data. Journal of Business & Economic Statistics, 18(2):139, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524855.

Tsay:2016:SMA

[Tsa16] Ruey S. Tsay. Some methods for analyzing big dependent data. Journal of Business & Economic Statistics, 34(4):673–688, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1148040.

Tse:1987:DTM

[Tse87] Y. K. Tse. A diagnostic test for the multinomial logit model. Journal of Business & Economic Statistics, 5(2): 283–286, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509587.

Tse:1989:PRU

[Tse89] Y. K. Tse. A proportional random utility approach to qualitative response models. Journal of Business & Economic Statistics, 7(1):61–65, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509713.

Tsionas:2022:EMC

[Tsi22] Mike G. Tsionas. Estimating monotone concave stochastic production frontiers. Journal of Business & Economic Statistics, 40(3):1403–1414, 2022. CODEN ???? ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2021.1931240.

Tiao:1983:MTS

[TT83] George C. Tiao and Ruey S. Tsay. Multiple time series modeling and extended sample cross-correlations. Journal of Business & Economic Statistics, 1(1):43–56, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1983.10509323.

Tse:2002:MGA

[TT02] Y. K. Tse and Albert K. C. Tsui. A multivariate generalized autoregressive conditional heteroscedasticity model with timevarying correlations. Journal of Business & Economic Statistics, 20(3):351–362, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618496.

Todorov:2006:SML

[TT06] Viktor Todorov and George Tauchen. Simulation methods for Lévy-driven Continuous-Time Autoregressive Moving Average (CARMA) stochastic volatility models. Journal of Business & Economic Statistics, 24(4):455–469, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500106000000260.

Todorov:2011:VJ

[TT11] Viktor Todorov and George Tauchen. Volatility jumps. Journal of Business & Economic Statistics, 29(3):356–371, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2010.08342.

Tauchen:1995:DBN

[TTW95] George Tauchen, Ruey Tsay, and Mark Watson. Daniel B. Nelson, 1959–1995. Journal of Business & Economic Statistics, 13(4):361, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524610.

Taylor:1990:SNS

[TU90] John B. Taylor and Harald Uhlig. Solving nonlinear stochastic growth models: a comparison of alternative solution methods. Journal of Business & Economic Statistics, 8 (1):1–17, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509766.

Tucker:1992:RFI

[Tuc92] Alan L. Tucker. A reexamination of finite- and infinite-variance distributions as models of daily stock returns. Journal of Business & Economic Statistics, 10(1):73–81, 1992. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1992.10509888.

Tuzcuoglu:2022:CLE

[Tuz22] Kerem Tuzcuoglu. Composite likelihood estimation of an autoregressive panel ordered probit model with random effects. Journal of Business & Economic Statistics, 41(2):593– 607, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2044829.

Thimme:2015:ACS

[TV15] Julian Thimme and Clemens Völkert. Ambiguity in the crosssection of expected returns: an empirical assessment. Journal of Business & Economic Statistics, 33(3):418–429, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2014.958230.

Tsay:1998:EA

[TW98] Ruey S. Tsay and Mark W. Watson. Editorial announcement. Journal of Business & Economic Statistics, 16(1):1, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1998.10524729.

Turtle:2014:MCC

[TW14] H. J. Turtle and Kainan Wang. Modeling conditional covariances with economic information instruments. Journal of Business & Economic Statistics, 32(2):217–236, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2013.859078.

Tschernig:2013:LRI

[TWW13] Rolf Tschernig, Enzo Weber, and Roland Weigand. Long-run identification in a fractionally integrated system. Journal of Business & Economic Statistics, 31(4):438–450, 2013. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2013.812517.

Tang:2022:QCB

[TXLT22] Wenlu Tang, Jinhan Xie, Yuanyuan Lin, and Niansheng Tang. Quantile correlation-based variable selection. Journal of Business & Economic Statistics, 40(3):1081–1093, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1899932.

Trajtenberg:1989:DIM

[TY89] Manuel Trajtenberg and Shlomo Ystzhaki. The diffusion of innovations: a methodological reappraisal. Journal of Business & Economic Statistics, 7(1):35–47, 1989. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1989.10509711.

Tyrrell:1983:UPS

[Tyr83] Timothy J. Tyrrell. The use of polynomials to shift coefficients in linear regression models. Journal of Business & Economic Statistics, 1(3):249–252, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509348.

Tan:2019:EPC

[TZ19] Lili Tan and Yichong Zhang. M-estimators of U-processes with a change-point due to a covariate threshold. Journal of Business & Economic Statistics, 37(2):248–259, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2017.1319373.

Uhlig:1994:C

[Uhl94] Harald Uhlig. Comment. Journal of Business & Economic Statistics, 12(4):410–412, 1994. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1994.10524562.

Urga:2007:CFE

[Urg07] Giovanni Urga. Common features in economics and finance. Journal of Business & Economic Statistics, 25(1):2–11, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000602.

Uematsu:2023:ESI

[UY23a] Yoshimasa Uematsu and Takashi Yamagata. Estimation of sparsity-induced weak factor models. Journal of Business & Economic Statistics, 41(1):213–227, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
2008405.

Uematsu:2023:ISI

[UY23b] Yoshimasa Uematsu and Takashi Yamagata. Inference in sparsityinduced weak factor models. Journal of Business & Economic Statistics, 41(1):126–139, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.2003203.

Valliant:1988:ELP

[Val88] Richard Valliant. Estimation of Laspeyres price indexes using the prediction approach for finite population sampling. Journal of Business & Economic Statistics, 6(2): 189–196, 1988. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1988.10509653.

Valliant:1991:VEP

[Val91] Richard Valliant. Variance estimation for price indexes from a two-stage sample with rotating panels. Journal of Business & Economic Statistics, 9(4):409–422, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509868.

VanBiesebroeck:2008:SPE

[Van08] Johannes Van Biesebroeck. The sensitivity of productivity estimates. Journal of Business & Economic Statistics, 26 (3):311–328, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500107000000089.

Varneskov:2016:FTR

[Var16] Rasmus Tangsgaard Varneskov. Flat-top realized kernel estimation of quadratic covariation with nonsynchronous and noisy asset prices. Journal of Business & Economic Statistics, 34(1):1–22, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1005622.

Viscusi:1995:GLR

[VB95] W. Kip Viscusi and Patricia Born. The general-liability reform experiments and the distribution of insurance-market outcomes. Journal of Business & Economic Statistics, 13 (2):183–188, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524592.

vanderHoeven:1986:MSA

[vdHH86] H. van der Hoeven and A. J. Hundepool. A method for seasonally adjusting time series with variation in the seasonal amplitude. Journal of Business & Economic Statistics, 4 (4):455–471, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509545.

vanderKlaauw:2005:C

[vdK05] Wilbert van der Klaauw. Comment. Journal of Business & Economic Statistics, 23(2):154–157, 2005. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1198/073500104000000721. See [AV05b, AV05a].

vanderKnoop:1989:MIM

[vdKH89] Han S. van der Knoop and Frans C. Hooijmans. A multivariate intervention model for the Dutch mint circulation: Estimation and Monte Carlo simulation. Journal of Business & Economic Statistics, 7(2):179–189, 1989. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1989.10509726.

vanderKlaauw:2003:TNA

[vdKK03] Bas van der Klaauw and Ruud H. Koning. Testing the normality assumption in the sample selection model with an application to travel demand. Journal of Business & Economic Statistics, 21(1):31–42, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102288618739.

Vrontos:2000:FBI

[VDP00] I. D. Vrontos, P. Dellaportas, and D. N. Politis. Full Bayesian inference for GARCH and EGARCH models. Journal of Business & Economic Statistics, 18(2):187–198, 2000. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.2000.10524861.

Velasco:2022:IES

[Vel22] Carlos Velasco. Identification and estimation of structural VARMA models using higher order dynamics. Journal of Business & Economic Statistics, 41(3):819–832, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2075000.

Verbrugge:2012:DCP

[Ver12] Randal J. Verbrugge. Do the Consumer Price Index’s utilities adjustments for owners’ equivalent rent distort inflation measurement? Journal of Business & Economic Statistics, 30(1): 143–148, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic).

URL http://www.tandfonline.com/doi/full/10.1198/jbes.
2011.08016.

VanDijk:1999:TST

[VFL99] Dick Van Dijk, Philip Hans Franses, and André Lucas. Testing for smooth transition nonlinearity in the presence of outliers. Journal of Business & Economic Statistics, 17(2): 217–235, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524812.

VandenBerg:1997:JSC

[VG97] Gerard J. Van den Berg and Cees Gorter. Job search and commuting time. Journal of Business & Economic Statistics, 15(2):269–281, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524705.

vonHagen:1989:RCP

[vH89] Juergen von Hagen. Relative commodity prices and cointegration. Journal of Business & Economic Statistics, 7 (4):497–503, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509763.

vonHaefen:2004:EWA

[vHPP04] Roger H. von Haefen, Daniel J. Phaneuf, and George R. Parsons. Estimation and welfare analysis with large demand systems. Journal of Business & Economic Statistics, 22(2):194–205, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500104000000082.

Vinod:1985:MED

[Vin85a] H. D. Vinod. Measurement of economic distance between blacks and whites. Journal of Business & Economic Statistics, 3(1):78–88, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/

07350015.1985.10509429. See comment [Gas85] and reply [Vin85b].

Vinod:1985:R

[Vin85b] H. D. Vinod. Reply. Journal of Business & Economic Statistics, 3(4):408–409, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509480. See [Vin85a, Gas85].

ValleeAzevedo:2006:TBC

[VKR06] João Valle e Azevedo, Siem Jan Koopman, and António Rua. Tracking the business cycle of the Euro area. Journal of Business & Economic Statistics, 24(3):278–290, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500105000000261.

Van:2004:PGF

[VLK04] Phu Nguyen Van, François Laisney, and Ulrich Kaiser. The performance of German firms in the business-related service sector. Journal of Business & Economic Statistics, 22(3): 274–295, 2004. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000145.

Vanhonacker:1990:CRS

[VLS90] Wilfried R. Vanhonacker, Donald R. Lehmann, and Fareena Sultan. Combining related and sparse data in linear regression models. Journal of Business & Economic Statistics, 8 (3):327–335, 1990. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1990.10509803.

Vailiant:1989:CME

[VM89] Richard Vailiant and Stephen M. Miller. A class of multiplicative estimators of Laspeyres price indexes. Journal of Business & Economic Statistics, 7(3):387–394, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509747.

vanOphem:2000:MSC

[vO00] Hans van Ophem. Modeling selectivity in count-data models. Journal of Business & Economic Statistics, 18(4): 503–511, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524889.

Vogelsang:1998:TSM

[Vog98] Timothy J. Vogelsang. Testing for a shift in mean without having to estimate serial-correlation parameters. Journal of Business & Economic Statistics, 16(1):73–80, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1998.10524736.

Vogelsang:2014:C

[Vog14] Timothy J. Vogelsang. Comment. Journal of Business & Economic Statistics, 32(3):334–338, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.926818. See [Mül14a,

Mül14b].

Vogelsang:2018:CHI

[Vog18] Timothy J. Vogelsang. Comment on “HAR inference: Recommendations for practice”. Journal of Business & Economic Statistics, 36(4):569–573, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1497503.

Volpicella:2022:SIT

[Vol22] Alessio Volpicella. SVARs identification through bounds on the forecast error variance. Journal of Business & Economic Statistics, 40(3):1291–1301, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1927742.

Vossmeyer:2016:SST

[Vos16] Angela Vossmeyer. Sample selection and treatment effect estimation of lender of last resort policies. Journal of Business & Economic Statistics, 34(2):197–212, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1024837.

Vanhonacker:1992:UMA

[VP92] Wilfried R. Vanhonacker and Lydia J. Price. Using metaanalysis results in Bayesian updating: The empty-cell problem. Journal of Business & Economic Statistics, 10(4): 427–435, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509918.

Vlaar:1993:MWE

[VP93] Peter J. G. Vlaar and Franz C. Palm. The message in weekly exchange rates in the European monetary system: Mean reversion, conditional heteroscedasticity, and jumps. Journal of Business & Economic Statistics, 11(3):351–360, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509963.

VannOphem:1999:MLM

[VSV99] Hans Vann Ophem, Piet Stam, and Bernard Van Praag. Multichoice logit: Modeling incomplete preference rankings of classical concerts. Journal of Business & Economic Statistics, 17 (1):117–128, 1999. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1999.10524801.

vonUngern-Sternberg:1987:DSN

[vUS87] Thomas von Ungern-Sternberg. Does the Swiss national bank stabilize the Swiss franc exchange rates? Journal of Business & Economic Statistics, 5(1):105–113, 1987. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1987.10509565.

Vella:1999:EIM

[VV99] Francis Vella and Marno Verbeek. Estimating and interpreting models with endogenous treatment effects. Journal of Business & Economic Statistics, 17(4):473–478, 1999. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1999.10524835.

Vukina:2011:HHC

[VZ11] Tomislav Vukina and Xiaoyong Zheng. Homogeneous and heterogeneous contestants in piece rate tournaments: Theory and empirical analysis. Journal of Business & Economic Statistics, 29(4):506–517, 2011. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
jbes.2010.08345.

Waite:1985:CFA

[Wai85] Charles A. Waite. Comment: The future of administrative records in the economic programs of the Census Bureau. Journal of Business & Economic Statistics, 3(4):400–401, 1985. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1985.10509477.

Wallis:1998:C

[Wal98] Kenneth F. Wallis. Comment. Journal of Business & Economic Statistics, 16(2):164–165, 1998. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1998.10524747.

Wallis:2004:C

[Wal04] Kenneth F. Wallis. Comment. Journal of Business & Economic Statistics, 22(2):172–175, 2004. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/073500104000000055. See [PSW04a] and rejoinder

[PSW04b].

Wang:1996:PUE

[Wan96] Wenyu Wang. Prediction of the U.S. employment links: an application of an empirical Bayes procedure. Journal of Business & Economic Statistics, 14(2):243–250, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524651.

Wang:2003:SFA

[Wan03] Hung-Jen Wang. A stochastic frontier analysis of financing constraints on investment. Journal of Business & Economic Statistics, 21(3):406–419, 2003. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500103288619016.

Warga:1989:EDT

[War89] Arthur Warga. Experimental design in tests of linear factor models. Journal of Business & Economic Statistics, 7 (2):191–198, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509727.

Watson:1987:C

[Wat87] Mark W. Watson. Comment. Journal of Business & Economic Statistics, 5(4):451–453, 1987. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1987.10509612.

Watson:1988:C

[Wat88] Mark W. Watson. Comment. Journal of Business & Economic Statistics, 6(4):408–409, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509680.

Watson:1996:C

[Wat96] Mark W. Watson. Comment. Journal of Business & Economic Statistics, 14(3):394–396, 1996. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1996.10524667.

Watanabe:2000:BAD

[Wat00] Toshiaki Watanabe. Bayesian analysis of dynamic bivariate mixture models: Can they explain the behavior of returns and trading volume? Journal of Business & Economic Statistics, 18(2):199–210, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524862.

Watanabe:2003:EDB

[Wat03] Toshiaki Watanabe. The estimation of dynamic bivariate mixture models. Journal of Business & Economic Statistics, 21 (4):577–580, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619304.

Wolfson:2004:MW

[WB04] Paul Wolfson and Dale Belman. The minimum wage. Journal of Business & Economic Statistics, 22(3):296–311, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500104000000154.

Weisberg:1986:BR

[WCM^{+}86] Herbert I. Weisberg, Samprit Chatterjee, Mukul Majumdar, Edward L. Melnick, and Alan J. Oppenheim. Book reviews. Journal of Business & Economic Statistics, 4(2):

275–278, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509523.

Wedel:1996:EFM

[WD96] Michel Wedel and Wayne S. Desarbo. An exponential-family multidimensional scaling mixture methodology. Journal of Business & Economic Statistics, 14(4):447–459, 1996. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1996.10524674.

Weber:2000:RTC

[Web00] Christian E. Weber. “rule-of-thumb” consumption, intertemporal substitution, and risk aversion. Journal of Business & Economic Statistics, 18(4):497–502, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524888.

Wecker:1989:CAA

[Wec89] William E. Wecker. Comment: Assessing the accuracy of time series mode! forecasts of count observations. Journal of Business & Economic Statistics, 7(4):418–419, 1989. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1989.10509751.

Weiss:1986:ABT

[Wei86] Andrew A. Weiss. ARCH and bilinear time series models: Comparison and combination. Journal of Business & Economic Statistics, 4(1):59–70, 1986. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1986.10509494.

Werner:1999:AZD

[Wer99] Megan Werner. Allowing for zeros in dichotomous-choice contingent-valuation models. Journal of Business & Economic Statistics, 17(4):479–486, 1999. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1999.10524836.

West:2001:TFE

[Wes01] Kenneth D. West. Tests for forecasts encompassing when forecasts depend on estimated regression parameters. Journal of Business & Economic Statistics, 19(1):29–33, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/07350010152472580.

West:2012:C

[Wes12] Kenneth D. West. Comment. Journal of Business & Economic Statistics, 30(1):34–35, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2012.634350. See [PT12a, PT12b].

Westerlund:2014:HRP

[Wes14] Joakim Westerlund. Heteroscedasticity robust panel unit root tests. Journal of Business & Economic Statistics, 32(1):112–135, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2013.
857612.

Westerlund:2015:RUA

[Wes15] Joakim Westerlund. Rethinking the univariate approach to panel unit root testing: Using covariates to resolve the incidental trend problem. Journal of Business & Economic Statistics, 33(3):430–443, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.962697.

West:2018:DLL

[Wes18a] Kenneth D. West. Discussion of Lazarus, Lewis, Stock, and Watson, “HAR Inference: Recommendations for Practice”. Journal of Business & Economic Statistics, 36(4):560–562, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.

1505627. See [Mül18, LLSW18a].

Westerlund:2018:UGD

[Wes18b] Joakim Westerlund. On the use of GLS demeaning in panel unit root testing. Journal of Business & Economic Statistics, 36(2):309–320, 2018. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2016.1152969.

Weng:2022:DCC

[WF22] Haolei Weng and Yang Feng. Discussion of “Cocitation and Coauthorship Networks of Statisticians”. Journal of Business & Economic Statistics, 40(2):486–490, 2022. CODEN RFSCFJ.

ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2022.

2037432. See [JJKL21].

Wlazlowski:2012:PTE

[WGBM12] Szymon Wlazlowski, Monica Giulietti, Jane Binner, and Costas Milas. Price transmission in the EU wholesale petroleum markets. Journal of Business & Economic Statistics, 30(2):165–172, 2012. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2012.
672290.

Whited:1998:WDI

[Whi98] Toni M. Whited. Why do investment Euler equations fail? Journal of Business & Economic Statistics, 16(4):479–488, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1998.10524787.

Wiesenfarth:2014:BNI

[WHKCS14] Manuel Wiesenfarth, Carlos Matías Hisgen, Thomas Kneib, and Carmen Cadarso-Suarez. Bayesian nonparametric instrumental variables regression based on penalized splines and Dirichlet process mixtures. Journal of Business & Economic Statistics, 32(3):468–482, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.907092.

Wang:2022:SMF

[WHT22] Lijia Wang, Xu Han, and Xin Tong. Skilled mutual fund selection: False discovery control under dependence. Journal of Business & Economic Statistics, 41(2):578–592, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2044337.

Wu:2022:NGD

[WHW22] Shuyuan Wu, Danyang Huang, and Hansheng Wang. Network gradient descent algorithm for decentralized federated learning. Journal of Business & Economic Statistics, 41(3):806– 818, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2074426.

Wang:2022:RIN

[WHYC22] Shouxia Wang, Tao Huang, Jinhong You, and Ming-Yen Cheng. Robust inference for nonstationary time series with possibly multiple changing periodic structures. Journal of Business & Economic Statistics, 40(4):1718–1731, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1970574.

Wilcox:1983:DDU

[Wil83] James A. Wilcox. Disaggregating data using related series. Journal of Business & Economic Statistics, 1(3):187–191, 1983. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1983.10509338.

Wilson:1993:DOD

[Wil93] Paul W. Wilson. Detecting outliers in deterministic nonparametric frontier models with multiple outputs. Journal of Business & Economic Statistics, 11(3):319–323, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509959.

Wilson:2009:IBC

[Wil09] Daniel J. Wilson. IT and beyond: The contribution of heterogeneous capital to productivity. Journal of Business & Economic Statistics, 27(1):52–70, 2009. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.0005.

Winkler:1989:C

[Win89] Robert L. Winkler. Comment. Journal of Business & Economic Statistics, 7(4):419–422, 1989. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1989.10509752.

Winkelmann:1995:DDD

[Win95] Rainer Winkelmann. Duration dependence and dispersion in count-data models. Journal of Business & Economic Statistics, 13(4):467–474, 1995. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1995.10524620.

Winkler:2005:C

[Win05] William E. Winkler. Comment. Journal of Business & Economic Statistics, 23(2):153–154, 2005. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500104000000686. See [AV05b, AV05a].

Wang:2008:EFO

[WJS08] Shanshan Wang, Wolfgang Jank, and Galit Shmueli. Explaining and forecasting online auction prices and their dynamics using functional data analysis. Journal of Business & Economic Statistics, 26(2):144–160, 2008. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500106000000477.

Wilson:1991:HMC

[WK91] Jeffrey R. Wilson and Kenneth J. Koehler. Hierarchical models for cross-classified overdispersed multinomial data. Journal of Business & Economic Statistics, 9(1):103–110, 1991. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1991.10509832.

Wang:2023:BTH

[WKX23] Lengyang Wang, Efang Kong, and Yingcun Xia. Bootstrap tests for high-dimensional white-noise. Journal of Business & Economic Statistics, 41(1):241–254, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
2008407.

Wang:2021:MTS

[WL21] Jingli Wang and Jialiang Li. Multi-threshold structural equation model. Journal of Business & Economic Statistics, 41(2):377– 387, 2021. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.2023553.

Wang:2007:RRS

[WLJ07] Hansheng Wang, Guodong Li, and Guohua Jiang. Robust regression shrinkage and consistent variable selection through the LAD-lasso. Journal of Business & Economic Statistics, 25 (3):347–355, 2007. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000251.

Wu:2018:BIA

[WLL18] Jiexing Wu, Kate J. Li, and Jun S. Liu. Bayesian inference for assessing effects of email marketing campaigns. Journal of Business & Economic Statistics, 36(2):253–266, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2016.1141096.

Wang:1998:APD

[WlMCP98] Peiming Wang, lain M. Cockburn, and Martin L. Puterman. Analysis of patent data — a mixed-Poisson-regression-model approach. Journal of Business & Economic Statistics, 16 (1):27–41, 1998. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1998.10524732.

Wang:2020:IGS

[WLTL20] Zhanfeng Wang, Xianhui Liu, Wenlu Tang, and Yuanyuan Lin. Incorporating graphical structure of predictors in sparse quantile regression. Journal of Business & Economic Statistics, 39 (3):783–792, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1730859.

Wang:2019:STT

[WLW19] Feifei Wang, Jingyuan Liu, and Hansheng Wang. Sequential textterm selection in vector space models. Journal of Business & Economic Statistics, 39(1):82–97, 2019. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2019.1634079.

Wu:2022:ION

[WLZT22] Yujia Wu, Wei Lan, Tao Zou, and Chih-Ling Tsai. Inward and outward network influence analysis. Journal of Business & Economic Statistics, 40(4):1617–1628, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1953509.

Whiteside:1989:RRC

[WN89] M. M. Whiteside and A. Narayanan. Reverse regression, collinearity, and employment discrimination. Journal of Business & Economic Statistics, 7(3):403–406, 1989. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.1989.10509749.

Wurst:1989:ESS

[WNG89] John Wurst, John Neter, and James Godfrey. Efficiency of sieve sampling in auditing. Journal of Business & Economic Statistics, 7(2):199–205, 1989. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1989.10509728.

Wolff:1987:TVP

[Wol87] Christian C. P. Wolff. Time-varying parameters and the out-of-sample forecasting performance of structural exchange rate models. Journal of Business & Economic Statistics, 5 (1):87–97, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509563.

Wolf:2000:SRD

[Wol00] Michael Wolf. Stock returns and dividend yields revisited: a new way to look at an old problem. Journal of Business & Economic Statistics, 18(1):18–30, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524844.

Wooldridge:1998:AZT

[Woo98] Jeffrey M. Wooldridge. Announcement: Zellner Thesis Award. Journal of Business & Economic Statistics, 16(1):125, 1998. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07350015.1998.10524742.

Wood:2000:MMR

[Woo00a] Robert A. Wood. Market microstructure research databases: History and projections. Journal of Business & Economic Statistics, 18(2):140–145, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524856.

Wooldridge:2000:ER

[Woo00b] Jeffrey M. Wooldridge. Editor’s report. Journal of Business & Economic Statistics, 18(1):137, 2000. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07350015.2000.10524854.

Wooldridge:2001:EA

[Woo01] Jeffrey M. Wooldridge. Editorial announcement. Journal of Business & Economic Statistics, 19(1):1, 2001. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/07350010152472562.

Womer:1983:ETL

[WP83] N. Keith Womer and J. Wayne Patterson. Estimation and testing of learning curves. Journal of Business & Economic Statistics, 1(4):265–272, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509350.

Womer:1987:C

[WP87] Norman Keith Womer and J. Wayne Patterson. Comment. Journal of Business & Economic Statistics, 5(3): 336–337, 1987. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1987.10509596.

Wright:1983:MPS

[Wri83] Roger L. Wright. Measuring the precision of statistical cost allocations. Journal of Business & Economic Statistics, 1 (2):93–100, 1983. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1983.10509328.

Wright:1985:R

[Wri85] Roger L. Wright. Reply. Journal of Business & Economic Statistics, 3(3):286–288, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509459.

Wright:2000:AVR

[Wri00a] Jonathan H. Wright. Alternative variance-ratio tests using ranks and signs. Journal of Business & Economic Statistics, 18(1):1–9, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524842.

Wright:2000:CIU

[Wri00b] Jonathan H. Wright. Confidence intervals for univariate impulse responses with a near unit root. Journal of Business & Economic Statistics, 18(3):368–373, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524877.

Wright:2000:CSC

[Wri00c] Jonathan H. Wright. Confidence sets for cointegrating coefficients based on stationarity tests. Journal of Business & Economic Statistics, 18(2):211–222, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524863.

Wright:2009:C

[Wri09] Jonathan H. Wright. Comment. Journal of Business & Economic Statistics, 27(3):323–326, 2009. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.08274. See [KM09b, KM09a].

Wright:2015:C

[Wri15] Jonathan H. Wright. Comment. Journal of Business & Economic Statistics, 33(1):12–13, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.969429. See [AGO^{+}14a, AGO^{+}14b].

Wang:2020:SAD

[WS20] Xuexin Wang and Yixiao Sun. A simple asymptotically F - distributed portmanteau test for diagnostic checking of time series models with uncorrelated innovations. Journal of Business & Economic Statistics, 40(2):505–521, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1832505.

Wu:2022:TTS

[WSX22] Jilin Wu, Xiaojun Song, and Zhijie Xiao. Testing for trend specifications in panel data models. Journal of Business & Economic Statistics, 41(2):453–466, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2035227.

Weisberg:1984:PIL

[WTC84] Herbert I. Weisberg, Thomas J. Tomberlin, and Sangit Chatterjee. Predicting insurance losses under cross-classification: a comparison of alternative approaches. Journal of Business & Economic Statistics, 2(2):170–178, 1984. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1984.10509385.

Wu:2019:RLC

[Wu19] Ximing Wu. Robust likelihood cross-validation for kernel density estimation. Journal of Business & Economic Statistics, 37(4):761–770, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1424633.

Wuthrich:2020:CTQ

[Wüt20] Kaspar Wüthrich. A comparison of two quantile models with endogeneity. Journal of Business & Economic Statistics, 38(2):443–456, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1514307.

Wegmann:2011:BIS

[WV11] Bertil Wegmann and Mattias Villani. Bayesian inference in structural second-price common value auctions. Journal of Business & Economic Statistics, 29(3):382–396, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2011.08289.

Walsh:2022:LSM

[WV22] Christopher Walsh and Michael Vogt. Locally stationary multiplicative volatility modeling. Journal of Business & Economic Statistics, 41(2):497–508, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2036612.

West:1996:CAI

[WW96] Kenneth D. West and David W. Wilcox. A comparison of alternative instrumental variables estimators of a dynamic linear model. Journal of Business & Economic Statistics, 14 (3):281–293, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524657.

Wheelock:2009:RNQ

[WW09] David C. Wheelock and Paul W. Wilson. Robust nonparametric quantile estimation of efficiency and productivity change in U.S. commercial banking, 1985–2004. Journal of Business & Economic Statistics, 27(3):354–368, 2009. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.06145.

Wen:2020:TKE

[WW20] Kuangyu Wen and Ximing Wu. Transformation-kernel estimation of copula densities. Journal of Business & Economic Statistics, 38(1):148–164, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2018.1469999.

Wang:2000:BTS

[WZ00] Jiahui Wang and Eric Zivot. A Bayesian time series model of multiple structural changes in level, trend, and variance. Journal of Business & Economic Statistics, 18(3): 374–386, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524878.

Wu:2015:NIT

[WZ15] Weichi Wu and Zhou Zhou. Nonparametric inference for timevarying coefficient quantile regression. Journal of Business & Economic Statistics, 35(1):98–109, 2015. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07350015.2015.1060884.

Wooldridge:2020:IAS

[WZ20a] Jeffrey M. Wooldridge and Ying Zhu. Inference in approximately sparse correlated random effects probit models with panel data. Journal of Business & Economic Statistics, 38(1):1–18, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2019.

1681276. See comment [HW20a] and rejoinder [WZ20b].

Wooldridge:2020:R

[WZ20b] Jeffrey M. Wooldridge and Ying Zhu. Rejoinder. Journal of Business & Economic Statistics, 38(1):25–26, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2019.1681278.

Wang:2022:TMD

[WZS22] Guochang Wang, Ke Zhu, and Xiaofeng Shao. Testing for the martingale difference hypothesis in multivariate time series models. Journal of Business & Economic Statistics, 40(3):980– 994, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1889568.

Xu:2021:DTS

[XCK21] Xiaofei Xu, Ying Chen, and Steven Kou. Discussion on “Text Selection”. Journal of Business & Economic Statistics, 39 (4):883–887, 2021. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2021.1942890. See [KMM21].

Xu:2020:PEE

[XHL20] Wen Xu, Yanxi Hou, and Deyuan Li. Prediction of extremal expectile based on regression models with heteroscedastic extremes. Journal of Business & Economic Statistics, 40(2):522– 536, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2020.1833890.

Xie:2022:IGE

[Xie22] Erhao Xie. Inference in games without equilibrium restriction: An application to restaurant competition in opening hours. Journal of Business & Economic Statistics, 40(4):1803–1816, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1981914.

Xu:2021:LSQ

[XKZ21] Zhuying Xu, Seonjin Kim, and Zhibiao Zhao. Locally stationary quantile regression for inflation and interest rates. Journal of Business & Economic Statistics, 40(2):838–851, 2021. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2021.1874389.

Xu:2011:TNE

[XP11] Ke-Li Xu and Peter C. B. Phillips. Tilted nonparametric estimation of volatility functions with empirical applications. Journal of Business & Economic Statistics, 29(4):518–528, 2011. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/jbes.2011.09012.

Xu:2000:IGG

[Xu00] Kuan Xu. Inference for generalized Gini indices using the iteratedbootstrap method. Journal of Business & Economic Statistics, 18(2):223–227, 2000. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.2000.10524864.

Xu:2022:EEM

[XZ22] Zhanxiong Xu and Zhibiao Zhao. Efficient estimation for models with nonlinear heteroscedasticity. Journal of Business & Economic Statistics, 40(4):1498–1508, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1933991.

Xia:2020:MED

[XZL20] Yifan Xia, Ling Zhang, and Iris L. Li. Multidimensional economic dispersion index and application. Journal of Business & Economic Statistics, 39(3):729–740, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2020.1730185.

Xie:2014:VCE

[XZW14] Shangyu Xie, Yong Zhou, and Alan T. K. Wan. A varyingcoefficient expectile model for estimating value at risk. Journal of Business & Economic Statistics, 32(4):576–592, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07350015.2014.917979.

Xu:2016:DVC

[XZW16] Minya Xu, Ping-Shou Zhong, and Wei Wang. Detecting variance change-points for blocked time series and dependent panel data. Journal of Business & Economic Statistics, 34(2):213–226, 2016. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1026438.

Xu:2022:CAC

[XZW22] Shirong Xu, Yaoming Zhen, and Junhui Wang. Covariate-assisted community detection in multi-layer networks. Journal of Business & Economic Statistics, 41(3):915–926, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2085726.

Yamamoto:2016:FNF

[Yam16] Yohei Yamamoto. Forecasting with nonspurious factors in U.S. macroeconomic time series. Journal of Business & Economic Statistics, 34(1):81–106, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1004071.

Yang:2020:NCE

[Yan20a] Lu Yang. Nonparametric copula estimation for mixed insurance claim data. Journal of Business & Economic Statistics, 40(2):537– 546, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2020.1835668.

Yang:2020:SEC

[Yan20b] Xiye Yang. Semiparametric estimation in continuous-time: Asymptotics for integrated volatility functionals with small and large bandwidths. Journal of Business & Economic Statistics, 39(3):793–806, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1733583.

Yang:2022:ELE

[Yan22] Xiye Yang. Estimation of leverage effect: Kernel function and efficiency. Journal of Business & Economic Statistics, 41(3):939– 956, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2097910.

Yao:2008:DFS

[Yao08] Tong Yao. Dynamic factors and the source of momentum profits. Journal of Business & Economic Statistics, 26(2): 211–226, 2008. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500106000000648.

Yao:2014:Ca

[Yao14a] Qiwei Yao. Comment. Journal of Business & Economic Statistics, 32(2):165–166, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.887014. See [GPI13b, GPI13a].

Yao:2014:Cb

[Yao14b] Qiwei Yao. Comment. Journal of Business & Economic Statistics, 32(2):201, 2014. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.887015. See [FQX14a, FQX14b].

Yagi:2020:SCK

[YCJK20] Daisuke Yagi, Yining Chen, Andrew L. Johnson, and Timo Kuosmanen. Shape-constrained kernel-weighted least squares: Estimating production functions for Chilean manufacturing industries. Journal of Business & Economic Statistics, 38(1):43–54, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2018.
1431128.

Yu:2020:TRT

[YF20] Ping Yu and Xiaodong Fan. Threshold regression with a threshold boundary. Journal of Business & Economic Statistics, 39 (4):953–971, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1740712.

Yousuf:2022:TPP

[YF22] Kashif Yousuf and Yang Feng. Targeting predictors via partial distance correlation with applications to financial forecasting. Journal of Business & Economic Statistics, 40(3):1007– 1019, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1895812.

Yilmaz:2003:MPE

[Yil03] Kamil Yilmaz. Martingale property of exchange rates and central bank interventions. Journal of Business & Economic Statistics, 21(3):383–395, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619034.

Yitzhaki:1991:CJV

[Yit91] Shlomo Yitzhaki. Calculating jackknife variance estimators for parameters of the Gini method. Journal of Business & Economic Statistics, 9(2):235–239, 1991. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1991.10509849.

Yitzhaki:1996:ULR

[Yit96] Shlomo Yitzhaki. On using linear regressions in welfare economics. Journal of Business & Economic Statistics, 14(4): 478–486, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524677.

Yu:2014:C

[YL14] Philip L. H. Yu and Guodong Li. Comment. Journal of Business & Economic Statistics, 32(2):166–167, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.885436. See [GPI13b,

GPI13a].

Yang:2017:IHT

[YL17] Yaxing Yang and Shiqing Ling. Inference for heavy-tailed and multiple-threshold double autoregressive models. Journal of Business & Economic Statistics, 35(2):318–333, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1064433.

Yin:2019:FIC

[YLL19] Shou-Yung Yin, Chu-An Liu, and Chang-Ching Lin. Focused information criterion and model averaging for large panels with a multifactor error structure. Journal of Business & Economic Statistics, 39(1):54–68, 2019. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2019.1623044.

Yu:2017:GCA

[YLN17] Philip L. H. Yu, W. K. Li, and F. C. Ng. The generalized conditional autoregressive Wishart model for multivariate realized volatility. Journal of Business & Economic Statistics, 35(4):513–527, 2017. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1096788.

Yang:2020:UTD

[YLPC20] Bingduo Yang, Xiaohui Liu, Liang Peng, and Zongwu Cai. Unified tests for a dynamic predictive regression. Journal of Business & Economic Statistics, 39(3):684–699, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1714632.

Yu:2022:CFA

[YLYJ22] Xianshi Yu, Ting Li, Ningchen Ying, and Bing-Yi Jing. Collaborative filtering with awareness of social networks. Journal of Business & Economic Statistics, 40(4):1629–1641, 2022. CO-DEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07350015.
2021.1954527.

Yamauchi:2019:MSV

[YO19] Yuta Yamauchi and Yasuhiro Omori. Multivariate stochastic volatility model with realized volatilities and pairwise realized correlations. Journal of Business & Economic Statistics, 38 (4):839–855, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1602048.

Yogo:2008:APU

[Yog08] Motohiro Yogo. Asset prices under habit formation and referencedependent preferences. Journal of Business & Economic Statistics, 26(2):131–143, 2008. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000205.

Yogo:2009:C

[Yog09] Motohiro Yogo. Comment. Journal of Business & Economic Statistics, 27(3):326–328, 2009. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/jbes.2009.08275. See [KM09b, KM09a].

Young:1988:CNM

[You88] Kan H. Young. Comment on “A Neglected Method of Separating Demand and Supply in Time Series Regression,” by Stephen E. Haynes and Joe A. Stone. Journal of Business & Economic Statistics, 6(1):135–138, 1988. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1988.10509645.

Yasar:2008:FTT

[YP08] Mahmut Yasar and Catherine J. Morrison Paul. Foreign technology transfer and productivity. Journal of Business & Economic Statistics, 26(1):105–112, 2008. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1198/073500107000000197.

Yang:2018:IPP

[YQZ18] Yi Yang, Wei Qian, and Hui Zou. Insurance premium prediction via gradient tree-boosted Tweedie compound Poisson models. Journal of Business & Economic Statistics, 36(3):456–470, 2018. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2016.
1200981.

Yatchew:2003:EES

[YSD03] Adonis Yatchew, Yiguo Sun, and Catherine Deri. Efficient estimation of semiparametric equivalence scales with evidence from South Africa. Journal of Business & Economic Statistics, 21 (2):247–257, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288618936.

Yang:2007:CPM

[YT07] Z. L. Yang and Y. K. Tse. A corrected plug-in method for quantile interval construction through a transformed regression. Journal of Business & Economic Statistics, 25(3):356–376, 2007. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500106000000684.

Yu:2020:NEC

[YYX20] Xiufan Yu, Jiawei Yao, and Lingzhou Xue. Nonparametric estimation and conformal inference of the sufficient forecasting with a diverging number of factors. Journal of Business & Economic Statistics, 40(1):342–354, 2020. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/07350015.2020.1813589.

Yao:2019:SSC

[YZK19] Feng Yao, Fan Zhang, and Subal C. Kumbhakar. Semiparametric smooth coefficient stochastic frontier model with panel data. Journal of Business & Economic Statistics, 37(3):556–572, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1390467.

Zivot:1992:FEG

[ZA92] Eric Zivot and Donald W. K. Andrews. Further evidence on the great crash, the oil-price shock, and the unit-root hypothesis. Journal of Business & Economic Statistics, 10 (3):251–270, 1992. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1992.10509904.

Zivot:2002:FEG

[ZA02] Eric Zivot and Donald W. K. Andrews. Further evidence on the great crash, the oil-price shock, and the unit-root hypothesis. Journal of Business & Economic Statistics, 20 (1):25–44, 2002. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500102753410372.

Zarnowitz:1985:REM

[Zar85] Victor Zarnowitz. Rational expectations and macroeconomic forecasts. Journal of Business & Economic Statistics, 3(4): 293–311, 1985. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1985.10509462.

Zivot:2009:C

[ZC09] Eric Zivot and Saraswata Chaudhuri. Comment. Journal of Business & Economic Statistics, 27(3):328–331, 2009. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/jbes.2009.08305. See [KM09b, KM09a].

Zou:2017:EEI

[ZC17] Tao Zou and Song Xi Chen. Enhancing estimation for interest rate diffusion models with bond prices. Journal of Business & Economic Statistics, 35(3):486–498, 2017. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2015.1089773.

Zhang:2021:LHA

[ZCZ^{+}21] Jingfei Zhang, Biao Cai, Xuening Zhu, Hansheng Wang, Ganggang Xu, and Yongtao Guan. Learning human activity patterns using clustered point processes with active and inactive states. Journal of Business & Economic Statistics, 41(2):388– 398, 2021. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.2025065.

Zellner:1983:ES

[Zel83] Arnold Zellner. Editorial statement. Journal of Business & Economic Statistics, 1(1):1–4, 1983. ISSN 0735-0015 (print), 1537-

2707 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07350015.1983.10509314.

Zellner:1993:RMT

[Zel93] Arnold Zellner. Remarks on my term at JBES. Journal of Business & Economic Statistics, 11(4):425–426, 1993. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07350015.1993.10509977.

Zheng:2000:ION

[ZFSC00] Buhong Zheng, John P. Formby, W. James Smith, and Victor K. Chow. Inequality orderings, normalized stochastic dominance, and statistical inference. Journal of Business & Economic Statistics, 18(4):479–488, 2000. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.2000.10524886.

Zhang:2022:UFE

[ZG22] Fengqing Zhang and Jiangtao Gou. A unified framework for estimation in lognormal models. Journal of Business & Economic Statistics, 40(4):1583–1595, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.1952878.

Zellner:1992:CCC

[ZH92] Arnold Zellner and Chansik Hong. Comments on “Canonical Correlation in Multivariate Time Series Analysis With an Application to One-Year-Ahead and Multiyear- Ahead Macroeconomic Forecasting,” by P. W. Otter. Journal of Business & Economic Statistics, 10(4):470–471, 1992. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07350015.1992.10509924. See [Ott90].

Zhao:2019:CIB

[ZH19] Qingyuan Zhao and Trevor Hastie. Causal interpretations of black-box models. Journal of Business & Economic Statistics, 39(1):272–281, 2019. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2019.1624293.

Zhao:2015:ILA

[Zha15] Zhibiao Zhao. Inference for local autocorrelations in locally stationary models. Journal of Business & Economic Statistics, 33(2):296–306, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.948177.

Zhao:2020:DBP

[Zha20] Zifeng Zhao. Dynamic bivariate peak over threshold model for joint tail risk dynamics of financial markets. Journal of Business & Economic Statistics, 39(4):892–906, 2020. ISSN 0735-0015 (print),

1537-2707 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/07350015.2020.1737083.

Ziliak:2023:TEV

[ZHB23] James P. Ziliak, Charles Hokayem, and Christopher R. Bollinger. Trends in earnings volatility using linked administrative and survey data. Journal of Business & Economic Statistics, 41(1):12– 19, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2102023.

Zhang:2016:DNF

[ZHCB16] Junni L. Zhang, Wolfgang K. Härdle, Cathy Y. Chen, and Elisabeth Bommes. Distillation of news flow into analysis of stock reactions. Journal of Business & Economic Statistics, 34(4):547–563, 2016. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1110525.

Zhou:1996:HFD

[Zho96] Bin Zhou. High-frequency data and volatility in foreignexchange rates. Journal of Business & Economic Statistics, 14(1):45–52, 1996. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1996.10524628.

Ziliak:1997:EEP

[Zil97] James P. Ziliak. Efficient estimation with panel data when instruments are predetermined: an empirical comparison of momentcondition estimators. Journal of Business & Economic Statistics, 15(4):419–431, 1997. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1997.10524720.

Zhang:2005:IDG

[ZK05] Xibin Zhang and Maxwell L. King. Influence diagnostics in generalized autoregressive conditional heteroscedasticity processes. Journal of Business & Economic Statistics, 23(1): 118–129, 2005. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500104000000217.

Zhu:2022:DCC

[ZK22] Xiaojing Zhu and Eric D. Kolaczyk. Discussion of “Cocitation and Co-authorship Networks of Statisticians”. Journal of Business & Economic Statistics, 40(2):494–496, 2022. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2022.2044335. See [JJKL21].

Zhu:2015:NPT

[ZL15] Ke Zhu and Wai Keung Li. A new Pearson-type QMLE for conditionally heteroscedastic models. Journal of Business & Economic Statistics, 33(4):552–565, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.977446.

Zhang:2022:UPC

[ZL22] Haozhe Zhang and Yehua Li. Unified principal component analysis for sparse and dense functional data under spatial dependency. Journal of Business & Economic Statistics, 40(4):1523– 1537, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1938085.

Zhu:2023:TSC

[ZLLC23] Fukang Zhu, Mengya Liu, Shiqing Ling, and Zongwu Cai. Testing for structural change of predictive regression model to threshold predictive regression model. Journal of Business & Economic Statistics, 41(1):228–240, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.2008406.

Zhao:2019:SIP

[ZLS19] Shunan Zhao, Ruiqi Liu, and Zuofeng Shang. Statistical inference on panel data models: a kernel ridge regression method. Journal of Business & Economic Statistics, 39(1):325–337, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2019.
1660176.

Zheng:2020:SSS

[ZLWW20a] Zemin Zheng, Yang Li, Jie Wu, and Yuchen Wang. Sequential scaled sparse factor regression. Journal of Business & Economic Statistics, 40(2):595–604, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1844212.

Zhou:2020:AMS

[ZLWW20b] Jing Zhou, Jin Liu, Feifei Wang, and Hansheng Wang. Autoregressive model with spatial dependence and missing data. Journal of Business & Economic Statistics, 40(1):28–34, 2020. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://

amstat.tandfonline.com/doi/full/10.1080/07350015.2020.
1766471.

Zhu:2017:BAM

[ZLY17] Ke Zhu, Wai Keung Li, and Philip L. H. Yu. Buffered autoregressive models with conditional heteroscedasticity: an application to exchange rates. Journal of Business & Economic Statistics, 35(4):528–542, 2017. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2015.1123634.

Zhang:2004:DRR

[ZO04] Qiang Zhang and Masao Ogaki. Decreasing relative risk aversion, risk sharing, and the permanent income hypothesis. Journal of Business & Economic Statistics, 22(4):421–430, 2004. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1198/073500104000000406.

Zhu:2022:FSM

[ZPWW22] Xuening Zhu, Rui Pan, Shuyuan Wu, and Hansheng Wang. Feature screening for massive data analysis by subsampling. Journal of Business & Economic Statistics, 40(4):1892–1903, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2021.1990771.

Zheng:2014:RSV

[ZS14] Tingguo Zheng and Tao Song. A realized stochastic volatility model with Box–Cox transformation. Journal of Business & Economic Statistics, 32(4):593–605, 2014. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2014.918544.

Zhang:2015:AMP

[ZS15] Yongli Zhang and Xiaotong Shen. Adaptive modeling procedure selection by data perturbation. Journal of Business & Economic Statistics, 33(4):541–551, 2015. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2014.965307.

Zhao:2020:MMT

[ZSZ20] Zifeng Zhao, Peng Shi, and Zhengjun Zhang. Modeling multivariate time series with copula-linked univariate d-vines. Journal of Business & Economic Statistics, 40(2):690–704, 2020. CODEN RFSCFJ. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1859381.

Zimmer:2006:UTC

[ZT06] David M. Zimmer and Pravin K. Trivedi. Using trivariate copulas to model sample selection and treatment effects. Journal of Business & Economic Statistics, 24(1):63–76, 2006. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1198/073500105000000153.

Zhou:2015:ESA

[ZTCW15] Jing Zhou, Yundong Tu, Yuxin Chen, and Hansheng Wang. Estimating spatial autocorrelation with sampled network data. Journal of Business & Economic Statistics, 35(1):130–138, 2015. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2015.
1061437.

Zuehlke:2003:BCD

[Zue03] Thomas W. Zuehlke. Business cycle duration dependence reconsidered. Journal of Business & Economic Statistics, 21 (4):564–569, 2003. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1198/
073500103288619241.

Zaslavsky:1993:TSM

[ZW93] Alan M. Zaslavsky and Glenn S. Wolfgang. Triple-system modeling of census, post-enumeration survey, and administrativelist data. Journal of Business & Economic Statistics, 11 (3):279–288, 1993. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07350015.1993.10509955.

Zhang:2022:NEM

[ZWL22] Dixin Zhang, Yulin Wang, and Hua Liang. A novel estimation method in generalized single index models. Journal of Business & Economic Statistics, 41(2):399–413, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2027777.

Zhang:2022:NQR

[ZWLL22] Xiaoyu Zhang, Di Wang, Heng Lian, and Guodong Li. Nonparametric quantile regression for homogeneity pursuit in panel data models. Journal of Business & Economic Statistics, 41(4):1238– 1250, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/07350015.2022.2118125.

Zhang:2012:FIC

[ZWZ12] Xinyu Zhang, Alan T. K. Wan, and Sherry Z. Zhou. Focused information criteria, model selection, and model averaging in a Tobit model with a nonzero threshold. Journal of Business & Economic Statistics, 30(1):132–142, 2012. ISSN 0735-0015 (print), 1537-2707

(electronic). URL http://www.tandfonline.com/doi/full/10.
1198/jbes.2011.10075.

Zhong:2022:EIM

[ZWZ22] Wei Zhong, Chuang Wan, and Wenyang Zhang. Estimation and inference for multi-kink quantile regression. Journal of Business & Economic Statistics, 40(3):1123–1139, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2021.
1901720.

Zhu:2022:SFM

[ZWZL22] Rong Zhu, Haiying Wang, Xinyu Zhang, and Hua Liang. A scalable frequentist model averaging method. Journal of Business & Economic Statistics, 41(4):1228–1237, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2022.
2116442.

Zhu:2023:KAE

[ZZWZ23] Rong Zhu, Xinyu Zhang, Alan T. K. Wan, and Guohua Zou. Kernel averaging estimators. Journal of Business & Economic Statistics, 41(1):157–169, 2023. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2021.2006668.

Zhou:2020:PAC

[ZZZ20] Yeqing Zhou, Yaowu Zhang, and Liping Zhu. A projective approach to conditional independence test for dependent processes. Journal of Business & Economic Statistics, 40 (1):398–407, 2020. ISSN 0735-0015 (print), 1537-2707 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/07350015.2020.1826952.

Zhu:2022:STR

[ZZZ^{+}22] Xuehu Zhu, Qiming Zhang, Lixing Zhu, Jun Zhang, and Luoyao Yu. Specification testing of regression models with mixed discrete and continuous predictors. Journal of Business & Economic Statistics, 41(4):1101–1115, 2022. CODEN ???? ISSN 0735-0015 (print), 1537-2707 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07350015.2022.2110879.

Zhang:2019:GFT

[ZZZS19] Shulin Zhang, Qian M. Zhou, Dongming Zhu, and Peter X.-K. Song. Goodness-of-fit test in multivariate jump diffusion models. Journal of Business & Economic Statistics, 37(2):275–287, 2019. ISSN 0735-0015 (print), 1537-2707 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07350015.2017.
1321547.